A Bibliography of Publications in IEEE Software: 1984–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 2.02

Title word cross-reference

4 + 1 [Kru95]. ^{3} [Har87].

-ilities [Voa04b].

.NET [LL03, NV02].

02 [DP03]. 05 [AW06].

1 [BRS^{+}85, ES09]. 1.0 [San98e]. 12207 [FS98]. 130J [Con02]. 18 [Mat96]. 1986 [Har03c]. 1990 [Har03c]. 1990s [ME90]. 1994 [Har03c]. 1998 [Har03c]. 1st [Lew98].

2 [BHHM86, Gla09c, Gut86, LRO02, MPT86, Rov86, SM09b]. 2.0 [BK03, OHM^{+}07]. 2/MUPE [MPT86]. 2000 [Ano99d, Tho98]. 2002 [Har03c]. 20th [Har03c, You07]. 25th [Ano09a, GAO09]. 25th-Anniversary [Ano09a, GAO09].

3 [DNS^{+}06, Mas98]. 3D [DW90].

5 [Nus97]. 500-Language [LV01]. 5th [Tau95].

6.0 [Fer01b].

830 [Pos85].

9 [Kor96]. 90 [Rys95]. 9001 [Col94, Pau95]. 9126 [JKC04]. ’96 [Ano97h]. ’97 [Cha97d]. ’98 [Ano98o]. 982.1 [Sch07].

AADL [DDMS86]. AADL/S* [DDMS86]. AAS [BLP96]. Abbot [RP07]. Abstract [Cam88]. Abstraction [HPLH90, OCHW91, Ris07, Sha84, Spi07a, FKKM89, MMS88]. Abstractions [WB06d, Nav95]. Abundance [Boo09e, Boo09f]. Abuse [Ano93-39, Dav98a, Hum01]. Abusing [Spi08c]. Academia [Ano93l, BCKM97, CT90, McC97f, ZJ00]. Academic [Gla03a]. Accelerated [CB00]. Accelerating [LG03, Pre01]. Acceptable [Car88, Jon05]. Acceptance [MM05]. Access [Fow03a, JW09, KGRB03, SBK86]. Access-Oriented [SBK86]. Accidental [Boo06a]. Accommodate [Tho05]. according [Pre95]. Accountability [Par96b]. Accounting [Koc00, Lai08]. Accoutrements [Boo09c]. Accreditable [FTB06]. Accreditation [Eng99]. Accredited [DJSS97]. Accurate [BL92, Fai02]. Achieve [Byt99, KMY91, NO91]. Achievements [BV08]. Achieving [Abe07, Das94, GV94, HKPS96, McC97a, PDHT97, YD01]. Achilles [Voa99e]. Acquiescence [Erd08d]. Acquire [Mai09b]. Acquiring [MN98]. Acquisition [EM93, LG03, SJM02, US04]. Across [Bai98]. Act [Gra99a, WJS97]. Action [McC96f, Ada95]. Active [WSG^{+}05]. Active-Architecture [WSG^{+}05]. Ada [Ano96f, BR85, Bas87a, BFC00, Eme95, EMB^{+}99, GP87, HL85, LCS91, Lv85, OCM^{+}84, Ros85, Wit90, WCW85]. Ada-Based [WCW85]. Adaptability [FHS^{+}06]. Adaptable [OLPS04]. Adaptation [SDML^{+}06]. Adapting [FTR^{+}94, KKH95, Mel05]. Adaptive [ACM^{+}07, Lit05, MSK00]. Adding [Fon96, MMS88]. Addressing [CP02]. Adept [CTC07]. Adopt [VVM08]. Adopted [Rif01a]. Adopting [GE04, vvSO^{+}98]. Adoption [BG98, CK02, Gar05, Kem92, Ksh04, SV02, WW01]. Ads [Ano00l]. Adults [Lis97, SPZ08]. Advanced [CCRZ^{+}90]. Advancement [Spi06d]. Advances [PWS^{+}95, PC97]. Advancing [KCL93]. Advice [WB07d]. Advisory [Ano01i]. Aero [DHMR02]. Aerospace [Dor99b]. Affect [NV02]. Affecting [SS04, ZME00]. Affects [Got99b, Kem92, Tom87]. Affordably [BTC^{+}99]. Afraid [Rob01a]. After [Bro95, HDS05, McC99a, MC03b, Mil84, Moo02a, Wil99a, Pfl08]. After-the-Fact [HDS05]. Again [Dav96b, San98c, MMM^{+}01]. Against [Voa99e, And03]. Age [Ano93n, Boo94, CSR^{+}99, CS09, Erd07a, Mic99, SC06, LSB02]. Agent [Mub08]. Agent-Oriented [Mub08]. Agents [Gon02]. Ages [VD96]. Agile [Ano09b, Aoy98, BT05, CR08, CSSP05, DD09, Got09a, KR05, Mel05, NT06, Orr04, Rei02b, RME03, SA05, Sel09b, SMHT09, SC03, Tho05, UA03, WB09c, KHA^{+}03, Erd07a, Mug08, THDK06]. Agility [Amb02, Erd09a, Lit05]. Aging [BC86]. Agreement [Ano97d, RS01, Zan92]. Agrees [Sim95a]. Ahead [Ano93x, Got01e, Fis89]. AI [Ano85]. Aid [Mou01]. Aided [Ano92i, LK88, Luq92b, BD94, Luq92a]. Aids [HM02b, KM86]. Ailing [Mcl98]. Aim [Bec01]. Aims [Wal96]. Ain’t [DeM99]. Air [Got00, HJA08, JC00, TB97]. Aircraft [HSW91]. Ajax [SA07]. Al [Har03c]. Alan [Pat08d]. Alarm [LLH^{+}03]. alert [CGR94a]. Algorithm [BFJT05, BS85, DSS85]. Algorithms [Ano96t, NT92, SFG94, WOS94, For88]. Algres [CCRZ^{+}90]. Align [CC95a]. Aligning [BB99a, US09]. Alignment [vWHvW09]. Alleviating [CA01]. Allocation [DBL04, SJ86]. Allow [Mye89b]. Allows [Dak95c]. Alpha [Con96]. Alternative [Hor87, Pem87, SS06a]. Alternatives [MCL^{+}01]. Amaze [BC85]. Ambiguous [Pat08a]. Amended [Tom87].

Amendment [Bar97b]. Analogy [DDS^{+}97]. Analyses [KJC07]. Analysis [Ano93g, AHM^{+}08, Bas87a, Boe00, CH93, Car98, CE09, CG88, ELLM99, EL02, FHRK93, FKLW91, FLKMR91, GFP09, GdHN^{+}08, Gra90, HN90, Hor96, HSG^{+}94, JL06, Jun98, JML00, Kai99, Lou06b, MR90, MCL^{+}01, OBM^{+}90, PTA94, PTH^{+}09, vMC92, Deu88, EJW95, KABC96, MMN^{+}89, War89, Hou01]. Analyst [Mai09a]. Analyste [Gra03]. Analysts [Rob02]. Analytics [Sel09a]. Analytics-Driven [Sel09a]. Analyzers [Dya95]. Analyzing [Ano92a, BCH^{+}01, FCB00, GA06, KKJ91, LX99, Map85, PSC02, TP98, dSA95]. Anarchy [Gla04a]. Anatomy [BLP96, Ber94]. Anchoring [Boe96]. Ancient [Gla08b]. Angela [Dav96e]. Animation [BS85]. Anna [Lv85]. Anniversary [Ano09a, GAO09, Har03c, Wei99, You07]. Annual [Ano95a, Ano96a, Ano98a, Ano99o, Ano00a, Ano01a, Ano02a, Ano03a, Ano04b, Ano05a, Ano06a, Ano08a, Ano09c]. Annualized [McC97b]. Answer [CL00, Gla00a, Zve98b]. Answers [Rob09a]. Ant [SC04]. Antarctica [Ana99]. Antidotes [FW03]. Anup [GAS^{+}01]. Any [Kla95]. Anyone [Sel09b]. Anyway [Erd07f, Har03f, Shu07a]. Apache [DGC^{+}07]. APIs [MRTS98, Rob09a]. APL [TP86]. Apple [Mil84]. Apples [Per95]. Applet [Kin97]. Applet-Browser [Kin97]. Application [Ama07, Bal00, BS00, BHKR01, BKW09, DFL00, GV95, Gla00c, GB00, HM02a, HKN85, JS95, KT05, Les06, SRK^{+}09, SM97, Vis08, Vo90, Cle88, GMS^{+}96, Gue94]. Application-domain [GV95]. Application-Independent [SM97]. Application-Specific [BS00]. Applications [Alo03, Bar97b, BRS^{+}85, BS95, BFP95, Bol99, CL91, CCRZ^{+}90, Clo01, CL02, CAGR91, DLR07, Fer01b, FF08, Ger90, Kod04, LM07, LSG^{+}05, LL92a, MBL08, Mar04, Mar96, ME00, Off02, OHM^{+}07, PHRS07, Pou95, PAD^{+}98, RD96, Sch92, SC04, SFVFM08, Sta94, SH93, Wel00, YBBP05, ZLJW95, HSA89, KLR89, Sch96a]. Applicative [OC88]. Applied [Fei07, MF00]. Apply [Ben00]. Applying [Ano00k, CN95a, ELM90, Etl96a, GFB96, Har89a, JB00, Jul93, KMSW00, Les06, LN92, Mor00, Pal01b, Pfl97c, SK92]. Appreciation [Bri96]. Apprenticeship [SHM02]. Approach [ADS91, Ano93e, Ano93n, Ano93-30, Ano96p, Ano96q, CT93, CCD09, Cha97c, CDPP96, CP06, DDMS86, DS07, DFL00, Dut93, ELST95, FGSK03, Gil96b, Gli08, Got00, Gut86, Hor98a, HSG^{+}94, HB06, KR97, Kis96, KYM^{+}04, LM84, LX99, MH07, MW01, MSK00, MS99, Mye95d, NGBR05, RW90, Rog00, Sch07, She91, Sim04, SMHT09, Sto99, TP98, WSG^{+}05, WMH07, ZJ98, dLdCGR06, AO89, BAS89, RHK^{+}03, Zav89]. Approaches [CA01, HG00, Mel05, Veg86]. Approves [MHC03]. Apps [SA07]. Arachne [SDML^{+}06]. ARC [Pau85]. Archaeology [HT02a]. Architect [Boo08c, Bus09a, Fow03d]. Architecting [BK03]. Architects [Par05]. Architectural [Boo08a, Boo08b, Boo09c, BDHJ96, Cop97, Cop99, GAO95, GAO09, HAZ07, MKMG97, Sha95b, Sta06a, dLdCGR06, KC97]. Architecture [Ano93f, Ano07j, BS86, BH01, BSGF98, Boa95, BALV03, Boo06d, Boo06a, Boo07b, Boo07c, Boo07f, Boo08e, Boo09b, Boo09c, Boo09a, Boo09d, CS09, Dig98, Erd09a, FHS^{+}06, FO85, Fuj99, GC85, GPW97, KB02, KS96, KOS06, KCD09, LCM06, LSB02, MRZ^{+}05, NT06, NCC^{+}02, Pfl97b, RP06, San03, SC06, SM97, Sta06a, THV09, TA05, UA03, WSG^{+}05, KABC96, Kru95, Nic92, MBO93]. Architecture-Centric [NT06]. Architecture-First [Boo07b]. Architectures [Ano93p, DSS85, FNO92, HG99, MKHV06, MM05, Med05, MJ97, RKJ^{+}97, vM95a, Tic89]. Archives [GHH^{+}09, NZZ09]. Area [LC85]. Aren’t [JLLP03]. Argument [Shu07b]. Ariane [Nus97]. Arjuna [SDP91]. Army [Fen90, Nis00]. Art [AD03, Dav96d, DDS^{+}97, EMB^{+}99, HMP^{+}01, Hol92, HT03c, JL03, JM07, Mat00b, McC98c, McC01b, Pfl05, Rei03a, RS05, Spi05c, BPS^{+}03]. Article [Ano95b, Ano96b, Ano97a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99a, Ano99b, Ano99c, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano01b, Ano01c, Ano01d, Ano01e, Ano01j, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07a, Ano07b, Ano07c, Ano07d, Big96, Dav96a, Eve95, Fow02c, McC98f, McC02c]. Articles [Ano98m, Ano98n, Ano99h, Ano99e, Ano01g, Ano02i, Ano03h, Ano03i, Ano03j, Hor97, O’L09]. Artifacts [Boo07a]. Artificial [STSK92]. Artistry [Boa95]. ASDL [KS89]. Asia [GAS^{+}01, IS94]. Asimov [Fei07]. Ask [Mcs96, Nie97d, Rai07a]. Asking [GSB03]. Aspect [AK03a, BCC^{+}06, MS06, VV00]. Aspect-Oriented [AK03a, BCC^{+}06, MS06, VV00]. AspectJ [Les06]. Aspects [Ano99f, BCA^{+}06, TM04, dSSS^{+}09]. ASPIS [PTP^{+}88]. Assembling [Nau98]. Assembly [Fow04, MKHV06]. assertions [Voa97a]. Assess [Mos92, Moy97, vWAS06]. Assessing [CMH92, Cuk05b, HM92, Lin85, MDM03]. Assessment [Ano93b, Bau94, BR87, CTC07, CMH92, DAEE08, Dou85, GPW97, HMK^{+}94, Kän97, KN96, Mad97, WS00, Wil94, WMH07, KKL97]. Asset [Spi06e]. Assets [GG08, MBE05, Sal03, BMM05]. Assistance [KFP88, RRD07]. Association [DG09b]. Assurance [Frü99, HK94, KM02, MRA98, PL03, Voa99a, Voa99e, Voa03]. assure [Sie92]. Assuring [LH89, Voa03]. Asynchronous [Bur94, SFG94]. ATC [Hal96b]. ATM [JWKZ99]. Atomic [ADT91]. ATP [Bar95b]. Attack [Hou95, Ano97e]. Attacking [MM00]. Attaining [MD02]. Attitude [WB09c]. Attribute [Far84, OBSN08]. Attributes [Bøe06, BVB04, EG04a, NV02, Off02, OLPS04, URK01]. Audio [Ama07, Rei96a]. Author [Ano97a]. Authors [Erd07e]. Automated [EG04a, KSTM98, Lav00, MM05, McL06, SCCA04, WC86, Yac03]. Automatic [BFJT05, CDPP96, Gon02]. Automatically [RW04]. Automating [BCMV05, CC93, GdHN^{+}08, Ko09, SC04, Ste95a, VVS^{+}08]. Automation [HT02c, JRS^{+}09]. Automotive [TH02a, WW03]. Autosar [Hon09]. Available [NOK08, MCA00, RM00]. Avenues [HPG^{+}00]. Average [Ano93-31]. Aviation [HH02a]. Avoid [Gra99a, Jør09, Tur03]. avoidance [CGR94a]. Avoiding [Fow01a, JG08, McC96b, Pre01, Rai07b, Tho97]. Award [Big96, Dav96a]. Awards [Bal00, GAS^{+}01]. Aware [KH02]. Awareness [Row96]. Away [Ano96h, Wil99a]. Axiomatic [DWH86]. AXP [Con96].

B [Ano93m]. Baby [Ben00]. Bachelor [FTB06]. Back [Ano93x, Ano98o, Bar96a, Ben99, Boo08e, Har06d, PT09, VD96]. Backlash [You00]. Backtracking [ADS91]. Bacteriologic [BFJT05]. Bad [And98b, Gla97, Gla01a, HM02b, JOL^{+}00, Mac96, NCC^{+}02, Rei00b, Wil99b]. Badly [VCM^{+}97]. Bag [Ano92o]. Balance [BCH08, Dio93, Roy05, TB02]. Balanced [Mai02]. balances [Sie92]. Balancing [Ano97g, BB08, KHA^{+}03]. Ball [Ano93w]. Band [HM02b, Mou01]. Band-Aid [Mou01]. Band-Aids [HM02b]. Bandwagon [CCS^{+}03]. Bang [Ano08e, Boo09b]. Bank [Lew84]. Banking [NS99]. Barrier [CK02]. Barriers [Bel99, MDD09, TTT^{+}02]. Base [Ano93d, Che84, DFLS05, Sym88, SSGT86]. Based [AHZ02, Amb02, AO88, Ano93s, Aoy98, Bas87c, Bas07, BV08, BB08, BOS00, BS03, CSSP05, CSHA07, CR92, Dak96c, DRR^{+}07, Den91, DSD02, DW90, DKJ05, Ebe01, EFM^{+}91, FGSK03, GC85, Gli08, GB00, Got09b, HB06, JL06, Jor05, Jør09, KKM87, LANC07, LCB^{+}09, Lub91, MMSM09, MH06, NSV^{+}08, New95, Ols85, OYE06, PSC02, PS90, PTA94, PAD^{+}98, Pue97, PTP^{+}88, RFHM86, RBBC03, Sev87, SMR93, SS08, SRS^{+}86, TM97, Ter96, TP98, TM04, TB02, VNS94, Voa98a, Wey98, WSD^{+}09, WMH07, YBBP05, ZJ98, vvSO^{+}98, ABB^{+}93, ARW96, HN90, Kar88, KABC96, PR94, SSGT86, WCW85]. Bases [SCCA04]. Basic [Spi09a, Bol00, Nau98]. Basics [FJWC01, VD96]. Basili [Bur95a, SSZ06]. Basis [Bur95a]. Battle [Ban93, Cha97f]. Battling [You00]. Bazaar [DDPW09, Wes08]. Be [BPS^{+}03, Car97, Cur01, Fow01d, Jac98c, Lon02, MGR04, NCC^{+}02, PAD^{+}98, Rei95d]. Beating [Gla02e]. Beautiful [Iel96, WB07b]. Beauty [Fuj99]. Because [Boo07d, Har05a]. Becomes [Rys95]. Beethoven [Kla97]. Before [AW03, Fea04]. Begged [Zve98b]. Beginning [Gla05d]. Behave [VCM^{+}97, Voa01c]. Behavior [AHM86, And03, HPLH90, Kru05a, OCM^{+}84]. Behind [Ano93-37, Erd09c, SGM06]. Being [CK02, Iel96, Lar01a]. Believable [DFS93]. Believe [And95]. Bell [DKKK90, DKMS91, JL03]. belongs [Lou95]. Belt [Bud99]. Benchmarking [Ano01g, CZ01, FLN01, Got01e, Hei01, LWHS01, MF00, Max01, SM01, TS01]. Benchmarks [JLLP03]. Beneficial [GA06]. Benefit [Ebe09]. Benefits [Ano92a, MBO93, ME00, SSV99]. Best [Big96, BTC^{+}99, Dav96a, DLV99, Har04a, Ker08, McC96b, McC96c, McC96d, McC96e, McC97b, McC97a, McC97c, McC97d, McC97e, McC98c, McC98a, McC98b, McC98d, McC00d, Pos95, RBBP03, Rog00, Ros06, RE03, Tur03, US09, WB06f, You95]. Better [Ale06, Ber96b, BTC^{+}99, Ded09, DAS^{+}07, Etl96b, GD02, HM02b, Knu87, KC96, MH06, Nes98, Nie97b, VBM^{+}02, Voa01b]. Better-Quality [MH06]. Between [Dav97a, EMW^{+}97, KE88, CT90, Hoh03, MKCC03, McK02, San08, TB02, ZJ00]. Beware [Pal96]. Beyond [Acl88, Ano93a, Ano93q, AD03, Bil94, Erd07b, EAS94, HG99, KT05, Kic96, Mac98, MW96, Nak96, Tiw04, EL04, Got01e, Jur97]. Bias [GRS02]. Bidirectional [NT92]. Bids [CC95c, Jør09]. Big [Ano08e, Dav95c, Gla03a, RS08]. Bill [Cha99c]. Bind [SB95a]. Biological [Kru08a]. Bit [Rei02c]. bits [Cox97]. Black [Ano96d, Cla98a, Kic96, Tiw04]. Blacksmiths [Dav95d]. Blame [CC95c, DDS^{+}97, Gue95, Hei98]. blaming [MW96]. Blend [SLG^{+}00]. Blending [Bud91, MK03]. Blends [Fuj99]. Blessings [Kap00]. Blinn [GF99]. Block [Knu87]. Block-Structured [Knu87]. Blocks [KG87, LSW87, vM95a]. Blog [Gla07c]. Bloun [Dak96e]. Blown [DDS^{+}97]. Blue [KIB^{+}08]. Blueprint [CC96]. Blueprints [Aae03, Sta06a]. Blum [Jur97]. Blur [Ano97g]. Board [Dav96a, Ano96i, Ano01i, Big96]. Boards [Ano01i]. Body [BDA^{+}99, DB99, EMB^{+}99, JW09]. Bolts [Ano99l, Ano99r]. Book [Bei97b, Bro95, Hor96, OC90a, PMD07, WJS97, Jac95]. Books [BALV03]. Bookshelf [Ano96d, Ano96c, Ano97b, Ano98g, Ano98h, Ano99d, Ano01f, Ano02h, Ano06h, Ano06i, Ano06j, App06, Bei97b, BMGT01, BS04, BJ06, BALV03, BBJ04, BdFD07, BSFS07, Can95, Cha96a, Coo95, Coo96, DSBM99, DW00, DFL00, Dug98, DSB99, Ebe01, EM04, EA07, Etl96a, Etl96b, Fer95, Fer01b, Frü99, GS07, Gou97, Gou98, GF99, Gre96, HKPS96, Hor96, Hor98a, Hou01, JL03, JML04, Jur97, Kal94, Kus96, Lar01b, LSB02, LM03, Lew98, LRO02, Mas98, Mat00b, MRA98, MLL03, MSK00, MF04, Nav95, Nus96, Pal01a, Par96a, PWS^{+}95, Pfl97b, Pou99, Rog00, Ros98b, Rys95, Sal97a, Sal97b, STBC05, SC00, SDE^{+}05, SMP07, SJM02, Sta06b, VBM^{+}02, WJS97, WJSP05, Wil96, Wil97c, Wil97a, Wil97b, Ada95]. Boomers [Ben00]. Boone [Mil01b]. Bootstrap [Ano93b, HMK^{+}94]. Borders [SHZ^{+}98]. Both [Per95, KHA^{+}03]. Bots [Gra00b]. Bottom [Don01, Gla01a, Jak98, TM94, Wei98a]. Bottom-Up [TM94, Wei98a, Jak98]. Box [Ano96d, Gla03a, Kic96, Sha95a, Tiw04, MLL03]. BPEL [Lou08]. Brass [Rei04a]. Brazil [Dua02]. Break [Bel99, Mac99b]. Breakdown [DDS^{+}97]. Breaking [Bol02]. Breathing [Lot97]. Brick [Dro06]. Bridges [DW00]. Bridging [BHS85, KE88]. Brief [Ano92j, Ano96n, Ano97e, Bar95c, Bur94, Cha97a, Cha97d, Ebe08b]. Briefs [Ano96e, BLP96, Cha99c, Gla99c]. Bright [Ano96r]. Bringing [Ber96a, Bri96]. British [Bar98]. Broad [FCH^{+}08, Pou95]. Broadcast [LC85]. Broke [DeM99]. Broker [Har06f]. Brook [JOL^{+}00, McC99b, SLG^{+}00]. Brooklyn [DM93]. Browser [Kin97, PCS95]. Browsers [Jaa02]. Browsing [SSSW86]. Bruce [Har03c, Jur97]. BSA [Bar97a]. Budding [Ju98, Ju01]. Budget [Ano92i]. Bug [SC05, Spi06a, Tho98]. Bugs [Ano97g, AHM^{+}08, KAS93, Ano96d]. Bugzilla [SC05]. Build [CM96, DDS^{+}97, LRO02, Mag00, McC96c, OBR^{+}08, Rov86, Vo90]. builder [CFTZ89]. Builders [Spi08b]. Building [Ano93e, Ano01h, Bac95b, Cle88, CM07, CB00, DLR07, GHW02, Har06a, HM02b, KG87, KKM^{+}09, KL05, LSW87, Lev90, McC98e, MRTS98, Pfl93, PS84, SC04, VBM^{+}02, WSL^{+}99, vM95a, FKU97]. Built [Mat00b]. Buley [Pat09]. Bullet [Ano93s, Ber08, Dro06, FM08, Moo01c]. Bulletproof [VBM^{+}02]. Bullets [Spi07e]. Bumping [PM98]. Burgess [Dav96e]. Bus [MM05]. Business [Ano98o, Ano06k, Cha98a, Chi02, Dig98, Fav02, GK02, Har05f, Hec99, Hou01, Lei99, LSG^{+}05, Mat00b, McC02e, ME02, MLL03, MK03, NT02, New95, Pat08a, Ram00, Rob04, RP09, SRK^{+}09, Vic07, Voa04a, WEL02, Wes08, ZJ00, Sim96]. Business-Critical [Ram00]. Business-Driven [NT02]. Business-Process [New95]. Busters [Spi06a]. Buyers [Ano92g, Ano93c]. Bzip2 [PJT09].

C [Fer01b, Wil97c, Ano96j, Ano99d, Bec90, CG88, Con02, DSBM99, DSB99, EFM^{+}91, Oma89, Sho06, WC06, Wil97c]. C# [LL03]. C-130J [Con02]. C-Based [EFM^{+}91]. C/C [WC06]. Cache [SDML^{+}06, TM97, San98e]. Café [vdL02]. Calculating [BCM^{+}04]. Calendar [Cur95]. Call [Ano98i, Ano98j, Ano98m, Ano99e, Ano99g, Ano99f, Ano00i, Ano00j, Ano00h, Ano01g, Ano01h, Ano03h, Ano03i, Ano03j, Ano03k, Gra97, Hor97, KH97, Spe99]. Calls [Ano98n, Ano99h, Ano00k, Ano02i, Got02a]. Calmer [Rei94b]. Cambridge [San95a]. Campus [Hol04]. Can [Ano92b, Ano92c, Ano92h, Bin97, Boo08d, Ebe09, EG04a, EA03, Gau05, GK02, Har05a, Hay03, KC96, Lai08, Lon02, MSG^{+}07, McC97f, Pfl02, PAD^{+}98, Sim04, VV00, VMM91, VCM^{+}97, Voa97b, Voa00b, WES87, Voa97a]. Canadian [Str02]. Can’t [Gil96a, Rob99, Spe99, Yam98b]. Capabilities [AJM06, KGRB03]. Capability [BM91, BM09, HC91, PCCW93, Rug93]. Capers [Cha97d]. Capital [DL98, Weg84]. Capital-Intensive [Weg84]. Capitalizing [Dak95b]. CAPS [Luq92a]. Capstone [UHC02]. Capture [HAZ07, Lit04, ZSF91]. Capturing [BDA04, KSMS02, SSV99, Zdu09]. Card [Mai09b, URK01]. Cards [Dav97e]. Care [JB04]. Career [Hal09, Kru08b, Kru09a, Kru09b, Nau98, WOBP00, Wil09]. Careers [Cos03]. Careful [Daw95a]. Cares [McC96e]. Cargo [McC00a]. Carl [Har03c]. Cascade [ZF99]. CASE [Ano92i, Bas07, BFJT05, BBRTL90, CK00, Che01, CR88, CGR94b, CGR94a, CSR^{+}08, DLSZ05, DFL00, ELM90, ELLM99, GCR94a, GCR94b, Gue94, HG00, Har05b, Hei01, JD98, Kan01, KCF^{+}08, KAKG96, KPP95, Lei99, LJ88, Mey87, PJT09, Pau85, PWS^{+}95, Rob04, SRK^{+}09, Sne84, THV09, Tar09, WKCJ00, Win05, ZAV02, vdLLM09, Den91, LD95, WTMS95, Acl88, Big88, CN92, CC95b, Cur88, GLS99, HBP92, Jar92, Kem92, Mar88, MS92, NC92, OBM^{+}90, PTP^{+}88]. Cases [Ale03, JB04, LX99, TRW03, Yam98a]. Cash [Ado96]. Casting [Kru05a]. Catch [Erd07d, Ing08]. Catching [Rei04a]. Cathedral [Wes08]. Causal [Ano93g, Car98, FKN02]. Cause [Pes97]. Cautionary [JD98, Pfl96, Wey98]. CBSE [BW98]. CD [Rei95e]. CD-ROM [Rei95e]. Cedar [Emr85, Tei84]. Celebrating [You07]. Center [Cha98b, GB00, Got01c, Lie02, PM98]. Centered [Clo01, CL02, MSJ^{+}09, Rog00]. Centers [Mar85]. Centric [NZ07, NT06, Pat08d, Pat09, PHRS07, VZM^{+}07]. Centricity [Pat07]. Cents [Ano93-34]. Century [Ano98p, Lew98]. CEO [And02]. Certificate [Cha96c]. Certification [Ano98-27, Col94, CNL99, Mat00a, McC99f, NS05, RD99, Voa99b, WKW99]. Certifiers [Got99a]. Certifying [Bøe06, Got99a, PMM93, Voa99a]. Challenge [Ano99d, BC08, Cha97d, FPG94, Lev90, LO90, Mar88, Sha96, Lew98]. challenged [Dak96f]. Challenges [Ano99d, Ath98, BGM06, BT05, BV08, BCH^{+}02, Irv97, WW03, Fis89]. Challenging [DFL00, Sid94]. Chameleon [MBO93]. Champions [Gou98]. Chang [Har03c]. Change [BMM04, BMM05, Big96, Cha97b, Cox84, DC04, GFP09, MNA05, Mil01a, Pfl98, RG04, Tho05, WB08d, For97, MSK00]. Changed [Lus04]. ChangeDistiller [GFP09]. Changes [CCD09, Cos03, Cus03, HH02b, KL97a, Koh05]. Changing [Cha94, Coy00a, McC02b]. Chaos [DFL00, Par03]. Chaotic [Voa00b]. Characterizing [Hum88, WB06b]. Charlotte [ACF87]. charts [AMM^{+}89]. Charybdis [Dav97a]. Chasm [Kel07]. Cheap [And94, Ano98h]. Cheaper [Voa01b]. Check [Rei94a]. Checklist [Ano93c]. Checklists [Hat08]. checks [Sie92]. CHI [Cur95]. Chime [Rei96b]. Chimera [Dro96]. China [BHAW09, GA05, Ju98, Ju01, Sim95a, Zho89]. Chinese [Ksh05]. Chip [MW09]. Choice [Bri97, Voa98b]. Choose [Lov94]. Choosing [GN95, Spi06b, VB97]. Christopher [Mey96b]. Chunking [MW01]. Cisco [Gar00]. Cited [O’L09]. Civil [Joh95]. Civilizing [Gou97]. Claims [GB00]. CLAM [Ama07]. Clarity [Fea04]. Clash [Per99]. Class [PM98, WJS97, Wil97c]. Classes [GBdlHQCG^{+}00, WB06b, WB07a]. Classic [McC96b]. Classification [Gla08b, Gla09c, PS90, Sha97b]. Classifications [KS96]. Classified [Ano00l]. Classifying [PDF87, TM97]. Classroom [UHC02]. Clean [Voa97b]. Cleanroom [Ano93e, Bei97a, BTC^{+}99, Boy97, Bri97, DB97, HBP92, Lin94, Osh98, TB97, WKA96]. Clearer [Etl96b]. Clearing [PR08, TB97, WOBP00]. Client [Ano93p, Frü99, Pou99, Sch04, Ano97i]. Client/Server [Ano93p, Frü99, Pou99]. Climbing [Dro06]. Clinical [MMM^{+}09]. Closed [Lar01a]. Closer [Rei03a]. Closing [BCKM97, McC02a]. Closings [Dav97c]. Clothes [Spi98]. Cloud [Erd09c]. Cloudy [Rei95f]. Clueless [Har04b]. Clutter [BNST99]. CM [Sch95a]. Cm* [SJ86]. CM-Tool [Sch95a]. CMM [Car00a, Eic03, FS98, GA06, GE04, HPG^{+}00, JB00, MD02, MK03, Pau95, Pau01, Rei03b, TS01]. CMMI [vWHvW09]. CMMI-DEV [vWHvW09]. Coaches [BTC^{+}99]. Coaching [Hoh99, Kan05]. Cobol [Ano93d, AHL^{+}00, AP00, AAB^{+}00, BMGT01, CK00, Coy00b, Gla99a, HD00, HVF^{+}00, Hor00, Jal89, Sam00, Sch00]. Cocomo [BV08, HPSD00]. Cocomo-Based [BV08]. Code [Ano98n, Ano08c, BP94, BGL^{+}09, CE09, DDS^{+}97, DRW03, Fow01c, Got99b, Hat08, HS90, Hor87, HJA08, JP03, KCF^{+}08, Lai08, LY98, LFL^{+}90, Lou06b, MM00, Nor03, PWS^{+}95, RBBP03, Rai07e, RS08, Rem05, Rot99, SCCA04, VMMF00, WC06, WB07b, WB08a, Sme92, BAS89]. Code-Generation [Hor87]. Coding [Daw95a, HM02b, JP03, Lou05]. Coffee [Hoh05]. Cognitive [Ano93e, Fis87]. Coherence [TM97]. Coladas [Hay03]. Cold [Bak95]. Collaboration [CP06, FH07, Fro07, Got03, REM^{+}09, SSC06, WSD^{+}09, WC09]. Collaborations [BCKM97]. Collaborative [DHMR02, FPM01, LANC07, MDWTR93, MSK00, Pat09, PDHT97]. Collaborative-Learning [LANC07]. Colleagues [MLI04]. Collecting [Max01, PSC02]. Collection [Moo99, Rog00]. collision [CGR94a]. collision-avoidance [CGR94a]. Color [ON92]. Colors [GN95]. Combinatorial [CDPP96]. Combine [Pos95]. Combining [KR05, BHKR01]. Comdex [Ano97c, Ano92g, BLP96, Rei94b, Rei95a].

Come [Ano93n, Ano09e, DM06, DeM09, MG01, Mat96, Rei95a]. Comes [Cur98, PM99b, SM01]. Comfort [Ste95b]. Coming [Boo94, Erd07a]. Command [Luq92a, OMTHC94]. Command-and-Control [Luq92a, OMTHC94]. Comment [HC91]. Commentary [WJS97]. Comments [Cod88]. Commerce [Car00b, OBR^{+}08]. Commercial [Ano02j, Cha96d, Dor99b, HC02, MD04, MLL03]. Commit [Hoh05]. Commitment [Thi88]. Commitments [RP09]. Commodification [vdLLM09]. Common [Ano93d, KG07, McC01c, BDHJ96]. Commonality [CHW98]. Commonsense [Cha97c, Dut93, YNM^{+}91]. Communicate [ZSF91]. Communication [Ano96d, ACF87, DESG00, Etl96a, Fer98, FWB85, Hay03, HA00, LC85, NS92, PDK05, STSK92, Spi09a, Har89b, KKH95, WKUI89]. Community [Aha99, BCC^{+}08, DP03, McC98e, San08, Sca04, SRW00, WSL^{+}99, Wil99b, WC09]. Companies [CTC07, CSSP05, JBHD08, KMSW00, NC00, Rob03, vWAS06, Mat00b]. Company [Dav00, Gre01, MDM03]. Comparability [Max01]. Compares [Pau95]. Comparing [Dor99b, DGC^{+}07, Sha95b, Tic89]. Comparison [DRW03, SM95, WGM85, Wol89, KB88, Sca89]. Comparisons [Mir01, SO92]. compelling [Shn91]. Competence [Dak96b]. Competencies [BHKR01]. Competing [BCH08]. Competitive [Wal96]. Compilation [Gut86]. Compiler [BBE^{+}88, Far84, GAY93, Sal97a]. Compleat [Gra03]. complete [SA97]. Completing [RW92]. Complex [CCRZ^{+}90, MSK00, KLR89]. Complexities [YM94a]. Complexity [Boo08b, DFL00, HS90, Lit05, McC96a, MK92, NO91, SVSL08, Sha01]. Component [Bal00, BSGF98, Bøe06, BB08, CTW98, Dig98, DT03, DZB^{+}05, GK02, Kot98, LCB^{+}09, MP07, US04, Voa98a, Voa01a, Wey98, ZJ98, Hat97a, RHK^{+}03]. Component-Based [BB08, LCB^{+}09, Voa98a, Wey98, ZJ98]. Component-Oriented [DT03, RHK^{+}03]. Components [Ano98y, CFQ07, Cou99b, LCB^{+}09, MLL03, Par03, RIP^{+}01, Sta94, Sch96a]. Composing [KH02, Voa01a]. Composition [Bla07]. Compositional [Zav89]. Comprehend [RS08]. Comprehension [LS86]. Compromise [Lai08, Sme92]. Computation [Pra85]. Computational [OLPS04]. Computer [Ano84b, Ano92i, Ano92q, Ano93m, Ano01f, Bar98, BC85, DSBM99, DSS85, GBdlHQCG^{+}00, Irv97, KM86, Lam84, Lin85, LK88, Luq92a, Luq92b, Mil84, Par99, PWS^{+}95, Ree91, Spi98, Wei99, Fis89]. Computer-Aided [Ano92i, LK88, Luq92b, Luq92a]. Computers [FWB85, FHMU85, KT88, Kal94, Kus96, Mou01]. Computing [Ano92d, Ano96d, Ano98p, BCC^{+}08, Cha99c, CT02, Chw01, Erd09c, Gla00b, Kel07, Kle98, Rei95b, Rei95f, Wil96, Zad94]. Con [WGK^{+}02]. Concept [Got06a, Got02c]. Concepts [BLSS08, DSBM99, KMSW00, SSSW86, Sny93]. Conceptual [Bol09, Bor85, DG09b, LSS94, Rei87]. Concerns [Got01d, KR03, VVJ06]. Concurrency [Ano93e, Ano96f, ZSM92]. Concurrent [Aoy93, CT91, For88, HR95, LSM84, PH93, RP95, Ros91, San97a, Wol88, FKU97]. Concurrent-development [Aoy93]. Cone [Ano06p, Lit06]. Conference [Ano96g, Ano96l, Bol04, Cha98a, GLS99, Ree06b]. Conferences [Erd09b, Voa00a]. Confession [Gla08f]. Confessions [Che97]. Confidence [Ano04p]. Configurable [LM07]. Configuration [Ber03, BH95b, LC87, SO97]. Conflict [Mac99a]. Conflicts [BI96, EG04a]. Confronting [Tru94, YM94a]. Congress [Bar95d, San95a, SB96]. congruence [MW96]. Congruent [Ada95]. Conjurer [HJA08]. Connecting [Rei90, WB08a]. Connection [Kra95, Moh01, Hat97a]. Connections [Cha98a]. Connectivity [Har05a, San98f]. connector [BCM88]. Conscious [Mac99a]. Consciousness [McC01f, Kal94]. Consciousness-Raising [Kal94]. Consider [Pat08b]. Consistency [BG95]. Console [Got01a]. Consoles [GAS^{+}01]. Constant [Har05a]. Constraint [TBY96, Bal89]. Construct [SHZ^{+}98]. Constructing [SMR93, Sta94, Ano93f]. Construction [Ano02i, Ano02j, CJK^{+}02, DH93, DBL04, DSB99, HC02, HG00, HSA89, HT02b, McC96e, McC06, Sal97a]. Consul [Bal89]. Consultancy [Kee00]. Consumer [GB00, RAv96]. Consumption [Mat01]. Contact [TB02]. Contemporary [GV95]. Content [Cha98a, Har03a, Har06b, IAJR97]. Context [BDA04, DZB^{+}05, Gau05, Lit05, Mai09c, Mau90, SO90]. Context-Adaptive [Lit05]. Context-Free [Mau90]. Contexts [Mye98]. Continue [MHC03]. Continued [CK00]. Continues [MHC03]. Continuing [Ano99d, Sha09]. Continuous [AAR09, DJH^{+}08, KM02, Sho04a]. Contract [Dak96a]. Contractor [Rug93]. Contribute [Ano92h]. Contributions [SSZ06]. Control [Ano92i, Ban93, BMM05, BDKP99, BALV03, Bri96, Car94, CM90, DJH^{+}08, EA03, FH07, FCB00, Got00, HK94, JC00, JP03, KGRB03, Lev90, LLH^{+}05, Lou06d, Luq92a, OMTHC94, PHRS07, Sha01, Spi05e, Spi05f, TV99, VNS94, Wel00, WGFC86, Har89b, Put91, dSA95, MSK00]. Controller [CG08, DG07]. controlling [Smi91]. Controversy [Ano92d]. Conundrum [Ber08]. Convenience [Ano05k]. Conversation [Pat08d]. Conversions [TV00]. Converting [BPS94].

Conway [HG99]. Cool [Rei96c]. Cooper [Pat08d]. Cooperation [DVG09, EG04a]. Cooperative [DAEE08, Dua02, dSSS^{+}09]. coordinate [BLA88]. Coordinated [CT93]. Coordinating [LSG^{+}05]. Coordination [AAR09, HG99]. Coping [Ben95, CM96]. Copyright [Ano93-32, Bar95e, Cos94, Sim95a, Tom87, Sam88]. Copyrighting [Mil84]. CORBA [Ano01f, Frü99, Pud04]. Corner [GF99]. Cornering [Dro96]. Corporate [Dio93, Pfl93, Ros00, Sal03]. Corporations [REM^{+}09, San08]. Correct [TL98]. Correction [Ano02j, Ano03l, Dav95b, SHZ^{+}98]. Correctness [Ano02j, BH01, HC02]. CoScripter [Ko09]. Cosmic [MHC03]. Cosmic-FFP [MHC03]. Cost [AH90, Ano96v, BB91, Boe00, CBMP05, EPSJ93, Gla02a, HF95, HM02b, Jør09, Jun98, Kan01, Kän97, KR97, KT05, Mad97, MH06, PM97, She91]. Cost-Effective [BB91, HM02b, She91]. Cost-Value [KR97]. Cost/Schedule [AH90]. Costs [Ano93-40, Voa99b, Smi91]. COTS [BOS00, CL00, DZB^{+}05, EMP05, GK02, KT05, Koh05, LSG^{+}05, LG03, MN98, MP07, PWW05, RBBC03, TM04, US04, Voa98b, WSG^{+}05, YBBP05]. COTS-Based [BOS00, RBBC03, TM04, YBBP05]. COTS-Intensive [Koh05]. Could [MGR04]. Council [Ree00, MM00, Moo01b, Moo01a]. Count [Ano92h]. Countdown [McC99f]. Counterpoint [AW03, Bac97, Bec07, Car97, CK02, CHJ^{+}06, CB00, DeM99, EM09, Fen96, FV07, FPK07, GC08, HvG09, HM02b, Hor00, Jac96, Kis96, LB00, Law98, LS07, MC94, McC97f, Nie97b, PC09, RM04, RC95, RH05, SY99, SS06b, SVSC09, SN07, SM01, TB02, UA03, UE07, WCCR08, Zve98a, vGB08, CC95b, Gil96b]. Counterproductive [Ano96p]. Counting [VK07]. Countries [Ksh04]. Country [BHKR01, Coc01, Ju01, Moi01]. Coupling [Fow01b, OYE06]. course [Sam88]. Courses [CT02]. Coverage [WGM85]. Cow [Ado96]. Cowboy [Cur01]. Cowboys [HMP^{+}01]. CPM [Cha96c]. CPU [HKR^{+}90]. Cracking [LV01, Spi07b]. Craft [EMW^{+}97, HMP^{+}01, JL03, Wil97c]. Crafting [Gou98]. Cray [AGK^{+}87]. Create [Ano92b, Ano92d, Kot98]. Creating [BHAW09, DDS^{+}97, FTB06, Mac98, Sch96a, SL97, SB95a, SC96, Sym88, WB09a, vM95a, Lew98]. Creation [Lit04]. Creative [Coo96]. Creativity [Gla01c, MGR04, OGE08]. Creators [Sam00]. Credible [WKW99]. Crisis [Ano99d, Gla94b, Gla98c, MC94, Lew98]. Criteria [Ano97g, CFQ07, Eng99]. Critical [Ano98k, BM91, BM09, CMKC09, DNR04, GCR94c, HL99, HC91, Jac95, KL94, KYM^{+}04, NK04, RBBP03, Ram00, Ree99, Sef99, Wil94, YJC09, ZME00, Bei97a, Duk89]. Criticism [WB07e]. Critiquing [FN88]. CRM [Kim04]. Cross [Ano98l, ME00]. Cross-Pollinating [Ano98l]. Cross-Pollination [ME00]. Crosscutting [GSS^{+}06, VVJ06]. Crossing [Nak96, TTT^{+}02]. Crouching [Fer01a]. Crypto [Kus96]. Cryptography [Pfl95a]. Cryptology [DSB99]. Crystal [Ano93w]. CS [Ano96h, Gla05b, San97b]. CS-IS [San97b]. CSDP [NS05]. Culling [MLI04]. Cult [McC00a, MJP^{+}00]. Cults [Gla00d]. cultural [Nak96]. Culture [Ben00, Cla98a, Dor99b, Eme95, Fer98, Gar00, Got08a, Hei98, Hen99, Jac98a, Mac98, Mac99a, Mac99b, Mac00, Mcl98, Mic00, Nis00, Per99, Ros00, Rot99, Rot00, SRW00, Lew98]. Cultures [Dor99b, Gla00d, Gla06a, Mac98]. Cure [Ano09b]. Current [BW98, Jon96, WPJH98]. Currents [GC07, GDD07, Got09a, Ing08, Sha07]. Curricula [Ano07i, LLS^{+}06]. Curriculum [BCC^{+}06, LB06, SS95]. Curses [Ano93f]. Curtain [Wil99a]. Curve [Kem92]. Custom [NV02]. Customer [Chr94, DJS08, Frü99, Har89a]. customer-oriented [Har89a]. Customization [DVG09]. Cusumano [Mye95c]. Cyber [Got02c]. Cyber-UL [Got02c]. Cybersecurity [PR08]. Cyberspace [GD02, Gou97]. Cycle [Bas87a, Ebe06, GS90a, HF95]. Cyclomatic [McC96a]. Cyclone [Got02a]. Czar [Rot99].

D [ECW09, San95a]. Daily [CM96, McC96c]. Dance [Pet99]. Danger [Got01a]. Dangerous [BK05, Mar98, Pon98, Rei02c]. Dangers [Har04d, Str96]. Dark [Rob01a, VD96]. Darlington [CGR94b]. Dashboards [Sel09a]. Data [AW91, BLP96, BMGT01, Bie00, Bif00, BBGK07, CBMP05, CC93, Cuk05a, DCH04, Fow03a, Gon02, Got02d, HS91, JW09, JMVS09, LH01, LL87, LWHS01, MR95, MBO93, Mau90, Max01, Oat98, QH90, Ros98a, Var88, WC06, Wel93, Cod88, MMS88, Vis94, Win90a]. Data-Driven [DCH04]. Data-Gathering [Gon02]. Data-Parallel [QH90]. Database [Ano93e, Bas87b, CCRZ^{+}90, CFW^{+}98, Fon96, Gai94, Lin87, MDM^{+}94, PHRS07, TL98, Wie84, YM94a, YM94b]. Database-Centric [PHRS07]. Databases [Alo03, Amb07, AEP96, DY93, Pou95, RH89, KR88]. Dataflow [FL86]. Dave [Ano96h]. David [Ano96m, Pfl97b]. Davidow [Ano97f]. Davis [Har03c]. Days [Gla02b]. DCE [RD96]. Dead [Gla02e, Rei95e]. Dead-End [Rei95e]. Deadline [Lew98]. Deadly [McC01e, Rei01, Rei02a, SGW02]. deal [Dak95a]. Dealing [SK08]. Dear [Spi05a]. Death [Cla98a]. Debatable [And95]. Debate [MHC03, McL05, SLH01, GA05]. Debates [MMM^{+}01]. Debating [Coo96, Cur95, KS96]. Debugger [ISO87, SFG94]. Debuggers [Spi06c]. Debugging [ADS91, Ano96d, AFC91, CBM90, HL85, LST91, MW91, OCHW91, PWS^{+}95, PHK91, PT09, Sev87, SI91, TBY96]. Debunking [Gra00a]. Decade [DSBM99, Dun89]. Decision [Ano92i, CDD08, Don06, EPSJ93, Gla08a, KCD09, MMM^{+}09, Mey96a, Str96]. Decision-Support [MMM^{+}09]. Decisionmakers [Dav95e]. Decisions [ABCF91, Bar96b, CMKC09, Erd05, FCH^{+}08, HAZ07, ROL90, TA05]. Deck [Got01b]. Declarative [Lou07, Spi03]. Decline [Ros00]. Declining [Gro04]. Decompiler [San98f]. Decomposing [Bla07]. Decrypted [DSB99]. Deductive [DY93]. Deep [IKBM08]. Defect [Ano93g, Bif00, Car98, Gra04, KT04, KL05, McC97c, RAT^{+}06]. Defect-Prediction [KL05]. Defects [Eme95, Pre01, Rai07b]. Defend [McC96d]. Defending [MCA00]. Defenestration [Boo09c]. Defense [AAB^{+}00]. Deferred [Sho06]. Define [MKH98]. Defined [CDF96, Hum96]. Defines [BLP96]. Defining [DB96, EEM01, FKKM89, Gla98b, Jac98b, OO92, SM88]. Definition [Jon05]. Definitions [TN92]. Degree [DJM04]. Degrees [PTH^{+}09]. Déja [Dav96b]. Delayed [Got01d]. Delaying [Thi88]. Delegation [ADT91]. Delivered [Got02b]. delivering [GSB03]. delivery [McC97b]. Delocalized [LS86]. Delphi [Gor95]. Demand [Erd07d, Jon94, Kle98]. Demise [Nie96a]. Demo [Ara95]. Demonstrating [LSM84]. Demonstration [RGR85]. Demystifying [TA05]. Denning [Bar97c]. Density [Wit90, Hat97a]. Departments [You00]. Dependability [Ano99f, BDA04, BHJM04, DZB^{+}05]. Dependable [KL94, dLdCGR06]. Dependencies [Lub91]. Depends [Mai07]. Deployment [DNS^{+}06]. Describing [Sim06]. Description [Jac98b, LCM06, RP06]. Descriptions [Lau03]. Design [AW91, Ano99g, Ano04o, Ano08d, AMM^{+}89, ABCF91, Bai98, Ber93, Boe84, Boy97, BC86, Bud99, CN07, CC90, CS90, CSR^{+}99, CDPP96, DDMS86, DG09a, Dic05, DSS85, DB96, DSB99, Etl96a, EMB^{+}99, Fac01, Fow01a, Fow01b, Fow01c, Fow05, GP05, Gla99b, GYvdR06, Got02a, GSS^{+}06, HLS90, Hat97b, HS90, Hol92, HKPS96, Hou01, JC00, JS08, JOL^{+}00, Jur97, Kai99, Kan01, KMY91, KM86, KS96, Kru05a, Kru05b, Kuh98, LANC07, Lam84, LCM06, Lar01a, LM84, LH01, Leh86, Lub91, MSJ^{+}09, MMSM09, MS04, Mey87, Mey04, MKMG97, Ner01, NV02, Nie95b, Pat08d, PWS^{+}95, Ray06, RW90, RG02, Rog00, Rom90, Ros85, RW92, ROL90, Sha95b, STSK92, SM97, Sho04a, Sim06, SO92, Sto99, THV09, Ter96, Tho05, WSL^{+}99, Wan95, WHC02, WB06c, WB07b, WB07c, WB07d, WB07e]. Design [WB07f, WB08a, WB08b, WB08c, WB08e, WB08f, WB09a, WB09c, WB09b, WB09e, WB09d, Yam98a, YKIY93, ZAV02, ZZ93, Zdu09, BDHJ96, Kar88, MMN^{+}89, OBM^{+}90, Sch96a, Son97, TK89, War89, DeM95b, Jur97, Pou99, Rog00]. Design-Method [SO92]. Design/Requirement [Wan95]. Designer [WB09f]. Designers [Law98, VBM^{+}02]. Designing [Ano08d, Clo01, CN95b, DSBM99, DV89, DDPW09, ES92, Hou01, Jef97, Nie96c, SB95b, SMK^{+}00, WB06a, WB07a, WB08b, WB09c, WB09b]. Designs [PM99a, Sel09a, WB09a]. Desk [SGM06]. Desktop [Got09b, HAK^{+}96]. Destroy [Lin93]. Destroys [Fav96]. Detecting [EGP92]. Detection [KH97, MP07, MCA00, PCOM97, RAT^{+}06]. Determination [KL88]. Determining [CFQ07, EPSJ93, MSG^{+}07, Sim96]. DEV [vWHvW09]. Develop [Hal96b]. Developer [Ano92r, Ano96r, Ano04p, Gee06, HSB^{+}08, Har03e, Har04e, McL06, PSHH04]. Developers [Ano97d, Ano06k, Cha95, FJWC01, Got01a, Har05f, Ing08, Jac95, KL97a, MC03b, MHC03, McL05, MKF06, Mye01, PWS^{+}95, PM98, PCS95, Rob09a, Sef99, US09, Dak96f, PSV94, Lew98]. Developing [Ano02j, Ano03h, Ano03i, BGB01, BR99, BOS00, Con02, DJSS97, FK04, Ger90, Gon05, HC02, Jaa02, KS89, Ksh04, MBL08, MH07, MS95, Mub08, PWS^{+}95, RBIM97, SM08b, SM09b]. Development [ABF05, AH90, AKM^{+}08, AJM06, Ama07, Amb02, Amb07, AFGD01, And95, Ano92i, Ano93e, Ano93o, Ano93z, Ano98g, Ano99d, Ano00k, Ano00i, Ano00h, Ano05o, Aoy98, AK03a, AK03b, ASC07, Bal00, Bas90, BRL^{+}03, Bas95, BKS98, BCKS01, Blu87, BT05, BB08, Bor85, BCC^{+}06, Bur94, BPD02, CA01, Che95, Cri06, CM07, CMKC03, CMKC09, DM06, DWH86, DM93, DCH04, DAB06, DG07, Dua02, DD09, Ebe01, ED01, EMP05, Erd09d, FH07, FTR^{+}94, FNY09, Gla99d, Gla04e, Got01c, GAS^{+}01, Got02e, Got07, GH00, Gum06, Hal09, Har04c, Hei01, HM01a, HBP92, Hsi96, JBR99, Jan04, JBHD08, JS08, JKP^{+}05, JHMW07, Jon96, Jon03, JG08, JBR09, JM02, KFP88, Kan09, Kem90, KCF^{+}08, Kim04, Kru08b, Kru09a, Kru09b, KR03, LM07, LF05, LNR98, LHB^{+}90, Les06, LO90, LCB^{+}09]. Development [LR00, Lit04, LVS92, Lou06c, LB06, MKCC03, MSG96, MSJ^{+}09, Mac99c, Mar96, Mar00, Mar07, MMSM09, MF00, Max01, McC98c, McC99e, MR09, MCF03, MD90, MJ88, MSK00, Mot06, Mug08, Mul86, Mye02, NOK08, NS99, NT06, NK04, NC00, OGE08, OY95, Pal01a, Pal01b, PWS^{+}95, Pit93, PS90, Pue97, RKK05, RRD07, Rei00c, Rei02a, RIP^{+}01, RJ00, RE04, RP07, RM00, Sal97b, SK08, San95b, Sca04, SA05, Sch07, Sel03, SK03, SCSC04, Sim06, SMHT09, Sot01, SS04, TM04, UA03, Uhl08, VB97, VE05, Vis08, VK07, WEL02, WW03, WPJH98, Wil09, YJ07, ZAV02, Zha04, ZS06, vdLLM09, AO89, Aoy93, CW89, For97, LD95, MMN^{+}89, MW96, Mey91, Per90, VT88, Zho89]. Developments [Ano00j, Rus91]. Device [MTKE94, SGM06]. Devices [DW00]. Devoted [Ano96h]. Diagnosis [Hat01, TC86, SSGT86]. Diagrams [Mil02, RW04, Nic92]. dialects [KB88]. Dialog [Mar00]. Dialogue [SK95]. Diana [Ros85]. Dictionary [Ano99f]. Did [Hei01, JD97]. Died [WJS97]. Dies [Ano96m, Dav95b]. Difference [DeM90, Hoh03]. Different [BRL^{+}03, DESG00, DSS85, Gla06a, RvW07]. Difficult [Spi09d]. Difficulties [Kai99, MBE05, Mey91]. Diffusing [RC89]. Diffusion [GH00, SS08]. Digital [Di 05, Gue94]. Dijkstra [Che02, DCF^{+}02]. Dilbert [Mac98, Wei98b]. Dilemma [Ano92p, HT03b, Wan95]. Dilemmas [BM04b]. Dimension [Bag99]. Dimensional [CG88, LST91]. Dimensions [Gum06]. Diminishing [Gla09d]. Direct [Ano97e]. Directions [Ano93e, BR87, BLWG99]. Directives [Ano92d]. Dirty [GF99, Voa97b]. Disasters [Har05c, Wil97c]. Discipline [DDS^{+}97, Gon05, Jac98b, LSB02, Mat00b, MAMY99, Mye95b, Sha90, Sha09, TR05, Was96, Pfl97b]. Disciplines [Ano98l, Bai98, Rob05a]. Discomfort [Gla04c]. discount [CN95a]. Discover [BGL^{+}09, PSV94]. Discovering [BCA^{+}06, WB07c, ZMT05]. Discovery [BBGK07, LM04, WMGM08]. Discussion [HVF^{+}00]. Dispatching [Sho06]. Dispelling [SBM02]. Display [SSSW86]. Dispute [SLG^{+}00]. Distance [CA01, GSB03, HG99, QBN^{+}06]. Distilled [MBE05]. Distinguish [Gla04b]. Distributed [AAR09, BPD02, CV88, Che84, CBM90, CW84, DV89, EG04b, Fer95, Fer01b, FO85, FLR04, GC85, HKPS96, KKM87, KYM^{+}04, Lin87, MBL08, MDWTR93, ND05, RS84, RKK05, RS01, RIP^{+}01, SDP91, SSC06, SFG94, SAG^{+}97, Fid96, Fer95]. Distribution [Gum06, WMGM08]. Diverse [JB00]. Diversity [Ano98j, Ano98w, Dec01, Erd09d, Gla00b, Kan01, KMY91, KL98, LR00, MTE00]. Divide [Kin97, MKMH07, San97b, Nak96]. Divided [Erd08e]. Diving [IKBM08]. DLLs [San98f]. DMCA [Got01d, MC03b]. DNS [Cha97e]. Do [Ano93a, Ano96p, Ano97g, Bat98, Boo08d, CL00, Chr94, Dak95c, Dak96b, DJS08, DBS09, DD09, Fen96, Gau05, Gla08c, Gor95, HSB^{+}08, HVF^{+}00, Har05b, Har05f, Hay03, Jon07, Kla95, LB00, Law96, Law98, MR95, MG01, McC97f, OG07, OBR^{+}08, RAT^{+}06, Sch96b, Spe99, Zve03]. Doctrine [Ano93-35]. Document [LCS97, Ste95a, Yac03]. Documentation [Kot98, LSF03, Sel09b]. Documents [GS90a, Ros06, SFVFM08]. DoD [Cha96d, Fer01a]. Does [Aha99, Ale95, Ano96v, Coy00b, EM96, Erd09c, Gla94a, Hat98, JS08, Mye95c, RC95, Spr95, Voa99d, WB07b, WLC01]. Doesn’t [Har05a, Hoh05, Kla95]. Dog [Har06c]. Doing [Cha98a, Gla02a]. Dollars [Ano93-34]. Domain [CRR09, CM07, FNY09, Fow09, Gla00c, HTLPS09, JRS^{+}09, KP09, LM07, MSG96, MKH98, Mil02, SMTS09, TH02d, WHG^{+}09, GV95]. Domain-Configurable [LM07]. Domain-Specific [CRR09, CM07, Fow09, HTLPS09, JRS^{+}09, KP09, SMTS09, WHG^{+}09, MSG96]. Domains [MA00]. Domination [Per99]. Don [And02]. Done [Mai05, RR08]. Don’t [AW03, Bar95a, DeM99, EA03, Mag00, Mel95, San98c, RHK^{+}03]. Dorothy [Bar97c]. Dose [Ano96r]. Double [Mat00a]. Double-Edged [Mat00a]. Doubt [Gla09a]. Down [Ano93g, LB00, LCE^{+}01, Par96a, PM99b, TM94]. Down-to-Earth [Par96a]. Dragon [Fer01a, GA05]. Drawer [Ano92n, Ano93c, Ano96u, Feo97, Sha97a, Ano97i]. Drawing [Moe90, Spi09c]. Drill [Ano98p]. Driven [ABF05, Amb07, Ano05o, AK03b, Cri06, CM07, DS07, DCH04, DSLB03, DG07, DFL00, Jan04, JS08, JHMW07, KYM^{+}04, KR03, LX99, Mar07, MMSM09, MMM^{+}09, MR09, MCF03, Mug08, NT02, RS04, RP07, Sel09a, Sel03, SK03, SH02, UA03, Uhl08, VK07, WB07c, Zha04, ZS06]. Driver [UA03]. Drivers [KT05]. Drives [Ano93g, Cos03, Ebe07]. Driving [Cri06]. DRY [HT04b]. DSR [Bar96a, SB96]. DTV [Ano97d]. Dunnit [Nus97]. during [Ano00k, DL03, MCL^{+}01, RBBC03, Sim06]. Dwarf [WJS97]. Dynamic [BCMV05, DLR07, EFM^{+}91, FLR04, MP07, Moe90, MK92, RS84, SF93, SDML^{+}06, Shn94, SFG94, WC86]. Dynamically [Ano08b, KKH95, TW07]. Dynamics [Ber96b, Chi02].

E-Commerce [Car00b, OBR^{+}08]. E-science [TSP^{+}09]. e-Service [GYvdR06]. E-Voting [Sto03]. e3value [GYvdR06]. Each [Whi96]. Early [BCA^{+}06, Cuk05b, FCH^{+}08, Got01e, KAKG96]. Earth [Par96a]. Easier [CRR09, Xu03]. Easily [BBE^{+}88]. Easy [Ros98a]. Eating [Har06c]. Ebert [Ano96i]. EC [Ano92d, Cha98a]. Eclipse [YJ07, DGC^{+}07, Fro07, MKF06]. ecological [TK89]. Economic [Ano98g, Boo09e, Boo09f, Erd05, PR08, SV02]. Economical [Voa98b]. Economics [Boo07b, Ksh04, Lim94, Cox97]. Ecosystems [TH04b]. Edge [Wal96]. Edged [Mat00a]. Edgewatch [San98a, San98b, San98c]. edited [Bei97b]. Edition [Ano01f, DW00, Wei99, San98e, DW00]. Editor [McC99a, McC99c, McC99e, McC00b, McC00e, McC00c, McC00f, McC01b, McC01c, McC01a, Spi05a, Ano92h, Big96, Bol02, Byt99, Car95, Cha94, Cuk05a, Cur00, Dav95c, Dav95a, Dav95d, Dav95b, Dav96b, Dav96c, Dav97a, Dav97b, Dav97e, Dav97c, Dav98d, Dav98b, Dav98c, Dav98a, Ebe09, Erd08b, Erd09c, Erd09d, Erd09e, Erd09b, Gla03c, Har03a, Har03d, Har03c, Har03e, Har03b, Har03f, Har04a, Har04b, Har04d, Har04c, Hor98b, Hum00, Kru05b, McC98e, McC98f, McC99b, McC99d, McC99f, McC00d, McC00a, McC01d, McC02e, McC02b, McC02c, McC02d, Mel05, Mil01a, PC97, Rif09, Str99, Str01, Voa99b, Voa01c, Voa03, Wil03, Zdu09]. Editorial [Ano01i, Dav95a, Dav96a, Ano96i, Big96]. Editors’ [JWC01, AP00, BD97, BF00, BB99b, CBM98, CW00, DM06, DP94a, DP03, ES09, EMP05, EFS04, EH95, GW07, HF02, HM01a, JM07, JMS06, KV07, KABM09, KB98, LFW00, LR00, LB06, MRE05, ME00, MNJP02, MCL97, MCF03, ME02, MS06, NZZ09, NZ07, NOK08, PL03, PT05, RvW07, RL02, SE98, SM09b, SS04, SMTS09, TW07, VM00, WE04, dSSS^{+}09]. Edsger [Che02]. Educating [CP02]. Education [Ano96r, Bac97, CK00, DJM04, FPM01, HGK09, Hil97, HH02b, HM92, Irv97, McC97f, MCL97, MJP^{+}00, Moo02b, PDHT97, TR05, vV06, KCC^{+}01]. Effect [Kam87]. Effective [Ano92i, Ano93-36, BB91, BS03, DSBM99, HM02b, KL05, SL97, She91, ABL89, HF97]. Effectively [LN92]. Effectiveness [Che01, DAS^{+}07, Voa97a]. Effects [BM04b, Cla00, Gla04a, Lim94]. Efficient [BR85, Gut86]. Efficiently [Mos92]. Effort [Cla00, Got01e, JMDB00, Jor05, JG08, JBR09, PRSV97, VT88]. Effortlessly [Rai07f]. Efforts [DAB06, FS98, Got02d]. Egoless [Wei99]. EIA [FS98]. EIC [Har03c]. Eight [Cla02, RBBC03, Voa99f]. ElcomSoft [MC03b]. Electronic [Ano93-39]. Electronically [Kus95]. Electronics [RAv96]. element [FKLW91]. Elements [Ano92e, DeM95b]. Elicit [URK01]. Elicitation [DJS08, SPZ08]. Eliminating [CK02]. Ellen [Ano98q]. Else [And94, Sam90]. Eludes [Ano92g]. Elusive [Gla03b, KL96]. Email [Rei95c]. Embedded [Ano99p, CE09, ES09, GLT03, LSB02, LT09, Med05, MW09, SMHT09, TYZP05]. Embrace [MSK00]. Embracing [FLR04]. Emerald [HAK^{+}96]. Emerging [Pfl97b, SS96, US04, VZM^{+}07, WC89]. Emotional [CNS08]. Emphasizing [AJM06]. Empirical [CR08, CDM^{+}05, Pfl05]. Empirically [JMDB00, PS90]. Employees [Yam99]. Employment [Dak95a]. Empowering [DG09a]. Emulator [San98f]. Enable [DC04, Gar05, Sel09a]. Enabling [SSC06, WB08d]. Enbugging [HT03c, BPS^{+}03]. Encourage [WOS94]. Encourages [Got08a]. Encryption [Ban93, Bar97a, Bar96b]. End [Ano92f, Ano92g, DDPW09, Gra04, Har04d, IKBM08, KABM09, Rei94a, Rei94b, Rei95b, Rei95c, Rei95d, Rei95e]. End-to-End [Gra04]. End-User [Har04d, KABM09]. Ending [Gla05a, ZJ00]. Ends [San97b]. Enduring [Pfl97b]. Engine [Daw95b]. Engineer [BCC^{+}08, BS95, CC96, Dak97a, GAS^{+}01, Hal09, Har05e, Kla95, Pae08, Spe99]. Engineered [PAD^{+}98]. Engineering [ANB99, Ano84a, Ano92i, Ano93-34, Ano93t, Ano03k, Ano04o, Ano07i, Ano09e, AW06, BHS85, BLP96, BHAW09, Bas87c, Bas07, Bat98, BL98, Ber08, BGM06, BHJ^{+}03, Bol02, BDA^{+}99, BD94, Bur95a, CR08, CDD08, Cha94, Cha97d, Cha98b, CW00, CT02, CR88, CC90, Chi02, CM90, CL02, CHW98, Cos01, Cou99a, CMA^{+}02, Cuk05a, CP06, CR92, Dak96a, Dam07, Dan04, DJSS97, DP94a, Dav96d, DeM09, DRR^{+}07, DJM04, DSD02, DBS09, DVG09, DT03, Dor99b, DAEE08, DP03, DB99, DSB99, DLV99, DKJ05, Ebe01, Ebe06, Eng99, Erw09, FPM01, Fra95, Fra99, FM08, FTB06, Gla01b, Gla03a, Gla03c, Gla03d, Gla04d, Gla06b, Gla07a, Gla07b, GW07, Gom06, Got08b, Got99b, GLT03, GRS02, HL05, Har03b, HBB01, HF02, HL01, IS94, Jac98c, Jac99, Joh95, JMS02, KLD02, Kel07, KABM09]. Engineering [Koh05, Kru08a, Lab94, LFW99, LFW00, LM03, LWE01, LLS^{+}06, Lew98, LT09, LL92b, LB06, McC98c, McC99d, MT99, McC00d, McC00a, McC01b, McC01d, MAMY99, ME02, MLL03, Moo99, MJP^{+}00, Moo01b, Moo01a, Moo02a, Moo02b, Moo03, Mub08, Mus85, ME90, Mus93, Mye95d, NGBR05, Nav95, NL03, NC00, PDK05, PF92, PJT09, Par99, PDK93, PWS^{+}95, Pfl02, Pfl05, PM98, PH93, PT05, QBN^{+}06, RP95, Ree00, Ree05, Ree06b, Rei03a, Rob05b, Ros05, Row96, RS04, RL02, SBM02, SSV99, Sch95b, Sch99, SRW00, Sha90, Sha09, SD88, Sid94, SS96, Som05, SRS^{+}86, SMK^{+}00, SHM02, Sym88, Tha03a, Tha03b, TSP^{+}09, TR05, Toc97, Tot06, TB02, Voa04a, WKA96, WW03, WA02, WE04, Zve03, dSSS^{+}09, vV06, Ano93-27, CN95a, Pot93, RC89, Sam90, SA97, Tho89, Was96]. Engineering [Bei97b, Mye95b, PTH^{+}09]. Engineers [Bla07, Con02, DN97, DHMR02, Dor99a, FPG94, Har03a, Hum01, Kru08b, LSF03, Mag00, McC02f, NS05, Nor07]. English [Gla08d]. Enhanced [Mau90]. Enhancing [HPSD00, HS00, vWHvW09]. Enough [Bac95a, CJK^{+}02, HB06, JOL^{+}00, Mcs96, Ree03, UA03, WA02, You95, Zve03]. Enrollment [Con96]. Ensuring [CE09]. Enter [GA05, Ano85]. Enterprise [DDPW09, Erd07d, Hou01, Mur95, Par05, PHRS07, Uhl08]. Entertainment [GAS^{+}01]. Entity [San03, SM88]. Entity-Life [San03]. entity-relationship [SM88]. Entrepreneurial [Ano92i]. Envelope [Ben99]. Environment [AK85, AO88, BRS^{+}85, Blu87, CAGR91, CR92, DLR07, EM91, GGJ^{+}89, LM07, LVS92, MPT86, Map85, OCM^{+}84, PF92, Pem87, PH01, Pra85, Pue97, PTP^{+}88, Rei87, Ros85, San03, Sch92, SR85, SD88, SBK86, WMWM87, vvSO^{+}98, BDHJ96, LC87, MMN^{+}89, Rei90, SA97, WKUI89].

Environments [CL91, CNL99, HKR^{+}90, Jar92, JB00, KKM87, MS92, PDK05, SB95a, STM88, SAG^{+}97, TN92, VZM^{+}07, Voa99a, KS89, Oma89, Tho89, Pou99]. Envisioning [And07, Ano98o]. Epidemiology [TSP^{+}09]. Epochs [SVSL08]. Equality [Hal95]. EQUITY [Ree06a]. Era [Got02e, Jur97, Kru05b, MLL03]. Eras [Dav96a, Ano96q]. Ergonomic [Bil94]. Eric [Mil01b]. EROS [SH02]. ERP [Dan04, DVG09, HL99, WC09]. ERPs [SS06a]. Errata [Dav97c]. Erring [KCC^{+}01]. Error [Ano93g, Gla03b, Gla08f, ON92, PWS^{+}95, SHZ^{+}98, TV99, Wit90]. Error-Free [Gla03b, Gla08f, TV99]. Errors [Pos85]. Esaps [vdL02]. Escort [WKUI89]. escrow [Sme92]. ESI [Wal96]. ESPRIT [Ano89, Ano93h]. Essay [Ano99j]. Essence [Sny93]. Essentials [Erd08a, Rin97]. Establishing [Dak95d, OJ97]. Estimates [Bie00, Fai02, Jør09, KL97b, Mir01]. Estimating [Cha96c, JG08, Rei00c, VT88]. Estimation [AHM86, Ano00j, Bif00, BF00, BV08, HPSD00, Jor05, JBR09, Kän97, Lit06, MH06, PM97, Rif09, Put91]. Ethical [Hal09]. Ethics [Got99b, KCC^{+}01, KHA^{+}03]. Etiquette [Spi09a]. Eureka [Rob02]. Euro [Ano98p, Jon99]. Europe [Ano93b, DLV99, KCL93, vdL02]. European [Ano93e, Ano96r, Car00b, JMS02, MDM03, Sch99, Sku99]. Evaluate [Sch98a]. Evaluating [Ano92i, BKS98, BW96, DZB^{+}05, Oat98, PS92, ZZ93]. Evaluation [BDKP99, KPP95, Rug93, Sch95a, Tia04, WHC02]. Evaluations [BM91, BM09, HC91]. Event [LSG^{+}05]. Events [EG04b]. Ever [Jac98c, LB00, Mai05]. Every [SM01]. Everyday [VVS^{+}08]. Everyone [Car97, Can95]. Everything [Car95, Cus03, HTLPS09, MRA98]. Everywhere [Ana99]. Evidence [DBS09, DKJ05, JMVS09, MH06, NSV^{+}08, Pfl08, Shu07a, SS08, vWS09]. Evidence-Based [DKJ05, MH06, SS08]. Evidential [Pfl05]. Evolizer [GFP09]. Evolution [BG94, FLR04, Kam87, Kee00, Leh98, SV02]. Evolutionary [Ano93z, Cox84, FH07]. Evolving [Ano93i, ND05, SS95]. Exactly [BMGT01]. Examination [Fer95, Bei97a]. Example [FHMU85, SS08, Wil99b, WC86]. Examples [Ano93j, Rai07a, Veg86]. Excellence [Ano92i, Bal99]. Exception [BMM05, Knu87, WB06f]. Exception-Handling [Knu87, WB06f]. Exceptional [Ano98h]. Excerpt [Ben99, Bro95, DL98, JBR99, KP99, Wei99, Jac95]. Executable [JB04]. Executing [ACM^{+}07]. Execution [ADS91, KC90, MS99, TS06]. Execution-Backtracking [ADS91]. Executive [SM99, Wil97c]. Executives [SJM02]. Exercise [JC00, HMP^{+}01]. Exist [Coy00b]. Existing [DSLB03]. Exists [HVF^{+}00]. Expanding [Ano97e, Gee06]. Experience [Ano97g, Ano02i, BGS01, DSBM99, Ebe01, FPM01, FLN01, GCR94c, HL05, KH00, KCC^{+}01, MRZ^{+}05, NR89, Pat09, Pre01, Rus91, SvHB02, SR91b, Tau95, UHC02, WKA96]. Experienced [CP02, Moy97]. Experiences [GJPD99, Kan09, Str01, TSP^{+}09, Tot06, TH02d, WW03]. Experiment [OCM^{+}84, SSSW86]. Experimental [Bur95a, DM87, DRW03, JWKZ99, Ree06b, Sca89]. Experimentally [JMVS06]. Experiments [FNO92, PSV94]. Expert [ABG^{+}97, Ano92i, Ano96j, Dou85, DJ05, Jor05, JBR09, Nar85, WGFC86, CFTZ89]. Expert-Judgment-Based [Jor05]. expert-system [CFTZ89]. Expertise [ZSF91]. Experts [AS03, FNY09]. Explained [Ano93f, MSK00]. Explaining [WB06c]. Explanation [Ray06]. Explanation-Oriented [Ray06]. Explicit [Fow01d, NCC^{+}02]. Exploration [Kuh98]. Exploratory [Ano93n]. Explore [JKM^{+}09, MJ97, WSD^{+}09]. Explores [Mye94, Ber99]. Exploring [Hat01, Hud88, HPG^{+}00, LNR98, MCL^{+}01, Sim95c, WK03]. Explosion [Got99a]. Exports [Bar96b]. Extended [Mil02]. Extending [DP94b, Kor86, Nic92, Rov86, SH93]. Extends [Mur95]. Extensible [EL02, Sel90, WB07a, KR88]. Extensions [CS05, ZHL^{+}89]. Extracting [CS90]. Extreme [Dan03, Lon02, MSK00, Gla01a, Gre01, Pau01, PH01, Sch01, Str01, Mar00]. Extremely [Gla03b, SGW02]. Eye [MHC03].

FAA [Bar96a]. FAC [TP86]. Face [AS03, Boo09e, Boo09f, Cha94, Sim95b]. Facilitating [Fer98]. Facility [AKM^{+}08, SM88]. Fact [HDS05, MT99, San98b]. Factor [HL01]. Factories [Ano92l]. Factors [Ano98k, DC04, Faf94, HL99, Mad97, McC00b, Ree99, ZME00, FI94]. Factory [Con02, FNO92, Cus89]. Facts [App02, BALV03, Gla01b, LCB^{+}09]. Fad [Car95, Fra98]. Fail [Lan96, Sho04b]. Failed [Ros04]. Failure [Bus09b, Gla02b, GT05, Got01e, Hat01, TTT^{+}02]. Failure-Early [Got01e]. Failures [EK08, Rob03, GF99]. Fair [CS96, Dak95d, For92, Dak95a]. Fall [Ros00]. Families [KM99, vdL02]. Family [Ebe01, GHW85]. Family-Based [Ebe01]. Fantasies [BALV03]. Fantasy [San98b]. Far [DM06]. Farewell [Dav96e]. Farewells [Dav98c]. Fashion [Boa97]. Fast [And94, DY93, HM01b, JS95, LL87, Sho04b, TS01]. Fast-Moving [HM01b]. Fast-Track [TS01]. Faster [HM02a, Pre01, Voa01b]. fastest [Ols95]. Father [Got09b]. Fault [Ano98h, DHK^{+}01, HKPS96, KMY91, LCS91, Voa01c, YD01, dLdCGR06, Hat97a, SSGT86]. Fault-Tolerant [YD01, dLdCGR06]. Faults [And03, VMM91]. Faust [GGJ^{+}89]. Fear [Pre98]. Fearless [JM07]. Feasibility [McC98a]. Feature [KLD02, NT02, Sav08, HGK09]. Feature-Oriented [KLD02]. Feature-Rich [Sav08]. Features [Bor85]. Federal [Got02a]. Feedback [Che95]. Feels [Dav96b]. Feet [Lis98]. Few [Wil03]. Few-Minutes [Wil03]. FFP [MHC03]. Fiction [MT99]. Field [PSHH04, Str01, Rei90]. Fifth [Mat00b]. Fight [Sim95a]. File [FO85, Gai94, Nie96a]. Files [Gut86]. Filing [GTF^{+}87]. Fillin [WP86]. Filling [CSR^{+}99, Dav95c, Phi96]. Finally [CL00]. finance [GMS^{+}96]. Find [AHM^{+}08, Hen94]. Finding [Bur95a, CCS^{+}03, CBK85, CBMP05, Dum95, KP99, RW04, VMMF00]. Fine [ARTZ03, HMK^{+}94]. Fine-Grained [ARTZ03]. Fine-Tuning [HMK^{+}94]. Finite [FKLW91]. Finite-element [FKLW91]. Fire [Ano98p, Bec01]. Firewalls [ABG^{+}97]. Firms [MD04, Ano85]. Firmware [DDMS86, DWH86]. First [Ano06l, Boo07b, Cur98, Got03, OCM^{+}84, PM98, Wil97c]. First-Year [Wil97c]. Fishing [Fri97]. Fit [Boo06c]. Fitness [MHB08]. Fits [Ano93o, Col94, RHK^{+}03]. Fitting [OC90b]. Five [Lin93, Ros98a]. Fix [DeM99]. Fixes [Got01d]. Flaws [SV03]. Flexibility [Erd05]. Flexible [Ano92l, CG08, DLR07, IAJR97, Mub08, OHM^{+}07]. Flight [Kan09, Lie02, Duk89]. Flip [SB95b]. Flooded [Rei95c]. flowcharts [Sca89]. fly [SF93]. Focus [Ano93-31, Ano98y, Ano98z, Ano99i, Ano99j, Ano99l, Ano99k, Bar95e, Bar96a, Bau94, Daw95a, MHC03, Moo01a, Nie97a, ES09, Rif09, SMTS09]. Focused [Gil96b]. Focuses [Ano95c]. Focusing [Deu88]. Fold [Ber96a]. Follow [Gla05a, GAS^{+}01, vS09]. Follow-the-Leader [Gla05a]. Follow-Up [vS09]. Food [Har06c]. Foolishness [Boa97]. Fools [Cha95]. Force [Pfl05]. Forecasting [KCF^{+}08]. Forecasts [Rei95f]. Forges [REM^{+}09]. Forgotten [Gla01b, Ner01, Spi05c, Tho98]. Form [Kis96, WP86, Jac96]. Form-Oriented [WP86]. Formal [BBRTL90, BH95a, Cha98a, DG90, Ger90, GCR94c, GFB96, Hal90, Hal96b, HA00, HSG^{+}94, JKM^{+}09, JBR09, Jul93, Kem90, LG97, MD86, ND90, Pal01b, SB90, Win90b, YJC09, FKU97, Law96, LST91]. Formalism [CS05, Mey85]. Formalisms [HGK09]. formality [Jac96]. Formalizing [HKK93]. Fortran [Ano93f, Rys95, MMS88, KB88, Rys95]. Fortunes [KCL93]. Forward [RE03]. Foundation [AK03b]. Four [DNR04, Ebe06]. Fourth [MJ88, VT88]. Fourth-Generation [MJ88, VT88]. Fplus [CAGR91]. Fragility [FH07]. Fragments [Har87]. Frame [Bas87c, Bas07]. Frame-Based [Bas87c, Bas07]. Framework [Ama07, AFC91, ACM^{+}07, BR87, BW96, Car00b, CN92, CBM90, DCF^{+}02, Fow09, Har87, HS00, Hum88, Kru05a, LL03, MMW86, OHM^{+}07, RKK05, Vit84, Sch96a, Shn97]. Frameworks [AK03a, Bas95, DFL00, Fac01, SA07, Spi06c, Sel90]. Frank [HEH06]. Franklin [Mil84]. Free [Gla03b, Gla08f, Kis96, Mau90, Sca04, TV99]. Free-Form [Kis96]. Frequently [Gla01b, Zve98b]. Fresh [AP00, Fle99]. Friend [Ano96h]. front [WB08e]. Fruit [Per95]. full [SM88]. full-screen [SM88]. Fun [Gla07a, Gla07b, GAS^{+}01]. Function [HPLH90, Kru05a, Fur97]. Function-Behavior-Structure [Kru05a]. Functional [And07, Ano96d, BK89, DG09b, How85, Lau03, MMW86, TP86]. Functions [CBK85]. Fundamental [Gla01b]. Fundamentals [Ano01f, Fid96, Pfl97a]. Funding [Cha99b, DCH04]. Funds [Cha98b, PM98, San95a]. Future [Ale99, Ano98o, Ath98, BLSS08, Bar95b, Bol04, Boo08e, BHS07, Cur88, Dav97e, Dug98, Gla99a, Gla02a, GHH^{+}09, Har06a, Kay98, KOS06, Leh98, Mic00, Mus85, Mye95c, Pes97, PT09, Ree05, Rei95b, Rei95c, San98a, WB09b, Lou95]. Futures [You98]. Fuzzy [Zad94]. Fuzzyspeak [Rei95b].

Galaxy [BB89]. GAMB&IT [LM84].

Gamble [LJ97]. Game

[Bal00, BC85, Got01a, Hei98, LM84, MM08a, Sca04, vWS09]. Games [GAS^{+}01]. Gap [BCKM97, KE88, McC02a, San08]. Gaps [CSR^{+}99, Phi96, TK89]. Garden [Rei87]. Gargantuan [Boo06b]. Garlan [Pfl97b]. Gas [DM93]. Gate [KR05]. Gatekeeper [Far06]. Gates [HVF^{+}00]. Gateway [GCR94a]. Gathering [Gon02]. Gauging [McC97c]. Gender [Kra95]. Gender-Quality [Kra95]. Genera [WMWM87]. General [Ano98m, AFC91]. Generalization [Gla00c]. Generalized [TL96]. Generating [Alo03, Far84, Mau90, CGR94b]. Generation [Ale99, And07, Bal87, CDPP96, Den91, DDPW09, Got01a, GAS^{+}01, Hor87, LJ88, Mar88, MJ88, Ste95a, Voa00a, VT88]. Generations [Hix90]. generator [Ano97i, Gue94]. Generators [BST^{+}94, HKN85, LFL^{+}90, Cle88]. Generic [RH89, RKJ^{+}97]. Gentle [Can95, LSB02]. Gentler [Ano92g]. GenVoca [BST^{+}94]. German [RRD07]. Germany [BHKR01]. Get [AW03, Ber93, BL92, Gil96a, Gla05d, Ano97e]. Gets [AS03, SA05]. Getting [Ano93l, Ano93k, Ano98h, E^{+}94, Got03, MVS^{+}99, Pat08c, Rei03a]. Ghosh [GAS^{+}01]. Giant [Got08a]. Gilb [Bur94]. GISs [CB98]. Give [Mac99b]. Giving [DP94a, Har06d, WB07d]. Glanville [Hor87]. Glimpse [Cha96a]. Global [Ano00i, Ano00h, BCKS01, BM01, Bel99, CA01, Cha96d, CW84, Cur00, DM06, Dam07, DAB06, ED01, EAS94, HKNS01, HM01a, Mat00b, Nie96d, RKK05]. Globalization [Cos03, Jon94, MW01]. Glossary [Ano04o, Tha03a, Tha03b]. GNU [Ana99]. GNU/Linux [Ana99]. Go [AS03, Bil94, Cha96d, HVF^{+}00, IH90, Sch04].

Goal [LX99, Rob09b, Ven05]. Goal-Driven [LX99]. Goals [CC95a, HKPS96, MK03, NE07]. Goddard [Lie02]. Goes [EL04]. Going [HM02a, Nie96d]. Gold [GB00, McC99a]. Golden [CS09, Dav98e, SC06]. Gone [Ano09e, Cod09, DeM09]. Gonna [Gra97]. Good [Ada95, And98b, Ano97g, Ano97d, BTC^{+}99, CJK^{+}02, Erd07f, Gla97, Gla01a, Hat97b, JOL^{+}00, Lon03, Lov94, Mac96, McC02c, McL05, PR08, Ree03, Rei00b, Rei02b, SLG^{+}00, Shu07b, UA03, VCM^{+}97, You95, Ber99, HMP^{+}01, LH89]. Goodbye [Gla09b]. Goodies [Ano92o]. Goodness [Boo06c]. Googling [Got08a]. Gordon [DKKK90, DKMS91]. Goto [Che02]. Goto-less [Che02]. Government [SV03, Voa99c, Got02a]. Government-Regulated [Voa99c]. Governments [AS03]. GQM [Lav00, vvSO^{+}98]. GQM-Based [vvSO^{+}98]. Grab [Ano92o]. Grade [Fer01b, Pal01a]. Graduated [ON92]. Graham [Hor87]. Graham-Glanville [Hor87]. Grain [KL88]. Grained [ARTZ03, Mye94]. Grammar [Far84]. Grammars [Mau90]. Grand [GB00]. Grant [Bar95b, Cha98b, PM98]. granularity [CH93]. Graph [RW90]. Graph-Parsing [RW90]. Graphical [HLS90, RGR85, vMC92]. Graspin [MMN^{+}89]. Great [Ano93-37]. Greatest [Voa99f]. Greece [Gla06a]. Greenfield [Ras03]. Grid [NE07]. Ground [Ano01h, GHW02, SH02]. Grounded [BLP96]. Groundwork [Ahu99]. Group [Mai02, vGvDSV01, OO92]. groups [Nie97a]. Groupware [BGB01]. Grow [Got01c]. Growing [BTC^{+}99, Got09a, MKMH07]. Grows [Fra99]. Growth [AvGCS07, BLP96, Cop99]. Guard [McC02b]. Guards [BFC00]. Guesstimation [JMDB00]. Guest [AP00, BD97, BF00, BB99b, Bol02, Byt99, Car95, CBM98, CW00, Cuk05a, Cur00, DM06, DP94a, DP03, ES09, Ebe09, EMP05, EFS04, EH95, Gla03c, GW07, HF02, HM01a, Hor98b, Hum00, JM07, JWC01, JMS06, KV07, KABM09, KB98, LFW00, LR00, LB06, MRE05, ME00, MNJP02, MCL97, MCF03, Mil01a, ME02, MS06, NZZ09, NZ07, NOK08, PL03, PC97, PT05, RvW07, Rif09, RL02, SE98, SM09b, SS04, SMTS09, Str99, Str01, TW07, Voa99b, VM00, Voa01c, Voa03, WE04, Wil03, Zdu09, dSSS^{+}09]. GUI [Lou07, RP07]. Guidance [Ano98n, GH00]. Guide [Ano92i, Ano96d, Ano99d, BDA^{+}99, DB99, Far06, Mas98, Par96a, Rog00, Ros98b, Wil97c, Wil97b]. Guided [JMDB00, PS90]. Guideline [MMM^{+}09, Mey04]. Guideline-Driven [MMM^{+}09]. Guidelines [Bar97a, Cha98a, FNY09, HH02a, Jor05, Pfl97c, PCS95, RD96]. Guy [Che97, HT04b].

Habitation [JRS^{+}09]. Habits [Mat00b]. Hackers [MC03b]. Half [Kru08a]. Half-Life [Kru08a]. Hallucinations [Lis98]. Hammer [SMTS09]. Hand [Spi05d]. Handbook [DW00, LRO02, Ano93m, Bei97b]. Handhelds [Got01b]. Handling [Knu87, KT04, Mai09c, NCC^{+}02, Ros98a, WB06f, WB07e]. Hands [Ano97g, Ebe01]. Hands-On [Ebe01]. Handwritten [Dya95]. Happen [Gla02d, Mac96]. Hard [GAO09, LB00, RS84, Rif01b, Rob09a, Whi00, ZLJW95, GAO95]. Hardly [Bas95]. Hardware [Bil94]. Hare [Gra00d]. Harissa [MS99]. Harlan [Bri96]. Harmony [Dav98d]. Harms [Pat08a]. Hash [CBK85]. hashing [KR88]. Haystack [Whi96]. Head [Ano99n]. Heads [Cha99b, DAS^{+}07]. Heal [Dav96f]. Healing [Mcl98, Mou01]. Health [Gla06c, JB04]. Healthy [Hen99]. hear [MR95]. Heart [SK03]. Heaven [Med05]. Heel [Voa99e]. Held [Bac95c]. Helix [FO85]. Help [Ano96r, EG04a, Got01c, Hef95, MSG^{+}07, MC03b]. Helped [DS97]. Helping [Hef95, vWAS06]. Here [Gil96a, HVF^{+}00]. Heresy [Gla04d]. Hero [Gla08e]. Heroes [Bac95a, Bak95]. Heterogeneous [SN89]. Heuristic [Mad97]. Heuristics [Sot01]. Hewlett [Pfl95b]. Hidden [Fer01a, Voa99b]. Hide [Erd09c, VMM91]. Hiding [McC96f]. Hierarchical [MDM^{+}94]. High [BCC^{+}08, Dav86, FVHF95, Har05c, HKPS96, LC85, LM04, NO91, OCHW91, Pit00, PRSV97, RP06, Ros00, SJM02, Voa99a, Wil96, FKKM89, Ano92i, Mat00b, Pou99]. High-Assurance [Voa99a]. High-Integrity [LM04, RP06]. High-Level [Dav86, LC85, SJM02, FVHF95, HKPS96]. High-Payoff [PRSV97]. High-performance [Wil96]. High-Performance-Computing [BCC^{+}08]. High-Tech [Har05c, Ros00, Ano92i]. Higher [Das94, DJM04, HKPS96]. Higher-Education [DJM04]. Hill [Har03b]. Hints [Lam84]. Hire [Cos94]. Historian [Bas87b]. Historic [Gla99a]. Historical [Bie00, Cus89]. Histories [PWS^{+}95]. History [Ano08e, Ebe08b, JL03, SS08, Tar09]. Hit [Mye95a, Spr95]. Hobson [Bri97]. Hold [Got02c]. Holy [ZJ00]. Home [JRS^{+}09, Nie95a, Nie95b, Spr95]. Home-Page [Nie95a, Nie95b]. Homeopathic [Jac98a]. Hood [Cha98b]. Horse [Gla02e, MA07]. Hostile [Ale03]. Hot [GB00, Rei04b, Rei96c]. Hour [Gla94a, Ols93]. House [Ano93l, BFC00, O’D98, VE05]. HSDPA [MMSM09]. Huge [Mat01]. Hughes [HSW91]. Human [AJM06, Ano92q, DSBM99, DL98, Fis89, Lin85, MW91, PWS^{+}95, dSSS^{+}09]. Human-Computer [Ano92q, DSBM99, Lin85, PWS^{+}95, Fis89]. Humane [Ano92b]. Humphrey [Mye95d]. Hunt [Tur03]. Hurdles [NCC^{+}02]. Hurt [EA03]. Husband [Gla00b]. Hybrid [MS99]. Hype [Gla99c, HPG^{+}00, MVS^{+}99, RC95]. hypercube [RWS88]. Hypermedia [LF05]. Hypertext [Big88, CR92, GS90a]. Hypertext-Based [CR92]. Hyperview [MR90].

i* [GYvdR06]. i-Mate [LG03]. IC [KM86]. ICASE [Ano93c]. Icon [Har03e, Kar88]. ICON-based [Kar88]. Iconic [IH90, RH89]. ICSE [Mat96, Mye95a]. ICSM [Ano97h]. iDAVE [BHJM04]. IDE [MKF06]. Idea [Ano09e, DeM09, Mat00b]. Ideal [CC96]. Ideas [Ano97g, Kru08a, SJM02]. Ideate [BGL^{+}09]. Ideawork [Ano84b]. Identifying [BI96, EG04a, HG00]. Idioms [Cop97]. IEC [JKC04]. IEEE [Ano84a, Ano96h, Ano97e, Ano98n, Ano99o, Ano01i, Ano02a, Ano03a, Ano05a, Ano09a, Erd07b, FS98, Har03a, Hor97, O’L09, Pos85, Sch07]. IEEE/EIA [FS98]. If [Mel95, Dak97b]. Ifs [Vo90]. ilities [Voa01a, Voa04b]. Ills [Ano09b]. Image [PCS95, ZSF91]. Image-browser [PCS95]. Image-Processing [ZSF91]. Imaginate [HT04a]. Imagine [MGR04]. Imaging [FHMU85]. Immersive [CNL99]. Impact [AHM86, BMJH96, Erd05, Kal94, SV02]. Impacts [Cri06, PSHH04]. Impatient [Ano93-33]. Impediment [Pre96a]. Impending [HH02b, Nie96a]. Imperative [Bud91, MM05]. Implantable [MTKE94]. Implement [Rob99]. Implementation [ANB99, BR85, BM04a, HL99, HKPS96, Kic96, Pud04, WC09]. Implementing [AGK^{+}87, BT05, CS97, DFLS05, DAEE08, Ebe01, Fen90, HF97, KMY91, Kil01, Nar85, Par96b, SvHB02, SK95, SO97]. Implications [Dak96e, DBS09, Gra99b, GRS02, PSHH04]. Implied [Dak96b]. Imply [WB07b]. Importance [BH02, BMM05, DH93, Iel96, Lar01a, LM03]. Important [Gla00a, Mey04]. Improbable [Dak96d]. Improve [Ale95, Ano93-31, Che95, EM96, Fen96, Hay03, JP03, JS08, Mai06a, Mye95b, Sch96b, SJM02, Tar09, Voa00b, WLC01, KC97]. Improved [Hay03, OC90a, THV09]. Improvement [Aae03, AH96, Ano92a, Ano92e, Ano93-38, BZM^{+}95, Cla00, CF02, DLSZ05, DS97, Dio93, DAEE08, FO99, Got08a, HG00, HSW91, Hum95, Jak98, JB96, KM02, MNA05, McC02e, PSV94, PR94, Rob95, SM08a, TM94, TS01, Wei98a, Wie99, Yam99, vS04, vS09, Hal96a, Vis94, Coo95]. Improvements [KP99, Nie96e]. Improving [Ale01, Bie00, BBGK07, Che01, CF02, Dec01, EL02, Fin96, HDS05, HHM94, JKP^{+}05, Kan09, KCC^{+}01, Lev91, MKHV06, Mir01, Mot06, NSV^{+}08, PDK05, Smi91, TP98, WSL^{+}99, WHG^{+}09, Ols95]. Improvisation [Dyb00]. Improvisers [MM08a]. Impurity [Gla08d]. In-House [Ano93l, VE05]. Incompatibility [MP07]. Incorporating [JHMW07, Kod04]. increase [Voa97a]. Increasing [GG08]. Incremental [DCH04, Got00, Raj00, RG04]. Increments [NT02]. Incubator [Moo01b]. Incubators [DGC^{+}07]. Indecent [Cha98a]. Independence [SK95, Var88, Cod88]. Independent [CCD09, SM97]. Index [Ano95a, Ano96a, Ano97a, Ano98a, Ano99o, Ano00a, Ano01a, Ano02a, Ano03a, Ano04b, Ano05a, Ano06a, Ano08a, Ano09c]. India [Ano98g, Bag99, GAS^{+}01, KRF01, MHC03, Moi01]. Indicators [Sel09a]. indigenous [KCC^{+}01]. Industrial [Ano98g, Bro96, Cox90, DNS^{+}06, End93, HF95, Jac93, Osh98, SH93, vvSO^{+}98, vdLLM09]. Industrial-strength [Bro96]. Industrializing [AO89]. Industries [BM04b]. Industry [Ano92s, Ano93l, Ano93w, Ano95c, Ano96o, Ano98g, Ano02i, Ano03j, Bag99, BCKM97, BC08, CP02, CT90, Cha97d, Cur95, DFL00, Dua02, Ebe08a, Ebe09, GCB^{+}06, GFB96, Gro04, HG00, Ju98, Ju01, JMS02, JMS06, Ksh05, Mat96, Moi01, Mor00, Rei96a, Rei96b, RJI02, Str99, Str02, Ter01, Voa99c, KCC^{+}01, WC89, Ano01i]. Industry-Academia [Ano93l, BCKM97]. ineffective [HMP^{+}01]. Infamous [Erd08c]. Infeasible [Bha96]. Inferred [PM99a]. Influence [DJM04, MD97]. Influences [McC00d]. Informal [HKPS96, WEF01]. Information [AW91, ABB^{+}00, Ano93p, Ano99e, BP94, BMM04, BLWG99, Blu87, Bor85, CR88, CC95a, CSR^{+}99, Dav95e, FK04, GPW97, Hal96b, Har03f, Has05, JG08, KM02, McC96f, OBSN08, SCSC04, Shn94, Voa99e, Pfl97a]. Information-System [AW91]. Infosec [MM00]. Infosys [RJ02]. Infrastructure [Bol04, FK04]. Infrastructures [BCH^{+}01, Voa99d]. Inheritance [Ano93-30, DT94, HO87]. Initiating [MNJP02]. Initiatives [Lie02]. Injection [Ano98h]. Innovation [Boo09d, BFC00, Ebe07]. Innovations [GH00, Rif01a]. Innovative [JBHD08, KIB^{+}08]. Input [PM98]. Inquiry [PTA94]. Insecure [Kus95]. Insecurity [Lai08, Mic99]. Insider [Ano92k, Boa97, Coc97, Eic03, Ju98, KC96, Mat95b, Var95, Wat97, Ano92l, Chr94, Gri95]. Insight [DSBM99, Mat00b]. Insights [KNR^{+}01]. Inspect [Xu03]. Inspecting [SS08]. Inspection [ARTZ03, BP94, Bif00, DRW03, GV94, Hat95, HKPS96, HPG^{+}00, MDWTR93, PL03, RBBP03, Ter96, Wel93, O’N97, Rus91]. Inspections [DS07, Hat08, PV97, SS08, TRW03, vGvDSV01, ABL89]. Installshield [San98e]. Institute [BHS85, Mye95d]. Instructor [BTC^{+}99]. Instrumentation [SR85]. Instruments [JML00]. Integrate [NC92, War89]. Integrated [BSGF98, CN92, CFTZ89, Dow87, GGJ^{+}89, HBP92, MMM^{+}09, Moo99, PF92, PWS^{+}95, Rov86, SB95a, Som05, KS89, RSE^{+}90, SSGT86, MBO93]. Integrating [AFGD01, CW00, CT02, EMP05, Har87, Jar92, Kän97, Kem90, SN89, SBK86, WEL02, Mey91]. Integration [Ano93i, Bol09, Coy00a, DSLB03, HS91, MS92, Mur95, OHM^{+}07, TN92, WSG^{+}05, YJ07, Yin94, Sel90, Son97]. Integrity [CE09, Gla06c, LM04, RP06]. Intellectual [Bri96]. Intelligence [STSK92]. Intelligent [Ano93a, Ano93p, KFP88, LLH^{+}03, RRD07, TR88]. Intensional [DW90]. Intensive [CS97, Gre01, Koh05, Weg84]. Intent [Ale03]. Interaction [Ano92q, DSBM99, LANC07, Pat08d, PWS^{+}95, Fis89]. Interaction-Based [LANC07]. Interactions [MPT86, RP09]. Interactive [CBK85, FLKMR91, SA07]. Interconnections [DW00]. Interests [Cha97e]. Interface [BHHM86, Bil94, CN07, CN95b, DSBM99, Dum95, Etl96a, Fon96, Fow01c, Got09b, GN95, HLS90, Hix90, Hol04, Iel96, Jef97, Kin97, LH01, Lov94, Ner01, Nie95a, Nie96c, Nie96d, Nie96b, Nie97d, Nie97c, Nie97a, OO92, PWS^{+}95, PP03, PRSV97, Pue97, Rog00, RW92, Shn91, SR91b, VB97, Har89b, LHB^{+}90, Mye89a, Nie96e, Per90, TK89, Nie95b]. Interfaces [Ano93j, Ano93q, kCCP^{+}95, CN95b, ES92, ELST95, Fow02a, GSS^{+}06, Jef97, Lin85, MGG^{+}95, PWS^{+}95, SL97, FKKM89, Shn97, Tho89]. Interfacing [FF08]. interior [ALMS92]. interior-point [ALMS92]. Interleave [LSM84]. Intermedia [Rei94a]. Intermediate [Ros98b]. Internal [Wes08]. International [Cha99c, Moh01, Got02a]. Internationalizing [RP95]. Internet [Pou99, Amb02, Ano93r, Ano95c, Ano98i, Ano98t, BRL^{+}03, BS03, Cha97d, Cha98a, Col02, DW00, HF02, IAJR97, LSB02, Mic99, MDM03, PM98, PAD^{+}98, Rei02a, RG02, SM99, Sim95b]. Internet-Based [Amb02, BS03, PAD^{+}98]. Internet-Speed [BRL^{+}03]. Internet2 [Cha97b]. Internetwork [Kus95]. Internetworking [Ano93f, DW00]. Interop [Cha97d]. Interoperability [MHC03]. interpretation [Cus89]. Interprocess [ACF87]. Interrupts [vBv98]. Interview [Ahu99, Ano97f, Ano98q, Bar97c, Bur94, Bur95a, Cha97c, Cha97d, Dav98e, HEH06, Kay98, Ker99, Ko09, Mey96a, Mey96c, Mey96b, Mil01b, Mye95c, Mye95d, Mye01, Ray99, San95a, Ano92m]. Interviews [Mil01b]. Intranet [Rei02a]. Introducing [Ano01i, BMGT01, Bus09a, DSBM99, DN97, Jan04, JWC01, NS99, PHRS07, Ras03]. Introduction [AP00, BB89, BL90, BD97, BF00, BB99b, Bol02, Byt99, Can95, Car95, CBM98, CW00, Cuk05a, Cur00, DM06, DP94a, DC04, DP03, ES09, Ebe09, EMP05, EFS04, EH95, Gla03c, GW07, HF02, HM01a, Hor98b, Hum00, JM07, JWC01, JMS06, KV07, KABM09, KB98, Kru05b, LFW00, LSB02, LR00, LB06, MRE05, ME00, MNJP02, MCL97, MCF03, Mel05, MRA98, Mil01a, ME02, MS06, Mye89a, NZZ09, NZ07, NOK08, PL03, PC97, PT05, RvW07, Rif09, RL02, SE98, SM09b, SC00, SS04, SMTS09, Str99, Str01, TW07, Voa99b, VM00, Voa01c, Voa03, WE04, Wil03, Win90b, Zdu09, dSSS^{+}09, Wil97c]. Introductory [CT02]. Intrusion [KH97, MCA00, PCOM97]. Intrusions [BS03]. Intuition [Gla08a]. Intuitive [CKMV95, Str96]. Intuitively [Gla98b]. Invariants [Pai86]. Invent [Rob02]. Inventing [Kay98]. Inventive [RS08]. Inventorying [BMM04]. Investigating [AH90]. Investment [Ano03j, EFS04, HB06, Ree06a]. Investments [GA06]. Invisible [Whi01]. Involved [Gla01c]. Involvement [Cla98b]. IOGen [LJ88]. IP [Ano93f]. IPSE [BM92]. Ireland [Coc01, KCC^{+}01]. Iron [Bud99]. Irrelevance [Boo07c]. Irrelevant [JG08, Spi05b]. ISBSG [LWHS01]. isn’t [EJW95]. ISO [BFC00, Col94, JKC04, MHC03, Pau95, PWS^{+}95]. ISO/IEC [JKC04]. ISSRE [Tau95]. Issue [Hor97, JD97, Lub91]. Issue-Based [Lub91]. Issues [Bar95e, BLWG99, Cha98a, Col02, Coo96, DAB06, GP87, MS04, O’N97, PR08, Zan94, Zan92]. IStar [Dow87]. Itch [SLG^{+}00]. Iterative [Hen94, Sot01]. ITracker [SC05]. Ivar [Mye02].

J [MRA98]. J2EE [AHZ02, Les06]. Jackson [Gre96]. Jacobsohn [Ano96h]. Jacobson [Mye02]. James [San95a]. Japan [Ano85, AMM^{+}89, Mat01, Tom87]. Japanese [AO89, Fuj99]. Jargon [Ano97g]. Java [DFL00, MRA98, San98e, Frü99, Ano97i, Ano98g, Ano98h, BALV03, Fra98, Hor00, Kod04, Lai08, LRO02, Mar98, MRA98, MS99, MKF06, Nau98, Pon98, San98f, Spi05b, VMMF00, Wil97c]. JavaCC [Kod04]. Jazz [Fro07]. Jetpack [HRS08]. Jim [GF99]. Job [CCS^{+}03, LD95, Mat00b, RW92, Spi09b]. Job-shop [LD95]. John [ECW09]. Johnny [Yam98b]. Join [Par05]. Joining [CSR^{+}99]. Joint [CT90, Ano85]. Joke [SM01]. Jolt [Cha99a]. Jones [Cha97d]. Journey [Pit00]. Journeyman [SC00]. Joy [HBB01]. Judgement [PWS^{+}95]. Judgment [Jor05, JBR09]. Jump [Ano98h]. Jump-Start [Ano98h]. JUnit [Lou05]. Just [Ano93a, Ano96p, App02, Cha95, Dom98, GM09, Har05a, Har05b, Rai07c, vBv98].

Karel [Etl96b]. Keep [Bar97b, Gra00b, HT04b]. Ken [Cha97c]. Kernel [Che84, Kor96, Spi90, dAdS99, SR91a]. Kerth [Mey96c]. Key [Ano97g, Bar96b, Cla98b, DDS^{+}97, DC04, Ebe06, GV94, Mey96a, MA00, Rob95]. Keynotes [Mye95a]. Keys [Dak96a, LL87, Mat00b]. Kill [Far06]. Kinder [Ano92g]. Kinds [SMTS09]. King [WJS97]. Know [Ano93l, Dak95c, DJS08, DBS09, DD09, HSB^{+}08, Har03b, Har05f, JMVS06, MM08a, McC02d, NE07, RAT^{+}06, Zve03]. Knowing [Hay04, MJP^{+}00]. Knowledge [AO88, Ano93s, BDA^{+}99, Cha96a, DAB06, DBS09, DVG09, DFLS05, DB99, EM93, EMB^{+}99, GP05, HN90, JM02, KS04, KNR^{+}01, KSMS02, Lie02, PTP^{+}88, RJ02, Ram02, RFHM86, Rei02c, RL02, Sev87, SMR93, SRS^{+}86, Sym88, Tiw04, WHC02, Wie84, Zdu09, SSGT86]. Knowledge-Based [AO88, Ano93s, PTP^{+}88, RFHM86, Sev87, SMR93, HN90]. Known [And03]. Korea [WC89]. KosovoNet.org [Ber99]. Kudos [CCS^{+}03].

Lab [End93, For92, Bur95a]. Labor [Jon99, San97b, Str99]. Labs [JL03]. LaFleur [Mey96a]. Lands [Sto99]. Language [BL90, CCD09, CDM^{+}05, FL86, GS90b, Gla08d, HTLPS09, HKPS96, JRS^{+}09, KKM87, Kod04, KSS84, LV01, Lv85, MPT86, MC03b, Pem87, Rys95, SHZ^{+}98, SLG^{+}00, Spi06b, TV00, TP86, WHG^{+}09, Wie98b, Bal89, BB89, kCCP^{+}95, Got02a, Nak96]. Language-Based [KKM87]. Language-Independent [CCD09]. Languages [Ano96s, Ano08b, Bor85, CL91, Cha87, CM07, Erd08f, Fow09, GHW85, HR95, JDT01, Jon07, Kor86, RP06, Sha84, SFVFM08, Spi05b, TW07, Wol89, WKH92]. Larch [GHW85]. Large [AHZ02, AEP96, BRS^{+}85, Cha96b, CS90, DM96, FHMU85, HF95, KT04, LL92a, Mar96, MA00, Mye94, NGBR05, NS99, PS84, Rov86, SRK^{+}09, Sel09a, THDK06, WCW85, ZZ93, GMS^{+}96, MW96, Mye89b]. Large-Grained [Mye94]. Large-Scale [BRS^{+}85, LL92a, MA00, NGBR05, SRK^{+}09, Sel09a, THDK06, ZZ93, Cha96b, HF95, Mye89b]. Last [Hei01, Mat00b]. Lasting [Mat00b]. Later [PT05]. Latest [Gre96]. Lau [Ko09]. Launching [Gre01]. Law [Big96, Cos01, Cos94, Dak95b, Dak95c, Dak95d, Dak96b, Dak96a, Dak96e, Dak96c, HG99, Spr95, SMK^{+}00, Tom87, Tru94, Zan92, Zan94, Sam88, JOL^{+}00, McC99b, SLG^{+}00]. Laws [Fei07]. Layer [LSG^{+}05]. Laying [Ahu99]. Lazy [BCMV05]. Lead [Gla94a, VE05, VV00]. Leader [Gla05a]. Leadership [Ath98, BTC^{+}99, HVF^{+}00, Nis00]. Leading [Mey96c, Sel09a]. Leah [Pat09]. Lean [DeM95a]. Leaner [McC97a]. Learn [Har05b, KC96, Pfl02, Rob09a, Sim04]. Learned [BGB01, BCH^{+}02, CFQ07, Dam07, Dan04, Di 05, DC04, End93, GB95, GF99, Gri95, Ker99, Mat00b, NK04, Pfl93, Pit93, Ras03, RBBC03, Rom90, RW98, Fis89]. Learning [BGS01, BM92, Bus09b, Car98, Erd05, Gla04b, HD00, Har04e, Hsi93, KAS93, Kem92, LANC07, Mat00b, Nie97c, Nol99, Pre01, QBN^{+}06, Rob05a, SvHB02, VBM^{+}02, VK07, GSB03, NR89]. Leave [BDS02]. Leg [Bud99]. Legacy [Ado96, Ano93d, Ben95, BLWG99, Che02, Coy00a, Dav95c, Got01e, RW98, SE98, Sch98a, Sne95]. Legal [Car00b, Car88, Dak96e, Gra99b, Gra00d, Sam90]. Legislation [Bar95e, Gra00c]. Legitimacy [Fra95]. Legs [HT04c]. Lemmingineering [Ano93-29]. Less [Ano93-32, Gla98d, Mat96, MC00, Che02]. Lesson [SY99, Tho98]. Lessons [Amb02, Ano06k, BGM06, BGB01, BCH^{+}02, CFQ07, Coo95, Dam07, Dan04, Di 05, DC04, End93, Fis89, GB95, GF99, GV94, Gri95, Hsi93, Joh95, KC96, Ner01, NK04, Pfl93, Pit93, Ras03, RBBC03, Rom90, RW98, Sta99, Wel93]. Let [Jak00, Nie97d, SHZ^{+}98]. Letters [Ale95, Ano96q, Ano96r, Ano97g, Ano98r, Ano98s, Ano99q, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano05k, Ano05h, Ano05i, Ano05j, Ano06p, Ano06m, Ano06n, Ano06o, Ano07i, Ano07e, Ano07f, Ano07g, Ano07h, Ano08c, Ano09d, Bac95c, Bak95, Bas95, BHR^{+}98, BPS^{+}03, BTC^{+}99, Boy97, Bri97, CCS^{+}03, CJK^{+}02, CSR^{+}99, CM96, DMN^{+}98, DDS^{+}97, DB97, Dic95, DB96, DCF^{+}02, Eme95, EMB^{+}99, FV97, Fri97, GSB03, GN95, Gue95, HVF^{+}00, HMP^{+}01, Hou95, HPG^{+}00, JLLP03, Joh95, JOL^{+}00, JD97, KCC^{+}01, KS96, Kla95, Kla97, Kle98, KHA^{+}03, Kra95, LCE^{+}01, Mar98, MVS^{+}99, MMM^{+}01, MJP^{+}00, Mye98, NCC^{+}02, PB98, RHK^{+}03, RP98, Rin97, SLH01, SHZ^{+}98, SLG^{+}00, SM95, SS95, SGW02, Sta97, SB96, SB95b, SMK^{+}00, TTT^{+}02, TB97, VD96, WSL^{+}99, WGK^{+}02, Wan95, WEF01, Wei98a]. Letters [WAGT95, CC95c]. Level [Bor85, Dav86, Eic03, Gil96a, LC85, MD02, OCHW91, Ros98b, SJM02, FKKM89, FVHF95, HKPS96]. Levels [Das94]. Lever [Sto03]. Leverage [Sta99]. Leveraging [BCKS01, FS98, RKK05]. Lewis [Har03c]. Lexical [Dya95]. Lexicon [Gre96]. Liability [Spr95, KCC^{+}01]. Liable [Ano93-39, Tru91]. Liars [Cha95]. Liberalisation [Ano98g]. Libraries [Bac95b, Has05]. Library [BHHM86, BAB^{+}87, Di 05, Win88]. Libre [GBdlHQCG^{+}00]. License [Ano98n, Dak95c, Dor99a, JOL^{+}00, Mel95, Moo03]. Licensing [CTW98, Kru08b, Sch99]. Life [Ano98i, Ano98t, BLP96, Bas87a, Bud99, Ebe06, GS90a, Kru08a, Lau98, Lot97, San03, Sim95b]. Life-Cycle [GS90a]. Light [Ano99p, Hir07, Mat95b, Rei96a, SHZ^{+}98]. Lightweight [EL02, HA00, JDT01]. Like [Boo09a, Dav96b, Hay03, MGR04, Shu07b, Gla02b, SGW02]. Likely [Bar96b]. Limitations [MW91]. Limiting [Str96]. Limits [Lar01b]. Line [BHJ^{+}03, CJNM05, Don01, Ebe01, Gla01a, Jaa02, KLD02, KMSW00, Nor02, SV02, TH02a, WHG^{+}09, Ano08e, MSG96]. linear [ALMS92]. Lines [BCM^{+}04, MNJP02, VK07]. Linguistic [JML00, NGBR05]. Linguistic-Engineering [NGBR05]. Link [Gra02, Toc97]. Linked [SI91]. Linked-List [SI91]. Linking [Cha97f, LH01]. links [Ano97e]. Linux [Ana99, Ano98u, Ano99i, Ano99j, Ano99l, Ano99k, AS03, BR99, BB99b, Bol99, Ded09, Gla99c, HHJC99, JWKZ99, Ksh04, Lei99, Mar04, Ray99, Ros98b, San98a, WSL^{+}99]. Lips [Wie98a]. Lisp [AGK^{+}87, GG89, ZHL^{+}89]. List [SI91]. Listen [Pes06]. Literate [Ram94]. Literature [Cop97]. Little [Erd08f, HKPS96, KS04, Rei02c, Rei96a]. Live [Jak00, RC95, Zve98a]. Lives [Mun98]. Living [Ale99, Bal87]. Loadable [dAdS99]. Local [LC85, SN89]. Localization [Col02]. Lock [CH93]. Lock-granularity [CH93]. Lockheed [DHMR02]. Logging [Spi06c]. Logic [Dav85, DW90, HKPS96, HR95, KE88, Zad94, Tic89]. logic-programming [Tic89]. Logical [Var88, Cod88]. Loglisp [Nar85]. Long [Got02d, Mun98, Rok03]. Long-Term [Got02d, Rok03]. Longer [Etl96b, Zve98a]. Longitudinal [FO99]. Look [Ano98h, BM91, BM09, Gla08b, HKPS96, HC91, JMVS09, Lie02, ND90, NSV^{+}08, Pfl95a, Rin97, Sch95b, Oma89]. Looking [Acl88, Ano93x, Gla02b, WB06d]. Looks [Cha98a]. Loops [Wol88]. Loral [SB96]. Losing [Gra99a]. Lost [GM09, Lon03]. Lot [Hum01]. Lots [Rei96a]. Low [PRSV97]. Low-Effort [PRSV97]. Lowball [CC95c]. Loyal [Bud99, Cur01, Ded09, Gla98b, Gla98d, Gla98e, Gla98c, Gla99b, Gla99c, Gla00c, Gla00a, Gla00b, Gla00d, Gla00e, Gla01b, Gla02d, Gla08b, Gla09c, Gla09d, Jac98b, Jac99, Rif01a, Rob01b, Spi98, ZJ00]. Lua [Hir07]. Lucky [BTC^{+}99]. Lure [Pal96]. Lying [GRM08]. Lyn [Bei97b].

M [Cod88]. Mach [BRS^{+}85]. Mach-1 [BRS^{+}85]. Machiavelli [Pre95]. Machine [Dav00, Fra98]. Machines [Ano93a, Ano98q, TH02b]. Macintosh [MP88]. MacWorld [Rei96a, Ano98o]. Mad [Bur95b]. Made [Cos94, Med05]. Madness [Ano98q]. Mail [SN89]. Mailbox [Gla94a]. Maintain [CRR09]. Maintainable [ES92]. Maintaining [SLG^{+}00, Voa98a, WMH93]. Maintenance [Ano97h, Ano05m, ABFP86, Bas90, CAW97, Gla98d, Gla02a, KFP88, OC90a, ONR^{+}90, OC90b, PH01, Ram00, RBBC03, Sch98a, Sha96, Tar09, ZAV02, HHM94]. Maintenance-Oriented [ZAV02]. Major [DP03]. Make [BPS^{+}03, CKMV95, Fow03c, Hor96, Jon07, KKC93]. Makes [Ano93-36, Chu93, PV97, Rif01b, Rob09a, Spi05b]. Making [Ano92i, AD03, BBE^{+}88, BB91, BFC00, CDD08, Dav97b, DeM90, Fai02, FCH^{+}08, Fer01b, Gla08a, Gla09d, Hsi96, Hum95, Kau99, KB02, Lon03, Mey96a, NOK08, OBSN08, Pal01a, Pfl95b, Rot00, Str96, Voa01c, Xu03]. Malaysia [Sch99]. Malicious [Ano99e, MM00, VM00]. Mammoth [Bol00]. Man [Bro95, Mye95a]. Man-month [Mye95a]. Manage [GS90a, Tho94]. Management [Ada95, And98a, Ano92c, Ano93p, Bas87b, Ber03, BD97, BT05, BD94, BH95b, Bro96, BS03, CT93, Car97, CSSP05, CC99, Cha96b, CSHA07, CJNM05, Con96, CS97, Coo95, CG08, CSR^{+}08, CP06, DL03, DBS09, Don06, Fai94, Fai05, Feo97, Gar05, GPW97, HVF^{+}00, HS97, Hix90, JKP^{+}05, KR05, KNR^{+}01, LM03, Lie02, Lin87, Lis97, LLH^{+}03, Lon02, MC94, MR09, McK02, MA00, MDD09, MSK00, NGBR05, ND05, PWS^{+}95, Phi96, PT05, RJ02, Ram02, RKP08, Rei00a, Rei00b, Rei01, RSE^{+}90, Roy00, Roy05, RL02, SM08a, SCCA04, SSC06, SJM02, SM01, Tom03, VE05, WHC02, Wie84, WWD97, YNM^{+}91, Boe91, Fen90, For88, Har89b, LC87, Ros90, SO97, Lew98, Mat00b]. Management-aided [BD94]. Management-Speak [McK02]. Manager [And94, And98a, Aus99, Bak01, Ber96a, Boe00, Cla02, Cos01, Cox97, DeM95a, Fer01a, Gil96a, HS97, Hoh99, Hum95, LJ97, Lis98, LRO02, Lot97, McC00e, ND90, Pet99, Phi96, Pre96a, Pre96b, Pre98, Pre00, PM97, Rei00a, Rei00b, Rei01, Rob99, Ros00, Roy00, San98f, Smi91, Sta99, Tho98, Voa98b, Wei98b, You95, Cha95]. Managers [BTC^{+}99, Moy97, Pre98]. Managing [BP94, BG98, CAW97, Cha97c, Con96, DM96, Dec01, DAB06, Fav02, For97, Gil96b, HdLGG03, Kis96, Leh98, Lit05, MSG96, MNA05, MRA98, MSK00, Mug08, Nes98, NJJ^{+}96, Nor03, Pit93, RKK05, SM08a, Sch98b, Wil97c, Win90a]. Managment [Mey96a]. Maniacs [Cha97c]. Manufacturing [BHR^{+}98, Bin97, Lev91, LRO02, Sch96a]. Many [Erd08f, GA04, Mye98, STM88]. Mapping [CS05, HDS05, Mil02]. Marietta [HHM94]. Mark [Dav97b, JD97, TR05, WEF01]. Market [Dak95b, GB00, GAS^{+}01, Got09a, Gue95, Mat01, Rei00c, US04]. Marketecture [Hoh03]. Marketing [PM00]. Marketplace [MS04]. Markets [Sch99, Wes08]. Markup [SFVFM08]. Marriage [Chi02, Har03d]. Mars [Mac00, Sto99]. Martin [DHMR02, HHM94]. Mary [Mil01b, Pfl97b]. Mashups [GG08]. Masses [Lew84]. Massive [GF99]. Master [PTH^{+}09, SC00]. Mastering [Bro99, RS01, Rog00]. Match [Med05]. Matching [And03]. Mate [LG03]. Material [HT03a]. Materials [Pes06]. Math [HMP^{+}01, MMM^{+}01]. Mathematical [FVHF95, Kan93]. Mathematical-Modeling [Kan93]. Mathematics [BH02, Gla00a, LCE^{+}01]. Matters [Gau05]. Mature [DP03, Lai93]. Maturing [Cur95, LSB02]. Maturity [Ano93u, Cur00, EMW^{+}97, EL04, EM96, Hum88, MD02, PCCW93, Pit00, RKK05]. Maxims [DSB99]. May [Bar97b, Car97, Rei95d]. maybe [Jac96]. McConnell [Har03c, Dav96a, SMK^{+}00]. MDA [Tho04]. Me [Bat98, CJK^{+}02, Dés08, Gla05d, Mac99b]. Mead [Big96]. Mean [App02, CL00, DeM95a, Har05a, Sei03, Spe99]. Meaning [Gla02c, Got01d]. Meanings [GA04]. Means [DeM95a]. Measure [EA03, Erd08c, RHK^{+}03]. Measured [Hum96]. Measurement [Ano93-36, BB99a, BFC00, Bur95b, Cla02, DG09b, ELM90, Erd08d, FKN02, JD97, Kau99, Lav00, LK03, MRA98, PR94, Rif09, Rom90, SKT^{+}92, SM92, vvSO^{+}98, OJ97, PJCK97, Put91, Sel90]. Measurement-based [PR94]. Measurements [Tia04]. Measures [Ano92i, Ano99f, BL92]. Measuring [Ano97g, Ano98g, Boo08b, EL04, JKC04, KKJ91, MK92, Pfl96, Rif01b, SM01, US09, vS04]. Mechanism [CM96, ZSF91]. Media [Ano93e, Ano93p, Ano93v, DESG00, ZMT05, Ano92i, Ano93f]. Medical [FHMU85, HdLGG03, Hol04, MTKE94, Sim04, TH02d]. Medium [KG07, KMSW00, KT04]. Medium-Sized [KMSW00]. Meet [Gla05b, MK03, Rei95a]. Meeting [Sch04, Sha96]. Meets [Erd09a, Tom03]. Megaproject [Ven05]. Megaprojects [Gil96b, Kis96]. Melding [KG87]. Mellor [FHRK93]. Members [Ano01i]. Memorial [Che02]. Memory [Boo08f, OC88, San98f, Bec90, For88]. Men [Hou95]. Menace [Aus99]. Mentoring [Ram01]. Merger [Ano93-35, Mic00]. Message [Ano93r, Cox84, Ols85, Rei90]. Message-Based [Ols85]. Message/Object [Cox84]. Metaclasses [DP94b]. Metadata [BBMH05, Fow02b]. Metadesign [FNY09]. Metamodeling [AK03b, FGSK03]. Metamodels [NJJ^{+}96]. Metamorphosis [WUG03]. Metaphor [Cop99, FV97, Gla94a, OO92]. Metaphors [EEM01, Lov94]. Metaprogramming [Spi08a]. Metatools [EM93]. Metaview [STM88]. Method [Ano93b, Bai98, BDKP99, CTC07, Dav98a, DCH04, FHRK93, FTR^{+}94, Hor87, KPP95, LL87, MHC03, NC00, Nus96, SO92, WS00, ALMS92, FKU97, Kar88]. Methodology [Coc00, McC99c, Ros85, THV09]. Methods [Ano97g, Ano98q, BBRTL90, BNST99, BH95a, Bud99, Cha98a, Col02, DG09b, DSB99, Ger90, GCR94c, Hal90, Hal96b, HA00, Jac94, JKM^{+}09, KR05, Kem90, KKH95, LG97, MJ97, MD86, NT06, Rei02b, RME03, Rog00, RAT^{+}06, SC03, Voa00b, Win90b, RC89, Son97]. Metric [PS90]. Metric-Based [PS90]. Metrics [Ano93k, CSHA07, Coo95, DK05, GJPD99, Kil01, KJC07, MD90, MLL03, OYE06, Pfl93, Pfl08, PSC02, WGM85, ZZ93, Fen90, Fur97, HF97, Har89a, HAK^{+}96, Nie96e, Ros90]. Metrics-Based [CSHA07]. Metro [GCR94b]. Meyer [Mey96b]. Michael [Bei97b, Mye95c]. MICO [Pud04]. Microarchitectural [GP05]. Microcode [MD86, Veg86, WC86]. Microprocessor [SB90]. Microprocessors [EM91]. Microprogramming [Dav86]. MicroScope [AO88]. Microsoft [Fer01b, Ano92n, Ano93l, Ano98z, Cha98a, Lou07, Mye95a, Mye95c, Nus96, San95a]. Microtool [EM91]. Middleware [Bro99, LG03, San98f, Tho97, VVJ06]. Migrates [CFW^{+}98]. Migrating [DSBM99, Hor97, Hor98b, Rys95]. Migration [MDM^{+}94, WC86]. Military [Got01b]. Millennium [Feo97, Ree00, Sch00]. Mills [Bri96]. MIMD [LVS92]. Mind [Bol02]. Mind-Mold [Bol02]. Mini [WS00]. Mini-Assessment [WS00]. Minimizing [TK89]. Mining [DJ05, GHH^{+}09, NZZ09, Tar09, WSD^{+}09, ZS06]. Minute [vBv98]. Minutes [Wil03]. Mirror [Kla97]. Mirroring [JS95]. MIS [SMK^{+}00]. Miscalculate [BHR^{+}98]. Miscellaneous [Ano01j]. Misconceptions [SBM02]. Misleading [JG08, KJC07]. Mismanagement [Har06b]. Mismatch [GAO09, GAO95]. Miss [JD97]. Missing [Gra02, McC96f, Toc97]. Missing-Link [Gra02]. Mission [DNR04, NK04, Duk89]. Mission-Critical [DNR04, NK04, Duk89]. Mistakes [Ano92c, BM92, Car98, Gla08f, McC96b]. Misuse [Ale03, Nie97a]. Mixing [AW91]. Mobile [Jaa02, LANC07, MSG^{+}07, SGM06]. Möbius [MM08b]. Mock [TH02c]. Mode [Bur94]. Model [ABB^{+}93, Aoy93, AK03b, BDA04, BST^{+}94, BO89, Bin97, BHJM04, BCC^{+}06, CS05, Car00b, CM07, CG08, DSLB03, DT03, FCH^{+}08, FGSK03, GGLW06, GGJZ00, HL99, KM99, Kim04, Koc00, KYM^{+}04, KR03, Lin94, Lit04, Lot97, MMSM09, MMM^{+}09, MCF03, NSV^{+}08, PCCW93, Pue97, RKJ^{+}97, RS04, Sel03, SK03, Sim06, UA03, Uhl08, Var88, WUG03, YNM^{+}91, Zha04, ZS06, Bei97a, Cod88, Kru95, LCS97, LD95, RHK^{+}03]. Model-Based [NSV^{+}08, Pue97, ABB^{+}93]. Model-Driven [AK03b, CM07, DSLB03, KYM^{+}04, KR03, MMSM09, MCF03, RS04, Sel03, SK03, Uhl08, Zha04, ZS06]. Model-Integrated [MMM^{+}09]. Model-View-Controller [CG08]. Modelers [Tho04]. Modeling [AW91, AR04, Bol09, CCY^{+}89, CBMP05, CFJ^{+}91, Dan02, DFL00, Dut93, FKN02, FVHF95, FLKMR91, Gla04c, GYvdR06, Gra04, HKK93, Hor98a, Hou01, IKBM08, Kan93, KP09, KSTM98, LK03, Law98, LSS94, PP03, RKP08, RF09, San97a, San03, SMTS09, TH02a, TC86, VNS94, YJC09, Zha04, ZF99, vMC92]. Models [AvGCS07, Ano98g, BHR^{+}98, BPS94, CRR09, DSLB03, FHS^{+}06, FC03, HPSD00, JBR09, KG07, KL05, LH01, LN92, Rog00, SK92, Sei03, Spi03, Tia04, WEL02, Wes08, Wie98a, HAK^{+}96]. ModelTalk [HTLPS09]. Modern [BC86, Sha84]. Modest [Hil97, Mil01c]. Modification [SF93]. Modula [BHHM86, Gut86, MPT86, Mas98, Mul86, Rov86]. Modula-2 [BHHM86, Gut86, MPT86, Rov86]. Modula-2/MUPE-2 [MPT86]. Modula-3 [Mas98]. Modula-Prolog [Mul86]. Modular [GSS^{+}06, Lea91, Pal01a, WS00]. Modularity [Kam87]. Modularization [SRK^{+}09]. Module [Fow04]. Modules [dAdS99]. Mold [Bol02]. Money [Gla06c]. Monitor [KC90]. Monitoring [LLH^{+}05, RP09, Rob09b]. Monoliths [GG08]. Month [Bro95, Mye95a]. Monument [Pfl97b]. Moore [Big96]. Morale [SLG^{+}00]. Morality [Boo08c]. morally [Dak96f]. Most [Mey04, O’L09, Spi09d]. Motif [SR91b]. Motivation [BLSS08, HSB^{+}08]. Motorola [BDY94, DS97, Joo94]. Mountain [Cur95]. Mousetrap [Coc97]. Mouth [MA07]. Move [BB99b, Lai93]. Movie [Har03e]. Movies [BPS^{+}03]. Moving [HM01b, JMS02]. Mozilla.org [Sch99]. MP [CH93]. MPD [PHK91]. MS [Bac95b]. MS-DOS [Bac95b]. Much [Ano93-40, Ano96v, Gla06b, HB06]. Muddle [Bro99, Tho97]. Mugwumpish [Gla04e]. Multi [Ano93p]. Multi-Media [Ano93p]. Multibox [Dya95]. Multicast [FWB85]. multicomputer [RWS88]. Multicomputers [QH90]. Multicore [PJT09]. Multicustomer [SCCA04]. Multidisciplinary [BPD02]. Multidomain [RD99]. Multikey [KR88]. Multimedia [GTF^{+}87, PM98, Rei94b, Sch92, HSA89]. Multimethods [Sho06]. Multinet [GCR94a]. Multiparadigm [HDR^{+}86, SBK86, Zav89]. Multiplayer [BC85]. Multiple [Dak95b, DT94, HKR^{+}90, NJJ^{+}96, Pat08b]. Multiple-CPU [HKR^{+}90]. Multiprocessing [ZHL^{+}89]. Multiprocessor [BRS^{+}85, Emr85, Map85, MW09, OC88, SJ86, Wol88]. Multiproject [PDK05]. multitasking [CCY^{+}89]. Multithreaded [PHK91, Wil97b]. Multithreading [Ano99d, RD96]. Multiuser [KKM87]. multiview [Mey91]. MUPE-2 [MPT86]. Musa [ECW09]. Muse [GTF^{+}87]. Music [Ama07]. Musings [EMB^{+}99]. Must [Ano09d, Erd08e, Law98]. Must-Reads [Ano09d]. My [Ano09d, HRS08, Lus04, Mai07, Wey03, Wie98a]. Mycin [Nar85]. Myself [Spe99]. MySQL [Di 05]. Mysteries [Spi07e]. Myth [Ros06]. Mythbusters [JKM^{+}09]. Mythical [Bro95]. Myths [BH95a, Gra00a, Gra02, Hal90, SBM02, Voa99f].

N [Hat97b]. N-Version [Hat97b]. Nails [SMTS09]. Naked [Paw02]. Name [Kla95, McC00f]. Named [Big96, Dav96a, GAS^{+}01, Ano96i]. Naming [CW84, Whi96, Win90a]. NASA [Lie02]. National [Cha98b, GD02, BFC00, Mar85, PM98]. NATO [Ano97e]. Natural [GS90b, SLG^{+}00]. Naturalness [Gla02e]. Nature [Pfl98]. Navigability [Gon02]. Navigating [Ano97h, EMW^{+}97, Kin97, NO91]. Nebula [Erd09c]. Necessity [Dav95b]. Need [AW03, Bac95a, Bar95a, Ben00, Dak96b, Har05f, HS97, Mye01, SS06a, SMTS09, Tho94, Jac96, Law96, Sie92]. Needle [Whi96]. Needs [CP02, Fow03d, GCB^{+}06, MSG^{+}07, McC01d, Mel05, Shu07a]. Negative [Gla08c]. Neglected [Dav95b]. Negotiation [BGB01, DESG00]. Net [Can95, CC99, CCY^{+}89]. Netscape [Ano99p, Cha98b]. Network [Ano96d, Ano98g, Ano98h, BSGF98, BM04b, Cha98b, ELLM99, FWB85, JWKZ99, Pau85, PM98, vMC92, LC87, Kus95]. Networks [KWM92, LC85, MC03b, WSD^{+}09]. Networld [Cha97d]. Neural [KWM92]. Never [BDS02, Gla02d, Gla05b, Mat00b, Sch04, vBv98]. News [Ano96k, Ano96n, Ano96l, Ano96m, Ano97e, Ano97d, Ano98o, Ano98p, Ano98n, Ano99p, Ano00m, AS03, AD04, Bal00, Bar97a, Bar97b, Bar95d, Bar95e, BLP96, Bar96a, Bar96b, Bar97c, BFC00, Bur94, Can95, Cha96c, Cha96d, Cha97a, Cha97b, Cha97c, Cha97d, Cha97e, Cha97f, Cha98a, Cha98b, Cha99a, Cha99b, Cha99c, Chw01, Col05, Cos03, EG96, Gla99c, Got99a, Got00, GB00, Got01a, Got01b, Got01c, GAS^{+}01, Got01d, Got02a, GD02, Got02c, GV04, GC05, GM05, Got06b, GLS99, Irv97, Kor96, Lan06, MM06, MG06, MC03a, MC03b, MHC03, MC04, MA04, McL05, MA05, Mey96b, Moh01, Mye95c, NCC^{+}02, PM98, Row96, San95a, San95b, Sch99, Sim95a, Sim95b, SV03, Sto03, VNS04, Wal96, Got02b, HH02a]. Next [And07, Ano98p, AP00, Bal87, DDPW09, GB00, Got01a, GAS^{+}01, McL06, Sch00, Dun89]. Next-Generation [Bal87, DDPW09, GAS^{+}01]. Nial [JGM86]. Niccolo [Pre95]. Nine [Boo08d, HDR^{+}86, McC01e]. Nineties [Rys95]. Nirvana [Rei00a, Erd09c]. NIST [GD02]. No [Ano92s, Ano93w, Dro06, Dya95, FM08, Hal95, Hou95, HT04c, Mat00b, VBM^{+}02, Wie98a, Dak97b, Zan92]. Nokia [Kil01]. Noncrises [Gla00e]. Nondeterminacy [EGP92]. Nonfiction [App02]. Nonhuman [Gou98]. Nonrelational [Fon96]. Norm [Mey96c]. Nothing [Nie97b, Shu07b]. Noting [EMB^{+}99]. Noun [TH03b]. Novel [Lew98]. Novelty [Erd07c]. NSDIR [Got01e]. NSF [Cha98b, PM98]. NSFNet [Sim95b]. nuclear [CGR94b]. Nuking [JOL^{+}00]. Numbers [Gla97]. Nurturing [TH04a]. Nuts [Ano99l, Ano99r].

Object [ADT91, Ale95, Ale01, Ano93p, Ano99d, Boo94, Cox84, DM93, DeM95b, Dig98, DFL00, Dug98, DRW03, DSB99, EFM^{+}91, FHRK93, FTR^{+}94, Gla02e, Got02b, HO87, HD00, Jac93, JML00, KKH95, Koc98, KE88, Lau98, Lea91, Mar96, MJ97, MRA98, Mey87, Mil02, Pal01a, PG96, Pit93, Pou99, RG04, Ram01, Sal97a, Sal97b, Str88, TC97, WMH93, Wol89, WKH92, YKIY93, LH89, War89, Pou95, Pou99].

Object-communication [KKH95]. Object-Oriented [ADT91, Ale01, Ano93p, Ano99d, Boo94, DeM95b, Dug98, DRW03, DSB99, EFM^{+}91, FHRK93, FTR^{+}94, Got02b, HO87, HD00, JML00, KE88, Lau98, Lea91, MRA98, Mey87, Pal01a, Pit93, Pou99, RG04, Ram01, Sal97a, Sal97b, Str88, WMH93, Wol89, WKH92, YKIY93, DM93, Mar96, LH89, Pou99, Pou95]. Object-Powered [Koc98]. Objectifying [Par96a]. Objective [Oma89, SO92]. Objectively [BKS98]. Objects [BLP96, DeM95b, Hor95, JDT01, KD02, MKMG97, Paw02, Sny93, TH02c]. Oblivious [Har04b]. Obscenity [Tru94]. observation [Fid96]. Observations [BBMH05]. Ocean [KIB^{+}08]. ODBMSs [DP94b]. Odds [Ale95]. Off [AH90, CK02, DD95, Got02a, LCB^{+}09, WSL^{+}99]. Off-Shore [DD95]. Off-the-Shelf [LCB^{+}09, Got02a]. Offbeat [Sch95b]. offer [Shn91]. Offers [Ano98n, Cha96c]. Office [SC96, Aha99, BFC00]. Offices [Ano93p]. offs [MKCC03]. Offsetting [MW91]. Offshore [BGM06, CP06]. Often [Ano92s]. Oi [Mai09a]. Old [BLP96, Boo08d, Fle99, Gla00b, Gla06b, Ker08, O’D98, Pos95, Rei96b, SS06a]. Old-Timers [Rei96b]. Older [SPZ08]. Omniscient [PT09]. On-Demand [Erd07d]. On-the-fly [SF93]. Onboard [FCB00]. Once [Ara95]. One [And02, Cha99a, Dav96d, DAS^{+}07, FLN01, Hat97b, HT04b, MMW86, Rei95a, SGW02]. Onerous [Ano93d]. Ones [Gla97]. Ongoing [LM04]. Online [San98b]. Onto [For92]. Ontology [GP05, OYE06]. Ontology-Based [OYE06]. OO [BFP95, BPS94, CFW^{+}98, DK05, EEM01, EJW95, Hat98, HS97, HT04b, JDT01, Kai99, KSTM98, KKC93, Nes98, NS99, Pou95, Wat97]. OOA [HKK93]. OOPSLA [BLP96]. OOT [BG98]. OPC [LLH^{+}05]. Open [Abe07, Ano03h, Ano03i, AS03, BNST99, Cha98b, DV89, DGC^{+}07, Ebe07, Ebe08a, Ebe09, Gla99c, Gla00d, GAS^{+}01, Got07, Hec99, Kic96, KHA^{+}03, MH07, McC99c, OG07, Ray99, REM^{+}09, RP09, San98a, San08, Sch92, SS06a, Spi06d, SV03, Tot06, VVM08, WW01, Wil99b, WLC01, vdLLM09, vWHvW09, HS00, DFLS05, FK04, GA04, Has05, KT04, Lus04, MD04, McL05, NK04, Pud04, RE04, Sca04, SCSC04, SS04, TH04b]. Open-Source [BNST99, Hec99, McC99c, Ray99, Wil99b]. Opening [Ano98p, GB00, HVF^{+}00]. Openings [Dav97c]. Operating [Ano98u, Bal87, BRS^{+}85, BHHM86, Emr85, Fer95, Gra99b, LRO02, Ols85, SH02, Wil97c, BDHJ96]. Operational [Ber96b, Mus93, Voa00c]. Operations [CC93, Gai94, KG07, TL98, Deu88]. Operations-Research [KG07]. Opportunistic [BGL^{+}09, GG08, JBHD08, NOK08, OGE08, Rob05b]. Opportunities [BB08, Dak95b, San98b]. Opportunity [Ano98p, Orr04]. Opposition [Bud99, Cur01, Ded09, Gla98b, Gla98d, Gla98e, Gla98c, Gla99b, Gla99c, Gla00c, Gla00a, Gla00b, Gla00d, Gla00e, Gla01b, Gla02d, Gla08b, Gla09c, Gla09d, Jac98b, Jac99, Rif01a, Rob01b, Spi98, ZJ00]. Optimism [BHR^{+}98]. Optimization [BFJT05, Fow02c, PWS^{+}95, RKP08, SM06a, Veg86]. Optimizer [DBL04]. Optimizing [AGK^{+}87, Gai94, Jun98]. options [Dak95a]. Oracle8 [Pou99]. Oranges [Per95]. Orchestrating [Lou08]. Order [HKPS96]. Ordinary [Ano93q]. Organization [ASC07, CDD08, CJNM05, CN95b, Cur98, DSBM99, GE04, MH07, NS99, SvHB02, Mat00b]. Organizational [Ano99g, Boo08a, Faf94, LWHS01, Mil01a, EM96]. Organizations [Ano01g, BT05, CZ01, CB00, DLSZ05, Dyb00, GH00, Har04e, HM01b, Kau99, PSV94, RvW07]. Organizing [HAJ07]. Orientation [Gla02e, War89].

Orientations [Hor95]. Oriented

[ADT91, Ale01, Ano93p, Ano99d, AK03a, ASC07, Bas90, Bla07, Boo94, BCC^{+}06, DeM95b, DT03, Dug98, DRW03, DSB99, EFM^{+}91, FHRK93, FTR^{+}94, Frü99, Gli08, GKMM04, Got02b, HO87, HD00, JML00, KLD02, KE88, LM07, Lau98, Lea91, MKHV06, MRA98, Mey87, Mub08, MS06, Pal01a, PG96, Pit93, Pou99, RG04, Ram01, Ray06, Sal97a, Sal97b, Sta06a, SBK86, Str88, VV00, WP86, WMH93, Wol89, WKH92, YKIY93, ZAV02, DM93, Har89a, LH89, Mar96, RHK^{+}03, Pou99, Pou95]. Orienting [Boy97]. Origin [Dak95e, Pat08d]. Origins [Ale99]. Orwell [Ano93-28]. Oscilloscope [DG90]. Other [Gla00e, HT04b, Nie95a, NE07, Rob05a, SC05, Spi07e, Str01, US09]. Ounce [McC01a]. Our [Ano93-31, Ano03m, Ano06q, Ano07k, Bac97, Car98, Dav97d, Har03c, Jon96, McC00g, McC02g, Str09, WEF01, KC97]. Ourselves [Mel95]. Outcome [WMH07]. Outcome-Based [WMH07]. outcomes [Sim96]. Outdated [Wil97c]. Outer [Kor96]. outperform [Sca89]. Output [Mat00b]. Outside-the-Box [Gla03a]. Outsourcing [BGM06, HKNS01, KRF01, MHC03, Rei04b, Tiw04]. Outwitting [Ano05l]. Over-the-Transom [Ano92o]. Overcoming [BGM06, MDD09]. Overhaul [Nie95a]. Overlaps [Tiw04]. Overlooked [TM04]. Overoptimistic [Jør09]. Override [Bar97a]. Overshadow [SM06b]. Overview [AHL^{+}00, Bol99, Lv85, SDP91, Wil96, WC86, WF89]. OVID [RBIM97]. Own [Bac97, Gra97, Har06c]. Ownership [Nor03]. Owns [Gem85]. Oxymoron [Orr04].

Pacific [GAS^{+}01]. Package

[Ano93o, FC03]. Packard [Ano96m, Pfl95b]. Packets [PM98]. Packing [HT03b]. Pad [RGR85]. Page [Nie95a, Nie95b, URK01]. Pain [GRS02]. Painless [LVS92]. Pair [DCF^{+}02, DAS^{+}07, GSB03, WKCJ00].

Paired [Mir01]. Pairs [PSHH04]. Palm [SY99]. Panacea [MLL03]. Pandora [MLL03]. Panelists [Mcs96]. Paper [Nie95b, Spi07d]. Paperless [SC96]. Papers [Ano00k, Ano00i, Ano00j, Ano00h, Ano01h, Ano03k]. Parable [Kla97]. Paradigm [Ano93-37, BMM04, BMM05, Coc97, DCF^{+}02, Jur97, KL04, OC90a, San95b, Ven05, SR91a]. Paradigms [PWS^{+}95]. Parafunctional [Hud88]. Parallel [AK85, Ano96s, BO89, Bec90, CK91, Cla98a, EGP92, GG89, GGJ^{+}89, HE91, KT88, KL88, KC90, LC87, Ols85, Pra85, QH90, Ree91, SR85, WJS97, ASM89, Bal89, BAS89, KB88, Tic89]. Parallel-Execution [KC90]. Parallel-Programming [BO89, ASM89]. Parallelism [CV88, Dou85, OC88, WKH92]. Parallelizing [CC94, GAY93, PJT09]. Parameterized [TS06]. Paring [BALV03]. Paris [GCR94b]. ParseMe. [Lew98]. Parsers [Dya95]. Parsing [RW90]. Part [Boo09e, Boo09f, Bus09b, CDD08, Gla07b, Gom06, McC06, Spi09d, Ano04o, Gla09c, SM09b, Spi05e]. Participation [Ano99f, DRR^{+}07]. Partly [Rei95f]. Partners [San95a]. Partnerships [San08]. Party [Cou99a, Cou99b]. Pascal [DSBM99, GAY93]. Passes [Ano96h]. Passing [Fra98, Rei90]. Passion [Har06e]. Passport [Bar95a]. Passwords [Har06e]. Past [BHS07, GA06, Gla99a, KOS06, Pes97, Ree05, VBM^{+}02]. Pat [ASM89]. Patch [BS03]. Patent [Bar97b, Bar95e, Gra00a, Gra00c, Aha99]. Patents [Gra99a, GK02, McL05, Sku99, WGK^{+}02]. Path [AH96, PR08, WOBP00]. Pathfinder [Sto99]. Pathological [Bak95]. Patrick [DMN^{+}98]. Pattern [Ale99, CG08, FGSK03, WC02]. Pattern-Based [FGSK03]. Patterns [BGS01, BALV03, Boo08a, BHS07, Cop97, DeM95b, Fac01, Fow03b, HAJ07, HL05, HAZ07, HM01b, Hou01, KM99, KV07, Kot98, MKMH07, MJ97, MKMG97, Ris07, Sta06a, TYZP05, WB06e, WB06f, KC97]. Pauses [Rei94a]. PAWS [ACM^{+}07]. Payoff [PRSV97]. Pays [CK02]. PC [FKLW91]. PCs [Pem87]. PCTE [Tho89]. PDAs [Got01b]. Pearl [SH93]. Pearls [Ben99]. Pedagogical [Fow09]. Peer [Chw01]. Peer-to-Peer [Chw01]. Peering [BNST99]. Penguins [Ana99]. People [Bag99, GDD07, Lar01b, Mur95, PSV94, Sim95c, VBM^{+}02, Wil97c, Ano99e]. Peopleware [DL98, You07]. Perceptions [MD97, Pre96b]. Perceptual [GRS02]. Perfect [CBK85]. Performability [CFJ^{+}91]. Performance [Ano93f, BCC^{+}08, DSD02, FLKMR91, GMS^{+}96, HE91, JHMW07, KKJ91, KP99, LM03, MR90, Map85, NDJ^{+}90, NO91, Pou99, PWW05, WSL^{+}99, Jal89, Wil96]. Perilous [Cla98a]. Perl [Dom98]. Perry [HEH06]. Persistence [OLPS04]. Persistent [BVB04]. Personal [Ano92p, Hum95, Hum00, JD98, KH00, Mye95b, Mye95d, Rai07d, Tom03, ZME00, Hum96]. Perspective [And02, Ano96r, Bar98, BCC^{+}08, BM01, BC08, Fle99, HD00, Kal94, Pau01, MRA98]. Perspectives [Ano92q, Ano93e, Ano96d, BF00, Coy00a, Mil01b, NJJ^{+}96, Pfl97b, US09, Lar01b]. Pervasive [JB04]. Phantom [Aus99]. Phase [Hoh05, Ano89]. Phenotypes [Dav03]. Philosopher [Gra90]. Philosophy [Got09a]. Physics [MJP^{+}00]. Physiological [JW09]. Pick [RHK^{+}03]. Picks [Ano09a, GAO09]. Picture [Gla98e, JLLP03]. Pictures [Boc03]. PID [DG07]. PIE [SR85]. Piecemeal [Cop99]. Pieces [OC90b]. Pilgrim [Lar01b]. Pilot [SY99]. Piloting [DNR04]. Piña [Hay03]. pinpointing [LST91]. Pipe [FV97]. Pipelined [SB90]. Pipelines [TH03a]. Pipes [Voa97b]. piracy [Dak97b]. Pisces [Pra85]. Pit [Ado96]. Pitfalls [Sal97b, Tur03]. Pixels [GF99]. Pizza [Bak95]. Place [Boo09d, DDS^{+}97]. Plan [CSSP05, Kor96]. Plan-Based [CSSP05]. Planning [Ano93x, AW03, BDKP99, BALV03, Cox90, DDPW09, JB00, McC01e, MS04, NT02, PMM93, Rai07d, RS05, Sne95]. Plans [LS86]. Platform [Jac93, LRO02, PCOM97, YJ07]. Play [JLLP03, San98c]. Please [Voa00c]. Pleased [SGW02]. Pleasure [GRS02]. plenty [Mye89b]. Plight [Pre98]. Plural [BK05]. Poetry [Ano08c]. Point [Ano93u, AW03, Bac97, Bec07, CC95b, Cla98b, CK02, CHJ^{+}06, Coy00b, CB00, EM09, FV07, Fra99, FPK07, GC08, Gil96b, Gra97, HvG09, HM02b, LB00, Law96, Lis97, LS07, MC94, McC97f, Mun98, PC09, RM04, RC95, RH05, Sch96b, SY99, SS06b, SVSC09, SN07, Spr95, SM01, TB02, UA03, UE07, WCCR08, vGB08, ALMS92]. Point-Counterpoint [MC94, CC95b]. Point/Counterpoint [AW03, Bac97, Bec07, CK02, CHJ^{+}06, CB00, EM09, FV07, FPK07, GC08, HvG09, HM02b, LB00, LS07, McC97f, PC09, RM04, RH05, SS06b, SVSC09, SN07, SM01, TB02, UA03, UE07, WCCR08, vGB08]. points [Fur97]. Poised [GAS^{+}01]. policies [GA05]. Policy [ABB^{+}00, Ano98g, Kus96]. Political [Ros04]. Politics [And98a, SM06b]. Pollinating [Ano98l]. Pollination [ME00]. Poor [Voa99b]. Pops [Cha98b]. Popular [MLI04]. Portability [Spi06e]. Portable [BBE^{+}88, BHHM86, GAY93, JS95, WC06]. Portend [Mye95c]. Portfolios [CSHA07]. Porting [Sch92]. Pose [KL97a]. Position [Mey96c]. Positive [Ano96q, EG96, Hir05]. Possibilities [DSD02]. Postmodern [Kru05b]. Postmortem [BDS02, CDF96]. Postwaterfall [CCS^{+}03]. potatoes [Ano97e]. Potential [BHS85, Gla03a, Jal89]. Potholes [Gra99c]. Power [Boo07e, McL06, Ris07, Ste95b]. Powered [Koc98]. Powerful [WB06d]. Powering [DW00]. Powerless [Mey96c]. powers [Got02a]. PQ [CCY^{+}89]. Practical [Cha98a, CP06, DSBM99, Dav96c, DS07, DBS09, DFL00, DRW03, Jor05, KSMS02, LSB02, Mye98, OY95, OBSN08, Pet85, Rog00, SKC93, SM92, Wel00, WE04, Wol89, Yam98a, Coo95]. Practice [AvGCS07, Ano03k, Ano05n, BHS85, Big96, BHJ^{+}03, Bol99, BTC^{+}99, CC99, CLB03, CMKC03, CMKC09, Dam07, Dan04, DK05, DSB99, Gar05, Gla98a, Gla99d, Gla03c, Gla05c, GGLW06, GLT03, Har03d, How85, Hsi93, KP99, KL05, KCD09, LCM06, LSF03, MRZ^{+}05, Mat00b, Moo03, NL03, PR08, Rei03a, Rob01b, Ros05, SKT^{+}92, TH04c, TB02, WPJH98, Whi00, WWD97, WB09e, Sch95b]. Practices [Ano98g, BC86, CR08, Dec01, DAB06, DLV99, Got08a, Har04a, JB00, Jon96, Jon03, JMS06, KP09, Kle98, Koc00, LF05, MKCC03, McC96b, McC96c, McC96d, McC96e, McC97a, McC97c, McC97d, McC97e, McC98c, McC98a, McC98b, McC98d, RBBP03, Rei04b, Rog00, Ros06, Run06, Sca04, SJM02, Str99, Tur03, US09, VE05, WB06f, Boe91, McC97b, Kus95]. Practitioner [BH02, CN07, Dav96f, Gla00a]. Practitioners [Ano96r, Ano98x, Ano98y, Ano98z, DKJ05, LCE^{+}01, PM00, San98d, San98e, San98f]. Pragmatic [Ber03, Bus09a, JBHD08, Pat09, SC00]. Pragmatics [Sel03]. Praise [BHR^{+}98, Gla98a, Kla97]. Predictable [Ano93f, MKHV06]. Predicting [Gla02a, HS90, VMM91, VCM^{+}97]. Prediction [KWM92, KAKG96, KL05]. Predictions [Dav98c, HDS05, Lan96, Voa00b]. Preparing [HT03a]. Prescription [Ada95]. Prescriptions [BBMH05, TH02d]. Present [BHS07, KOS06]. Preservation [Got02d]. Preserve [SE98]. Preserving [KYM^{+}04]. Pressure [Gla04a]. pretty [Cam88].

Preventing [Pos85, Wil97c]. Prevention [McC01a]. Price [Dak95d]. Pricing [Dak96c]. Primavera [SA05]. Prime [McC99c, UA03]. Primer [CTW98, RE04, Wil97b]. Principle [KL04, LSM84, SH02]. Principle-Driven [SH02]. Principles [Ano93f, Ano98g, Boe91, Etl96a, FNO92, Kus95, McC99d, RG02, Sch95b, WB09e]. printer [Cam88]. Priorities [BCH08, Cha99b, Pet85]. Prioritization [ASC07]. Prioritized [TRW03]. Prioritizing [KR97]. Priority [Ano93-30]. Privacy [Ano92p, Ano93-28, Got01d, Kus96, PC97]. Private [Ano96d]. Prize [DKKK90, DKMS91]. Pro [WGK^{+}02]. Proactive [CK02]. Problem [AW03, Gla03a, Gla04b, LV01, McC98b, PM97, Ree06a, Rob05b, RD99, San03, Sim95c, Str99, Win88]. Problems [BLP96, FW03, Fle99, FHMU85, Jac94, JMS02, KL97a, LG97, Pes97, RS01, Vic07]. Proceedings [Cur95, Mye94, Sim95c]. Process [Aae03, Ano92a, Ano92e, Ano93-38, Ano98j, Ano98w, Ano00m, Aoy93, Bac95a, Bai98, BG94, BZM^{+}95, Bau94, BO89, BMGT01, Boe96, BALV03, Boo07a, BM04a, Car94, Cla00, Col94, CDF96, CF02, Coo95, Cur00, CSR^{+}08, DLSZ05, Dec01, DS97, Dio93, DT03, DC04, Dut93, EMP05, EA03, EL04, Erd05, Erd08a, Erd09f, FH07, FO99, FCB00, FLN01, Gla00b, Got08a, Gre01, HMK^{+}94, HK94, HS97, Hum88, HSW91, Hum95, Hum96, Hum00, JP03, JBR99, Jak98, JB96, JD98, KM02, KH00, Kem90, Kim04, Lav00, Lev91, Lev90, LR00, Lin94, Lit04, LG03, LLH^{+}05, MNA05, Mat00a, MD02, Mcs96, MS92, MSK00, Mye95b, Mye95d, New95, PSV94, PR94, PWS^{+}95, Pre96a, RKK05, Ram02, Rif01a, RJ00, RM00, SM09a, SC04, SC03, Sut00, TV99, TM94, TS01, TC86]. Process [Tom03, UHC02, Vis94, Wei98a, Wel00, Wie99, WS00, Yam99, dSA95, vS04, vS09, ABL89, Bei97a, BCM88, Hal96a, Pre96b, RHK^{+}03, JBR99, Rog00, Ebe01, MRA98, SC00]. Process-Control [Lev90, dSA95]. Process-Improvement [Ano92e]. Process-Intensive [Gre01]. Processes [Ano05o, ACM^{+}07, BT05, BOS00, Cur98, DAEE08, EEM01, Jak00, Kan09, Kee00, Lai93, MTE00, RP09, Wil97c, YBBP05, ZME00, vWAS06]. Processing [Ano96s, DY93, GG89, Kod04, KL88, Lew84, Ols85, SR85, ZSF91]. Produce [EG96, Voa97b]. Product [BHAW09, BHJ^{+}03, BCM^{+}04, CSHA07, CJNM05, Ebe01, Ebe06, FH07, Fen96, Jaa02, JKC04, KLD02, KM99, KMSW00, MSG96, MNJP02, NT02, Nor02, RE04, SV02, Sch96b, Sim06, TH02a, WHG^{+}09, vdL02, Gra90, KCC^{+}01]. Product-Line [Ebe01, MSG96]. Production [Far84]. Productivity [AH96, BMJH96, Cha99a, Gla08c, Gro04, HG00, HS00, Lim94, MKCC03, MF00, Max01, PSHH04, PS84]. Products [Ano93c, Ano98x, Ano98y, Ano98z, AW06, Gem85, HdLGG03, Pat08a, Rei96b, RAv96, San98d, San98e, San98f, Pou95]. Profession [Gee06, JL03, Koc00, Mus85]. Professional [BTC^{+}99, McC99e, MT99, NS05, Row96, Spi06d]. Professionalism [Mar07, Wat97]. Professionals [CP02, Cus03]. Profile [Ano92r, Voa00c]. Profiles [Mus93]. Profound [Cha96a]. Program [And03, Ano92e, Bar95d, BZM^{+}95, Con96, DJSS97, DHMR02, Eng99, FTB06, GA06, GdHN^{+}08, HN90, HPLH90, Kil01, Leh86, LS86, MRA98, MK92, NS05, OCM^{+}84, Pfl93, RW90, SF93, Sev87, SSSW86, Spi08d, TL98, Fen90]. Programmer [Fri97, McC97d, MP88, SC00, Wil03]. Programmers [Ano93f, Gor95, Jef97, McC98b, WES87]. Programming [AK85, AO88, Ano93e, Ano96j, Ano96s, Ano98h, BO89, Bec90, BGL^{+}09, Bri96, Bud91, CK91, Che02, CDM^{+}05, Coo96, Cox84, CAGR91, DSBM99, Dav85, Dom98, DCF^{+}02, DAS^{+}07, EM91, EG04b, FVHF95, Frü99, GSB03, Got02b, GGJ^{+}89, HO87, Har04d, Hoa84, Hud88, HR95, IH90, JM07, JGM86, Jon07, Jur97, KT88, KL04, KG07, KE88, KSS84, Lou07, MMW86, Mar00, McL06, MW09, Mil86, MSK00, MS06, Pai86, QH90, RG04, Ram94, RWS88, Ree91, Rei87, RGR85, SR85, Sha84, SDP91, Spi06b, SBK86, Str88, Tho05, VV00, WMWM87, WJS97, WKCJ00, Wil97c, Wil97a, Wil97b, WCW85, ASM89, ALMS92, BAS89, BK89, Tic89, Zav89, Ben99, Gla01a, Gre01, KP99, Pau01, PH01, Sch01, Str01, Wei99, Ano01f]. Programming-In-The-Large [WCW85]. Programs [Ale01, Bar95a, CT91, CC94, DJM04, EGP92, Eng99, HE91, HL85, Lea91, LCS91, LSM84, Moo01b, Par99, PHK91, PTH^{+}09, RD96, ROL90, SBM02, HF97, LH89, OJ97, Pal01a]. Progress [Dav86, Pfl08, San95a, SRS^{+}86, YM94b, Lar01b]. Project [AHM86, AR04, AHZ02, Ano92i, Ano99d, Bas87a, Bau94, BB99a, BDS02, BALV03, CSSP05, CC99, Cha96b, CJNM05, Coc00, CDF96, Coo95, Don06, Dow87, DGC^{+}07, DB99, EK08, Gar05, Gla02d, GF99, Han06, HS97, JB00, JKP^{+}05, KR05, LM03, LJ97, Lew98, Lin93, Lis97, Lon02, McC97e, McC01e, Mel05, Mey96a, Moy97, Mug08, ND05, Phi96, PSC02, Pre98, PT05, RKP08, Rob03, SM08a, Sch98b, Spi06e, SM01, THDK06, VE05, Wil97c, UHC02]. Project-Processes [Wil97c]. Project-Starting [Sch98b]. Projects [Ano93h, Ano93y, AW03, Byt99, Cla98a, CS97, DM96, DNR04, Fai94, FW03, GRM08, GDD07, GJPD99, Gum06, HDR^{+}86, HL01, KSMS02, KT04, Lar01b, Mac96, Nes98, Ram00, Ram01, RBH07, Ree99, Ros04, RM00, SM08a, MW96, Ras03, vdL02]. Prolog [Dav85, Kar88, CV88, Den91, GS90b, Mul86].

Prolog-Based [Den91]. Promise [Cha96a, Cuk05a]. Promises [Cha97b, LG97]. Promising [Dak96d, PC97]. Propaganda [Har04c]. Proper [CC95b]. Properties [KYM^{+}04, Xu03]. Proposal [Cha97e, Cha98a, Hil97, Mil01c]. Proposed [Bar97b, GD02]. Proprietary [SCSC04]. Prose [Ano96d]. Prospect [Mil86]. Prospects [Sha90, Sha09]. Protect [MC03b]. Protected [Lar01a]. Protecting [AK99, HM02b, Voa99e]. Protection [Ano84b, Ano93-32, Bol04]. Protectionist [GA05]. Protocol [Hol92, LL92a, LL92b]. Protocols [Ano93f, NS92, DW00]. Prototype [And94, BGL^{+}09, Sch04]. Prototypes [Nie95b]. Prototyping [AW91, DM87, GB95, Hol04, KSS84, LM84, LK88, Luq92a, SAG^{+}97, LB88]. Protoyping [Luq92b]. Proud [Dav95c]. Prove [Mye98]. Providers [GB00, Sim95b]. Providing [Lav00]. Proving [Fra95]. Provoking [MGR04]. pseudocode [Sca89]. PSP [BCH^{+}02, MMM^{+}01, Mor00, SC03]. Psychology [Rup02, Wei99]. Public [Ano96d, Cuk05a, Fow02a, Fin96]. Published [Fow02a]. Pull [Sto03]. Pulled [Wil99a]. Pulling [HJA08]. Purpose [Dan02, MHB08]. Pursuit [Cur00, Fav96, SLG^{+}00]. Pushes [Bar97a]. Pushing [Mye95d]. Put [Hsi93, OG07]. Putting [DDS^{+}97, Nie96b, RE03, Ste02, WWD97]. Puzzle [OC90b].

Q&A [Che97]. QA [Yac03]. Qlisp [GG89]. quadtrees [BK89]. Qualitative [Dou85, KKL97, TC86]. Quality [Abe07, Ale01, Ano93e, Ano93o, Ano97g, Ano00m, Bar98, BBD^{+}04, Bau94, BHR^{+}98, Bin97, BCH08, BDKP99, Bøe08, BI96, Car94, Car00a, CM90, Col94, CKMV95, Cou99a, Cou99b, Cri06, DJH^{+}08, DS07, EG04a, EM96, EMB^{+}99, Fav96, Fen96, FC03, Frü99, Gla98b, Gli08, Hal95, Hat95, Hay03, HDS05, HEH06, Hir05, Hou95, HB06, Jac95, JS08, JD98, JKC04, JM02, Kee00, KAKG96, KL96, KL98, Kra95, LA03, Lan96, LY98, Lev91, Lim94, LSS94, MKCC03, McC02f, MVS^{+}99, MH06, MRA98, Mil01c, MSK00, Off02, OY95, OBSN08, PL03, Pfl97a, Pfl97c, Pfl98, PV97, PS84, Put91, RSO08, Rob95, Rob09b, RAv96, Ros98a, Sch96b, SLG^{+}00, SSZ06, Sim96, SM06b, Str96, Tar09, Ter96, Tia04, URK01, VV00, Voa97b, Voa99b, Voa99d, Voa99e, Voa99f, Voa99c, Voa00b]. Quality [Voa00a, Voa00c, Voa01a, Voa01b, Voa03, Zve98b, Har89a, Ros90, Sie92, Ada95]. Quality-Driven [DS07]. Quality-Evaluation [Tia04]. Quantify [Bur94, Mai06a, MD90]. Quantifying [Cla00, McC00b, MA89]. Quantitative [CSR^{+}08, FCH^{+}08, MVS^{+}99, MW01, RKP08, Sch07, KKL97]. Quantitatively [Mos92]. Quantity [Ano97g]. Queries [Shn94]. Query [CB98, DY93]. Question [DDS^{+}97]. Questioning [Gla03d]. Questionnaires [AK99]. Questions [GSB03, JOL^{+}00, Kus96, Zve98b]. Quick [Ano98h, Rei00c, San98f]. Quick-to-Market [Rei00c]. Quickly [Gra99a]. Quiet [Ano00m]. Quit [MJP^{+}00]. Quo [BLP96]. Quotes [Cur95].

R [Bei97b, SSZ06, San95a]. R&D [Bel99]. Race [Ano93y]. RAD [Car95, MS95, RC95]. Radical [Cha97b, DCF^{+}02]. Radio [MLL03]. Rails [BK07, Vis08]. Raising [JOL^{+}00, McC01f, Kal94]. Ramping [Got02d]. ranking [LCS97]. Rapid [Ano93z, GB95, Hol04, RIP^{+}01, SAG^{+}97, TYZP05, Vis08, For97]. Rapidly [LB88]. Rates [Ano93g, GT05]. Ratio [Erd08c]. Rational [BMGT01, Spi08a]. Rationale [Fuj99]. Rattle [MRE05]. Raw [HT03a]. Raytheon [Hal96a]. RDBMS [BPS94]. Re [WSL^{+}99, AW06, DP03]. RE’03 [WE04]. Reactive [Sav08]. Read [Kru09b, McC98f, Wie98a]. Readers [FF08]. Readies [San95b]. Readiness [McC97c]. Reading [Dan03, Jac95]. Reads [Ano09d]. Ready [Got01a, LL03, McC99c, UA03, Bre89]. Real [And98b, Ano93f, Ano93-30, BS86, Coo95, DMN^{+}98, DN97, EH95, FL86, Gom06, Gor95, Kan01, KB02, KKJ91, Lau98, LLH^{+}05, LL03, Mar04, McC02f, Nie97c, Pal01b, Par96a, PH93, RP06, RS84, Sam00, Spi90, SH93, SMK^{+}00, Voa00c, WSL^{+}99, WGFC86, ZLJW95, CCY^{+}89, Deu88, FKU97, For88, LB88, SR91a]. Real-Life [Lau98]. Real-Time [Ano93f, Ano93-30, BS86, FL86, Gom06, KKJ91, LLH^{+}05, Mar04, PH93, RP06, RS84, Spi90, SH93, WGFC86, Par96a, ZLJW95, CCY^{+}89, Deu88, FKU97, For88, LB88, SR91a]. Real-World [Coo95, Kan01, Pal01b]. Realism [Ano96r]. Realities [DD95, JB96, TV00]. Reality [Rei94a]. Realizing [Jal89, NZ07]. Really [Ano93t, Car95, DeM95a, Gla98c, GSB03, JS08, LB00, Law96]. Reasoning [THV09]. Reasons [Ros04]. Reassurance [Kus95]. Recasting [WOS94]. Recently [Tom87]. Rechoirments [Dav98d]. Recipes [Aae03]. Recipients [Bar95b]. Recognition [GAS^{+}01]. Recognizing [Bal99, RW90, ROL90]. Recommendations [LLS^{+}06]. Recompiling [MBL08]. Reconfiguration [YD01]. Reconnaissance [BFC00]. Record [JOL^{+}00, MVS^{+}99]. Recovering [Dav98a]. Recovery [CC90, Sch01, WB06a]. Recursive [SM97, Sho06]. Red [GB00]. Red-Hot [GB00]. Redefining [Got08b, Hol92]. Redemption [Sch01]. Redesign [Fis87, SE98]. Redesigning [JC00]. Redocumentation [Raj00, WTMS95]. Redoing [Vit84]. Reduce [HF95]. Reducing [BS03, Fow01b, Voa99b]. Reengineering [Ado96, BH95b, DD95, EK02, MGG^{+}95, Moo02a, New95, PRSV97, Sne95]. Reentering [DB97]. Reevaluating [Cop99]. Reexamining [Hat97a]. Refactoring [FGSK03, MHB08]. Referees [Ano03m, Big96, Dav97d, McC00g, McC02g]. Reference [GGJZ00, LM03, San98f]. References [Har06a, MMM^{+}01, KHA^{+}03]. Refinement [Hen94]. Reflecting [Rob01b]. Reflection [OLPS04, vS09, GA05]. Reflections [KLR89, vV06]. Refocus [SMK^{+}00]. Reform [Gra00c]. Refreshing [Cha99a, WB06e]. Regarding [Gla04d]. Region [GAS^{+}01]. Regulated [Voa99c]. Reinforcement [Hir05]. Rejuvenation [YD01]. Relation [Var88, Cod88]. Relational [Amb07, AEP96, Bud91, CC93, Fon96, Kor86, MMW86, MDM^{+}94, KR88, SM88]. relationship [SM88]. Relationships [HKNS01]. Release [RS05, Bre89]. Releases [San95a]. Relevance [Ano07j, Hon09]. Relevant [Gla09d]. Reliability [AvGCS07, ARW96, Bei97b, BG95, BL92, CR88, Cuk05b, ELM90, Eve95, EH95, Ham92, KWM92, Lan96, LN92, Moo01a, ME90, Mus93, PMM93, SK92, SKT^{+}92, SM92, TL96, TP98, Var95, KKL97]. Reliable [Ano92i, JDT01, JM02, Tau95, VV00]. Reloaded [FM08, TS06]. Remains [Gla03b, Gla06b]. Remediation [BS00]. Remedies [Jac98a]. Remote [Gon02, San98f]. Renaissance [Cur88, Roy00]. Rendezvous [GC85]. Rendezvous-Based [GC85]. Renewal [Sne84]. Repair [Cha99c, WB08f]. Repealed [McC99b]. Repertory [NE07]. Repetition [Fow01a]. Replay [CT91]. Replicated [EK08]. reply [Cod88]. Report [BZM^{+}95, Bol04, BHKR01, Coc01, Dav95a, DKKK90, DKMS91, GD02, Ju01, KH00, Luq92b, MM00, MS95, Moi01, PDK93, PD93, WW01, WS00, PJCK97]. Reporting [ON92]. Reports [Ano02i, Str01]. Repositories [Cuk05a, Yin94, ZS06]. Repository [LWHS01]. repossession [Zan92]. Representing [GS90b, Lub91]. Requirement [BI96, Wan95]. Requirements [Ale06, And94, And98a, Ano08d, AW03, App02, AD03, ASC07, BHAW09, BL98, BGM06, BCH08, Bøe08, Boe84, BGB01, Bol09, Bus09b, CNS08, CR08, CW00, CC96, Cod09, DESG00, Dam07, Dan04, DP94a, DL03, DRR^{+}07, DJS08, DAEE08, DP03, Ebe06, EG04a, Fav02, FCH^{+}08, GW07, Gli08, Gon05, GM09, Got03, Gra02, Gra03, GGJZ00, HL05, HdLGG03, HL01, HA00, Jac96, JLLP03, JL06, JB04, Jun98, JMS02, KR97, KIB^{+}08, Koh05, Lab94, Lau03, LWE01, LX99, LM04, MM08a, MN98, MG01, MGR04, MRE05, Mai05, Mai06a, Mai06b, Mai07, Mai08a, Mai08b, Mai09b, MKH98, MM08b, Mug08, MCL^{+}01, NGBR05, NL03, NJJ^{+}96, Nor07, Orr04, PDK05, Pae08, PF92, PDK93, PG96, Pos85, PTA94, RSO08, Rei00a, Rob01a, Rob02, Rob04, RP09, Rog00, Ros06, Rup02, Sal03]. Requirements [SSV99, SM09a, Sel09a, SGM06, Sid94, SS96, Sim04, SSC06, Som05, SJM02, SPZ08, TSP^{+}09, TH04a, Tho94, TJM08, TH02d, WW03, WE04, ZMT05, dSA95, Law96, WKUI89, Rog00]. Requires [Ano93x]. Requiring [Ano08d]. Rescuing [Rai07e]. Research [AKM^{+}08, Ano92s, Ano93e, Ano93l, Bar98, CT90, Cha99b, Erd08e, Gla94b, Gla09d, Got02e, HDR^{+}86, Har03d, HM92, KG07, KHA^{+}03, Mai05, Mar85, Mat95b, Mat96, MF00, MM00, Pfl97c, Pot93, Rob01b, San08, SRS^{+}86, WK03, Ano89, Zho89]. Researchers [MC03b]. Resolutions [BBD^{+}04]. Resource [BV08, DDPW09, GB00, SJ86]. Resources [BCKS01, RKK05]. Respect [Aha99, AS03]. Response [BNST99]. Responses [Ano09e]. Responsible [WB09f]. Rest [BC08, TJM08]. Resting [Boo09d]. Restoring [RW98]. Restructuring [CS90]. Result [Eme95]. Results [Ano97g, EG96, MF00, MVS^{+}99, Pfl97c]. Rethinking [SHZ^{+}98]. Retrain [DHMR02, Rot99]. Retrieval [Has05]. Retrofitting [MTKE94]. Retrospect [Mil86]. Retrospectives [Gla02d]. Return [Ano93-38, Ano03j, EFS04, Ree06a]. Reusability [BR87, GP87, Mey87, PDF87, PD93]. Reusable [BAB^{+}87, DeM95b, ES92, Hen94, HJA08, KG87, RW04, Sta94, WP86]. Reuse [Ale95, Ano98g, Bar95d, BB91, Bas90, DSBM99, Fac01, Faf94, Fis87, GAO09, Gla98e, Gri95, HF95, JBHD08, Joo94, LSW87, Lim94, MD04, MAMY99, Moo01c, MTE00, Mye94, OMTHC94, PB98, PWS^{+}95, Pfl96, RBH07, SHZ^{+}98, Spi07b, WOS94, WHG^{+}09, WES87, FI94, GAO95, Pfl95b, Vis94, Wat97]. Reuse-Oriented [Bas90]. Reusing [BFP95]. Revalidation [HR90]. Revenge [Tho04]. Reverse [ANB99, Cha97d, CC90, Moo02a, RS04, Sam90]. Reverse-Engineering [ANB99, Sam90]. Review [BALV03, CDF96, Cos94, JP03, Lew84, Rem05, SHZ^{+}98, SLG^{+}00, Sta97, Tru94, Zan92, Zan94, Ada95, Can95, Coo95, DeM95b, Fer95, Kal94, Kus95, Mye95b, Nav95, Pou95, Rys95, Sch95b]. Reviewer [Ano07k, McC01g, McC02g, Str09]. Reviewers [Ano98n, Ano99h, Ano99e, Ano00h, Ano04a, Ano06q, Ano07k, Str09]. Reviews [CLB03, KB02, MRZ^{+}05]. Revising [Ano93-28]. Revisited [CS09, Pot93, Sta97, Wan95]. Revolution [Ano00m, Cox90, GAS^{+}01]. Revolutionary [Mat00b]. Rewards [JB96]. RF [MLL03]. RFID [FF08]. Rhetoric [Leh86]. Rich [Sav08, ZMT05]. Rich-Media [ZMT05]. Right [Ber93, CBMP05, GSB03, Got03, Sam88]. Righting [LBD^{+}04]. Rights [GB00]. Rigor [Gla09d]. Rigorous [Eme95, Kle98]. Ring [Rei04a]. RISC [Daw95b]. Rise [LL92b]. Risk [Bai98, BD97, Car97, Cha96b, CAW97, CS97, DL03, Fai94, Fai05, Feo97, GPW97, Gli08, GB00, KM02, Kän97, KL97b, Lis97, Mad97, MR09, Moy97, SK08, WWD97, WMH07, Boe91]. Risk-Based [Gli08]. Risks [BB08, HG00, LWE01, Rei02a]. RITE [Pat08c]. River [Boo09a]. Road [EMW^{+}97, Per99]. Roadmapping [RSO08]. Robert [Kru09a]. Robotics [Fei07]. Rocky [Cur95]. ROI [BCM^{+}04, BHJM04, vS04, vS09]. Role [Bas87b, CC95b, Gla08a, Hat01, KCD09, MCA00, PL03, Pfl05, Sut00, Voa04a]. Roles [ZF99]. ROM [Rei95e]. Rome [Gla06a]. Ron [Mey96a]. Rookie [Hoh99, Pre98]. Roundtable [ABG^{+}97, ABB^{+}00, AAB^{+}00, DHK^{+}01, GHH^{+}09, MIZ^{+}98, McC01d, PAD^{+}98]. Roundup [GLS99]. Routers [DW00]. Routinely [McC98d]. Routines [Fow03a, McC98d]. Rover [San98e]. Royce [Dav95b, RHK^{+}03]. RPDE [Har87]. Ruby [BK07, Vis08]. Rule [ARW96, VNS94]. Rule-Based [VNS94, ARW96]. Rules [KL98, Kru09a]. Ruminations [DSBM99]. Runaway [BLP96]. Runaways [GF99]. Runtime [BS00, FHS^{+}06]. RUP [NCC^{+}02]. Rush [Gla94a, GB00, McC99a, Ols93]. Russian [Ter01].

S* [DDMS86]. Saboteur [Har05d]. SAC [Gla05c]. Sad [Gla05c, WJS97]. Safe [Boe00]. Safety [Bha96, Eve95, EH95, HK94, LCS91, MTKE94, RBBP03, RD99, Tri02, Wil94, YJC09, dSA95]. Safety-Critical [RBBP03, Wil94, YJC09]. Sailing [KIB^{+}08]. Saluting [Ano03m, McC02g]. Sam [San98c]. Same [Bac95c]. Sameness [Erd07c]. Sank [FW03]. Satisfies [Yam99]. Satisfying [Vic07]. Sauce [Voa04b]. Savvy [CB00, Rei95d]. Says [Ano92s, Har04a]. Scalable [RG02]. Scale [BRS^{+}85, LB00, LL92a, MA00, NGBR05, SRK^{+}09, Sel09a, THDK06, ZZ93, Cha96b, HF95, Mye89b, Rus91]. Scaled [Bar96a]. Scaled-Back [Bar96a]. Scaling [LB00, LCE^{+}01, RME03]. Scanning [VMMF00]. Scarcity [Boo09e, Boo09f]. Scenario [HSG^{+}94, KABC96]. Scenario-based [KABC96]. Scenarios [Mye98, WPJH98, ZMT05]. Schedule [AHM86, AH90, Gla04a, Lit06, McC96d, San98e]. Schedules [Dak96e, DDS^{+}97]. Scheduling [RS84, ZLJW95]. Schema [BPS94, Hon09]. schemas [SM88]. Scheme [Sha97b]. Schmidt [Mil01b]. Schools [DeM90]. Science [Ano93e, Ano93m, Che97, DFL00, EMW^{+}97, FPG94, GBdlHQCG^{+}00, Hoa84, JL03, McC98c, McC01b, Par99, RS05, TSP^{+}09]. Scientific [AKM^{+}08, CAGR91, Kel07, MSJ^{+}09, Pra85, SK08, SM08b, SM09b, Tom03, VVS^{+}08, WK03, WMGM08]. Scientist [VVS^{+}08]. Scientists [DG09a]. SCM [Ano98h]. Scope [Kor86]. Scoping [Bus09b, LJ97]. Scorecard [Mai02]. Scratching [SLG^{+}00]. screen [SM88]. screens [Shn91]. Screenwriting [Nor07]. Scripting [Spi05b]. Scrum [RJ00]. Sculpting [Jur97]. Scylla [Dav97a]. SD [Ano98o]. SDM [Rob99]. SE [Bac97, Hil97, McC97f, MCL97]. SE2004 [LLS^{+}06]. Search [Alo03, Gla02c, Got06a, JMVS06, LL87, NT92, Rei00a, YM94b, Shn97]. Seaside [DLR07]. Second [Ano01f, DW00, Mar88, Ano89]. Second-Generation [Mar88]. Secondary [Ano92p]. Secondary-Use [Ano92p]. Secret [Voa04b]. Secrets [Cla02, DSB99, Mat00b]. Secure [Ano02j, BC08, Got02a, GD02, HC02, HM02b]. Secure-Software [BC08]. Securely [Ano01h, GHW02]. Security [ABB^{+}00, Ano96d, Ano98g, Ano98i, Ano01h, Ano05k, Bar98, BC08, BS00, BS03, EL02, Got01d, HAJ07, HEH06, HM02b, Irv97, IAJR97, KH02, Kus95, LRO02, MA00, PC97, SM99, Spi09b, TJM08, VMMF00, WLC01, Pfl97a]. Security-Aware [KH02]. SED [GJPD99]. SEDOS [DV89]. Seduce [Nie96c]. see [MR95]. Seeing [Jac04]. Seek [PM98]. Seeking [Rob09b, Shn94]. Seen [Rei95c]. SEI [Ano93-27, Das94, Nor02]. SEL [BG94, BZM^{+}95]. Select [Rug93]. Selecting [CFQ07, Coc00, HS91, Jør09, MKCC03, NM96, PS92]. Selection [BCMV05, FC03, MN98, Sch95a]. Selective [DT94, HR90]. Selenium [BKW09]. Self [Ano96r, CG08, DAEE08, FN99, MDD09, Mou01]. Self-Assessment [DAEE08]. Self-Healing [Mou01]. Self-Help [Ano96r]. Self-Management [CG08, MDD09]. Self-Tuning [FN99]. Selling [DFS93]. Semantic [OHM^{+}07, SBBK08]. Semantics [BR85]. Sense [Dan02, Kau99, McC01c]. Sensors [JW09]. Sentence [HT04b]. Separate [Gut86]. Separating [Fow01c, Hud88, OLPS04]. Separation [KL04, KR03]. SEPG [Cur95]. Sequence [RW04]. Sequential [CC94, OCHW91]. Serious [SM01]. Server [Ano93p, Frü99, Pou99, RG02, San98f]. Server-Side [RG02]. Service [Ano92l, ACM^{+}07, BBGK07, Bla07, GKMM04, GYvdR06, GB00, Jon05, LM07, MKHV06, NZ07, Sta06a, VZM^{+}07, Ols95, Smi91]. Service-Centric [NZ07, VZM^{+}07]. Service-Oriented [Bla07, GKMM04, LM07, MKHV06, Sta06a]. Services [Erd07d, Lou06a, Lou08, SM09a, VVJ06, Kee00]. Servicing [Mai06b]. SESA [VZM^{+}07]. Set [Ven05, HSA89]. SETI [Ano97e]. Setl [KSS84]. Setting [Hec99, MVS^{+}99, Wil99b]. Seven [BH95a, Erd09e, Gra02, Hal90, Rei01, Rei04b, Tur03]. SEW [Mcs96]. SGML [Lew98]. Shadows [Mat95b]. Shake [MRE05]. Shall [Gla05b]. Shallow [Pou95]. Shalt [BHR^{+}98]. Shaping [Mye02]. Share [Ano97g]. Shared [Bec90, OC88]. Shared-Memory [OC88, Bec90]. Shareware [Got99a]. Sharing [DVG09, HL05]. Sharpe [Hor98a]. Shaw [Pfl97b]. Shed [Mat95b]. Shedding [Ano99p, SHZ^{+}98]. Sheet [Dio93]. Shelf [For92, LCB^{+}09, PMD07, Got02a]. Shell [ABCF91, TC86]. Shelves [Wil09]. Shift [Ano93-37, DCF^{+}02]. Shifts [Bar96a, Ksh05]. Ship [Rai07f]. Shirley [Ano92m]. Shlaer [FHRK93]. Shlaer-Mellor [FHRK93]. Shoes [Dav95c]. Shop [Hec99, Hoh05, LD95]. Shore [DD95]. Short [Rok03]. Short-Term [Rok03]. Shortage [Jon99, San97b, Str99]. Shortfalls [Ano07i]. Shoudl [MC03b]. Should [And95, AS03, Har05a, McC98d, Rob02, Sto03, VVM08, Zve98a, Fur97]. Shoulders [KKM^{+}09]. Shriver [Har03c]. Shuttle [SK92, FCB00]. Shy [HT04b]. Side [RG02, SB95b, Ber99]. Sigma [Bie04, Car00a, MK03]. signaling [GCR94b]. Signpost [And03]. Signs [BLP96, Hen99]. Silicon [OCM^{+}84, Sch99]. Silver [Ano93s, Ber08, Dro06, FM08, Moo01c, Spi07e, Wei99]. Simple [Boe00, Dav95b, Pem87]. Simplicity [Sha01, WB07f]. Simplified [Ram94]. Sims [San95a]. Simulating [SHM02, TL96]. Simulation [RFHM86]. sin [SGW02]. Singapore [CW89]. Single [Ven05]. Sinking [HKNS01]. Sins [McC01e, Rei01]. Site [BM01]. Sites [Nie95a, OBR^{+}08]. Sitting [McC00c]. Situated [BLSS08]. Six [Bie04, Car00a, McK02, MK03]. Size [Bie00, DG09b, KL88, Wit90, Hat97a, VT88]. Sized [KMSW00]. Skills [Sef99]. Skinner [Har05e]. Skipping [CCS^{+}03]. Slack [Mye01]. Slashes [San95a]. Sleaze [Tru94]. Slicing [TL98]. Slightly [Rei94b]. Slippery [AH96]. Slogans [Pal96]. Slowdown [SM08a]. Small [Ale06, ASC07, Boo06b, CTC07, DLSZ05, Dyb00, FHMU85, GJPD99, GE04, HPG^{+}00, Kau99, KMSW00, LFW99, LFW00, Mai02, MH07, NC00, PHRS07, Rei96c, RvW07, RJ00, Rob03, RM00, vWAS06]. Smalltalk [DM87, NR89]. Smart [Rei95d]. Smiling [PM98]. Smoke [McC96c]. Smooth [ANB99]. Snapshot [Gla99d]. Snapshots [Cha97d]. soaked [Ano97e]. SOAP [Lou06a]. Soapbox [Bat98, Bel99, Bol00, Cur98, Dav00, Dor99a, Fuj99, Gla97, Gor95, Kic96, Mar00, Mat95a, Mat00a, Mel95, Nus97, O’N97, Pal96, Pes97, Pon98, San97b, Wie98a, Wil99b, Yam99, Yam98b, You00]. Soccer [Pfl02]. Social [BM04b, Kal94, WSD^{+}09]. Society [Kal94]. Sociology [AR04, Gla00d]. Soft [Dak95b, Dak95c, Dak95d, Dak96b, Dak96a, Dak96e, Dak96c, Jon07, McC00b, Spr95, Zad94]. Softbot [ELST95]. Softlaw [Dak96f, Dak97a, Dak97b, Gra99a, Gra99b, Gra99c, Gra00d, KL97a, Sku99, AK99, Car00b, Gra00a, Gra00b, Gra00c]. Software [Aae03, ABF05, ABB^{+}93, AHM86, AH90, Abe07, ABL89, AKM^{+}08, AJM06, Ada95, Aha99, AvGCS07, AFGD01, ARTZ03, And98b, Ano84a, Ano89, Ano92i, Ano93e, Ano93k, Ano93w, Ano93-29, Ano93-28, Ano93-27, Ano93-40, Ano95c, Ano96d, Ano96v, Ano97d, Ano98g, Ano98i, Ano98-27, Ano99d, Ano99f, Ano00i, Ano00j, Ano00h, Ano00m, Ano01g, Ano01h, Ano01i, Ano02i, Ano03h, Ano03i, Ano03j, Ano03k, Ano04o, Ano04p, Ano05m, Ano06k, Ano07i, Ano08e, Ano09a, Ano09e, AW03, Aoy98, ABFP86, AD03, Bac95c, Bag99, BS86, BLSS08, BH02, BHS85, BH01, Bar95d, BLP96, BHAW09, Bas90, BG94, BZM^{+}95, BBD^{+}04, BCC^{+}08, BRL^{+}03, Bas87c, Bas07, BKS98, BST^{+}94, BCKS01, Bei97b, Ber03, Ber96b, Ber08, Ber94, BC08, Bie04, Bin97, BCH08, Bla07, Boa95, BCM^{+}04, BDKP99]. Software [Bøe06, Boe84, Boe91, Boe96, BD97, Boe00, BF00, BHJM04, BV08, BB08, BM91, Bol02, BALV03, Bol04, BVB04, BM09, Boo08c, Boo08d, Boo09e, Boo09f, BDA^{+}99, BD94, BCC^{+}06, BC86, BW96, BHKR01, BMJH96, BCH^{+}01, Bud99, Bur95a, BPD02, BAB^{+}87, BHS07, Byt99, CTC07, Car94, CZ01, CA01, CFQ07, Car88, CT90, Cha94, CC99, CAW97, Cha96c, Cha99b, CTW98, Che95, CE09, CBMP05, Che84, CJK^{+}02, CT02, Chi02, CW89, CSHA07, Chu93, CLB03, Cla02, CJNM05, CS09, CM90, Coc01, Col94, Col02, Con02, CF02, CS97, Coo95, CHW98, Cos01, Cos03, Cou99a, Cou99b, CMA^{+}02, Cox90, Cri06, Cuk05a, Cuk05b, Cur95, Cus03, Cus89, CMKC03, CMKC09, CR92, Dak95d, Dak97a, DM06, DSBM99, DLSZ05, DJSS97, Dav96a, Dav97e, Dav98e, Dav00, DN97, DM96, DeM09, DDS^{+}97]. Software [DG09a, Ded09, DCH04, DAB06, DH93, DS97, DJM04, DBS09, DVG09, DT03, Don06, Dor99a, Dor99b, DFL00, Dua02, DBL04, DB99, DSB99, DLV99, Dyb00, DKJ05, DD09, Ebe01, ED01, Ebe08a, Ebe08b, ES09, Ebe09, EK08, End93, Eng99, Erd07e, Erd07f, Erd08a, Erd08b, Erd08e, Erd09d, Erw09, EAS94, Fai94, FW03, Fai05, FPM01, Fei07, FPG94, FKN02, Fer01a, FNO92, FJWC01, FNY09, FO99, FK04, FCB00, Fra95, Fra99, FC03, FM08, FHMU85, Frü99, GC85, GS90a, Gee06, Gem85, Ger90, GHW02, Gla85, Gla94b, Gla98c, Gla00a, Gla00e, Gla01b, Gla01c, Gla03a, Gla03b, Gla03c, Gla03d, Gla04d, Gla04e, Gla06a, Gla06b, GCB^{+}06, Gla07a, Gla07b, GRM08, Gla08e, Gla08f, Gla09a, GdHN^{+}08, GHH^{+}09, GKMM04, Gom06, Gon05, GBdlHQCG^{+}00, Got00, Got01c, GAS^{+}01, Got02a, Got08b, Got99b]. Software [GF99, GLT03, Gra90, Gra00a, GH00, Gri95, GSS^{+}06, Gro04, GRS02, Gum06, GK02, Hal96a, Hal09, Han06, HG00, Har03e, Har03b, Har04c, Har04e, Har05e, Har05f, Hat95, Hat01, Hay03, HDS05, HEH06, HH02a, Hec99, HKNS01, Hei01, HF02, Hen94, HM01a, HH02b, Ho89, HL01, HM02b, HM01b, Hon09, HKPS96, Hor97, Hor98b, Hou01, Hsi96, HB06, HAK^{+}96, HJA08, Hum88, HC91, HSW91, Hum00, HT02a, HT02b, IS94, Jac95, Jac98c, Jac99, JC00, Jac93, JBR99, Jak00, Jan04, JBHD08, JS08, JKM^{+}09, JB96, JD98, JMDB00, JOL^{+}00, JKP^{+}05, Jon94, Jon99, Jon03, Jon07, Joo94, Jor05, JBR09, Ju98, Ju01, Jul93, JKC04, JW09, Jur97, JM02, JMS06, KG87, KFP88, KH00, Kan09, Kan93, KN96, KS04, Kel07, KH02, Kil01, KV07, KL96, KL94, KABM09]. Software [Koc00, KSMS02, KSTM98, KSS84, Kru05a, Kru05b, KOS06, Kru08a, Kru08b, KCD09, Ksh05, Kuh98, LANC07, LFW99, LFW00, LCM06, Lan96, Lar01b, LSB02, LM03, LBD^{+}04, LY98, LNR98, Leh98, LSW87, LSF03, LLS^{+}06, Lev90, LCS91, LO90, Lew98, LCB^{+}09, LT09, Lin93, LR00, Lin87, Lit04, LRO02, Lou06c, LCE^{+}01, LB06, MKCC03, MSJ^{+}09, MN98, Mai02, Map85, MH07, MDWTR93, MNA05, Mat95b, Mat00b, MF00, Max01, McC96e, McC97a, McC97c, McC98c, McC99d, MT99, McC00d, McC00a, McC00e, McC01d, McC01f, McC02e, McC06, MNJP02, McK02, MC03b, McL05, Med05, MH06, MRA98, MS04, MLI04, Mil84, MAMY99, Mil01c, ME02, MJ88, MDD09, Moh01, MLL03, MSK00, Moi01, MTKE94, Moo99, MJP^{+}00, Moo01c, Moo01b, Moo01a, Moo02b, Moo03, MD97, Mot06]. Software [Mub08, Mul86, Mun98, MKF06, Mus85, ME90, Mus93, Mye95b, Mye95d, Mye02, NZZ09, NZ07, Nav95, NS05, NOK08, NSV^{+}08, NM96, NT06, NK04, Nor02, NC00, Nus96, OG07, OGE08, Off02, Ols93, Pal96, Pal01b, PJT09, Par99, PL03, Pat08a, Pat08c, PWS^{+}95, Pfl97b, PM00, Pfl02, Pfl05, Pfl08, PM98, PMM93, PS90, Pos85, Pot93, PS84, Pre95, Pre96a, Pre96b, PDF87, PD93, PCOM97, PT05, PTH^{+}09, QBN^{+}06, RBBP03, RP95, RRD07, RBH07, Ree00, Ree05, Ree06b, Ree99, Rei00b, Rei00c, Rei01, Rei02a, Rei03a, RIP^{+}01, RvW07, RF09, REM^{+}09, Rif01a, Rif01b, Rif09, Rin97, RJ00, RBIM97, Rob05b, RD99, Rog00, RAv96, RKJ^{+}97, Ros04, Ros05, Row96, Roy00, Roy05, RJI02, RE04, RS05, RL02, RM00, SBM02, Sam90, San97a, San95a, San98a, San98g, SK08]. Software [SVSL08, Sch95b, SvHB02, Sch07, Sch99, SM06a, SM08b, SM09b, SK03, SS06a, SKC93, Sha96, SRW00, Sha90, SC06, Sha09, SD88, SM99, SSZ06, Sku99, SS95, SO90, Sne84, Sot01, Spe99, SC00, SS04, Spi07b, Spi08b, SJM02, SC03, Ste02, SV03, SM06b, Str02, SHM02, Sut00, Sym88, SM01, TV99, THDK06, THV09, TM97, Tar09, TL96, Ter01, Tha03a, Tha03b, TRW03, TR05, Tiw04, Toc97, Tom03, Tom87, Tot06, Tri02, TB02, UHC02, VBM^{+}02, VVM08, VE05, VV00, VM95b, VCM^{+}97, Voa98b, Voa99a, Voa99d, Voa99f, Voa99c, Voa00b, Voa00a, Voa01a, Voa01c, Voa03, Voa04a, Voa04b, WGK^{+}02, WF89, WEL02, WW01, WP86, WKA96, Weg84, Wey98, WOBP00, WHG^{+}09, Whi00, Whi01, WA02, Wie99, WS00, WMH93, Wil97c, Wit90, WSD^{+}09, WES87, WMGM08, WC89, Xu03]. Software [YKIY93, YJ07, YJC09, You95, YD01, ZZ93, ZJ98, ZS06, ZME00, Zho89, Zve98a, Zve03, dSSS^{+}09, vGvDSV01, vS04, vS09, vV06, vdL02, vdLLM09, vWAS06, vWHvW09, AO89, Bre89, Dak97b, Duk89, FN88, FKU97, Fis89, Fur97, GA05, HF97, HHM94, Hum96, KABC96, LC87, McC97b, MMS88, MA89, Mye89b, O’N97, OJ97, PJCK97, RC89, SA97, Tho89, Tru91, Was96, Ano99e, Ano99o, Ano02a, Ano03a, Ano05a, DeM95b, Erd07b, FTB06, GF99, Har03a, Hor97, Nav95, O’L09, Pfl97b, Rog00, Sch95b, Ano01f]. Software-Based [TM97]. Software-Development [MF00]. Software-Engineering [BLP96, CR92, Sym88, Ano93-27, Pot93, RC89, Tho89]. Software-Reliability [ME90, Mus93]. Software-Research [Gla94b]. Software-Speak [McK02]. Software-System [BST^{+}94]. Solid [Fer95, Nav95]. Solution [Ano93l, CM96, Gla03a, Gla04b, KSMS02]. Solutions [Ano98n, MRA98, Pat08b, Pes97, Sim95c, TM97, San98e, WE04]. Solve [AW03, MC94]. Solved [PM97]. Solving [JMS02, Rob05b]. Some [AS03, FW03, Gla04d, Kap00, McL05, Mye01, Rom90]. Someone [Sam90]. Something [Nie97b]. Sometimes [Ker08]. sonification [MR95]. Sorcery [Hoa84]. Sorting [Car00a]. Sorts [Mai09b, URK01]. Soul [SK03]. Sound [Rei96a]. Sounding [WSL^{+}99]. Soup [Pfl05]. Source [Abe07, Ano03h, Ano03i, AS03, BNST99, Cha98b, DGC^{+}07, Ebe07, Ebe08a, Ebe09, Gla99c, Gla00d, GAS^{+}01, Got07, Hec99, HS90, KHA^{+}03, MH07, McC99c, OG07, Ray99, Rem05, San98a, San08, SS06a, Spi06d, SV03, Tot06, VVM08, WW01, Wil99b, WLC01, vdLLM09, vWHvW09, Sme92, DFLS05, FK04, GA04, Has05, KT04, Lus04, MD04, McL05, NK04, Pud04, RE04, Sca04, SCSC04, SS04, TH04b]. Source-Code [HS90, Sme92]. Sources [Gla04e]. South [WC89]. Space [Got02a, Kor96, Lie02, SK92, LCS97, FCB00]. Spammers [Ano05l]. SPC [GB00]. Speak [McK02]. Speaking [Boo07e]. Spec [PM98, BL90]. Special [BZM^{+}95, DKKK90, DKMS91, Hor97, MS95]. Specialization [Jac94]. specialized [Lou95]. Specializing [Jac99]. Species [Dak95e]. Specific [BS00, CRR09, CM07, Fow09, HTLPS09, JRS^{+}09, KP09, SMTS09, WHG^{+}09, MSG96]. Specification [BL90, DG90, GFB96, GHW85, HLS90, HKK93, Lea91, Lv85, ND90, Pos85, RP09, STM88, Spi03, AMM^{+}89]. Specifications [BFP95, BK05, Boe84, DWH86, Den91, GGJZ00, Mey85, Pal01b, Wil94, FN88, LST91, TR88, Win88]. Specifier [Win90b]. Specifying [RP09, Ros91, SJ86, Spi90, WKUI89]. Speculating [Zan94]. Speculation [Ano97g]. Speed [BRL^{+}03, Ste95b]. spend [PSV94]. SPI [BGS01, DCF^{+}02]. Spice [EG96]. Split [KGRB03, San97b]. Sport [TB02]. spot [KHA^{+}03]. Spotlight [Ano96g, Ano96l]. Spotlights [Cur95]. Spreadsheet [DW90, MR09]. Spreadsheets [Erw09]. Spring [SR91a]. Sprinting [Got07]. Spur [ZHL^{+}89]. Spy [Spi07c]. SQA [Bak01]. SQml [San98e]. Squid [SDML^{+}06]. Staff [Ram00]. Stage [HS90, KR05]. Stage-Gate [KR05]. Stages [Mac99c, Mot06]. Stake [GB00]. Stakeholder [DRR^{+}07, MSG^{+}07, WMH07]. Stakeholders [AR04, Dam07, GW07]. Stand [Erd08e, Voa00c]. Standard [Ano99f, Bha96, Bøe08, LLH^{+}05, MS95, BFC00]. Standards [Ano84a, Bac95c, Bil94, BBRTL90, BBMH05, Car88, Cha96d, Fen96, Gar05, Gla09a, GB00, MIZ^{+}98, Moo99, PM98, Sch96b, GA05, Win90a]. Standards-Based [GB00]. Stars [Rei94b]. Start [Ano98h, Ber94, JBHD08, Spi09d, Sut00, ASM89]. Start-Up [Ber94, JBHD08, Sut00]. Start/Pat [ASM89]. Started [Ano93k]. Starting [Sch98b]. Starts [Ano99p]. Startup [Hen99]. State [Ano98g, Ano03k, Ano05n, BHJ^{+}03, BW98, CMH92, CLB03, CMKC03, CMKC09, Fow05, Gla03c, Gla05c, GLT03, Hol92, LSF03, MCL97, NL03, Rei03a, TH02b, Wey03]. Static [AHM^{+}08, CE09, DH93, EL02, Hat95, Lou06b]. Statically [VMMF00]. station [CGR94b, Got02a]. Statistical [Car94, CM90, EA03, FCB00, JP03, Wel00]. Statistics [CDD08]. Status [Hum00, Luq92b, PDK93, PJCK97, PD93, WW01]. Steering [Roy05, Sam88]. Stencils [Cod09]. Step [DP03, GD02]. Steps [Ale06, OY95, Ros98a, SM92]. Steve [Ano92m, Har03c]. Still [GAO09, Gla06b, Gla08b, MC03b, Rei96c, Zan92]. Stimulating [OGE08]. stir [GA05]. Stirrings [Cou99a]. Stochastic [RKP08]. Stone [Gra90]. Stop [EPSJ93, MA89]. Stop-Test [EPSJ93]. Stories [WSL^{+}99]. Story [Ano05o, Coc01, Dés08, Gla01c, Gla05a, Gla05c, TV99]. Story-Driven [Ano05o]. Storytest [Mug08]. Storytest-Driven [Mug08]. Straight [MVS^{+}99]. Straightening [JOL^{+}00]. Strange [Gla05a]. Strategic [BB99a, Kap00]. Strategies [Bro96, DSBM99, Jar92, SSSW86, Var95, VBM^{+}02, YM94b]. Strategy [Ano92i, BCMV05, GD02, Rei96a, WB08c]. Stratified [AK03a]. Streamlined [Sto99]. Street [Lew84]. strength [Bro96]. Strengthening [WKCJ00]. Strikes [BTC^{+}99]. Strip [MM08b]. Structural [Ksh05, SVSL08, WGM85, WTMS95]. Structure [AEP96, Kru05a]. Structured [Ano92i, Che02, Knu87, Mil86, Sca89, TS01, AMM^{+}89, MMN^{+}89, War89]. Structures [CC93, DH93]. Structuring [Ano98n]. Students [RP98]. Studies [Bur94, GA06, KPP95, McC98a]. Studio [Kuh98]. Study [CR08, CDM^{+}05, DLSZ05, DLV99, ELM90, ELLM99, FO99, HG00, Hei01, JD98, Kan01, KCF^{+}08, KAKG96, NE07, PJT09, PSHH04, Pau85, SRK^{+}09, Sne84, Tar09, Win88, ZAV02, CGR94b, CGR94a, GCR94a, GCR94b, Gue94, LD95, WTMS95]. Style [Ano92b, Bri97, Lub91, Nis00, Roy05, LH89]. Styles [JGM86, MKMG97, Sha95b]. Stylish [Mas98]. Subcontractors [NM96]. Subjective [Mir01]. Subroutines [CC94]. subscription [MCA00, RM00]. Substance [FPG94]. Subtleties [Lai08]. Success [Ahu99, Ano92i, Ano98k, Ben95, BTC^{+}99, Cla98b, Coc01, DC04, FI94, Gla02b, HL01, HL99, Mat00b, Nol99, Nus96, Ray99, Ree99, TV99, TTT^{+}02, Ven05, VE05, WSL^{+}99]. Successful [Ano92r, AW03, AW06, BM04a, Byt99, Chu93, GH00, Hou01, Kim04, Mat00b, Mey96a, Roy05, SA05]. Successfully [SO97]. Suitcase [McC00c]. Summaries [Ano96b, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99a, Ano99b, Ano99c, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano01b, Ano01c, Ano01d, Ano01e, Ano01j, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07a, Ano07b, Ano07c, Ano07d, Ano95b, Eve95]. Summary [SM99]. Summer [Hei01]. Supercharge [Daw95b]. Supercomputer [Mar85]. Superdistribution [Cox97]. Superfluous [Boo09c]. Superhighway [Gra99c, Tru94]. Superprofessionals [Erd09e]. Supply [Jon94]. Support [AAR09, ARTZ03, Ano92i, DJH^{+}08, Don06, Dow87, GK02, KSTM98, Lav00, MM05, MMM^{+}09, McC97e, San98f, THV09, Ter96, WCW85, vGvDSV01, TR88]. Supporting [BMM04, BMM05, FNY09, HDS05, JL06, RSO08, Tho89, VVS^{+}08]. Surrounding [Lit06]. Survey [Ano95c, Cha87, EK08, HDR^{+}86, HKPS96, HKN85, JKC04, LF05, RRD07, Run06, TJM08, WKH92, Mye89a]. Surveys [PM98]. Survivable [ELLM99]. Survival [Ano99d, HM01b, Ols95]. Survive [Ano92c]. Surviving [Ano99d, ED01]. Sustainable [WB09a]. SVR4 [CH93]. SVR4/MP [CH93]. SW [GE04]. SW-CMM [GE04]. Swap [Rei95a]. sweet [KHA^{+}03]. Switches [DW00]. Sword [Mat00a]. Symbol [Gut86]. Symbolic [TS06]. Symbolics [WMWM87]. Symposium [Ano93-27]. Sync [Hat98]. Synching [HKNS01]. Synchronization [Ano93-30, Wol88]. Synergies [San98g]. Synergy [PDHT97]. Syntactically [BK05]. Synthesis [ABB^{+}93, Jul93, Kan93, MD86, SKC93]. System [Ado96, AW91, Ano93f, Ano98u, Bal87, Bas87b, BST^{+}94, BHHM86, BH95b, CV88, CBK85, CCRZ^{+}90, kCCP^{+}95, Dav03, Don06, ELLM99, Emr85, EFM^{+}91, FO85, GC85, GS90a, GTF^{+}87, Gla09c, Got06a, Hal96b, HC02, HLS90, IAJR97, JB04, Kam87, Kim04, KH97, KYM^{+}04, Lam84, LLH^{+}03, LRO02, LK88, Luq92a, Mai08b, MMM^{+}09, Mic99, Nar85, Nie96a, Ols85, Pet85, Pfl98, PMM93, PS84, RFHM86, RGR85, Sch92, Sch98a, SH02, STSK92, SDP91, SMHT09, STM88, TYZP05, VBM^{+}02, WHC02, WPJH98, WLC01, WGFC86, CCY^{+}89, CFTZ89, CGR94a, Fid96, GCR94b, SSGT86, GCR94a, Ano02j]. Systematic [CKMV95, Ros98a, SO92, Son97, FI94]. Systems [ANB99, Ano92i, Ano92l, Ano93e, Ano93i, Ano93-30, Ano96d, Ano99p, Ano01h, ARW96, BS86, Bat98, BB99a, BMM04, Ben95, BLWG99, BK03, Blu87, Bol09, Bor85, BC86, BOS00, Che84, CR88, CC95a, CS90, CW84, DWH86, DV89, DJM04, DDPW09, Dou85, DRW03, EH95, FN99, Fer01a, Fer95, FK04, FLR04, GCR94c, Gla99d, Gon05, Got08b, Gra99b, Har06b, Hix90, HKPS96, JBHD08, KG87, Kan05, Knu87, Koh05, LF05, LSB02, Lau98, LM04, MCA00, Med05, MA00, MW09, NZ07, NOK08, OMTHC94, OYE06, Par96a, PCOM97, PH93, PWW05, RP06, RS84, RS01, RBBC03, Rem05, RG02, Ros91, Rov86, Sav08, SE98, SF93, Sel09a, SCSC04, Sev87, SMR93, Sne95, Spi05f, SN89, SRS^{+}86, TH02a, VNS94, Vo90, Voa98a, Wil97c, Yac03, dSA95, dLdCGR06, vGvDSV01, Deu88]. systems [EJW95, LB88, MW96, Mey91, SR91a, MSK00, Fer95, Gar00]. Systems-on-Chip [MW09].

Table [ERJI02]. Tablog [MMW86]. Tackle [RBH07]. Tactical [CA01]. Tailoring [BTC^{+}99, Osh98]. Taiwan [Ho89]. Take [Got01d, Got02c, SGM06]. Takes [Got00, GD02, KKC93]. Tale [Erd09b, Gla00b, Gla02c, Pfl96, Wey98, You98]. Taligent [San95b]. Taming [Bol00]. Tandem [Lou05]. Tangle [Cha97e]. Tangled [Pre00]. Tapestry [Dav98e]. Tapping [Hat95]. Tar [Ado96].

Tarchitecture [Hoh03]. Target [KL96]. Tarpit [PWS^{+}95]. Task [DBL04, Got09b, KL94, Lau03, RS84, WSD^{+}09]. Task-Based [Got09b, WSD^{+}09]. Tasking [BR85, HL85]. Tasks [LH01]. Tata [Kee00]. Taxonomies [GV95]. Taxonomy [AHL^{+}00, CC90, Gum06, PCS95]. TCP [Ano93f]. TCP/IP [Ano93f]. TCSE [Moo01b, Moo02b, Ree05]. TDD [JM07]. Teach [RP98, Zve03]. Teaching [BCH^{+}02, BPD02, HGK09, HH02c, Per90, SD88]. Team [Ano93l, Ano00m, DW00, EEM01, E^{+}94, FLN01, HMP^{+}01, Jac98a, Lus04, Mac99c, Par05, SC00]. Teams [AAR09, MDD09, RJ00, WSD^{+}09]. Teamwork [HH02c, Nic92]. Tech [CB00, Har05c, Har06f, Ros00, Ano92i]. Tech-Savvy [CB00]. Technical [McC98f, McC02c, Ree00, Spi09a, WEF01, Moo01b, Moo01a]. Technique [TL96]. Techniques [AFGD01, Ano92i, Ano96d, Ano00k, BS85, DSD02, DRW03, Ebe06, HR90, PWS^{+}95, PWW05, RF09, Sha84, WHG^{+}09]. Technologies [BS95, MSG^{+}07, NSV^{+}08, Pit00]. Technology [Ale95, Ano92b, Ano99e, BLSS08, Ban93, BDY94, BMM04, BW96, BM04b, CE09, CW00, CMA^{+}02, Cox84, Dav96a, DFS93, DNS^{+}06, Ebe08b, GGLW06, Gra00d, Har06f, Hon09, Jac93, JW09, LK03, MW91, ME90, PM00, Pou95, Rei95e, RF09, San95a, SS08, VZM^{+}07, Weg84, Ber99, CW89, Ho89, WC89]. Technospeak [Ano97g]. Ted [Har03c]. Tedium [Blu87]. Telcordia [Pit00]. Telecommunication [YKIY93, ZJ98]. Telecommunications [CT93, KAKG96]. Telecommuting [Rot00]. Telemetry [JKP^{+}05]. Telephony [Rei96b]. Tell [Dak97a, Dés08, HT04b, SV03]. Telling [Gla97]. Tempered [Boo07f]. Temporal [San98e]. Ten [Ano99d, Rei02a]. Tenets [Nor02]. Term [Got02d, Rok03]. Terminal [San98f]. Terminology [DB96, Ste02, Win05]. Terms [Sny93]. Tessa [Ko09]. Test [ABF05, Amb07, Ano92n, BFJT05, Che01, CDPP96, Cri06, Den91, DG07, EPSJ93, Got08a, Jan04, JS08, JHMW07, LJ88, MM05, Mar07, Mau90, McC96c, MR09, MC00, RP07, SM06a, Ste02, Var95, VK07, WGM85, WB09d, Yam98a, Yam98b, Zha04, Voa97a]. Test-Case [Che01, LJ88, Den91]. Test-Driven [ABF05, Amb07, Cri06, DG07, Jan04, JS08, JHMW07, Mar07, MR09, RP07, VK07, Zha04]. Testability [SB95b, VM95b]. Tester [Ano97g]. Testers [WOBP00]. Testing [And07, Ano96d, Ano01f, ARW96, Ber96a, CT91, CKMV95, Cou99a, Cou99b, DNS^{+}06, E^{+}94, Feo97, GCB^{+}06, Gla08b, Gla09c, GdHN^{+}08, Gon02, GD02, Gra02, Ham92, HS91, Hat08, HMP^{+}01, HM02a, HM92, How85, JDT01, JHMW07, JMS06, JMVS06, KL98, Lou05, LWN^{+}90, MM05, Mil01c, NS92, NSV^{+}08, Nie96d, Nie96b, Pet85, PS92, Pos95, PCOM97, Rok03, Run06, Sav08, She91, SR91b, THDK06, TS06, TYZP05, VMM91, WKW99, Wey98, Whi00, Wil97c, GMS^{+}96, KCC^{+}01, MA89]. TestNG [RP07]. Tests [BKW09, MM08b, RE03, TS06]. Texans [Ano97e]. Text [Alo03, Ano96t, DJ05, HKPS96, Nav95, Wil97c]. Thanks [Ano92s, Ano07k, Dav97d, McC00g, McC01g, McC02g, Str09]. Theater [MM08a]. Theft [Gla85]. Their [GB00, PSV94]. Them [Byt99, Mai06a, Zve98b]. Themselves [Got01a, Hef95]. Theoretical [Ano93m]. Theories [CSR^{+}99]. Theory [Ale99, DSB99, How85, Ros05, SKT^{+}92, Tau95, TB02]. There [Ara95, Gil96a, Gla98c, Jac98c, Mic00, Shu07b, Dak97b]. These [Gla02b]. Thin [HJA08]. Thing [Gla07c, JLLP03, Rei02c]. Things [Ano96r, Boo08d, Mac96]. Think [Hat98, NE07, RGR85]. Third [Cou99a, Cou99b, MJ88]. Third-Generation [MJ88]. Third-Party [Cou99a, Cou99b]. Thorough [HKPS96]. Thou [BHR^{+}98]. Thoughtful [DeM95b]. Thread [Dav98e, San03]. Threads [Wil97a, Wil97b]. Threat [IKBM08]. Threaten [Voa99d]. Threatens [Bar95d, Rei95b]. Three [HT04c, KT88, Mot06, Wel93]. Thumbs [Wei98a]. Thyself [Dav96f]. Ties [Ano97g]. Time [Ano92i, Ano93f, Ano93-30, Ano08e, Ano09e, Ara95, BS86, Bar98, Bau94, Bin97, Car94, Car00a, Cha99a, Clo01, Col94, CKMV95, Cou99a, Dak97a, Dav96d, DeM09, EM96, Erd08f, EH95, FL86, Fav96, Gom06, Got01e, Got03, Har05b, Hat95, HF95, Jac95, JD98, JM02, KKJ91, KL98, LA03, Lan96, LLH^{+}05, Lon03, LL03, Mar04, McC99c, Mil01c, Pfl97a, Pfl97c, Pfl98, PV97, PH93, RP06, RS84, Rob95, Ros98a, SVSL08, Spi90, SH93, Str96, SMK^{+}00, UA03, VV00, Voa97b, Voa99d, Voa99e, Voa99f, Voa99c, Voa00b, Voa00a, Voa00c, Voa01a, Voa01b, Wie99, WGFC86, Zve98b, CCY^{+}89, Deu88, FKU97, For88, LB88, Mye89b, Par96a, PSV94, Put91, SR91a, ZLJW95]. timed [CCY^{+}89]. Timers [Rei96b]. Timing [Car95, TBY96, Xu03]. Tips [Erd07e]. TMM [ABFP86]. Today [Ano84b, Ker99]. Together [Mat96, NC92, OG07]. Tolerance [DHK^{+}01, HKPS96, HT02b, KMY91, Voa01c]. Tolerant [YD01, dLdCGR06]. Tolerate [Hum01]. Tom [Bur94]. Tomorrow [Ano84b, Ano92r, Dav95d]. Too [Jon07, Lon02, Rei96c, Mun98]. Tool [ARTZ03, Ano97g, ABCF91, BLA88, DJH^{+}08, Far06, Got09a, GLS99, Har87, Hir05, Kem92, KKM^{+}09, KPP95, MHJ91, McC97e, Mul86, PG96, RH89, RS08, Sch95a, Sha95a, Sha97b, Spi05c, SM01, TN92, VB97, Vo90, WP86, WS00, YJ07, Yin94, vMC92, Lou95]. Tool-Integration [YJ07]. Toolbox [CC95a, Feo97, McC96a, New95, Oat98, Sch95a, SO97, Sha96, Sha97a, Sha97b, Ste95a, Tho97, Yin94]. Toolkit [McC00e, ASM89]. Toolkits [Oat98]. Tools [ABF05, And07, AHL^{+}00, Bac95b, Bai98, BS95, BMJH96, CH93, CN07, CMH92, CC95b, CC95a, CS96, Chu93, Dum95, Ebe09, For92, Fow09, HKR^{+}90, Hon09, HM92, KN96, Ker08, KKM^{+}09, LHB^{+}90, LWN^{+}90, Mar88, McL06, Mos92, MHB08, NDJ^{+}90, ONR^{+}90, Pes06, PS92, Pos95, RF09, Ros85, SCSC04, SB95a, SO90, Spi05d, Spi05g, Spi07f, Spi09c, Spi09d, Tot06, ZSM92, BLA88, Mye89a, OBM^{+}90, RSE^{+}90, Rei90, Tho89]. Top [Ano92n, Ano93c, Ano96u, Ano97i, Ano99d, Ano09a, Feo97, GAO09, LWE01, Sha97a, TM94]. Top-down [TM94]. Topics [Ano98v]. Topology [TC97]. Tortoise [Gra00d]. Touch [Shn91]. Tour [Tei84]. Toxicity [Jac98a]. Trace [MHJ91]. Traceability [Dic05, EG04a, Ram02]. Traces [KH97]. Traceview [MHJ91]. Tracing [Dak95e, Dav95b, EK02, HDS05, PG96]. Track [TS01]. Trackers [SC05]. Tracking [ABCF91, CCD09, McC97c, McC97e, Nie96e]. Trade [AH90, Fow09, Ker08, KKM^{+}09, MKCC03, Spi09c, Spi09d]. Trade-Off [AH90]. Trade-offs [MKCC03]. Trademarks [Gra99c]. Traditional [BT05, Fuj99]. Traffic [Got00, JC00, CGR94a]. Train [MVS^{+}99]. Training [GB00, HMP^{+}01, Mat00b, MCL97, Sef99]. Traits [Erd09e]. Transaction [Dak96c]. Transaction-Based [Dak96c]. Transactional [LSB02]. Transactions [ADT91]. Transfer [BDY94, Bur94, Dav96a, GGLW06, Har06f]. Transformation [ABFP86, Ros85, SK03, TR88]. Transforming [OCM^{+}84]. Transition [Kai99, SA05]. Translation [GM09, MBO93, Nak96]. Translations [McK02]. Transom [Ano92o]. Traveling [Hir07]. Treasure [VBM^{+}02]. Treatise [DeM95b]. Treatment [Pou95]. Tree [TP98, AMM^{+}89]. Tree-Based [TP98]. tree-structured [AMM^{+}89]. Trees [LCS91, Moe90, PS90]. Trenches [Ano96r, Ano99k, Ano99m, Dav95a, KNR^{+}01]. Trend [Ano93-32, Dak96c]. Trends [Bur94, BHS07, CDM^{+}05, GDD07, Hum00, LT09, Put91, SC03, Ste95a, Var95]. Triage [Sim04]. Trials [EG96]. Tribal [Boo08f]. Tricks [Dum95, Jak98, Lus04]. Tricky [JLLP03]. Trilogy [LNR98]. Trip [HT03b]. Trip-Packing [HT03b]. Triplex [KC90]. Troubles [HMP^{+}01]. troubleshooter [Got02a]. Truce [DB97]. True [Ham92]. Trust [AAR09, Gla06c, OBR^{+}08]. Truth [Boo07e]. Truths [Sid94]. Try [Bur94, Rai07c, SGW02]. Trying [Pre98]. TSL [Ros91]. TTCN [DNS^{+}06]. TTCN-3 [DNS^{+}06]. Tuning [Ano93f, FN99, HMK^{+}94]. Tutorial [Cha87, Dav85, Kod04, Som05, Vis08, Whi00]. Twain [Gla05b]. Twisted [Hou95]. Two [BALV03, DAS^{+}07, Erd09b, Gla02c, Gla06a, Gla08f, Hoh05, Hor95, LST91, Mil01b, PSV94, Wol89, You98, Tic89]. Two-dimensional [LST91]. Two-Phase [Hoh05]. Type [BPS^{+}03, Fow03c]. Typed [Ano08b, EG04b, TW07]. Types [BCMV05, HO87]. typo [ERJI02].

Ubiquitous [Cha99c, HT02c, Kle98]. UCC [KL97a]. UCITA [Bal00]. Ugly [Rei00b]. UIMS [KLR89]. UK [Car00b]. UL [Got02c]. Ullman [Ano98q]. Ultralarge [Got08b, Rus91]. ultralarge-scale [Rus91]. UltraSAN [CFJ^{+}91]. UML [Hou01, BK03, Boc03, CRR09, DSD02, DFL00, LCM06, MMSM09, MBE05, Mil02, RW04, Tho04]. UML-Based [DSD02, MMSM09]. UMLi [PP03]. Uncertain [Bar95b]. Uncertainty [Ano06p, FKN02, KL97b, Lit05, Lit06]. Unconferences [Ing08]. Uncorked [Cha99c]. Uncryptic [Pfl95a]. Undergraduate [LLS^{+}06, TR05]. Understand [HPLH90, ZSM92]. Understandable [BFC00].

Understanding [AR04, Ano05o, BCC^{+}08, BG98, BALV03, DJS08, Ebe06, GKMM04, LSS94, Pat07, Ris07, San08, Yac03, Fin96]. Underused [Lab94]. Unidirectional [NT92]. Unified [BDA04, BMGT01, BALV03, CTC07, JBR99, TC97, BAS89]. Uninitiated [Wil96]. Union [DM93]. Unique [SM01]. Unit [JMVS06, Lou05, OCM^{+}84, Run06, TS06]. Universal [Ano93j, Bha96, Sid94, Var88, Cod88]. Universal-Relation [Var88, Cod88]. Universe [MRA98]. University [PTH^{+}09, Wei98b]. Unix [Ano93f, BCM88, Ano93-33, Cur88, EFM^{+}91, Gai94, Spi05g]. Unpopular [McC96d]. Unquestionables [Gla03d]. Unravel [Cha97f]. Unraveling [VVJ06]. Unsound [KJC07]. Untangling [CS96]. Untested [Voa99d]. Up-front [WB08e]. Update [Ano98p, Ano98n, Bal00, BM09, JWKZ99, McC99e]. Updates [BMGT01]. Updating [CMKC09, SF93]. Upgrade [BC86]. Upon [Ara95]. Upstarts [Rei96b]. US&R [Vit84]. Usability [AFGD01, Ano93-34, Ano93t, Ano00k, BM01, Don01, E^{+}94, FJWC01, JWC01, Lin85, MSJ^{+}09, Nie96e, CN95a]. Usable [MRTS98]. Usage [BBGK07, CL02, RF09, Rog00, Sim06]. Usage-Centered [CL02, Rog00]. Use [Ale03, Ano92p, Ano95c, DFL00, Gor95, GV94, Hoh05, JB04, JML00, LX99, LSF03, Lon03, McC98d, MD97, Osh98, OBSN08, RS08, Rog00, SM09a, Spi07f, TRW03, Win05, Fur97, Nie97a]. Useful [CL00, Mye98, SJM02]. Usenet [Can95]. User [Ano92f, Ano92g, Ano93u, Ano04p, CN07, kCCP^{+}95, Cla98b, Clo01, DSBM99, Fow01c, HLS90, Har04d, Har89b, Hix90, KABM09, LH01, LHB^{+}90, LX99, MSJ^{+}09, Mai08b, MGG^{+}95, MD97, Mye89a, Ner01, Nie96c, Nie96d, Nie96b, Pat07, Pat08d, Pat09, PWS^{+}95, PP03, PRSV97, Rei94a, Rei94b, Rei95b, Rei95c, Rei95d, Rei95e, Rob09b, Rog00, SL97, VB97, Deu88, Per90]. User-Centered [Clo01, MSJ^{+}09]. User-Experience [Pat09]. User-Goal [Rob09b]. User-Interface [CN07, HLS90, Hix90, Nie96b, Har89b, LHB^{+}90, Mye89a, Per90]. Users [DDPW09, Hef95, Nie97d, Whi01]. uses [Shn91]. Using [ABF05, AvGCS07, AHZ02, Ano99d, AHM^{+}08, Bie00, Bif00, BC86, CS05, Che95, CNL99, CG08, DESG00, DSLB03, DP94b, DAEE08, EEM01, EL02, FHRK93, FHS^{+}06, Fow02b, FC03, GdHN^{+}08, GYvdR06, HL05, HO87, Hal96b, HAZ07, HS91, HPLH90, Hen94, Hir05, HA00, Hum96, IAJR97, KGRB03, KWM92, KM99, KC97, Kot98, KSS84, LCS91, LLH^{+}05, LWHS01, Lou06c, Luq92a, Mad97, MBO93, MKH98, Mar04, Mil02, MD90, Mir01, Mub08, MKF06, NT02, Nie95a, Nie95b, PH01, PS90, Raj00, RBH07, RIP^{+}01, REM^{+}09, RBIM97, Ros90, RE04, Rug93, Sch07, SM06a, Sha01, STSK92, Sho06, SFVFM08, Spi08c, Sta06a, Str01, TH02a, TS01, TYZP05, URK01, WKA96, YJ07, ZSM92, vGvDSV01, KLR89, NR89, Rei90]. Utility [BHHM86]. Utopia [Tho04].

v [Mil84, Che84]. V&V [Duk89, Dun89]. Vadis [BLP96]. Vague [MMM^{+}01]. Valid [Gla08b, SM95]. Validating [Boe84, Che01, HA00]. Validation [KL98, LL92a, MA89, WF89]. Valley [Sch99]. Value [ASC07, BHAW09, Fav96, Fav02, Gli08, Hat08, HB06, Jun98, KR97, KT05, KIB^{+}08, Lit04, Pat08a, TB02, YBBP05, Sim96]. Value-Based [HB06, TB02, YBBP05]. Value-Innovative [KIB^{+}08]. Value-Oriented [ASC07, Gli08]. Values [WB07c]. Valuing [WB08f]. Vardi [Cod88]. Variability [CHW98, KM99, TH02a]. Variation [Lar01a, Spi07a]. Variations [Jon03]. Vasa [FW03]. vector [LCS97]. vector-space [LCS97]. Vectors [NT02].

Vehicle [TRW03]. velocity [Ols95]. Vendors [Ano92g, Mar88]. Ventures [Ano92i, Ano85]. Venus [Mac00]. Verbing [TH03b]. Verification [DDMS86, KL98, Lea91, LCS91, MD86, SB90, TYZP05, VM95b, YJC09, ABL89, WF89]. Verify [Xu03]. Verifying [Boe84, OCM^{+}84]. Version [Hat97b, Lou06d, PCCW93, PHRS07, Spi05e, Spi05f]. versus [Aae03, Fow02a, Hor00, JOL^{+}00, Mac00, MJ88]. Vertical [WC86]. Very [Gla06a]. Vest [VBM^{+}02]. via [KH97, LSG^{+}05]. viable [Sme92]. Victor [Bur95a, SSZ06]. Video [Bal00, JL06, LM84]. Video-Based [JL06]. View [Ano93u, CG08, Eic03, Fis87, Gla04e, KCD09, SJM02, Kru95]. Viewing [Bas90, RH89]. Viewpoint [San08]. Viewpoints [DeM09, JBR09, MKH98]. Views [Tru94]. Violations [Sim95a, TBY96]. VIPS [ISO87]. VIRE [KIB^{+}08]. Virtual [CNL99, Fra98, Hol04, LH01, SLH01, SAG^{+}97]. Virtualization [SM06a]. Virtues [Cuk05b, Hay04]. Virtuoso [Ger90]. Virus [Ano06l]. Viruses [Gla05d]. Visibility [Rob95]. Visible [Ano92i, Hsi96, LY98]. Vision [CL91, EAS94, Ree91, KC97]. Visionary [Mat00b]. Visual [BCM88, Cha87, kCCP^{+}95, Etl96a, Fer01b, FLKMR91, HGK09, ISO87, Shn94, Bol00]. Visual-language [kCCP^{+}95]. Visualization [Han06, MR90, MHJ91, SI91, ZSM92]. Visualizing [AEP96, BCH^{+}01, HE91]. Vital [Gla99a]. Voice [DP94a, DBS09, HA00, JMVS09, Pfl08, vWS09]. Vol [Ada95]. Volta [MBL08]. Volume [Ano93f, Ano01a, Ano02a, Ano03a, Ano05a, Ano06a]. Volumes [Ano93m]. Voting [Sto03]. vs [Ano99e, Ano05k, CB00, Dor99b, Gla06a, Gon05, Hat97b, LB00, MVS^{+}99, SM01, TM94]. Vu [Dav96b]. Vulnerabilities [BS00, VMMF00].

W [Che02]. Walker [RHK^{+}03]. Wall [Dro06]. Wanna [Cur01]. Want [BPS^{+}03, Dum95, LRO02]. War [ZJ00]. Warehouse [Oat98]. Warehousing [BMGT01]. Warranties [Dak96d]. Warranty [Dak96b]. Warren [Har03c]. Was [Ara95, Boo07d, Got01e, SM95, SB96, RHK^{+}03, Wil99a]. Wash [FV97]. Wasn’t [Har05e, SM95]. Watch [Ano96o, CMA^{+}02, Rei96a, Rei96b, Wie98b]. Watching [SB96]. Water [Voa97b]. Waterfall [Lot97]. Waters [Ano97h]. Watershed [Bei97b]. Watts [Mye95d]. Wave [McL06]. Way [Bak01, CB98, Fro07, Hir07, Lus04, Spi08d]. Ways [Ano92h, BL92, Ker08, Lin93]. Weapon [Fer01a]. Weather [KCF^{+}08]. Weave [Pre00]. Web [OHM^{+}07, San98e, AHZ02, Aoy98, ACM^{+}07, BM01, BBGK07, BBMH05, BKW09, Cha97e, Cha97f, CS96, Clo01, CFW^{+}98, CB98, CL02, Dom98, DLR07, Got08a, HM02a, Hor97, Hor98b, KD02, Ko09, Koc98, Lou06a, Lou08, MHC03, Ner01, New95, Nie95a, Off02, OBR^{+}08, PSC02, Pre00, Raj00, Rei00c, SM09a, SDML^{+}06, SA07, URK01, VVJ06, Vis08, Wie99]. Web-Based [AHZ02, Aoy98, New95, PSC02]. Web-Service [ACM^{+}07]. Weighing [MIZ^{+}98]. Well [Boo07f, Kla95, Mai07, RP98]. Well-Tempered [Boo07f]. We’re [Bac97, Gla02a, Dak97b]. West [Ano98o]. Whale [Hor96]. Wheels [Hat95]. Whence [BTC^{+}99]. Where [Ano93-38, Cod09, Erd07d, Gla94a, HVF^{+}00, HRS08, IH90, MG01, Mai09c, VMM91]. Which [Bak01, Cur98]. While [Dec01, Gla09d, Smi91]. Whitaker [Cha97c]. Who [Ano93-39, Bat98, CC95c, DMN^{+}98, DDS^{+}97, E^{+}94, Fow03d, Gem85, Gra97, Har04a, McC96e, McC01d, Mel95, Nus97, Shu07a]. Whole [E^{+}94, JLLP03]. Whose [Ano09e, DeM09, Har03f]. Wide [Hor97, Hor98b]. Widespread [GV94]. Wiki [DRR^{+}07]. Wiki-Based [DRR^{+}07]. Wikis [Lou06c, RBH07, SBBK08]. Will [Cha97f, Dak97a, Got01c, Got02c, Hum01, Jac98c, Kle98, Mel95, Voa00c]. William [Ano97f]. Wily [Che97]. WIMP [Cur95]. Windows [LH01, SLH01, Ano92n, Ded09, Lou07, San98f]. Winston [Dav95b, RHK^{+}03]. Winter [And02]. Wire [PM99b]. Wireless [JW09]. Wisdom [Dav97c, Dav98b, NC00, SS96]. Within [Ano92i, Har05d, REM^{+}09]. Without [Hal95, Sim95b, BDS02, Boc03, MCA00, RM00, Sch04, SHZ^{+}98]. Witnesses [DJ05]. Wives [Gla00b]. Wives’/Husband [Gla00b]. Wobble [HT04c]. Woes [Bus09b]. Women [Cha97d]. Wonderful [WJS97]. Woolly [Bol00]. Woos [GAS^{+}01]. Word [Lew84, PB98]. Words [Dav97c, Dav98b, Gla02c]. Work [Ano92b, Ben00, Bin97, Cla98a, Cos94, Cus03, Dor99b, Fer98, Gar00, Gla01c, Gra90, Hei98, Hen99, Hor95, Hor96, Jac98a, KB02, KKC93, Mac98, Mac99a, Mac99b, Mac00, Mcl98, Mic00, Pat09, PV97, PM99b, Rot99, Rot00, San95a, VVS^{+}08, Nis00, OO92, Pfl95b]. Work-product [Gra90]. Workable [Dak96a]. Workflows [VVS^{+}08, WMGM08]. Working [LM03, NC92, Rei87, Spi05g]. Workplace [HHJC99, MSG^{+}07, Mcl98]. Works [Lus04]. Workshop [BLP96, MP88, Mye94]. Workstation [SD88]. Workstyles [CN07]. World [Ale99, Ano96d, Ano98i, Ano98t, BR99, Boo94, Cha97d, Col94, Coo95, Dav97b, DN97, DC04, HD00, Jac04, Kan01, KB02, MKMH07, Nie97c, Pal01b, Par03, Per99, Rok03, SM99, SMK^{+}00, WEF01, Hor97, Hor98b]. Worlds [Dor99b, KHA^{+}03]. Worldwide [CMKC03, MRA98]. Worries [MC03b]. Worry [Dic95]. Worst [Ano93-31, Jon96, KP09]. Worth [Ano93t]. Worthy [Rog00]. Would [Chr94]. Wraps [Bar97b]. Write [McC02c]. Writer [App02, Lew84]. Writing [Ano96d, BGL^{+}09, KL94, McC97d, Spi05c, WEF01]. Wrong [Gla98e, Gla02a, PB98].

XML [Hon09, Spi08c, WC06]. XMS [FO85, GC85]. XP [DNR04, MDM03, NCC^{+}02, Ras03, Rei03b, Wil03, WK03]. Xylem [Emr85].

Y2K [Ano97e, Ano98p, Ano98n, AP00, AK99, Cha99c, Gla00e, Jon99, Kap00, McC99f, Moo02a, PM99b, San98c, Sch98b, You00]. Ya [Gra97]. Year [Bur94, Lew98, McC96a, Rei95f, Sha97b, Wil97c, Zan94, Ano99d, BBD^{+}04, GAS^{+}01, Tho98]. Years [Ano08e, Erd08b, JMVS09, Mai08a, Pfl08, PT05, Wel93, Bro95]. you’ll [SGW02]. Young [Ano96r, GAS^{+}01]. You’re [Gra97]. Yours [GN95, Wey03]. Yourself [AK99, Gra00b, MCA00].

Zen [PWS^{+}95]. Zero [Ano92l, Gla08e, HT02b]. Zero-Service [Ano92l]. Zero-Tolerance [HT02b].

References

Arranga:2000:RCD

[AAB^{+}00] Edmund Arranga, Ian Archbell, John Bradley, Pamela Coker, Ron Langer, Chuck Townsend, and Mike Wheatley. Roundtable: In Cobol’s defense. IEEE Software, 17(2):70–72, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2070.
pdf.

Aaen:2003:SPI

[Aae03] Ivan Aaen. Software process improvement: Blueprints versus recipes. IEEE Software, 20(5):86–93, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5086.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5086abs.htm.

Al-Ani:2009:TDT

[AAR09] Ban Al-Ani and David Redmiles. Trust in distributed teams: Support through continuous coordination. IEEE Software, 26(6):35– 40, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Abbott:1993:MBS

[ABB^{+}93] Ben Abbott, Ted Bapty, Csaba Biegl, Gabor Karsai, and Janos Sztipanovits. Model-based software synthesis. IEEE Software, 10 (3):42–52, May 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anderson:2000:RIS

[ABB^{+}00] Ross Anderson, Terry Bollinger, Doug Brown, Enrique Draier, Philip Machanick, Gary McGraw, Nancy Mead, Art Pyster, Howard Schmidt, and Tim Shimeall. Roundtable on information security policy. IEEE Software, 17(5):26–32, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5026.
pdf;   http://dlib.computer.
org/so/books/so2000/pdf/
s5mt_rt_web.pdf.

Arango:1991:TST

[ABCF91] Guillermo Arango, Laurent Bruneau, Jean-Francois Cloarec, and Albert Feroldi. A tool shell for tracking design decisions. IEEE Software, 8(2):75– 83, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Aberdour:2007:AQO

[Abe07] Mark Aberdour. Achieving quality in open source software. IEEE Software, 24(1):58–64, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Abacus:2005:UTD

[ABF05] Alex Abacus, Mike Barker, and Paul Freedman. Using test-driven software development tools. IEEE Software, 22(2):88–91, March/ April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/so/2005/02/s2088.
                                        pdf.

Arango:1986:TSM

[ABFP86] Guillermo Arango, Ira Baxter, Peter Freeman, and Christopher Pidgeon. TMM: Software maintenance by transformation. IEEE Software, 3(3):27– 39, May 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anderson:1997:FER

[ABG^{+}97] James P. Anderson, Sheila Brand, Li Gong, Thomas Haigh, Steven Lipner, Teresa Lunt, Ruth Nelson, William Neugent, Hilarie Orman, Marcus Ranum, Roger Schell, and Eugene Spafford. Firewalls: An expert roundtable. IEEE Software, 14(5):60– 66, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1997/pdf/
                                        s5060.pdf.

Ackerman:1989:SIE

[ABL89] A. Frank Ackerman, Lynne S. Buchwald, and Frank H.

Lewski. Software inspections: an effective verification process. IEEE Software, 6(3):31–36, May 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Artsy:1987:ICC

[ACF87] Yeshayahu Artsy, Hung-Yang Chang, and Raphael Finkel. Interprocess communication in Charlotte. IEEE Software, 4(1):22–28, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Acly:1988:LBC

[Acl88] Ed Acly. Looking beyond CASE. IEEE Software, 5(2):39–43, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ardagna:2007:PFE

[ACM^{+}07] Danilo Ardagna, Marco Comuzzi, Enrico Mussi, Barbara Pernici, and Pierluigi Plebani. PAWS: a framework for executing adaptive Web-service processes. IEEE Software, 24(6):39– 46, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Austin:2003:BRS

[AD03] Rob Austin and Lee Devin.

Beyond requirements: Software making as art. IEEE Software, 20(1):93–95, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s1093.
                                        pdf.

Applewhite:2004:N

[AD04] Ashton Applewhite and Alan Davis. In the news. IEEE Software, 21(2):94– 99, March/April 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Adams:1995:BPG

[Ada95] Tom Adams. Bookshelf: Prescription for good software management: review of Quality Software Management, Vol. 3: Congruent Action. IEEE Software, 12(6):114–115, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Adolph:1996:CCT

[Ado96] W. Stephen Adolph. Cash cow in the tar pit: Reengineering a legacy system. IEEE Software, 13(3):41– 47, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Agrawal:1991:EBA

[ADS91] Hiralal Agrawal, Richard DeMillo, and Eugene H. Spafford. An executionbacktracking approach to debugging. IEEE Software, 8(3):21–26, May 1991. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Aksit:1991:ADO

[ADT91] Mehmet Aksit, Jan Willem Dijkstra, and Anand Tripathi. Atomic delegation: Object-oriented transactions. IEEE Software, 8(2): 84–92, March 1991. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Antis:1996:VSL

[AEP96] Jacqueline M. Antis, Stephen G.

Eick, and John D. Pyrce. Visualizing the structure of large relational databases. IEEE Software, 13(1):72– 79, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Araki:1991:GFD

[AFC91] Keijiro Araki, Zengo Furukawa, and Jingde Cheng. A general framework for debugging. IEEE Software, 8 (3):14–20, May 1991. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anderson:2001:IUT

[AFGD01] Jean Anderson, Francie Fleek, Kathi Garrity, and Fred Drake. Integrating usability techniques into software development. IEEE Software, 18(1):46–53, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s1046abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s1046.pdf.

Anderson:1987:IOL

[AGK^{+}87] J. Wayne Anderson, William F. Galway, Robert R. Kessler, Herbert Melenk, and Winfried Neun. Implementing and optimizing Lisp for the Cray. IEEE Software, 4(4): 74–83, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Abdel-Hamid:1990:ICS

[AH90] Tarek K. Abdel-Hamid. Investigating the cost/ schedule trade-off in software development. IEEE Software, 7(1):97–105, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Abdel-Hamid:1996:SPP

[AH96] Tarek K. Abdel-Hamid. Slippery path to productivity improvement. IEEE Software, 13(4):43–52, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Aharonian:1999:DPO

[Aha99] Gregory Aharonian. Does the Patent Office respect the software community? IEEE Software, 16(4):87– 89, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4087.
pdf.

Arranga:2000:CTO

[AHL^{+}00] Edmund C. Arranga, Thane Hubbell, Alden C. Lorents, Steve Shiflett, and Jon Wessler. Cobol tools: Overview and taxonomy. IEEE Software, 17(2):59– 69, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2059abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2059.pdf.

Abdel-Hamid:1986:ISE

[AHM86] Tarek K. Abdel-Hamid and Stuart E. Madnick. Impact of schedule estimation on software project behavior. IEEE Software, 3(4):70–75, August 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ayewah:2008:USA

[AHM^{+}08] Nathaniel Ayewah, David Hovemeyer, J. David Morgenthaler, John Penix, and William Pugh. Using static analysis to find bugs. IEEE Software, 25(5):22–29, September/ October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ahuja:1999:ILG

[Ahu99] Sanjiv Ahuja. Interview: Laying the groundwork for success. IEEE Software, 16(6):72–75, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s6072.
                                        pdf.

Altendorf:2002:UJL

[AHZ02] Eric Altendorf, Moses Hohman, and Roman Zabicki. Using J2EE on a large, Web-based project. IEEE Software, 19(2):81– 89, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s2081.
                                        pdf;   http://www.computer.

org/software/so2002/s2081abs.
htm.

Acuna:2006:EHC

[AJM06] Silvia T. Acuna, Natalia Juristo, and Ana M. Moreno. Emphasizing human capabilities in software development. IEEE Software, 23(2):94–101, March/ April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Allen:1985:PPE

[AK85] John R. Allen and Ken Kennedy. Parallel programming environment. IEEE Software, 2(4):21–29, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Arriola-Kern:1999:SPY

[AK99] Trinidad Arriola-Kern. Soft-Law: Protecting yourself from Y2K questionnaires. IEEE Software, 16(3):103– 104, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3103.
pdf.

Atkinson:2003:AOD

[AK03a] Colin Atkinson and Thomas Kühne. Aspect-oriented development with stratified frameworks. IEEE Software, 20(1):81–89, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1081.
pdf;   http://www.computer.
org/software/so2003/s1081abs.
htm.

Atkinson:2003:MDD

[AK03b] Colin Atkinson and Thomas Kühne. Model-driven development: a metamodeling foundation. IEEE Software, 20(5):36–41, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5036.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5036abs.htm.

Ackroyd:2008:SSD

[AKM^{+}08] Karen S. Ackroyd, Steve H. Kinder, Geoff R. Mant, Mike C. Miller, Christine A. Ramsdale, and Paul C. Stephenson. Scientific software development at a research facility. IEEE Software, 25(4): 44–51, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Alexander:1995:LDO

[Ale95] Bruce Alexander. Letters: Does object technology improve reuse odds? IEEE Software, 12(3):9, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Alexander:1999:OPT

[Ale99]  Christopher  Alexander.

The origins of pattern theory: The future of the theory, and the generation of a living world. IEEE Software, 16(5):71–82, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5071abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5071.pdf.

Alexander:2001:IQO

[Ale01] Roger T. Alexander. Improving the quality of object-oriented programs. IEEE Software, 18(5):90– 91, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s5090abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5090.pdf.

Alexander:2003:MCU

[Ale03] Ian Alexander. Misuse cases: Use cases with hostile intent. IEEE Software, 20(1):58–66, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s1058.
                                        pdf;   http://www.computer.
                                        org/software/so2003/s1058abs.
                                        htm.

Alexander:2006:SSB

[Ale06] Ian Alexander. 10 small steps to better requirements. IEEE Software, 23(2):19–21, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Astfalk:1992:IPM

[ALMS92] Greg Astfalk, Irvin Lustig, Roy Marsten, and David Shanno. The interior-point method for linear programming. IEEE Software, 9(4): 61–68, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Alonso:2003:GTS

[Alo03] Omar Alonso. Generating text search applications for databases. IEEE Software, 20(3):98–105, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3098.
                                        pdf;   http://www.computer.
                                        org/software/so2003/s3098abs.
                                        htm.

Amatriain:2007:CFA

[Ama07] Xavier Amatriain. CLAM:

a framework for audio and music application development. IEEE Software, 24(1):82–85, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ambler:2002:LAI

[Amb02] Scott W. Ambler. Lessons in agility from Internetbased development. IEEE Software, 19(2):66–73, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2066.
pdf;   http://www.computer.
org/software/so2002/s2066abs.
htm.

Ambler:2007:TDD

[Amb07] Scott W. Ambler. Testdriven development of relational databases. IEEE Software, 24(3):37–43, May/ June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Aoyama:1989:DSJ

[AMM^{+}89] Mikio Aoyama, Kazuyasu Miyamoto, Noritoshi Murakami, Hironobu Nagano, and Yoshihiro Oki. Design specification in Japan: tree-structured charts. IEEE Software, 6(2):31–37, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anandakrishnan:1999:PEG

[Ana99] Sridhar Anandakrishnan. Penguins everywhere: GNU/ Linux in Antarctica. IEEE Software, 16(6):90–96, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6090abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6090.pdf.

Aiken:1999:REN

[ANB99] Peter Aiken, Ojelanki K. Ngwenyama, and Lewis Broome. Reverse-engineering new systems for smooth implementation. IEEE Software, 16(2):36–43, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s2036abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s2036.pdf.

Andriole:1994:MFC

[And94] Stephen J. Andriole. Manager: Fast, cheap requirements: Prototype, or else! IEEE Software, 11(2):85– 87, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Andriole:1995:DDW

[And95] Stephen J. Andriole. Debatable development: What should we believe? IEEE Software, 12(4):13–18, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Andriole:1998:MPR

[And98a]  Steve Andriole.  Man-

ager: The politics of requirements management. IEEE Software, 15(6):82– ??, November/December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1998/pdf/
s6082.pdf.

Andriole:1998:SGB

[And98b] Steve Andriole. Software:

The good, the bad, and the real. IEEE Software, 15(1):45–47, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1045abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s1045.pdf.

Andresen:2002:DWO

[And02]  Scott Andresen.  Don

Winter: One CEO’s perspective. IEEE Software, 19(6):53–55, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6053.
pdf.

Andrews:2003:SMP

[And03]  Mike Andrews.  Sign-

post: Matching program behavior against known faults. IEEE Software, 20(6):84–89, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6084.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6084abs.htm.

Andrea:2007:ENG

[And07] Jennitta Andrea. Envisioning the next generation of functional testing tools. IEEE Software, 24 (3):58–66, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1984:ISE

[Ano84a] Anonymous. IEEE software engineering standards. IEEE Software, 1(2):

94–98, April 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1984:PCI

[Ano84b] Anonymous. Protection of computer ideawork — today and tomorrow. IEEE Software, 1(2):74–82, April 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1985:UJA

[Ano85] Anonymous. US, Japan AI firms enter joint ventures. IEEE Software, 2(4): 101, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1989:SRE

[Ano89] Anonymous. Software research in ESPRIT’s second phase. IEEE Software, 6 (6):54–57, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:ABP

[Ano92a]  Anonymous.  Analyzing

benefits of process improvement. IEEE Software, 9(4): 83–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:CTC

[Ano92b] Anonymous. Can technology create a humane work style? IEEE Software, 9(6):8–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:CYS

[Ano92c] Anonymous. Can you survive your management mistakes? IEEE Software, 9 (4):72–??, July 1992. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:ECD

[Ano92d] Anonymous. EC computing directives create controversy. IEEE Software, 9(4): 75–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:EPI

[Ano92e] Anonymous. Elements of a process-improvement program. IEEE Software, 9(4): 83–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:EU

[Ano92f] Anonymous. End user. IEEE Software, 9(3):120– ??, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:EUK

[Ano92g] Anonymous. End user:

Kinder, gentler Comdex eludes vendors, buyers. IEEE Software, 9(1):112– ??, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:ECW

[Ano92h] Anonymous. From the Editor: Count the ways you can contribute in 1993. IEEE Software, 9(6):6–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:HTV

[Ano92i]  Anonymous.  High-Tech

Ventures: The Guide for Entrepreneurial Success; Measures for Excellence: Reliable Software On Time, Within Budget; CASE: Computer-Aided Software Engineering; Effective Structured Techniques: From Strategy to Case; Making Software Development Visible: Effective Project Control; Evaluating Decision Support and Expert Systems; New Media. IEEE Software, 9(4):86–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:B

[Ano92j]  Anonymous.  In brief.

IEEE Software, 9(4):75– ??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:I

[Ano92k] Anonymous. Insider. IEEE Software, 9(3):14–??, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:IFF

[Ano92l] Anonymous. Insider: Flexible factories and zeroservice systems. IEEE Software, 9(1):12–??, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:ISS

[Ano92m]  Anonymous.  Interview:

Steve Shirley. IEEE Software, 9(4):75–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:MTW

[Ano92n] Anonymous. Microsoft test for Windows; top drawer. IEEE Software, 9(4):92– ??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:TGB

[Ano92o] Anonymous. An over-the-transom grab bag of goodies. IEEE Software, 9(4): 96–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:PPS

[Ano92p] Anonymous. Personal privacy and secondary-use dilemma. IEEE Software, 9(4):73–??, July 1992. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:PHC

[Ano92q] Anonymous. Perspectives on human-computer interaction. IEEE Software, 9 (4):70–??, July 1992. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:PTS

[Ano92r] Anonymous. Profile of tomorrow’s successful developer. IEEE Software, 9(4): 6–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1992:WIO

[Ano92s] Anonymous. Why industry often says no thanks to research. IEEE Software, 9 (6):97–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:BIM

[Ano93a] Anonymous. Beyond intelligent machines: Just do it! IEEE Software, 10(1):100– ??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:BEA

[Ano93b] Anonymous. Bootstrap: Europe’s assessment method. IEEE Software, 10(3):93– ??, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:CBI

[Ano93c] Anonymous. Checklist for buyers of ICASE products; top drawer. IEEE Software, 10(6):108–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:CCB

[Ano93d] Anonymous. Cobol: Common base, onerous legacy. IEEE Software, 10(6):112– ??, November 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:CPD

[Ano93e]  Anonymous.  Concur-

rency in programming and database systems; building quality software; the cleanroom approach to quality software development; research directions in cognitive science: European perspectives; new media. IEEE Software, 10(6):105– ??, November 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:CPR

[Ano93f] Anonymous. Constructing Predictable Real-Time System; Unix for Fortran Programmers; Unix Curses Explained; Internetworking with TCP/IP, Volume I: Principles, Protocols, and Architecture; System Performance Tuning; New Media. IEEE Software, 10(3): 106–111, May 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:DCA

[Ano93g] Anonymous. Defect causal analysis drives down error rates. IEEE Software, 10 (4):98–??, July 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:EP

[Ano93h] Anonymous. Esprit projects. IEEE Software, 10 (6):45–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:ETS

[Ano93i] Anonymous. Evolving toward systems integration. IEEE Software, 10(6):12– ??, November 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:EUI

[Ano93j]  Anonymous.  Examples

of universal interfaces. IEEE Software, 10(5):76– ??, September 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:GSS

[Ano93k]  Anonymous.  Getting

started on software metrics. IEEE Software, 10(1):108– ??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:GKY

[Ano93l] Anonymous. Getting to know you: The industryacademia team; Microsoft’s solution: In-house research. IEEE Software, 10(6):96– ??, November 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:HTC

[Ano93m] Anonymous. Handbook of Theoretical Computer Science, Volumes A and B. IEEE Software, 10(1):118– ??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:EAC

[Ano93n] Anonymous. Has the exploratory approach come of age? IEEE Software, 10 (1):104–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:HQF

[Ano93o] Anonymous. How quality fits into package development. IEEE Software, 10 (5):85–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:IOO

[Ano93p]  Anonymous.  Intelligent

offices: Object-oriented multi-media information management in client/ server architectures. IEEE Software, 10(1):118–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:IOB

[Ano93q]  Anonymous.  Interfaces

for the ordinary and beyond. IEEE Software, 10 (5):76–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:IM

[Ano93r] Anonymous. The Internet message. IEEE Software, 10(1):118–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:SBK

[Ano93s] Anonymous. Is the silver bullet knowledge-based? IEEE Software, 10(1):10– ??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:UER

[Ano93t] Anonymous. Is usability engineering really worth it? IEEE Software, 10 (6):90–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:MUP

[Ano93u] Anonymous. Maturity from the user’s point of view. IEEE Software, 10(4):89– ??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:NM

[Ano93v] Anonymous. New media.

IEEE Software, 10(1):118, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:NCB

[Ano93w] Anonymous. No crystal ball in the software industry. IEEE Software, 10(4): 91–??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:PAR

[Ano93x]  Anonymous.  Planning

ahead requires looking back. IEEE Software, 10 (6):5–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:RP

[Ano93y] Anonymous. Race projects. IEEE Software, 10 (6):55–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:RED

[Ano93z] Anonymous. Rapid evolutionary development. IEEE Software, 10(1):118–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:SSE

[Ano93-27] Anonymous. SEI Softwareengineering Symposium. IEEE Software, 10(6):101– ??, November 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:SPR

[Ano93-28] Anonymous. Software and privacy: Revising Orwell. IEEE Software, 10(3):92– ??, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:SL

[Ano93-29] Anonymous. Software lemmingineering. IEEE Software, 10(5):79–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:SRT

[Ano93-30] Anonymous. Synchronization in real-time systems: a priority inheritance approach. IEEE Software, 10 (1):118–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:IOA

[Ano93-31] Anonymous. To improve our average, focus on the worst. IEEE Software, 10 (3):5–??, May 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:TTL

[Ano93-32] Anonymous. Trend is toward less copyright protection. IEEE Software, 10 (5):82–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:UI

[Ano93-33] Anonymous. Unix For The Impatient. IEEE Software, 10(1):118–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:UED

[Ano93-34] Anonymous. Usability engineering in dollars and cents. IEEE Software, 10 (3):88–??, May 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:WMD

[Ano93-35] Anonymous. What is the merger doctrine? IEEE Software, 10(5):82–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:WME

[Ano93-36] Anonymous. What makes for effective measurement? IEEE Software, 10(6):94– ??, November 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:WBG

[Ano93-37] Anonymous. What’s behind the great paradigm shift? IEEE Software, 10 (6):93–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:WRP

[Ano93-38] Anonymous. Where’s the return on process improvement. IEEE Software, 10 (4):12–??, July 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:WLE

[Ano93-39] Anonymous. Who’s liable for electronic abuse? IEEE Software, 10(1):106– ??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:WSC

[Ano93-40] Anonymous. Why software costs so much. IEEE Software, 10(3):90–??, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1995:AI

[Ano95a] Anonymous. Annual index 1995. IEEE Software, 12 (6):88–98, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1995:AS

[Ano95b] Anonymous. Article summaries. IEEE Software, 12 (4):14–15, July 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1995:SFS

[Ano95c] Anonymous. Survey focuses on software industry’s use of Internet. IEEE Software, 12(6):109, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:AI

[Ano96a] Anonymous. Annual index 1996. IEEE Software, 13 (6):84–96, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:AS

[Ano96b] Anonymous. Article summaries. IEEE Software, 13 (4):26–??, July 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:Ba

[Ano96c] Anonymous. Bookshelf. IEEE Software, 13(4):133– ??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:BBB

[Ano96d] Anonymous. Bookshelf: Black box testing: Techniques for functional testing of software and systems. computing perspectives. BUGS in writing: a guide to debugging your prose. network security: Private communication in a public world. IEEE Software, 13 (5):98–??, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:Bb

[Ano96e] Anonymous. Briefs. IEEE Software, 13(4):138–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:CA

[Ano96f] Anonymous. Concurrency in Ada. IEEE Software, 13(4):133–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:CS

[Ano96g] Anonymous. Conference spotlight. IEEE Software, 13(4):138–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:DJD

[Ano96h]  Anonymous.  Dave Ja-

cobsohn, devoted friend of the IEEE CS, passes away. IEEE Software, 13 (6), November 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:ENE

[Ano96i] Anonymous. Ebert named to Editorial Board. IEEE Software, 13(3):5, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:ECP

[Ano96j] Anonymous. Expert C programming. IEEE Software, 13(4):133–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:N

[Ano96k] Anonymous. In the news.

IEEE Software, 13(4):138– ??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:NCS

[Ano96l] Anonymous. In the news:

Conference spotlight. IEEE Software, 13(3):104, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:NDP

[Ano96m] Anonymous. In the news: David Packard dies. IEEE Software, 13(3):106, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:NB

[Ano96n] Anonymous. In the news: In brief. IEEE Software, 13(3):106, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:IW

[Ano96o]  Anonymous.  Industry

watch. IEEE Software, 13(4):138–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:JDI

[Ano96p] Anonymous. “just do it” approach is counterproductive. IEEE Software, 13(4): 6–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:LEP

[Ano96q]  Anonymous.  Letters:

“ERAS” and the “positive approach”. IEEE Software, 13(4):6–??, July 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:LMS

[Ano96r] Anonymous. Letters: More self-help for practitioners. European perspective on self-help. in the trenches with the “bright young things”. A dose of realism in developer education. IEEE Software, 13(5):6–14, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:PLP

[Ano96s] Anonymous. Programming languages for parallel processing. IEEE Software, 13(4):133–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:TA

[Ano96t] Anonymous. Text algorithms. IEEE Software, 13(4):133–??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:TD

[Ano96u] Anonymous. Top drawer.

IEEE Software, 13(3):117, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:WDS

[Ano96v] Anonymous. Why does software cost so much? IEEE Software, 13(4):133– ??, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1997:AAI

[Ano97a] Anonymous. 1997 author and article index. IEEE Software, 14(6):114–123, November/December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://computer.org/software/
                                        so1997/pdf/s6114.pdf.

Anonymous:1997:B

[Ano97b] Anonymous. Bookshelf. IEEE Software, 14(1): 86–??, January/February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1997:C

[Ano97c]  Anonymous.  Comdex.

IEEE Software, 14(1): 88–??, January/February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1997:NDA

[Ano97d] Anonymous. In the news: DTV agreement good for software developers. IEEE Software, 14(1):88–??, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1997:NBI

[Ano97e]  Anonymous.  In the

news: In brief: IEEE on Y2K; SETI potatoes?; Texans get soaked; direct attack; expanding NATO links. IEEE Software, 14(5):114, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5106.
pdf.

Anonymous:1997:IWD

[Ano97f]  Anonymous.  Interview:

William Davidow. IEEE Software, 14(1):88–??, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1997:LSY

[Ano97g] Anonymous. Letters: Share your results and your methods; do tool criteria ties testers’ hands? balancing experience and speculation; jargon, technospeak blur good ideas; quality, not quantity, key to measuring bugs. IEEE Software, 14(6):10, 12, 14, 16–17, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s6010.
                                        pdf.

Anonymous:1997:NMW

[Ano97h] Anonymous. Navigating the maintenance waters at ICSM ’96. IEEE Software, 14(1):88–??, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1997:TDJ

[Ano97i] Anonymous. Top drawer: Java client generator. IEEE Software, 14(2):125, March/ April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1998:AI

[Ano98a]  Anonymous.  1998 an-

nual index. IEEE Software, 15(6):114–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s6114.
                                        pdf.

Anonymous:1998:ASa

[Ano98b] Anonymous. Article summaries. IEEE Software, 15(2):4, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2004.
pdf.

Anonymous:1998:ASb

[Ano98c] Anonymous. Article summaries. IEEE Software, 15 (3):4, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3004.
pdf.

Anonymous:1998:ASc

[Ano98d] Anonymous. Article summaries. IEEE Software, 15 (4):4, July/August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1998/pdf/
s4004.pdf.

Anonymous:1998:ASd

[Ano98e] Anonymous. Article summaries. IEEE Software, 15(5):4–5, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5004.
pdf.

Anonymous:1998:ASe

[Ano98f] Anonymous. Article summaries. IEEE Software, 15(6):4–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6004.
pdf.

Anonymous:1998:BJN

[Ano98g] Anonymous. Bookshelf: Java network security: Measuring software reuse: Principles, practices, and economic models: India’s software industry: State policy, liberalisation, and industrial development. IEEE Software, 15(6):124–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6124.
pdf.

Anonymous:1998:BSQ

[Ano98h] Anonymous. Bookshelf: SCM: The quick and the cheap; getting a jumpstart in Java network programming; an exceptional look at fault injection. IEEE Software, 15 (5):115–118, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so1998/pdf/s5115.
pdf.

Anonymous:1998:CLI

[Ano98i] Anonymous. Call: Life in an Internet world: Software security. IEEE Software, 15(6):36–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6036.
pdf.

Anonymous:1998:CPDb

[Ano98j] Anonymous. Call: Process diversity. IEEE Software, 15(6):17–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6017.
pdf.

Anonymous:1998:CSF

[Ano98k] Anonymous. Critical success factors. IEEE Software, 15(5):69, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5069.
pdf.

Anonymous:1998:CPDa

[Ano98l] Anonymous. Cross-pollinating disciplines. IEEE Software, 15(5):22, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5022.
pdf.

Anonymous:1998:GCA

[Ano98m] Anonymous. General call for articles. IEEE Software, 15(6):21–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6021.
pdf.

Anonymous:1998:NLC

[Ano98n] Anonymous. In the news: a license to code; structuring solutions: IEEE offers Y2K guidance; Y2K update; calls for articles and reviewers. IEEE Software, 15(5):119–124, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5119.
pdf.

Anonymous:1998:NEF

[Ano98o] Anonymous. In the news; envisioning the future at SD and ’98 West; Macworld: Back to business. IEEE Software, 15(2):116–118, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2116.
pdf.

Anonymous:1998:NEY

[Ano98p] Anonymous. In the news:

The Euro: Y2K fire drill or is opportunity?: Opening up computing’s next century: Y2K update. IEEE Software, 15(6):107–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6107.
pdf.

Anonymous:1998:IIE

[Ano98q] Anonymous. Interview: Interview with Ellen Ullman: Of machines, methods, and madness. IEEE Software, 15(3):42–45, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3042.
pdf.

Anonymous:1998:La

[Ano98r]  Anonymous.  Letters.

IEEE Software, 15(5):6–9, September/October 1998.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s5006.pdf.

Anonymous:1998:Lb

[Ano98s]  Anonymous.  Letters.

IEEE Software, 15(6):6–??, November/December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s6006.pdf.

Anonymous:1998:LIW

[Ano98t] Anonymous. Life in an Internet world. IEEE Software, 15(5):110, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s5110.
                                        pdf.

Anonymous:1998:LOS

[Ano98u] Anonymous. The Linux operating system. IEEE Software, 15(5):17, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s5017.
                                        pdf.

Anonymous:1998:MT

[Ano98v] Anonymous. More topics. … IEEE Software, 15(5):125–127, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5125.
pdf.

Anonymous:1998:PD

[Ano98w] Anonymous. Process diversity. IEEE Software, 15(5):110–109, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5110.
pdf.

Anonymous:1998:PP

[Ano98x]  Anonymous.  Products

for practitioners. IEEE Software, 15(2):124–127, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2124.
pdf.

Anonymous:1998:PPFa

[Ano98y]  Anonymous.  Products

for practitioners: Focus on components. IEEE Software, 15(5):111–114, September/October 1998.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1998/pdf/
s5111.pdf.

Anonymous:1998:PPFb

[Ano98z] Anonymous. Products for practitioners: Focus on Microsoft. IEEE Software, 15(6):111–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6111.
pdf.

Anonymous:1998:SC

[Ano98-27] Anonymous. Software certification. IEEE Software, 15(5):53, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5053.
pdf.

Anonymous:1999:ASa

[Ano99a] Anonymous. Article summaries. IEEE Software, 16(1):4–??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1004.
pdf.

Anonymous:1999:ASb

[Ano99b] Anonymous. Article summaries. IEEE Software, 16 (4):4, July/August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1999/pdf/
s4004.pdf.

Anonymous:1999:ASc

[Ano99c] Anonymous. Article summaries. IEEE Software, 16(6):4, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6004.
pdf.

Anonymous:1999:BST

[Ano99d] Anonymous. Bookshelf:

Surviving the top ten challenges of software development; the Year 2000 crisis; the continuing challenge; software project survival guide; objectoriented multithreading using C++. IEEE Software, 16(1):114–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1114.
pdf.

Anonymous:1999:CARb

[Ano99e] Anonymous. Call for articles & reviewers: Malicious information technology: The Software vs. the People. IEEE Software, 16(6):89, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s6089.
                                        pdf.

Anonymous:1999:CPS

[Ano99f] Anonymous. Call for participation:: Standard dictionary of measures of the software aspects of dependability. IEEE Software, 16(1):57–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s1057.
                                        pdf.

Anonymous:1999:COD

[Ano99g] Anonymous. Call: Organizational design. IEEE Software, 16(1):64–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s1064.
                                        pdf.

Anonymous:1999:CARa

[Ano99h] Anonymous. Calls for articles and reviewers. IEEE Software, 16(1):??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FL

[Ano99i] Anonymous. Focus: Linux.

IEEE Software, 16(1):??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLE

[Ano99j] Anonymous. Focus: Linux — essay. IEEE Software, 16(1):??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLT

[Ano99k] Anonymous. Focus: Linux — from the trenches. IEEE Software, 16(1):??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLN

[Ano99l] Anonymous. Focus: Linux — nuts & bolts. IEEE Software, 16(1):??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:T

[Ano99m] Anonymous. From the trenches. IEEE Software, 16(1):??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:HH

[Ano99n]  Anonymous.  Head to

head. IEEE Software, 16 (1):??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:ISA

[Ano99o] Anonymous. IEEE Software 1999 annual index. IEEE Software, 16 (6):115–118, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6115.
pdf.

Anonymous:1999:NSL

[Ano99p] Anonymous. In the news: Shedding light on embedded systems; Netscape starts over. IEEE Software, 16(1):122–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1122.
pdf.

Anonymous:1999:L

[Ano99q] Anonymous. Letters. IEEE Software, 16(1):9–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1009.
pdf.

Anonymous:1999:NB

[Ano99r] Anonymous. Nuts & bolts.

IEEE Software, 16(1):??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2000:AI

[Ano00a] Anonymous. Annual index. IEEE Software, 17 (6):130–134, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6130.
pdf.

Anonymous:2000:ASa

[Ano00b] Anonymous. Article summaries. IEEE Software, 17(1):4–5, January/ February 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s1004.
pdf.

Anonymous:2000:ASb

[Ano00c] Anonymous. Article summaries. IEEE Software, 17(2):4, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s2004.
                                        pdf.

Anonymous:2000:ASc

[Ano00d] Anonymous. Article summaries. IEEE Software, 17 (3):4, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s3004.
                                        pdf.

Anonymous:2000:ASd

[Ano00e] Anonymous. Article summaries. IEEE Software, 17 (4):4, July/August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s4004.pdf.

Anonymous:2000:ASe

[Ano00f] Anonymous. Article summaries. IEEE Software, 17(5):6, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5006.
pdf.

Anonymous:2000:ASf

[Ano00g] Anonymous. Article summaries. IEEE Software, 17(6):4, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6004.
pdf.

Anonymous:2000:CPRb

[Ano00h] Anonymous. Call for papers & reviewers: Global software development. IEEE Software, 17(3):6, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3006.
pdf.

Anonymous:2000:CPG

[Ano00i] Anonymous. Call for papers: Global software development. IEEE Software, 17(2):69, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2069.
pdf.

Anonymous:2000:CPRa

[Ano00j]  Anonymous.  Call for

papers: Recent developments in software estimation. IEEE Software, 17(2): 89, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s2089.
                                        pdf.

Anonymous:2000:CPA

[Ano00k]  Anonymous.  Calls for

papers: Applying usability techniques during development. IEEE Software, 17(2):10, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s2010.
                                        pdf.

Anonymous:2000:CA

[Ano00l] Anonymous. Classified ads. IEEE Software, 17(1):61, January/February 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s1061.pdf.

Anonymous:2000:NTS

[Ano00m] Anonymous. In the news: The team software process: Quiet quality revolution. IEEE Software, 17(6):125–129, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6125.
pdf.

Anonymous:2001:AIV

[Ano01a] Anonymous. Annual index:

Volume 18, 2001. IEEE Software, 18(6):106–110, November/December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s6106abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6106.pdf.

Anonymous:2001:ASa

[Ano01b] Anonymous. Article summaries. IEEE Software, 18(1):4, January/ February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1004.
pdf.

Anonymous:2001:ASb

[Ano01c] Anonymous. Article summaries. IEEE Software, 18 (3):4, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so2001/pdf/s3004.
                                        pdf.

Anonymous:2001:ASc

[Ano01d] Anonymous. Article summaries. IEEE Software, 18(5):4, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s5004abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5004.pdf.

Anonymous:2001:ASd

[Ano01e] Anonymous. Article summaries. IEEE Software, 18(6):4, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s6004abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s6004.pdf.

Anonymous:2001:BTC

[Ano01f] Anonymous. Bookshelf: Testing Computer Software, second edition: CORBA 3 Fundamentals and Programming, second edition. IEEE Software, 18(3):104– 106, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/

books/so2001/pdf/s3104.
pdf.

Anonymous:2001:CAB

[Ano01g] Anonymous. Call for articles: Benchmarking software organizations. IEEE Software, 18(1):37, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s1038abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s1038.pdf.

Anonymous:2001:CPS

[Ano01h] Anonymous. Call for papers: Software security: Building systems securely from the ground up. IEEE Software, 18(2):95, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2095.pdf.

Anonymous:2001:IBI

[Ano01i] Anonymous. Introducing the boards: IEEE Software Editorial Board and Industry Advisory Board members. IEEE Software, 18(1):9–12, January/ February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so2001/pdf/s1009.
                                        pdf.

Anonymous:2001:MAS

[Ano01j]  Anonymous.  Miscella-

neous: Article summaries. IEEE Software, 18(4):4, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s4004.
                                        pdf.

Anonymous:2002:ISA

[Ano02a] Anonymous. 2002 IEEE Software annual index, volume 19. IEEE Software, 19(6):105–110, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

                                        tronic). URL http://www.
                                        computer.org/software/
                                        so2002/s6105.pdf.

Anonymous:2002:ASa

[Ano02b] Anonymous. Article summaries. IEEE Software, 19(1):2, January/ February 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2002/s1002abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2002/pdf/
                                        s1002.pdf.

Anonymous:2002:ASb

[Ano02c] Anonymous. Article summaries. IEEE Software, 19(2):4, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2004.
pdf;   http://www.computer.
org/software/so2002/s2004abs.
htm.

Anonymous:2002:ASc

[Ano02d] Anonymous. Article summaries. IEEE Software, 19 (3):4, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3004.
pdf.

Anonymous:2002:ASd

[Ano02e] Anonymous. Article summaries. IEEE Software, 19 (4):4, July/August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2002/pdf/
s4004.pdf.

Anonymous:2002:ASe

[Ano02f] Anonymous. Article summaries. IEEE Software, 19(5):4, September/ October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5004.
pdf.

Anonymous:2002:ASf

[Ano02g] Anonymous. Article summaries. IEEE Software, 19(6):4, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6004.
pdf.

Anonymous:2002:B

[Ano02h] Anonymous. Bookshelf. IEEE Software, 19(2):104– 108, March/April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2002/pdf/
s2104.pdf;   http://www.
computer.org/software/
so2002/s2104abs.htm.

Anonymous:2002:CAS

[Ano02i]  Anonymous.  Calls for

articles: Software construction and industry experience reports. IEEE Software, 19(3):96, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so2002/pdf/s3096.
pdf.

Anonymous:2002:CCC

[Ano02j] Anonymous. Correction:

Correctness by Construction: Developing a Commercial Secure System. IEEE Software, 19(2):10, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See [HC02].

Anonymous:2003:ISA

[Ano03a] Anonymous. 2003 IEEE Software annual index volume 20. IEEE Software, 20(6):102–107, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6102.
pdf.

Anonymous:2003:ASa

[Ano03b] Anonymous. Article summaries. IEEE Software, 20(1):4, January/ February 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1004.
pdf.

Anonymous:2003:ASb

[Ano03c] Anonymous. Article summaries. IEEE Software, 20(2):4, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2004.
pdf.

Anonymous:2003:ASc

[Ano03d] Anonymous. Article summaries. IEEE Software, 20 (3):4, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3004.
pdf.

Anonymous:2003:ASd

[Ano03e] Anonymous. Article summaries. IEEE Software, 20 (4):4, July/August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://www.computer.org/
software/so2003/s4004.
pdf.

Anonymous:2003:ASe

[Ano03f] Anonymous. Article summaries. IEEE Software, 20(5):4, September/ October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5004.
pdf.

Anonymous:2003:ASf

[Ano03g] Anonymous. Article summaries. IEEE Software, 20(6):4, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6004.
pdf.

Anonymous:2003:CADa

[Ano03h] Anonymous. Call for articles: Developing with open source software. IEEE Software, 20(3):1, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3001.
pdf.

Anonymous:2003:CADb

[Ano03i] Anonymous. Call for articles: Developing with open source software. IEEE Software, 20(4):1, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

tronic). URL http://www.
computer.org/software/
so2003/s4001.pdf.

Anonymous:2003:CAR

[Ano03j]  Anonymous.  Call for

articles: Return on investment in the software industry. IEEE Software, 20(5):1, September/ October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5001.
pdf.

Anonymous:2003:CPS

[Ano03k] Anonymous. Call for papers: The state of the practice of software engineering. IEEE Software, 20(2):14, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

tronic). URL http://www.
computer.org/software/
so2003/s2014.pdf.

Anonymous:2003:C

[Ano03l] Anonymous. Correction. IEEE Software, 20(6):11, November/December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/dl/mags/so/2003/06/

s6008.pdf. See [RE03].

Anonymous:2003:SOR

[Ano03m] Anonymous. Saluting our 2002 referees. IEEE Software, 20(1):100–102, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1100.
pdf.

Anonymous:2004:R

[Ano04a]  Anonymous.  2003 re-

viewers. IEEE Software, 21(1):107–110, January/ February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
01/s1107.pdf.

Anonymous:2004:AI

[Ano04b] Anonymous. Annual index. IEEE Software, 21(6):95–99, November/ December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
06/s6095.pdf.

Anonymous:2004:ASa

[Ano04c] Anonymous. Article summaries. IEEE Software, 21(1):4, January/ February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
01/s1004.pdf.

Anonymous:2004:ASb

[Ano04d] Anonymous. Article summaries. IEEE Software, 21 (2):4, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        02/s2004.pdf.

Anonymous:2004:ASc

[Ano04e] Anonymous. Article summaries. IEEE Software, 21(3):4, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        03/s3004.pdf.

Anonymous:2004:ASd

[Ano04f] Anonymous. Article summaries. IEEE Software, 21 (4):4, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        04/s4004.pdf.

Anonymous:2004:ASe

[Ano04g] Anonymous. Article summaries. IEEE Software, 21(5):4, September/October 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        05/s5004.pdf.

Anonymous:2004:ASf

[Ano04h] Anonymous. Article summaries. IEEE Software, 21(6):4, November/ December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
06/s6004.pdf.

Anonymous:2004:La

[Ano04i] Anonymous. Letters. IEEE Software, 21(1):5–7, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
01/s1005.pdf.

Anonymous:2004:Lb

[Ano04j]  Anonymous.  Letters.

IEEE Software, 21(2):8–12, March/April 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
02/s2008.pdf.

Anonymous:2004:Lc

[Ano04k]  Anonymous.  Letters.

IEEE Software, 21(3):8– 10, May/June 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
03/s3008.pdf.

Anonymous:2004:Ld

[Ano04l]  Anonymous.  Letters.

IEEE Software, 21(4):8–10, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        04/s4008.pdf.

Anonymous:2004:Le

[Ano04m]  Anonymous.  Letters.

IEEE Software, 21(5):8–16, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        05/s5008.pdf.

Anonymous:2004:Lf

[Ano04n]  Anonymous.  Letters.

IEEE Software, 21(6):9–11, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        06/s6009.pdf.

Anonymous:2004:SEG

[Ano04o] Anonymous. Software engineering glossary: Software design, Part I. IEEE Software, 21(5):110, September/October 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
05/s5110.pdf.

Anonymous:2004:UCS

[Ano04p] Anonymous. User confidence— and the software developer. IEEE Software, 21(6):5–8, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
06/s6005.pdf.

Anonymous:2005:ISA

[Ano05a] Anonymous. 2005 IEEE Software annual index, volume 22. IEEE Software, 22(6):106–110, November/ December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
06/s6106.pdf.

Anonymous:2005:ASa

[Ano05b] Anonymous. Article summaries. IEEE Software, 22(1):4, January/ February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
01/s1004.pdf.

Anonymous:2005:ASb

[Ano05c] Anonymous. Article summaries. IEEE Software, 22 (2):4, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2005/
                                        02/s2004.pdf.

Anonymous:2005:ASc

[Ano05d] Anonymous. Article summaries. IEEE Software, 22(3):4, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2005/
                                        03/s3004.pdf.

Anonymous:2005:ASd

[Ano05e] Anonymous. Article summaries. IEEE Software, 22 (4):4, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2005/
                                        04/s4004.pdf.

Anonymous:2005:ASe

[Ano05f] Anonymous. Article summaries. IEEE Software, 22(5):4, September/October 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.

org/comp/mags/so/2005/
05/s5004.pdf.

Anonymous:2005:ASf

[Ano05g] Anonymous. Article summaries. IEEE Software, 22(6):2, November/ December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
06/s6002.pdf.

Anonymous:2005:La

[Ano05h] Anonymous. Letters. IEEE Software, 22(1):8–9, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
01/s1008.pdf.

Anonymous:2005:Lb

[Ano05i]  Anonymous.  Letters.

IEEE Software, 22(2):8–10, March/April 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
02/s2008.pdf.

Anonymous:2005:Lc

[Ano05j]  Anonymous.  Letters.

IEEE Software, 22(5):8–12, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
05/s5008.pdf.

Anonymous:2005:LCV

[Ano05k]  Anonymous.  Letters:

Convenience vs. security. IEEE Software, 22(4):8, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
04/s4008.pdf.

Anonymous:2005:OS

[Ano05l] Anonymous. Outwitting spammers. IEEE Software, 22(3):8–9, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
03/s3008.pdf.

Anonymous:2005:SM

[Ano05m]  Anonymous.  Software

maintenance. IEEE Software, 22(4):103, July/ August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2005/04/s4103.
pdf.

Anonymous:2005:SP

[Ano05n]  Anonymous.  State of

the practice. IEEE Software, 22(6):6–8, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
06/s6006.pdf.

Anonymous:2005:USD

[Ano05o] Anonymous. Understanding story-driven development processes. IEEE Software, 22(6):103–105, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
06/s6103.pdf.

Anonymous:2006:AIV

[Ano06a] Anonymous. Annual index, volume 23, 2006. IEEE Software, 23(6):106–110, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
06/s6106.pdf.

Anonymous:2006:ASa

[Ano06b] Anonymous. Article summaries. IEEE Software, 23(1):4, January/ February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        01/s1004.pdf.

Anonymous:2006:ASb

[Ano06c] Anonymous. Article summaries. IEEE Software, 23 (2):4, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        02/s2004.pdf.

Anonymous:2006:ASc

[Ano06d] Anonymous. Article summaries. IEEE Software, 23(3):4, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        03/s3004.pdf.

Anonymous:2006:ASd

[Ano06e] Anonymous. Article summaries. IEEE Software, 23 (4):4, July/August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        04/s4004.pdf.

Anonymous:2006:ASe

[Ano06f] Anonymous. Article summaries. IEEE Software, 23(5):4, September/October 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
05/s5004.pdf.

Anonymous:2006:ASf

[Ano06g] Anonymous. Article summaries. IEEE Software, 23(6):4, November/ December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
06/s6004.pdf.

Anonymous:2006:Ba

[Ano06h] Anonymous. Bookshelf.

IEEE Software, 23(1):116– 118, January/February 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
01/s1116.pdf.

Anonymous:2006:Bb

[Ano06i] Anonymous. Bookshelf.

IEEE Software, 23(2):109– 111, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
02/s2109.pdf.

Anonymous:2006:Bc

[Ano06j] Anonymous. Bookshelf. IEEE Software, 23(3):100– 102, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        03/s3100.pdf.

Anonymous:2006:BLS

[Ano06k] Anonymous. Business lessons for software developers. IEEE Software, 23 (2):8, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        02/s2008.pdf.

Anonymous:2006:FV

[Ano06l] Anonymous. First virus? IEEE Software, 23(4):8–10, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        04/s4008.pdf.

Anonymous:2006:La

[Ano06m] Anonymous. Letters. IEEE Software, 23(1):10–12, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2006:Lb

[Ano06n]  Anonymous.  Letters.

IEEE Software, 23(3):8, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
03/s3008.pdf.

Anonymous:2006:Lc

[Ano06o]  Anonymous.  Letters.

IEEE Software, 23(6):8–11, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
06/s6008.pdf.

Anonymous:2006:LCU

[Ano06p]  Anonymous.  Letters:

The cone of uncertainty. IEEE Software, 23(5):8–10, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
05/s5008.pdf.

Anonymous:2006:OR

[Ano06q] Anonymous. Our 2005 reviewers. IEEE Software, 23(2):116–118, March/ April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/so/2006/02/s2116.
                                        pdf.

Anonymous:2007:ASa

[Ano07a] Anonymous. Article summaries. IEEE Software, 24(1):4, January/ February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2007/
                                        01/s1004.pdf.

Anonymous:2007:ASb

[Ano07b] Anonymous. Article summaries. IEEE Software, 24 (2):4, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2007/
                                        02/s2004.pdf.

Anonymous:2007:ASc

[Ano07c] Anonymous. Article summaries. IEEE Software, 24(3):4, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2007/
                                        03/s3004.pdf.

Anonymous:2007:ASd

[Ano07d] Anonymous. Article summaries. IEEE Software, 24 (4):4, July/August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
04/s4004.pdf.

Anonymous:2007:La

[Ano07e]  Anonymous.  Letters.

IEEE Software, 24(2):8, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
02/s2008.pdf.

Anonymous:2007:Lb

[Ano07f]  Anonymous.  Letters.

IEEE Software, 24(3):8– 9, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
03/s3008.pdf.

Anonymous:2007:Lc

[Ano07g]  Anonymous.  Letters.

IEEE Software, 24(4):8–11, July/August 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
04/s4008.pdf.

Anonymous:2007:Ld

[Ano07h]  Anonymous.  Letters.

IEEE Software, 24(6):6–8, November/December 2007.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
06/mso2007060006.pdf.

Anonymous:2007:LSE

[Ano07i] Anonymous. Letters: Software engineering curricula shortfalls. IEEE Software, 24(1):8, January/ February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
01/s1008.pdf.

Anonymous:2007:RA

[Ano07j] Anonymous. The relevance of architecture. IEEE Software, 24(5):8–9, September/October 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
05/s5008.pdf.

Anonymous:2007:RTO

[Ano07k]  Anonymous.  Reviewer

thanks: Our 2006 reviewers. IEEE Software, 24(2):100–102, March/ April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/

mags/so/2007/02/s2100.
pdf.

Anonymous:2008:AI

[Ano08a] Anonymous. Annual index. IEEE Software, 25 (6):0, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2008:DTL

[Ano08b]  Anonymous.  Dynam-

ically typed languages. IEEE Software, 25(2):7–10, March/April 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2008/
02/mso2008020007.pdf.

Anonymous:2008:LPC

[Ano08c] Anonymous. Letters: The poetry of code. IEEE Software, 25(6):107, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2008:RDD

[Ano08d]  Anonymous.  Requiring

design, designing requirements. IEEE Software, 25 (3):8–9, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2008/
03/mso2008030008.pdf.

Anonymous:2008:TLB

[Ano08e] Anonymous. Time line: The big bang: 25 years of software history. IEEE Software, 25(6):6–14, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2009:AIS

[Ano09a]  Anonymous.  25th an-

niversary: IEEE software’s 25th-anniversary top picks. IEEE Software, 26(1):9–11, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2009:ACA

[Ano09b] Anonymous. An agile cure for all ills? IEEE Software, 26(6):8, November/ December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2009:AI

[Ano09c] Anonymous. Annual index. IEEE Software, 26(6), November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2009:LMM

[Ano09d] Anonymous. Letters: My must-reads. IEEE Software, 26(1):8, January/ February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2009:RSE

[Ano09e]  Anonymous.  Responses

to “Software Engineering: An Idea Whose Time Has Come and Gone?”. IEEE Software, 26(5):5, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ambras:1988:MKB

[AO88] James Ambras and Vicky O’Day. MicroScope: a knowledge-based programming environment. IEEE Software, 5(3):50–58, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Akima:1989:ISD

[AO89] Noboru Akima and Fusatake Ooi. Industrializing software development: a Japanese approach. IEEE Software, 6(2):13–21, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Aoyama:1993:CDP

[Aoy93] Mikio Aoyama. Concurrentdevelopment process model. IEEE Software, 10(4):46– ??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Aoyama:1998:WBA

[Aoy98]  Mikio Aoyama.  Web-

based agile software development. IEEE Software, 15(6):56–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s6056abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s6056.pdf.

Arranga:2000:GEI

[AP00] Edmund C. Arranga and Wilson Price. Guest Editors’ introduction: Fresh from Y2K, what’s next for Cobol? IEEE Software, 17(2):16–20, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2016.
pdf.

Applewhite:2002:JFW

[App02] Ashton Applewhite. Not just the facts: What “requirements” mean to a nonfiction writer. IEEE Software, 19(1):87–89, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1087abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1087.pdf.

Appari:2006:B

[App06] Anthony Akins Ajit Appari. Bookshelf. IEEE Software, 23(5):99–101, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
05/s5099.pdf.

Alexander:2004:UPS

[AR04] Ian Alexander and Suzanne Robertson. Understanding project sociology by modeling stakeholders. IEEE Software, 21(1):23–27, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Arango:1995:OUT

[Ara95] Guillermo Arango. Once upon a time, there was a demo. IEEE Software, 12 (2):8, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anderson:2003:TSF

[ARTZ03] Paul Anderson, Thomas Reps, Tim Teitelbaum, and Mark Zarins. Tool support for fine-grained software inspection. IEEE Software, 20(4):42–50, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4042.
pdf;   http://www.computer.
org/software/so2003/s4042abs.
htm.

Avritzer:1996:RTR

[ARW96] Alberto Avritzer, Johannes P. Ros, and Elaine J. Weyuker. Reliability testing of rulebased systems. IEEE Software, 13(5):76–82, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Applewhite:2003:NSG

[AS03] Ashton Applewhite and Dale Strok. In the news: Should governments go open source?; Linux gets some respect; face to face with the experts. IEEE Software, 20(4):88– 93, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4088.
pdf.

Azar:2007:VOR

[ASC07] Jim Azar, Randy K. Smith, and David Cordes. Valueoriented requirements prioritization in a small development organization. IEEE Software, 24(1):32–37, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Appelbe:1989:SPP

[ASM89] Bill Appelbe, Kevin Smith, and Charlie McDowell. Start/Pat: a parallelprogramming toolkit. IEEE Software, 6(4):29–38, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Athey:1998:LCF

[Ath98] Tim Athey. Leadership challenges for the future. IEEE Software, 15(3):72– 77, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3072abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3072.pdf.

Austin:1999:MPM

[Aus99] Robert D. Austin. Manager: The phantom menace. IEEE Software, 16(5):15–18, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so1999/pdf/s5015.
                                        pdf.

Almering:2007:USR

[AvGCS07] Vincent Almering, Michiel van Genuchten, Ger Cloudt, and Peter J. M. Sonnemans. Using software reliability growth models in practice. IEEE Software, 24(6):82–88, November/ December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Alavi:1991:MPD

[AW91] Maryam Alavi and James C. Wetherbe. Mixing prototyping and data modeling for information-system design. IEEE Software, 8(3): 86–91, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anton:2003:PCS

[AW03] Annie I. Antón and Don Wells. Point/counterpoint: Successful software projects need requirements planning; don’t solve a problem before you get to it. IEEE Software, 20(3):44– 47, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3044.
                                        pdf.

Atlee:2006:RES

[AW06] Joanne Atlee and Roel Wieringa. RE 05: Engineering successful products. IEEE Software, 23 (3):16–18, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Burton:1987:RSL

[BAB^{+}87] Bruce A. Burton, Rhonda Wienk Aragon, Stephen A. Bailey, Kenneth D. Koehler, and Lauren A. Mayes. Reusable software library. IEEE Software, 4(4):25–33, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bach:1995:EAP

[Bac95a] James Bach. Enough about process: What we need are heroes. IEEE Software, 12(2):96–98, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bach:1995:BTL

[Bac95b] Richard Bach. Building tools and libraries in MS-DOS. IEEE Software, 12 (2):115–116, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Baczynski:1995:LSH

[Bac95c] George M. Baczynski. Letters: Software not held to same standards. IEEE Software, 12(3):7, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bach:1997:PCS

[Bac97]  James Bach.  Point/

counterpoint: SE education: We’re on our own. IEEE Software, 14(6):26, 28, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6026.
pdf;   http://dlib.computer.
org/so/books/so1997/pdf/
s6026abs.htm.

Bagchi:1999:ISI

[Bag99]  Subroto Bagchi.  In-

dia’s software industry: The people dimension. IEEE Software, 16(3):62– 65, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s3062abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s3062.pdf.

Baines:1998:ADR

[Bai98]  Robin Baines.  Across

disciplines: Risk, design, method, process, and tools. IEEE Software, 15(4):61– 64, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
cg1998/s4061abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s4061.pdf.

Baker:1995:LCP

[Bak95]  Trace Baker.  Letters:

Cold pizza and pathological heroes. IEEE Software, 12(3):5–6, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Baker:2001:MWW

[Bak01] Emanuel R. Baker. Manager: Which way, SQA? IEEE Software, 18(1): 16–18, January/February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2001/pdf/
s1016.pdf.

Balzer:1987:LNG

[Bal87] Robert M. Balzer. Living in the next-generation operating system. IEEE Software, 4(6):77–85, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). Also published in/as: ‘Proceedings of the tri-annual IFIP Conf’, H.Kugler(ed), N-H (Amsterdam), 1986.

Baldwin:1989:CPC

[Bal89] Doug Baldwin. Consul: a parallel constraint language. IEEE Software, 6 (4):62–69, July 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Baltes:1999:RE

[Bal99]  Cheryl Baltes.  Recog-

nizing excellence. IEEE Software, 16(4):113, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s4109.
                                        pdf.

Baltes:2000:NUU

[Bal00] Cheryl Baltes. In the news: Update: UCITA: Awards: Video game development: Application and component development. IEEE Software, 17(3):96– 98, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s3096.
                                        pdf.

Bollinger:2003:BFF

[BALV03] Terry Bollinger, John J. Abbott, Robert C. Larrabee, and Carl Vieregger. Bookshelf: Facts and fantasies: a review of two books; understanding the unified process; software architecture: Project planning and control; paring Java patterns. IEEE Software, 20(2):80– 85, March/April 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2080.

pdf. Review of A New Kind of Science by Stephen Wolfram, Prey: A Novel by Michael Crichton, The Unified Process Explained by Kendall Scott, Effective Work Breakdown Structures by Gregory T. Haugan, Applied Java Patterns by Stephen Stelting and Olav Maassen.

Banisar:1993:BCE

[Ban93] David Banisar. Battle for control of encryption technology. IEEE Software, 10 (4):95–97, July 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barker:1995:YPD

[Bar95a] Michael Barker. Your programs don’t need a passport. IEEE Software, 12 (2):116–117, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barlas:1995:FUA

[Bar95b] Stephen Barlas. Future uncertain for ATP grant recipients. IEEE Software, 12(6):109, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barlas:1995:B

[Bar95c] Stephen Barlas. In brief. IEEE Software, 13(2):112, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barlas:1995:NNC

[Bar95d] Stephen Barlas. In the news: New congress threatens software reuse program. IEEE Software, 12 (1):108–114, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barlas:1995:NPC

[Bar95e] Stephen Barlas. In the news: Patent, copyright issues focus of legislation. IEEE Software, 12(6):107– 108, November 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barlas:1996:NFS

[Bar96a] Stephen Barlas. In the news: FAA shifts focus to scaled-back DSR. IEEE Software, 13(2):110, 114, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barlas:1996:NKD

[Bar96b] Stephen Barlas. In the news: Key decisions likely on encryption exports. IEEE Software, 13(6):102– 108, November 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barias:1997:NBP

[Bar97a] Stephen Barias. In the news: BSA pushes for override of encryption guidelines. IEEE Software, 14 (2):136, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barias:1997:NPA

[Bar97b] Stephen Barias. In the news: Proposed amendment may keep patent applications under wraps. IEEE Software, 14(4):104, July/August 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4100.
pdf.

Barlas:1997:NID

[Bar97c] Stephen Barlas. In the news: Interview with Dorothy Denning. IEEE Software, 14(5):108–109, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1997/pdf/
                                        s5106.pdf.

Barnes:1998:QTC

[Bar98] Bruce H. Barnes. Quality time: Computer security research: a British perspective. IEEE Software, 15(5):30–33, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s5030.
                                        pdf.

Baskette:1987:LCA

[Bas87a] Jerry Harvey Baskette. Life cycle analysis of an Ada project. IEEE Software, 4(1):40–47, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bass:1987:RHD

[Bas87b] Leonard J. Bass. Role of a historian in a database management system. IEEE Software, 4(2):35–40, March 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bassett:1987:FBS

[Bas87c] Paul G. Bassett. Framebased software engineering. IEEE Software, 4(4): 9–16, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Browne:1989:CUA

[BAS89] J. C. Browne, Muhammad Azam, and Stephen Sobek. CODE: a unified approach to parallel programming. IEEE Software, 6(4): 10–18, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Basili:1990:VMR

[Bas90] Victor R. Basili. Viewing maintenance as reuseoriented software development. IEEE Software, 7 (1):19–25, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bassett:1995:LFN

[Bas95] Paul G. Bassett. Letters:

Frameworks a new development? hardly. IEEE Software, 12(4):10, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bassett:2007:CFB

[Bas07] Paul G. Bassett. The case for frame-based software engineering. IEEE Software, 24(4):90–99, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bate:1998:SDS

[Bat98] Roger R. Bate. Soapbox: Do systems engineering? who, me? IEEE Software, 15(4):65–66, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4065.
pdf.

Baumert:1994:QTP

[Bau94] John H. Baumert. Quality time: Process assessment with a project focus. IEEE Software, 11(2):89– 91, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Beetem:1989:IGL

[BB89] Anne F. Beetem and John F. Beetem. Introduction to the Galaxy language. IEEE Software, 6 (3):55–62, May 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barnes:1991:MRC

[BB91] Bruce H. Barnes and Terry B. Bollinger. Making reuse cost-effective. IEEE Software, 8(1):13–24, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Becker:1999:ASP

[BB99a] Shirley A. Becker and Mitchell L. Bostelman. Aligning strategic and project measurement systems. IEEE Software, 16 (3):46–51, May/June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so1999/s3046abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s3046.pdf.

Bollinger:1999:GEI

[BB99b] Terry Bollinger and Peter H. Beckman. Guest Editors’ introduction: Linux on the move. IEEE Software, 16(1):30–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1030.
pdf.

Boehm:2008:BOR

[BB08] Barry Boehm and Jesal Bhuta. Balancing opportunities and risks in component-based software development. IEEE Software, 25(6):56–63, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Basili:2004:NYR

[BBD^{+}04] Victor R. Basili, Barry Boehm, Al Davis, Watts S. Humphrey, Nancy Leveson, Nancy R. Mead, John D. Musa, David Lorge Parnas, Shari Lawrence Pfleeger, and Elaine Weyuker. New Year’s resolutions for software quality. IEEE Software, 21(1):12–13, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ballarin:1988:MCE

[BBE^{+}88] Enrico Ballarin, Helmar Burkhart, Rudolf Eigenmann, Heinz Kindlimann, and Michael Moser. Making a compiler easily portable. IEEE Software, 5(3):30– 38, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Birukou:2007:IWS

[BBGK07] Aliaksandr Birukou, Enrico Blanzieri, Paolo Giorgini, and Natallia Kokash. Improving Web service discovery with usage data. IEEE Software, 24(6):47– 54, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bollinger:2004:B

[BBJ04] Terry Bollinger, Mike Barker, and Philipp K. Janert. Bookshelf. IEEE Software, 21(5):106–109, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
05/s5106.pdf.

Bodoff:2005:WMS

[BBMH05] David Bodoff, Mordechai Ben-Menachem, and Patrick C. K. Hung. Web metadata standards: Observations and prescriptions. IEEE Software, 22(1):78–85, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Blyth:1990:CFM

[BBRTL90] David Blyth, Cornelia Boldyreff, Clive Ruggles, and Nik Tetteh-Lartey. The case for formal methods in standards. IEEE Software, 7(5):65–67, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Berglund:1985:AMC

[BC85] Eric J. Berglund and David R. Cheriton. Amaze: a multiplayer computer game. IEEE Software, 2(3): 30–39, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Britcher:1986:UMD

[BC86] Robert N. Britcher and James J. Craig. Using modern design practices to upgrade aging software systems. IEEE Software, 3(3): 16–24, May 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Beznosov:2008:SRU

[BC08] Konstantin (Kosta) Beznosov and Brian Chess. Security for the rest of us: An industry perspective on the secure-software challenge. IEEE Software, 25(1):10–12, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2008/
                                        01/mso2008010010.pdf.

Baniassad:2006:DEA

[BCA^{+}06] Elisa Baniassad, Paul C. Clements, Joao Araújo, Ana Moreira, Awais Rashid, and Bedir Tekinerdogan. Discovering early aspects. IEEE Software, 23(1): 61–70, January/February 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brichau:2006:MCA

[BCC^{+}06] Johan Brichau, Ruzanna Chitchyan, Siobhán Clarke, Ellie D’Hondt, Alessandro Garcia, Michael Haupt, Wouter Joosen, Shmuel Katz, Jacques Noyé, Awais Rashid, and Mario Südholt. A model curriculum for aspect-oriented software development. IEEE Software, 23(6):53–61, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Basili:2008:UHP

[BCC^{+}08] Victor R. Basili, Jeffrey C.

Carver, Daniela Cruzes, Lorin M. Hochstein, Jeffrey K. Hollingsworth, Forrest Shull, and Marvin V. Zelkowitz. Understanding the high-performance-computing community: a software engineer’s perspective. IEEE Software, 25(4):29–36, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buchsbaum:2001:VAS

[BCH^{+}01] Adam Buchsbaum, Yih-Farn Chen, Huale Huang, Eleftherios Koutsofios, John Mocenigo, Anne Rogers, Michael Jankowsky, and Spiros Mancoridis. Visualizing and analyzing software infrastructures. IEEE Software, 18(5):62– 70, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://computer.org/software/
                                        so2001/s5062abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5062.pdf.

Borstler:2002:TPC

[BCH^{+}02] Jürgen Börstler, David Carrington, Gregory W. Hislop, Susan Lisack, Keith Olson, and Laurie Williams. Teaching the PSP: Challenges and lessons learned. IEEE Software, 19(5):42–48, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s5042.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s5042abs.
                                        htm.

Blaine:2008:SQR

[BCH08] J. David Blaine and Jane Cleland-Huang. Software quality requirements: How to balance competing priorities. IEEE Software, 25(2): 22–24, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2008/
                                        02/mso2008020022.pdf.

Beckman:1997:CCI

[BCKM97] Kathy Beckman, Neal Coulter, Soheil Khajenoori, and Nancy R. Mead. Collaborations: Closing the industry-academia gap. IEEE Software, 14(6):49– 57, November/December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s6049.pdf;   http://dlib.
computer.org/so/books/
so1997/pdf/s6049abs.htm.

Battin:2001:LRG

[BCKS01] Robert D. Battin, Ron Crocker, Joe Kreidler, and K. Subramanian. Leveraging resources in global software development. IEEE Software, 18(2):70–77, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2070abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2070.pdf.

Bhattacharyya:1988:VPC

[BCM88] Mitali Bhattacharyya, David Cohrs, and Barton Miller. Visual process connector for UNIX. IEEE Software, 5(4):43–50, July 1988. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bockle:2004:CRS

[BCM^{+}04] Gunter Bockle, Paul Clements, John D. McGregor, Dirk Muthig, and Klaus Schmid. Calculating ROI for software product lines. IEEE Software, 21(3):23–31, May/ June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Berzal:2005:LTA

[BCMV05] Fernando Berzal, Juan-Carlos Cubero, Nicolas Marin, and Maria-Amparo Vila. Lazy types: Automating dynamic strategy selection. IEEE Software, 22(5):98–106, September/ October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brady:1994:MAS

[BD94] Sheila Brady and Tom DeMarco. Managementaided software engineering. IEEE Software, 11 (6):25–32, November 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:1997:GEI

[BD97] Barry W. Boehm and Tom DeMarco. Guest Editors’ introduction: Software risk management. IEEE Software, 14(3):17– 19, May/June 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bourque:1999:GSE

[BDA^{+}99] Pierre Bourque, Robert Dupuis, Alain Abran, James W. Moore, and Leonard Tripp. The guide to the software engineering body of knowledge. IEEE Software, 16(6):35–44, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6035abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6035.pdf.

Basili:2004:UMD

[BDA04] Victor Basili, Paolo Donzelli, and Sima Asgari. A unified model of dependability: Capturing dependability in context. IEEE Software, 21(6):19–25, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brookhart:2007:Ba

[BdFD07] Scott Brookhart, Cristiano di Flora, and Carlos Denner dos Santos, Jr. Bookshelf. IEEE Software, 24(1):92–94, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.

                                        org/comp/mags/so/2007/
                                        01/s1092.pdf.

Butler:1996:ADC

[BDHJ96] Shawn Butler, David Diskin, Norman Howes, and Kathleen Jordan. Architectural design of a common operating environment. IEEE Software, 13 (6):57–65, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boegh:1999:MSQ

[BDKP99] Jørgen Bøegh, Stefano Depanfilis, Barbara Kitchenham, and Alberto Pasquini. A method for software quality planning, control, and evaluation. IEEE Software, 16(2):69–77, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s2069abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s2069.pdf.

Birk:2002:PNL

[BDS02] Andreas Birk, Torgeir Dingsøyr, and Tor Stålhane. Postmortem: Never leave a project without it. IEEE Software, 19(3):43– 45, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3043.
pdf;   http://www.computer.
org/software/so2002/s3043abs.
htm.

Basili:1994:TTM

[BDY94] Victor R. Basili, Michael K.

Daskalantonakis, and Robert H. Yacobellis. Technology transfer at Motorola. IEEE Software, 11(2):70– 76, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Beck:1990:SMP

[Bec90] Bob Beck. Shared-memory parallel programming in C++. IEEE Software, 7(4): 38–48, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Beck:2001:AF

[Bec01] Kent Beck. Aim, fire.

IEEE Software, 18(5):87– 89, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s5087abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5087.pdf.

Beck:2007:PC

[Bec07] Ian Alexander Kent Beck.

Point/counterpoint. IEEE Software, 24(2):62–65, March/ April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Beizer:1997:CPM

[Bei97a] B. Beizer. Cleanroom process model: a critical examination. IEEE Software, 14(2):14–16, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Beizer:1997:BWS

[Bei97b] Boris Beizer. Bookshelf: a watershed software reliability book: Handbook of Software Reliability Engineering, edited by Michael R. Lyn. IEEE Software, 14(4):114–115, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4114.
pdf.

Belady:1999:SGR

[Bel99] Laszlo A. Belady. Soapbox: Global R&D: How to break barriers. IEEE Software, 16(4):15–17, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4015.
pdf.

Bennett:1995:LSC

[Ben95] Keith Bennett. Legacy systems: Coping with success. IEEE Software, 12(1):19– 23, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bentley:1999:EPP

[Ben99]  Jon Bentley.  Excerpt

from Programming Pearls: The back of the envelope. IEEE Software, 16 (5):121–125, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5121abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5121.pdf.

Bennett:2000:CWB

[Ben00] Al Bennett. Culture at work: Baby boomers need not apply. IEEE Software, 17(4):119–121, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4119.
pdf.

Bernstein:1993:GDR

[Ber93] Lawrence Bernstein. Get the design right! IEEE Software, 10(5):61–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bersoff:1994:ASS

[Ber94] Ed Bersoff. Anatomy of a software start-up. IEEE Software, 11(1):92–94, 100, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Berard:1996:MBT

[Ber96a]  Ed Berard.  Manager:

Bringing testing into the fold. IEEE Software, 13 (3):91–92, 95, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bernstein:1996:BST

[Ber96b] Lawrence Bernstein. Better software through operational dynamics. IEEE Software, 13(2):107–109, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bergen:1999:KOE

[Ber99] Shani Bergen. KosovoNet.org explores technology’s good side. IEEE Software, 16 (4):113, July/August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s4109.pdf.

Berczuk:2003:PSC

[Ber03]  Steve Berczuk.  Prag-

matic software configuration management. IEEE Software, 20(2):15–17, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2015.
pdf.

Berry:2008:SES

[Ber08] Daniel M. Berry. The software engineering silver bullet conundrum. IEEE Software, 25(2):18–19, March/ April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:2000:SEP

[BF00] Barry Boehm and Richard E.

Fairley. Software estimation perspectives: Guest Editors’ introduction. IEEE Software, 17(6):22–26, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6022.
pdf.

Brukardt:2000:NIS

[BFC00] Randall Brukardt, Nancy A.

Forbes, and David N. Card. In the news: An ISO Standard guards the Ada hen house: Innovation and the National Reconnaissance Office: Making measurement understandable. IEEE Software, 17(1):92–96, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s1092.
pdf.

Baudry:2005:ATC

[BFJT05] Benoit Baudry, Franck Fleurey, Jean-Marc Jezequel, and Yves Le Traon. Automatic test case optimization: a bacteriologic algorithm. IEEE Software, 22(2):76–82, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bellinzona:1995:RSO

[BFP95] Roberto Bellinzona, Maria Grazia Fugini, and Barbara Pernici. Reusing specifications in OO applications. IEEE Software, 12(2):65– 75, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Basili:1994:SPE

[BG94] Victor Basili and Scott Green. Software process evolution at the SEL. IEEE Software, 11(4):58–66, July 1994. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Birman:1995:RTC

[BG95] Kenneth P. Birman and Bradford B. Glade. Reliability through consistency. IEEE Software, 12(3):29– 41, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bhattacherjee:1998:UMO

[BG98] Anol Bhattacherjee and James Gerlach. Understanding and managing OOT adoption. IEEE Software, 15(3):91–96, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3091abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3091.pdf.

Boehm:2001:DGR

[BGB01] Barry Boehm, Paul Grünbacher, and Robert O. Briggs. Developing groupware for requirements negotiation: Lessons learned. IEEE Software, 18(3):46–55, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s3046abs.htm;
http://dlib.computer.

                                        org/so/books/so2001/pdf/
                                        s3046.pdf.

Brandt:2009:OPW

[BGL^{+}09] Joel Brandt, Philip J. Guo, Joel Lewenstein, Mira Dontcheva, and Scott R. Klemmer. Opportunistic programming: Writing code to prototype, ideate, and discover. IEEE Software, 26(5):18–24, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bhat:2006:ORE

[BGM06] Jyoti M. Bhat, Mayank Gupta, and Santhosh N. Murthy. Overcoming requirements engineering challenges: Lessons from offshore outsourcing. IEEE Software, 23(5):38–44, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Blanco:2001:SPL

[BGS01] Marina Blanco, Pedro Gutiérrez, and Giuseppe Satriani. SPI patterns: Learning from experience. IEEE Software, 18(3):28– 35, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s3028abs.htm;
                                        http://dlib.computer.

org/so/books/so2001/pdf/
s3028.pdf.

Bowen:1995:SMM

[BH95a] Jonathan P. Bowen and Michael G. Hinchey. Seven more myths of formal methods. IEEE Software, 12(4): 34–41, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bray:1995:RCM

[BH95b] Olin Bray and Michael M.

Hess. Reengineering a configuration management system. IEEE Software, 12 (1):55–63, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barber:2001:SAC

[BH01] K. Suzanne Barber and Jim Holt. Software architecture correctness. IEEE Software, 18(6):64–65, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6064abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6064.pdf.

Baldwin:2002:IMS

[BH02] Doug Baldwin and Peter B. Henderson. The importance of mathematics to the software practitioner. IEEE Software, 19 (2):112, 110–111, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2112.
pdf;   http://www.computer.
org/software/so2002/s2112abs.
htm.

Bhansali:1996:USS

[Bha96] Praful V. Bhansali. Universal safety standard is infeasible — for now. IEEE Software, 13(2):8–10, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barney:2009:REC

[BHAW09] Sebastian Barney, Ganglan Hu, Aybüke Aurum, and Claes Wohlin. Requirements engineering: Creating software product value in China. IEEE Software, 26(4):84–90, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Biagioni:1986:POS

[BHHM86] E. Biagioni, K. Hinrichs, G. Heiser, and C. Muller. Portable operating system interface and utility library for Modula-2. IEEE Software, 3(6):18–26, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Birk:2003:PLE

[BHJ^{+}03] Andreas Birk, Gerald Heller, Isabel John, Klaus Schmid, Thomas von der Maßen, and Klaus Müller. Product line engineering: The state of the practice. IEEE Software, 20(6):52–60, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6052.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6052abs.htm.

Boehm:2004:RSD

[BHJM04] Barry Boehm, LiGuo Huang, Apurva Jain, and Ray Madachy. The ROI of software dependability: The iDAVE model. IEEE Software, 21(3):54– 61, May/June 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Broy:2001:CRG

[BHKR01] Manfred Broy, Susanne Hartkopf, Kirstin Kohler, and Dieter Rombach. Country report: Germany: Combining software and application competencies. IEEE Software, 18(4):93– 95, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s4093.
                                        pdf.

Berman:1998:LPO

[BHR^{+}98] Kevin J. Berman, Andy Huber, Jarrett Rosenberg, David N. Card, and Robert Binder. Letters: In praise of optimism; thou shalt not miscalculate; manufacturing quality models. IEEE Software, 15(2):10–12, 14, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s2010.
                                        pdf.

Barbacci:1985:SEI

[BHS85] Mario R. Barbacci, A. Nico Habermann, and Mary Shaw. Software engineering institute: Bridging practice and potential. IEEE Software, 2(6):4–21, November 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2007:PPF

[BHS07] Frank Buschmann, Kevlin Henney, and Douglas C. Schmidt. Past, present, and future trends in software patterns. IEEE Software, 24(4):31–37, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:1996:IQR

[BI96] Barry Boehm and Hoh In.

Identifying quality requirement conflicts. IEEE Software, 13(2):25–35, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bielak:2000:ISE

[Bie00] James Bielak. Improving size estimates using historical data. IEEE Software, 17(6):27–35, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6027abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6027.pdf.

Biehl:2004:SSS

[Bie04] Richard E. Biehl. Six sigma for software. IEEE Software, 21(2):68–70, March/ April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Biffl:2000:UID

[Bif00] Stefan Biffl. Using inspection data for defect estimation. IEEE Software, 17(6):36–44, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6036abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6036.pdf.

Bigelow:1988:HC

[Big88] James Bigelow. Hypertext and CASE. IEEE Software, 5(2):23–27, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Biggerstaff:1996:EML

[Big96]  Ted Biggerstaff.  From

the Editor — Moore’s law: Change or die. 1995 best practice article award. Mead named to Editorial Board. 1995 referees. IEEE Software, 13(1):4–6, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Billingsley:1994:IES

[Bil94] Patricia A. Billingsley. Interface: Ergonomic standards go beyond hardware. IEEE Software, 11(2):82– 84, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). Discusses ISO 9241, “Ergonomic Requirements for Office Work with Visual Display Terminals”.

Binder:1997:QTC

[Bin97] Robert V. Binder. Quality time: Can a manufacturing quality model work for software? IEEE Software, 14 (5):101–102, 105, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5101.
pdf.

Berzal:2006:B

[BJ06] Fernando Berzal and Stuart James. Bookshelf. IEEE Software, 23(4):100– 102, July/August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
04/s4100.pdf.

Burton:1989:FPQ

[BK89] F. Warren Burton and John G. Kollias. Functional programming with quadtrees. IEEE Software, 6(1):90–97, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bjorkander:2003:ASU

[BK03] Morgan Björkander and Cris Kobryn. Architecting systems with UML 2.0. IEEE Software, 20(4):57– 61, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4057.
pdf;   http://www.computer.
org/software/so2003/s4057abs.
htm.

Berry:2005:SDA

[BK05] Daniel M. Berry and Erik Kamsties. The syntactically dangerous all and plural in specifications. IEEE Software, 22(1):55–57, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bachle:2007:RR

[BK07] Michael Bächle and Paul Kirchberg. Ruby on Rails. IEEE Software, 24(6):105– 108, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bassin:1998:ESD

[BKS98] Kathryn A. Bassin, Theresa Kratschmer, and P. Santhanam. Evaluating software development objectively. IEEE Software, 15(6):66–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s6066abs.htm;
http://dlib.computer.

org/so/books/so1998/pdf/
s6066.pdf.

Bruns:2009:WAT

[BKW09] Andreas Bruns, Andreas Kornstadt, and Dennis Wichmann. Web application tests with Selenium. IEEE Software, 26(5):88–91, September/ October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Berzins:1990:ISL

[BL90] Valdis Berzins and Luqi.

An introduction to the specification language Spec. IEEE Software, 7(2):74– 84, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brocklehurst:1992:NWG

[BL92] Sarah Brocklehurst and Bev Littlewood. New ways to get accurate reliability measures. IEEE Software, 9(4):34–??, July 1992. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Berry:1998:RE

[BL98] Daniel M. Berry and Brian Lawrence. Requirements engineering. IEEE Software, 15(2):26–29, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2026.
pdf.

Bisiani:1988:TCT

[BLA88] Roberto Bisiani, Francois Lecouat, and Vincenzo Ambriola. Tool to coordinate tools. IEEE Software, 5(6):17–25, November 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Blake:2007:DCS

[Bla07] M. Brian Blake. Decomposing composition: Serviceoriented software engineers. IEEE Software, 24(6):68– 77, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barlas:1996:NAR

[BLP96] Stephen Barlas, Michael Lutz, and Shari Lawrence Pfleeger. In the news — anatomy of a runaway: What grounded the AAS. now that objects are old hat, quo vadis, OOPSLA? signs of life — and growth — at Comdex. workshop defines problems with software-engineering data. briefs. IEEE Software, 13(1):104–106, 108– 112, January 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Balasubramaniam:2008:SSC

[BLSS08] Sriram Balasubramaniam, Grace A. Lewis, Soumya Simanta, and Dennis B. Smith. Situated software: Concepts, motivation, technology, and the future. IEEE Software, 25(6):50– 55, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Blum:1987:TDE

[Blu87] Bruce I. Blum. Tedium development environment for information systems. IEEE Software, 4(2):25– 34, March 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bisbal:1999:LIS

[BLWG99] Jesús Bisbal, Deirdre Lawless, Bing Wu, and Jane Grimson. Legacy information systems: Issues and directions. IEEE Software, 16(5):103–111, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5103abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5103.pdf.

Bollinger:1991:CLS

[BM91] Terry B. Bollinger and Clement McGowan. A critical look at software capability evaluations. IEEE Software, 8(4):25–41, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See comment [HC91].

Brown:1992:LIM

[BM92] Alan W. Brown and John A. Mcdermid. Learning from IPSE’s mistakes. IEEE Software, 9(2):23– 28, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Becker:2001:GPW

[BM01] Shirley A. Becker and Florence E. Mottay. A global perspective on Web site usability. IEEE Software, 18(1):54–61, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s1054abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s1054.pdf.

Borjesson:2004:SPI

[BM04a] Anna Borjesson and Lars Mathiassen. Successful process implementation. IEEE Software, 21(4):36– 44, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Browne:2004:NES

[BM04b] Glenn J. Browne and Nirup M. Menon. Network effects and social dilemmas in technology industries. IEEE Software, 21(5): 44–50, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bollinger:2009:CLS

[BM09] Terry Bollinger and Clement McGowan. A critical look at software capability evaluations: An update. IEEE Software, 26(5): 80–83, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Berzal:2001:BDW

[BMGT01] Fernando Berzal, Nicolás Marín, Ted Graham, and Andrey A. Terekhov. Bookshelf: Data warehousing 101 — not exactly; introducing the rational unified process; Cobol updates. IEEE Software, 18(5):92–95, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5092abs.htm;

                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5092.pdf.

Bruckhauas:1996:ITS

[BMJH96] Tillmann Bruckhauas, Nazim H. Madhavji, Ingrid Janssen, and John Henshaw. The impact of tools on software productivity. IEEE Software, 13(5):29–38, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ben-Menachem:2004:IIT

[BMM04] Mordechai Ben-Menachem and Garry S. Marliss. Inventorying information technology systems: Supporting the “Paradigm of Change”. IEEE Software, 21(5):34–43, September/ October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ben-Menachem:2005:IAC

[BMM05] Mordechai Ben-Menachem and Garry S. Marliss. IT Assets—control by importance and exception: Supporting the “Paradigm of Change”. IEEE Software, 22(4):94–102, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bollinger:1999:ROS

[BNST99] Terry Bollinger, Russell Nelson, Karsten Self, and Stephen Turnbull. Response: Open-source methods: Peering through the clutter. IEEE Software, 16(4):8–11, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4006.
pdf.

Beck:1989:PPP

[BO89] Bob Beck and David Olien.

A parallel-programming process model. IEEE Software, 6(3):63–72, May 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boasson:1995:ASA

[Boa95] Maarten Boasson. The artistry of software architecture. IEEE Software, 12 (6):13–16, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boasson:1997:IFF

[Boa97] Maarten Boasson. Insider:

The foolishness of fashion. IEEE Software, 14(4):18– 19, July/August 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s4018.
                                        pdf.

Bock:2003:UP

[Boc03] Conrad Bock. UML without pictures. IEEE Software, 20(5):33–35, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/05/s5033.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/so/2003/05/
                                        s5033abs.htm.

Boehm:1984:VVS

[Boe84] Barry W. Boehm. Verifying and validating software requirements and design specifications. IEEE Software, 1(1):75–88, January 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:1991:SRM

[Boe91] Barry W. Boehm. Software risk management: Principles and practices. IEEE Software, 8(1):32–41, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:1996:ASP

[Boe96] Barry Boehm. Anchoring the software process. IEEE Software, 13(4):73– 82, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:2000:MSS

[Boe00] Barry Boehm. Manager:

Safe and simple software cost analysis. IEEE Software, 17(5):14–17, September/October 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5014.
pdf.

Boegh:2006:CSC

[Bøe06] Jørgen Bøegh. Certifying software component attributes. IEEE Software, 23 (3):74–81, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boegh:2008:NSQ

[Bøe08] Jørgen Bøegh. A new standard for quality requirements. IEEE Software, 25(2):57–63, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bollinger:1999:LPO

[Bol99] Terry Bollinger. Linux in practice: An overview of applications. IEEE Software, 16(1):72–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s1072abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s1072.pdf.

Bollinger:2000:SVB

[Bol00]  Terry Bollinger.  Soap-

box: Visual Basic: Taming the woolly mammoth. IEEE Software, 17(3):16– 17, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3016.
pdf.

Bollinger:2002:GEI

[Bol02] Terry Bollinger. Guest Editor’s introduction: Breaking out of the software engineering mind-mold. IEEE Software, 19(4):90– 91, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4090.
pdf;   http://www.computer.
org/software/so2002/s4090abs.
htm.

Bollinger:2004:CRF

[Bol04] Terry Bollinger. Conference report: The future of software infrastructure protection. IEEE Software, 21(6):59–61, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bolloju:2009:CMS

[Bol09] Narasimha Bolloju. Conceptual modeling of systems integration requirements. IEEE Software, 26(5):66–74, September/ October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:1994:CAO

[Boo94] Grady Booch. Coming of age in an object-oriented world. IEEE Software, 11 (6):33–41, November 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2006:AA

[Boo06a] Grady Booch. The accidental architecture. IEEE Software, 23(3):9–11, May/ June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2006:SG

[Boo06b] Grady Booch. From small to gargantuan. IEEE Software, 23(4):14–15, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2006:GF

[Boo06c] Grady Booch. Goodness of fit. IEEE Software, 23(6):14–15, November/ December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2006:A

[Boo06d] Grady Booch. On architecture. IEEE Software, 23(2): 16–18, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2007:AP

[Boo07a] Grady Booch. Artifacts and process. IEEE Software, 24(6):26–27, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2007:EAF

[Boo07b] Grady Booch. The economics of architecturefirst. IEEE Software, 24(5): 18–20, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2007:IA

[Boo07c] Grady Booch. The irrelevance of architecture. IEEE Software, 24(3):10– 11, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2007:IWI

[Boo07d] Grady Booch. It is what it is because it was what it was. IEEE Software, 24(1): 14–15, January/February 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2007:STP

[Boo07e] Grady Booch. Speaking truth to power. IEEE Software, 24(2):12–13, March/ April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2007:WTA

[Boo07f]  Grady Booch.  The

well-tempered architecture. IEEE Software, 24(4):24– 25, July/August 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2008:AOP

[Boo08a] Grady Booch. Architectural organizational patterns. IEEE Software, 25 (3):18–19, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2008:MAC

[Boo08b] Grady Booch. Measuring architectural complexity. IEEE Software, 25(4): 14–15, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2008:MSA

[Boo08c] Grady Booch. Morality and the software architect. IEEE Software, 25(1):8– 9, January/February 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2008:NTY

[Boo08d] Grady Booch. Nine things you can do with old software. IEEE Software, 25(5):93–94, September/ October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2008:ABF

[Boo08e] Grady Booch. On architecture: Back to the future. IEEE Software, 25(6):20– 21, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2008:TM

[Boo08f] Grady Booch. Tribal memory. IEEE Software, 25(2): 16–17, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2009:ALR

[Boo09a] Grady Booch. On architecture: Like a river.

IEEE Software, 26(3):10– 11, May/June 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2009:AB

[Boo09b] Grady Booch. On architecture: Not with a bang. IEEE Software, 26(1): 74–75, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2009:ADS

[Boo09c] Grady Booch. On architecture: The defenestration of superfluous architectural accoutrements. IEEE Software, 26(4):7–8, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2009:ARP

[Boo09d] Grady Booch. On architecture: The resting place of innovation. IEEE Software, 26(2):12–13, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2009:SAFa

[Boo09e] Grady Booch. Software abundance in the face of economic scarcity, part 1. IEEE Software, 26(5): 12–13, September/October 2009. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2009:SAFb

[Boo09f] Grady Booch. Software abundance in the face of economic scarcity, part 2. IEEE Software, 26(6):15– 16, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Borgida:1985:FLD

[Bor85] Alexander Borgida. Features of languages for the development of information systems at the conceptual level. IEEE Software, 2 (1):63–72, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brownsword:2000:DNP

[BOS00] Lisa Brownsword, Tricia Oberndorf, and Carol A. Sledge. Developing new processes for COTS-based systems. IEEE Software, 17(4):48–55, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4048abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4048.pdf.

Boyer:1997:LOC

[Boy97] Kenneth W. Boyer, Jr.

Letters: Orienting cleanroom design. IEEE Software, 14(4):12, 14, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4010.
pdf.

Barnard:1994:MCI

[BP94] Jack Barnard and Art Price. Managing code inspection information. IEEE Software, 11(2):59– 69, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Burnell:2002:TDM

[BPD02] Lisa J. Burnell, John W.

Priest, and John R. Durrett. Teaching distributed multidisciplinary software development. IEEE Software, 19(5):86–93, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5086.
pdf;   http://www.computer.
org/software/so2002/s5086abs.
htm.

Blaha:1994:COM

[BPS94] Michael Blaha, William Premerlani, and Hwa Shen.

Converting OO models into RDBMS schema. IEEE Software, 11(3):28–39, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bogosian:2003:LYW

[BPS^{+}03] James Bogosian, Bob Pedigo, Noah Silva, Geert Poels, Michael Wisniewski, John Selbie, John D. Verne, Francisco Gutierrez, Mikko Kurki-Suonio, Ilango Veerasingam, Jiann-Ming Su, Artem Pyatakov, Jimmy Rimmer, Jose Robles, Magdin Stoica, and Chris Pursglove. Letters: So you want to be in the movies? the art of enbugging; when to make a type. IEEE Software, 20(3):8– 13, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3008.
pdf.

Baker:1985:ATS

[BR85] T. P. Baker and G. A. Riccardi. Ada tasking: from semantics to efficient implementation. IEEE Software, 2(2):34–46, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Biggerstaff:1987:RFA

[BR87] Ted J. Biggerstaff and Charles Richter. Reusability framework, assessment, and directions. IEEE Software, 4(2):41–49, March 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bokhari:1999:LDW

[BR99] Shahid H. Bokhari and Rafeequr Rehman. Linux and the developing world. IEEE Software, 16(1): 58–??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so1999/s1058abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s1058.pdf.

Brettschneider:1989:YSR

[Bre89] Ralph Brettschneider. Is your software ready for release? IEEE Software, 6(4):100, 102, 108– ??, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Britcher:1996:AHM

[Bri96] Bob Britcher. In appreciation of Harlan Mills: Bringing programming under intellectual control. IEEE Software, 13(2):111, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Britcher:1997:LHC

[Bri97] Robert N. Britcher. Letters: Hobson’s choice, cleanroom style. IEEE Software, 14(4):12–13, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s4010.
                                        pdf.

Baskerville:2003:ISS

[BRL^{+}03] Richard Baskerville, Balasubramaniam Ramesh, Linda Levine, Jan Pries-Heje, and Sandra Slaughter. Is Internet-speed software development different? IEEE Software, 20(6):70–77, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/06/s6070.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/so/2003/06/
                                        s6070abs.htm.

Brooks:1995:BEM

[Bro95] Frederick P. Brooks, Jr. Book excerpt: “The Mythical Man Month After 20 Years”. IEEE Software, 12 (5):57–60, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brown:1996:ISM

[Bro96] Norm Brown. Industrialstrength management strategies. IEEE Software, 13(4): 94–103, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brown:1999:MMM

[Bro99] Alan W. Brown. Mastering the middleware muddle. IEEE Software, 16(4):18– 21, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4018.
pdf.

Baron:1985:MOE

[BRS^{+}85] Robert Baron, Richard Rashid, Ellen Siegel, Avadis Tevanian, and Michael Young. Mach-1: An operating environment for largescale multiprocessor applications. IEEE Software, 2 (4):65–67, July 1985. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brown:1985:TAA

[BS85] Marc H. Brown and Robert Sedgewick. Techniques for algorithm animation. IEEE Software, 2(1):28–39, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Baker:1986:ART

[BS86] Theodore P. Baker and Gregory M. Scallon. An architecture for real-time software systems. IEEE Software, 3(3):50–58, May 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bell:1995:TEN

[BS95] Rodney Bell and David Sharon. Tools to engineer new technologies into applications. IEEE Software, 12(2):11–16, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bowen:2000:RAS

[BS00] Thomas F. Bowen and Mark E. Segal. Remediation of application-specific security vulnerabilities at runtime. IEEE Software, 17(5):59–67, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so/s5059abs.htm;   http:
                                        //dlib.computer.org/so/
                                        books/so2000/pdf/s5059.
                                        pdf.

Brykczynski:2003:RIB

[BS03] Bill Brykczynski and Robert A. Small. Reducing Internetbased intrusions: Effective security patch management. IEEE Software, 20(1):50–57, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1050.
pdf;   http://www.computer.
org/software/so2003/s1050abs.
htm.

Berzal:2004:B

[BS04] Fernando Berzal and Diomidis Spinellis. Bookshelf. IEEE Software, 21(3):104– 105, May/June 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brookhart:2007:Bb

[BSFS07] Scott Brookhart, Todd Schultz, Paul Freedman, and Art Sedighi. Bookshelf. IEEE Software, 24(6):109–112, November/ December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
06/mso2007060109.pdf.

Ben-Shaul:1998:INC

[BSGF98] Israel Ben-Shaul, James W.

Gish, and William Robinson Feature. An integrated network component architecture. IEEE Software, 15(5):79–87, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s5079.
                                        pdf;   http://www.computer.
                                        org/computer/co1998/s5079abs.
                                        htm.

Batory:1994:GMS

[BST^{+}94] Don Batory, Vivek Singhal, Jeff Thomas, Sankar Dasari, Bart Geraci, and Marty Sirkin. The GenVoca model of software-system generators. IEEE Software, 11(5):89–94, September 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:2005:MCI

[BT05] Barry Boehm and Richard Turner. Management challenges to implementing agile processes in traditional development organizations. IEEE Software, 22(5):30–39, September/ October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bollinger:1999:LSB

[BTC^{+}99] Terry Bollinger, James E. Tomayko, Forrest J. Cavalier III, Lee Aydelotte, Herb Krasner, Robert Schaaf, Michael Deck, Wolfgang Strigel, and Robert Oshana. Letters: Success is the best instructor; good coaches, better managers; growing lucky strikes; what is professional practice?; whence leadership?; tailoring cleanroom affordably. IEEE Software, 16(2):9–13, March/April 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2009.
pdf.

Budd:1991:BIR

[Bud91] Timothy A. Budd. Blending imperative and relational programming. IEEE Software, 8(1):58–65, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Budgen:1999:LOS

[Bud99] David Budgen. Loyal opposition: Software design methods: Life belt or leg iron? IEEE Software, 16 (5):136, 133–135, September/October 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s5136.
pdf.

Burgess:1994:NST

[Bur94] Angela Burgess. In the news: Studies try to quantify development trends; 1994: Year of the asynchronous transfer mode?; interview: Tom Gilb; in brief. IEEE Software, 11 (1):101–102, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Burgess:1995:FEB

[Bur95a] Angela Burgess. Finding an experimental basis for software engineering: Interview with Victor Basili, Software Engineering Lab. IEEE Software, 12(3):92– 93, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Burgess:1995:MAM

[Bur95b]  Angela Burgess.  Mad

about or mad at measurement? IEEE Software, 12 (1):115–118, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2009:IPA

[Bus09a] Frank Buschmann. Introducing the pragmatic architect. IEEE Software, 26(5):10–11, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2009:LFP

[Bus09b] Frank Buschmann. Learning from failure, part 1:

Scoping and requirements woes. IEEE Software, 26(6):68–69, November/ December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boehm:2008:ACC

[BV08] Barry W. Boehm and Ricardo Valerdi. Achievements and challenges in cocomo-based software resource estimation. IEEE Software, 25(5):74–83, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bollinger:2004:PSA

[BVB04] Terry Bollinger, Jeffrey Voas, and Maarten Boasson. Persistent software attributes. IEEE Software, 21(6):16–18, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
06/s6016.pdf.

Brown:1996:FES

[BW96] Allan W. Brown and Kurt C. Wallnau. A framework for evaluating software technology. IEEE Software, 13(5):39– 49, September 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Brown:1998:CSC

[BW98] Alan W. Brown and Kurt C. Wallnau. The current state of CBSE. IEEE Software, 15(5):37– 46, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s5037.pdf;   http://www.
                                        computer.org/computer/
                                        co1998/s5037abs.htm.

Bytheway:1999:GEI

[Byt99] Andrew J. Bytheway. Guest Editor’s introduction: Successful software projects and how to achieve them. IEEE Software, 16(3):15– 17, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s3015.
                                        pdf.

Basili:1995:SRS

[BZM^{+}95] Victor Basili, Marvin Zelkowitz, Frank McGarry, Jerry Page, Sharon Waligora, and Rose Pajerski. Special report: SEL’s software process improvement program. IEEE Software, 12(6):83–87, November 1995. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Carmel:2001:TAA

[CA01] Erran Carmel and Ritu Agarwal. Tactical approaches for alleviating distance in global software development. IEEE Software, 18(2):22–29, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2022abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2022.pdf.

Cunto:1991:FPE

[CAGR91] Walter Cunto, Jorge Araujo, Federico Giovannetti, and Juan Rivero. Fplus: a programming environment for scientific applications. IEEE Software, 8(5):81–??, September 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cameron:1988:APP

[Cam88] Robert D. Cameron. Abstract pretty printer. IEEE Software, 5(6):61–67, November 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cannava:1995:BGI

[Can95] Marc N. Cannava. Bookshelf: Gentle introduction to Usenet (review of Usenet: Net News for Everyone). IEEE Software, 12(3):106, May 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carver:1988:ALS

[Car88] Doris Carver. Acceptable legal standards for software. IEEE Software, 5(3): 87–93, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Card:1994:QTS

[Car94] David Card. Quality time: Statistical process control for software? IEEE Software, 11(3):95–97, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See letters and responses in January 1995 issue, pp. 8, 10.

Card:1995:GEI

[Car95] David Card. Guest Editor’s introduction: The RAD fad: Is timing really everything? IEEE Software, 12 (5):19–22, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carr:1997:CRM

[Car97] Marvin J. Carr. Counterpoint: Risk management may not be for everyone. IEEE Software, 14 (3):21, 24, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Card:1998:LOM

[Car98] David N. Card. Learning from our mistakes with defect causal analysis. IEEE Software, 15(1):56–63, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1056abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s1056.pdf.

Card:2000:QTS

[Car00a] David N. Card. Quality time: Sorting out six sigma and the CMM. IEEE Software, 17(3):11– 13, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3011.
pdf.

Carter:2000:SCU

[Car00b] Stephen Carter. SoftLaw:

E-commerce: The UK’s model for a legal European framework. IEEE Software, 17(2):90–91, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2090.
pdf.

Charette:1997:MRS

[CAW97] Robert N. Charette, Kevin MacG. Adams, and Mary B. White. Managing risk in software maintenance. IEEE Software, 14(3):43– 50, May/June 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Coetzee:1998:NWQ

[CB98] Serena Coetzee and Judith Bishop. A new way to query GISs on the Web. IEEE Software, 15(3):31– 40, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3031abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3031.pdf.

Curtis:2000:PCB

[CB00] Bill Curtis and Terry Bollinger. Point/counterpoint: Building accelerated organizations vs. building tech-savvy organizations. IEEE Software, 17(4):72– 75, July/August 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4072.
pdf.

Cercone:1985:ISF

[CBK85] Nick Cercone, John Boates, and Max Krause. An interactive system for finding perfect hash functions. IEEE Software, 2(6):38–53, November 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cheung:1990:FDD

[CBM90] Wing Hong Cheung, James P.

Black, and Eric Manning. A framework for distributed debugging. IEEE Software, 7(1):106–115, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chang:1998:GEI

[CBM98] Carl Chang, Maarten Boasson, and Tomoo Matsubara. Guest Editors’ introduction. IEEE Software, 15(6):22–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6022.
pdf.

Chen:2005:FRD

[CBMP05] Zhihao Chen, Barry Boehm, Tim Menzies, and Daniel Port. Finding the right data for software cost modeling. IEEE Software, 22(6):38–46, November/ December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chikofsky:1990:RED

[CC90] Elliot J. Chikofsky and James H. Cross. Reverse engineering and design recovery: a taxonomy. IEEE Software, 7(1):13–17, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cohen:1993:ARO

[CC93] Donald Cohen and Neil Campbell. Automating relational operations on data structures. IEEE Software, 10(3):53–60, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chu:1994:PSS

[CC94] Chih-Ping Chu and Doris L. Carvier. Parallelizing subroutines in sequential programs. IEEE Software, 11 (1):77–85, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chmura:1995:TTA

[CC95a] Alan Chmura and Henry D.

Crockett. Toolbox: Tools to align goals and information systems. IEEE Software, 12(3):108–110, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chmura:1995:PCW

[CC95b] Alan Chmura and Henry David Crockett. Point-counterpoint: What’s the proper role for CASE tools? IEEE Software, 12(2):18–20, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Codina:1995:LWB

[CC95c] Jorge Codina and Bob Charette. Letters: Who’s to blame for lowball bids? IEEE Software, 12(6):4, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Christensen:1996:BIR

[CC96] Mark Christensen and Carl Chang. Blueprint for the ideal requirements engineer. IEEE Software, 13(2): 12, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chang:1999:NPS

[CC99] Carl K. Chang and Mark Christensen. A net practice for software project management. IEEE Software, 16(6):80–88, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6080abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6080.pdf.

Canfora:2009:TYC

[CCD09] Gerardo Canfora, Luigi Cerulo, and Massimiliano Di Penta. Tracking your changes: a language-independent approach. IEEE Software, 26(1):50–57, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ceri:1990:AAD

[CCRZ^{+}90] Stefano Ceri, Stefano Crespi-Reghizzi, Roberto Zicari, Gianfranco Lamperti, and Luigi A. Lavazza. Algres: An advanced database system for complex applications. IEEE Software, 7(4):68–78, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Casey:2003:LFJ

[CCS^{+}03] Dermot Casey, James Cain, Bob Schaefer, Robert Glass, Linda Rising, Jon Dehn, and Cody Powell. Letters: Finding a job; kudos; skipping the bandwagon; postwaterfall. IEEE Software, 20(5):8–10, September/ October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5008.
pdf.

Chang:1989:MRT

[CCY^{+}89] Carl K. Chang, Young-Fu Chang, Lin Yang, Ching-Roung Chou, and Jong-Jeng Chen. Modeling a real-time multitasking system in a timed PQ net. IEEE Software, 6(2):46– 51, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Card:2008:MSP

[CDD08] David N. Card, Kevin Domzalski, and Glyn Davies. Making statistics part of decision making in an engineering organization. IEEE Software, 25(3):37– 47, May/June 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Collier:1996:DPP

[CDF96] Bonnie Collier, Tom De-Marco, and Peter Fearey. A defined process for project postmortem review. IEEE Software, 13(4):65–72, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chen:2005:ESP

[CDM^{+}05] Yaofei Chen, Rose Dios, Ali Mili, Lan Wu, and Kefei Wang. An empirical study of programming language trends. IEEE Software, 22 (3):72–78, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cohen:1996:CDA

[CDPP96] David M. Cohen, Siddhartha R. Dalal, Jesse Parelius, and Gardner C. Patton. The combinatorial design approach to automatic test generation. IEEE Software, 13(5):83– 88, September 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chelf:2009:STE

[CE09] Ben Chelf and Christof Ebert. Software technology: Ensuring the integrity of embedded software with static code analysis. IEEE Software, 26 (3):96–99, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Conradi:2002:ISP

[CF02] Reidar Conradi and Alfonso Fuggetta. Improving software process improvement. IEEE Software, 19(4):92–99, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4092.
pdf;   http://www.computer.
org/software/so2002/s4092abs.
htm.

Couvillion:1991:PMU

[CFJ^{+}91] Joseph Couvillion, Roberto Freire, Ron Johnson, W. Douglas Obal, M. Akber Qureshi, Manish Rai, William Sanders, and Janet Tvedt. Performability modeling with UltraSAN. IEEE Software, 8(5):69–??, September 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carvallo:2007:DCS

[CFQ07] Juan Pablo Carvallo, Xavier Franch, and Carme Quer. Determining criteria for selecting software components: Lessons learned. IEEE Software, 24(3):84– 94, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cherubini:1989:IES

[CFTZ89] M. Alessandra Cherubini, Laura Fanti, Piero Torrigiani, and Massimo Zallocco. Integrated expertsystem builder. IEEE Software, 6(6):44–52, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cobb:1998:ODM

[CFW^{+}98] Maria A. Cobb, Harold Foley, III, Ruth Wilson, Miyi Chung, and Kevin B. Shaw. An OO database migrates to the Web. IEEE Software, 15(3):22–30, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1998/s3022abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s3022.pdf.

Cmelik:1988:DAC

[CG88] Robert F. Cmelik and Narain H. Gehani. Dimensional analysis with C++. IEEE Software, 5(3):21– 27, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Curry:2008:FSM

[CG08] Edward Curry and Paul Grace. Flexible selfmanagement using the model-view-controller pattern. IEEE Software, 25 (3):84–90, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Craigen:1994:CST

[CGR94a] Dan Craigen, Susan Gerhaet, and Ted Ralston. Case study: traffic alert and collision-avoidance system. IEEE Software, 11 (1):35–37, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Craigen:1994:CSD

[CGR94b] Dan Craigen, Susan Gerhart, and Ted Ralston. Case study: Darlington nuclear generating station. IEEE Software, 11(1):30– 32, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Campbell:1993:LGA

[CH93] Mark D. Campbell and Russ L. Holt. Lockgranularity analysis tools in SVR4/MP. IEEE Software, 10(2):66–??, March 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chang:1987:VLT

[Cha87] Shi-Kuo Chang. Visual languages: a tutorial and survey. IEEE Software, 4(1):29–39, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chang:1994:ECF

[Cha94] Carl K. Chang. From the Editor: Changing face of software engineering. IEEE Software, 11(1):4–5, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See letter and response in May 1994 issue, pp. 4, 110.

Charette:1995:MWD

[Cha95] Robert N. Charette. Manager: Are we developers liars or just fools? IEEE Software, 12(4):90– 92, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See letter and response in November 1995 issue, p. 5.

Chadwick:1996:BGP

[Cha96a] David M. Chadwick. Bookshelf: a glimpse into the promise of “profound knowledge”. IEEE Software, 13(3):109–110, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Charette:1996:LSP

[Cha96b] Robert N. Charette. Largescale project management is risk management. IEEE Software, 13(4):110–117, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Charles:1996:NCO

[Cha96c]  John Charles.  In the

news: CPM offers certificate in software estimating. IEEE Software, 13(6), November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Charles:1996:NDS

[Cha96d] John Charles. In the news:

DoD standards go global, commercial. IEEE Software, 13(3):98–99, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Charles:1997:NB

[Cha97a]  John Charles.  In the

news: In brief. IEEE Software, 14(4):106, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4100.
pdf.

Charles:1997:NIP

[Cha97b] John Charles. In the news:

Internet2 promises radical change. IEEE Software, 14 (6):108–109, 112, November/December 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6108.
pdf.

Charles:1997:NMM

[Cha97c]  John Charles.  In the

news: Managing maniacs: a commonsense approach: Interview with Ken Whitaker. IEEE Software, 14(4):102–103, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4100.
pdf.

Charles:1997:NSI

[Cha97d] John Charles. In the news: Snapshots from industry. interview: Capers Jones. women in engineering. Internet World ’97. reverse engineering challenge. Networld+Interop ’97. in brief. IEEE Software, 14(3):112– 121, May/June 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Charles:1997:NWI

[Cha97e] John Charles. In the news: Web interests tangle over DNS proposal. IEEE Software, 14(4):100, 101, 105, July/August 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4100.
pdf.

Charles:1997:NWL

[Cha97f] John Charles. In the news:

Will linking battle unravel the Web? IEEE Software, 14(5):106–107, September/ October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5106.
pdf.

Charles:1998:NIP

[Cha98a]  John Charles.  In the

news: Indecent proposal? doing business with Microsoft; formal methods conference looks for practical connections; EC issues guidelines for Internet content. IEEE Software, 15(1):113–119, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1113.
pdf.

Charles:1998:NOS

[Cha98b] John Charles. In the news:

Open source: Netscape pops the hood: NSF grant funds national network: Engineering center.

IEEE Software, 15(4):79– 82, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s4079.
                                        pdf.

Charles:1999:NRP

[Cha99a] John Charles. In the news: Refreshing productivity — one jolt at a time. IEEE Software, 16(4):109–112, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s4109.
                                        pdf.

Charles:1999:NSR

[Cha99b]  John Charles.  In the

news: Software research heads US funding priorities. IEEE Software, 16(3):107– 110, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s3107.
                                        pdf.

Charles:1999:NUC

[Cha99c]  John Charles.  In the

news: Ubiquitous computing uncorked: The international Y2K repair bill: News briefs. IEEE Software, 16(2):97–100, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2097.

pdf. Discusses Sun’s plans for Jini.

Cheriton:1984:VKS

[Che84] David R. Cheriton. V kernel: a software base for distributed systems. IEEE Software, 1(2):19–38, 40– 42, April 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cheek:1995:UFI

[Che95] Martin Cheek. Using feedback to improve software development. IEEE Software, 12(3):89–90, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cheswick:1997:QCW

[Che97] Bill Cheswick. Q&A: Confessions of a wily science guy. IEEE Software, 14 (6):110–111, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chernak:2001:VIT

[Che01] Yuri Chernak. Validating and improving testcase effectiveness. IEEE Software, 18(1):81–86, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s1081abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s1081.pdf.

Chen:2002:MGL

[Che02] Peter P. Chen. Memorial: From goto-less to structured programming: The legacy of Edsger W. Dijkstra. IEEE Software, 19(5):21, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5021.
pdf.

Chillarege:2002:MBD

[Chi02]  Ram Chillarege.  The

marriage of business dynamics and software engineering. IEEE Software, 19(6):43–49, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6043.
pdf;   http://www.computer.
org/software/so2002/s6043abs.
htm.

Colyer:2006:PC

[CHJ^{+}06] Adrian Colyer, Rob Harrop, Rod Johnson, Alexandre Vasseur, Danio Beuche, and Cédric Beust. Point/ counterpoint. IEEE Software, 23(1):72–75, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Christensen:1994:IWY

[Chr94] Mark J. Christensen. Insider: Would you do this to your customer? IEEE Software, 11(1):12, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chusho:1993:WMS

[Chu93] Takeshi Chusho. What makes software tools successful? IEEE Software, 10 (5):63–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Coplien:1998:CVS

[CHW98] James Coplien, Daniel Hoffman, and David Weiss. Commonality and variability in software engineering. IEEE Software, 15(6):37–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s6037abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s6037.pdf.

Chweh:2001:NPP

[Chw01] Crystal R. Chweh. In the news: Peer-to-peer computing. IEEE Software, 18(1):103, January/ February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1100.
pdf.

Cherryh:2002:LHG

[CJK^{+}02] James Cherryh, Kevin Jameson, Charles Krueger, Brian Farquhar, Fred Rahmanian, James J. Besemer, James A. Whittaker, Steven Atkin, and Stephen Muires. Letters: How good is software construction?; what is enough?; what’s “in it” for me? IEEE Software, 19(5):8–11, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5008.
pdf.

Clements:2005:PMS

[CJNM05] Paul C. Clements, Lawrence G. Jones, Linda M. Northrop, and John D. McGregor. Project management in a software product line organization. IEEE Software, 22(5):54–62, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chandy:1991:PP

[CK91] K. Mani Chandy and Carl Kesselman. Parallel programming in 2001. IEEE Software, 8(6):11–??, November 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carr:2000:CCC

[CK00] Donald Carr and Ronald J.

Kizior. The case for continued Cobol education. IEEE Software, 17(2):33– 36, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2033abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2033.pdf.

Clements:2002:PCB

[CK02] Paul Clements and Charles Kreuger. Point/counterpoint: Being proactive pays off; eliminating the adoption barrier. IEEE Software, 19(4):28–31, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4028.
pdf;   http://www.computer.

                                        org/software/so2002/s4028abs.
                                        htm.

Cornelissen:1995:QTH

[CKMV95] Wieger Cornelissen, Ad Klaassen, Aart Matsinger, and Gerhard Van Wee. Quality time: How to make intuitive testing more systematic. IEEE Software, 12 (5):87–89, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cantoni:1991:LEV

[CL91] Virginio Cantoni and Stefano Levialdi. Languages and environments for vision applications. IEEE Software, 8(6):34–??, November 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carney:2000:WDY

[CL00] David Carney and Fred Long. What do you mean by COTS? finally, a useful answer. IEEE Software, 17(2):83–86, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s2083abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s2083.pdf.

Constantine:2002:UCE

[CL02] Larry L. Constantine and Lucy A. D. Lockwood. Usage-centered engineering for Web applications. IEEE Software, 19(2):42– 50, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2042.
pdf;   http://www.computer.
org/software/so2002/s2042abs.
htm.

Clark:1998:CWB

[Cla98a] Robert S. Clark. Culture at work: The black death: a parallel of perilous projects. IEEE Software, 15(2):60– 61, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2060.
pdf.

Clavadetscher:1998:PUI

[Cla98b] Carl Clavadetscher. Point:

User involvement key to success. IEEE Software, 15(2):30, 32, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2030.
pdf.

Clark:2000:QEE

[Cla00] Bradford K. Clark. Quantifying the effects on effort of process improvement. IEEE Software, 17(6):65–70, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6065abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6065.pdf.

Clark:2002:MES

[Cla02] Betsy Clark. Manager: Eight secrets of software measurement. IEEE Software, 19(5):12–14, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5012.
pdf.

Ciolkowski:2003:SRS

[CLB03] Marcus Ciolkowski, Oliver Laitenberger, and Stefan Biffl. Software reviews: The state of the practice. IEEE Software, 20(6):46–51, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6046.
pdf;   http://csdl.computer.

org/dl/mags/so/2003/06/
s6046abs.htm.

Cleaveland:1988:BAG

[Cle88] J. Craig Cleaveland. Building application generators. IEEE Software, 5(4):25– 33, July 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cloyd:2001:DUC

[Clo01] Molly Hammar Cloyd. Designing user-centered Web applications in Web time. IEEE Software, 18(1): 62–69, January/February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s1062abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s1062.pdf.

Cobb:1990:ESU

[CM90] Richard H. Cobb and Harlan D. Mills. Engineering software under statistical quality control. IEEE Software, 7(6):44–54, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cohen:1996:LDB

[CM96] Daniel S. Cohen and Steve McConnell. Letters: Daily build: Solution or coping mechanism?

IEEE Software, 13(6):6–10, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cuadrado:2007:BDS

[CM07] Jesús Sánchez Cuadrado and Jesús García Molina. Building domain-specific languages for model-driven development. IEEE Software, 24(5):48–55, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cowan:2002:SET

[CMA^{+}02] Robert David Cowan, Ali Mili, Hany Ammar, Alan McKendall, Jr., Lin Yang, Dapeng Chen, and Terry Spencer. Software engineering technology watch. IEEE Software, 19(4):123–129, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4123.
pdf;   http://www.computer.
org/software/so2002/s4123abs.
htm.

Chikofsky:1992:AST

[CMH92] Elliot J. Chikofsky, David A. Martin, and Chang Hugh. Assessing the state of tools assessment. IEEE Software, 9(3):18–??, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cusumano:2003:SDW

[CMKC03] Michael Cusumano, Alan MacCormack, Chris F. Kemerer, and Bill Crandall. Software development worldwide: The state of the practice. IEEE Software, 20(6):28–34, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6028.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6028abs.htm.

Cusumano:2009:CDS

[CMKC09] Michael A. Cusumano, Alan MacCormack, Chris F. Kemerer, and William (Bill) Crandall. Critical decisions in software development: Updating the state of the practice. IEEE Software, 26(5):84–87, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chen:1992:FIC

[CN92] Minder Chen and Ronald J.

Norman. A framework for integrated CASE. IEEE Software, 9(2):18–22, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Curtis:1995:ADU

[CN95a] Bill Curtis and Jakob Nielsen. Applying discount usability engineering. IEEE Software, 12 (1):98–100, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Curtis:1995:IDI

[CN95b] Bill Curtis and Jakob Nielsen. Interface: Designing interfaces for the organization. IEEE Software, 12(6):99–100, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Campos:2007:PTW

[CN07] Pedro Campos and Nuno Jardim Nunes. Practitioner tools and workstyles for userinterface design. IEEE Software, 24(1):73–80, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cruz-Neira:1999:UIV

[CNL99] Carolina Cruz-Neira and Robyn R. Lutz. Using immersive virtual environments for certification. IEEE Software, 16(4):26– 30, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s4026abs.htm;

http://dlib.computer.
org/so/books/so1999/pdf/
s4026.pdf.

Callele:2008:ER

[CNS08] David Callele, Eric Neufeld, and Kevin Schneider. Emotional requirements. IEEE Software, 25(1):43–45, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cochran:1997:IMP

[Coc97]  Robert Cochran.  In-

sider: The mousetrap paradigm. IEEE Software, 14(5):26, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5026.
pdf.

Cockburn:2000:SPM

[Coc00] Alistair Cockburn. Selecting a project’s methodology. IEEE Software, 17(4): 64–71, July/August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2000/s4064abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4064.pdf.

Cochran:2001:CRI

[Coc01] Robert Cochran. Country report: Ireland: a software success story. IEEE Software, 18(2):87–89, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s2087.
                                        pdf.

Codd:1988:CRU

[Cod88]  E. F. Codd.  Com-

ments, with reply, on ‘The universal-relation data model for logical independence’ by M. Vardi. IEEE Software, 5(4):4, 6–??, July 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Codephirst:2009:RWA

[Cod09] Colin Codephirst. Requirements: Where have all the stencils gone? IEEE Software, 26(4):91–92, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Collier:1994:QTH

[Col94] Francois Collier. Quality time: How ISO 9001 fits into the software world; certification process. IEEE Software, 11(1):98–100, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Collins:2002:SLI

[Col02] Rosann Webb Collins. Software localization for Internet software: Issues and methods. IEEE Software, 19(2):74–80, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2074.
pdf;   http://www.computer.
org/software/so2002/s2074abs.
htm.

Cole:2005:N

[Col05] Greg Goth Bernard Cole.

In the news. IEEE Software, 22(4):108–112, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
04/s4108.pdf.

Conklin:1996:EMM

[Con96] Peter F. Conklin. Enrollment management: Managing the Alpha AXP program. IEEE Software, 13 (4):53–64, July 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Conn:2002:DSE

[Con02]  Richard Conn.  Devel-

oping software engineers at the C-130J software factory. IEEE Software, 19(5):25–29, September/ October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5025.
pdf;   http://www.computer.
org/software/so2002/s5025abs.
htm.

Cook:1995:BRW

[Coo95] Ralph Cook. Bookshelf: Real-world lessons in software metrics: review of Practical Software Metrics for Project Management and Process Improvement. IEEE Software, 12(4):109– 110, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cook:1996:BDI

[Coo96] Ralph Cook. Bookshelf: Debating issues in creative programming. IEEE Software, 13(3):111, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Coplien:1997:IPA

[Cop97] James O. Coplien. Idioms and patterns as architectural literature. IEEE Software, 14(1):36–42, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Coplien:1999:RAM

[Cop99] James O. Coplien. Reevaluating the architectural metaphor: Toward piecemeal growth. IEEE Software, 16(5):40–44, September/October 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s5040.
pdf.

Costello:1994:LRC

[Cos94] John P. Costello. Law review: Copyright and work made for hire. IEEE Software, 11(3):93–94, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cosgrove:2001:MSE

[Cos01]  John Cosgrove.  Man-

ager: Software engineering and the law. IEEE Software, 18(3):14–16, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s3014.
pdf.

Costlow:2003:NGD

[Cos03] Terry Costlow. In the news: Globalization drives changes in software careers. IEEE Software, 20(6):14–16, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6012.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6012abs.htm.

Councill:1999:QTT

[Cou99a]  Bill Councill.  Quality

time: Third-party testing and stirrings of the new software engineering. IEEE Software, 16(6):76–79, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6076.
pdf.

Councill:1999:TPT

[Cou99b] William T. Councill. Thirdparty testing and the quality of software components. IEEE Software, 16(4):55– 57, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4055.
pdf.

Cox:1984:MOP

[Cox84]  Brad J. Cox.  Mes-

sage/object programming: an evolutionary change in programming technology.

IEEE Software, 1(1):50–61, January 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cox:1990:PSI

[Cox90] Brad J. Cox. Planning the software industrial revolution. IEEE Software, 7 (6):25–33, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cox:1997:MSE

[Cox97] Brad Cox. Manager: Superdistribution and the economics of bits. IEEE Software, 14(1):22–24, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Coyle:2000:LIC

[Coy00a] Frank P. Coyle. Legacy integration: Changing perspectives. IEEE Software, 17(2):37–41, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2037abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2037.pdf.

Coyle:2000:PDC

[Coy00b] Frank P. Coyle. Point:

Does Cobol exist? IEEE Software, 17(2):22, 24, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2022.
pdf.

Callahan:2002:EEI

[CP02] Dale Callahan and Bob Pedigo. Educating experienced IT professionals by addressing industry’s needs. IEEE Software, 19(5):57–62, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5057.
pdf;   http://www.computer.
org/software/so2002/s5057abs.
htm.

Cusick:2006:PME

[CP06] James Cusick and Alpana Prasad. A practical management and engineering approach to offshore collaboration. IEEE Software, 23(5):20–29, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chikofsky:1988:CRE

[CR88] Elliot J. Chikofsky and Burt L. Rubenstein. Case: Reliability engineering for information systems. IEEE Software, 5(2):11–16, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cybulski:1992:HBS

[CR92] Jacob L. Cybulski and Karl Reed. A hypertext-based software-engineering environment. IEEE Software, 9(2):62–68, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cao:2008:ARE

[CR08] Lan Cao and Balasubramaniam Ramesh. Agile requirements engineering practices: An empirical study. IEEE Software, 25(1):60–67, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Crispin:2006:DSQ

[Cri06]  Lisa Crispin.  Driving

software quality: How test-driven development impacts software quality. IEEE Software, 23(6):70– 71, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cao:2009:DSM

[CRR09] Lan Cao, Balasubramaniam Ramesh, and Matti Rossi. Are domain-specific models easier to maintain than UML models? IEEE Software, 26(4):19– 21, July/August 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Choi:1990:ERD

[CS90] Song C. Choi and Walt Scacchi. Extracting and restructuring the design of large systems. IEEE Software, 7(1):66–71, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chmura:1996:TFU

[CS96] Alan Chmura and David Sharon. Tools fair untangling the Web. IEEE Software, 13(5):61–72, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Conrow:1997:IRM

[CS97] Edmund H. Conrow and Patricia S. Shishido. Implementing risk management on software intensive projects. IEEE Software, 14 (3):83–89, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Caplat:2005:MMU

[CS05] Guy Caplat and Jean-Louis Sourrouille. Model mapping using formalism extensions. IEEE Software, 22(2):44–51, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Clements:2009:GAS

[CS09] Paul Clements and Mary Shaw. “The Golden Age of Software Architecture” revisited. IEEE Software, 26(4):70–72, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chulani:2007:MBM

[CSHA07] Sunita Chulani, P. Santhanam, Brent Hodges, and Kelley Blacksten Anders. Metrics-based management of software product portfolios. IEEE Software, 24(2): 66–72, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Church:1999:LTD

[CSR^{+}99] Vic Church, Philip Sargent, Andrew Raybould, Robert Glass, Alan Kaminsky, Larry Graham, and Larry L. Constantine. Letters: Theories on design; joining the information age; filling in the gaps. IEEE Software, 16 (5):10, 13–14, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s5010.
                                        pdf.

Curtis:2008:CQP

[CSR^{+}08] Bill Curtis, Girish V. Seshagiri, Donald Reifer, Iraj Hirmanpour, and Gargi Keeni. The case for quantitative process management. IEEE Software, 25 (3):24–28, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2008/
                                        03/mso2008030024.pdf.

Ceschi:2005:PMP

[CSSP05] Martina Ceschi, Alberto Sillitti, Giancarlo Succi, and Stefano De Panfilis. Project management in plan-based and agile companies. IEEE Software, 22 (3):21–27, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chang:1990:JSR

[CT90] Carl K. Chang and George B. Trubow. Joint software research between industry and academia. IEEE Software, 7(6):71–77, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carver:1991:RTC

[CT91] Richard H. Carver and Kuo-Chung Tai. Replay and testing for concurrent programs. IEEE Software, 8(2):66–74, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Callaghan:1993:CAT

[CT93] James Callaghan and Terry Turner. Coordinated approach to telecommunications management. IEEE Software, 10(6):28–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cheston:2002:ISE

[CT02] Grant A. Cheston and Jean-Paul Tremblay. Integrating software engineering in introductory computing courses. IEEE Software, 19(5):64–71, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5064.
pdf;   http://www.computer.
org/software/so2002/s5064abs.
htm.

Caffery:2007:AUA

[CTC07] Fergal Mc Caffery, Philip S.

Taylor, and Gerry Coleman. Adept: a unified assessment method for small software companies. IEEE Software, 24(1):24–31, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chavez:1998:SCL

[CTW98] Andrea Chávez, Catherine Tornabene, and Gio Wiederhold. Software component licensing: a primer. IEEE Software, 15(5):47–52, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5047.
pdf;   http://www.computer.
org/computer/co1998/s5047abs.
htm.

Cukic:2005:GEI

[Cuk05a] Bojan Cukic. Guest Editor’s introduction: The promise of public software engineering data repositories. IEEE Software, 22(6):20–22, November/ December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
06/s6020.pdf.

Cukic:2005:VAS

[Cuk05b] Bojan Cukic. The virtues of assessing software reliability early. IEEE Software, 22 (3):50–53, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cureton:1988:FUC

[Cur88] Bill Cureton. Future of Unix in the CASE renaissance. IEEE Software, 5(2):18–22, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Curtis:1995:PDW

[Cur95] Bill Curtis. Proceedings:

Debating WIMP at Rocky Mountain CHI. SEPG spotlights maturing software industry. in quotes. calendar. IEEE Software, 12(5): 100–102, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Curtis:1998:SWC

[Cur98]  Bill Curtis.  Soapbox:

Which comes first, the organization or its processes? IEEE Software, 15(6):10–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6010.
pdf.

Curtis:2000:GEI

[Cur00] Bill Curtis. Guest Editor’s introduction: The global pursuit of process maturity. IEEE Software, 17(4):76– 78, July/August 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4076.
pdf.

Curtis:2001:LOY

[Cur01] Tex Curtis. Loyal opposition: So you wanna be a cowboy. IEEE Software, 18(2):112, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2112.
pdf.

Cusumano:1989:SFH

[Cus89] Michael A. Cusumano. Software factory: a historical interpretation. IEEE Software, 6(2):10–73, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cusick:2003:HWS

[Cus03] James Cusick. How the work of software professionals changes everything. IEEE Software, 20(3):92– 97, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3092.

pdf;   http://www.computer.
org/software/so2003/s3092abs.
htm.

Carlton:1988:DPS

[CV88] Mike Carlton and Peter Van Roy. Distributed Prolog system with and parallelism. IEEE Software, 5(1):43–51, January 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Curtis:1984:GND

[CW84] Ronald Curtis and Larry Wittie. Global naming in distributed systems. IEEE Software, 1(3):76–80, July 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chin:1989:STD

[CW89] Tahn Joo Chin and Kai Yuen Wang. Software technology development in Singapore. IEEE Software, 6(2):61– 65, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cheng:2000:GEI

[CW00] Betty H. C. Cheng and David M. Weiss. Guest Editors’ introduction: Requirements engineering — integrating technology. IEEE Software, 17(3):18–20, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3018.
pdf.

Card:2001:BSO

[CZ01] David Card and David Zubrow. Benchmarking software organizations. IEEE Software, 18(5):16–17, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5016abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5016.pdf.

Desouza:2006:MKG

[DAB06] Kevin C. Desouza, Yukika Awazu, and Peter Baloh. Managing knowledge in global software development efforts: Issues and practices. IEEE Software, 23(5):30–37, September/ October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

deGoyeneche:1999:LKM

[dAdS99] Juan-Mariano de Goyeneche, Elena Apolinario, and Fernández de Sousa. Loadable kernel modules. IEEE Software, 16(1):65–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1065.
pdf.

Dorr:2008:IRE

[DAEE08] Jörg Dörr, Sebastian Adam, Michael Eisenbarth, and Michael Ehresmann. Implementing requirements engineering processes: Using cooperative self-assessment and improvement. IEEE Software, 25(3):71–77, May/ June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1995:EOW

[Dak95a] Karl J. Dakin. Employment options: What’s a fair deal? IEEE Software, 12 (1):103–104, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1995:SLC

[Dak95b]  Karl J. Dakin.  Soft

law: Capitalizing on multiple market opportunities. IEEE Software, 12(5):84– 86, September 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1995:SLD

[Dak95c] Karl J. Dakin. Soft law:

Do you know what your license allows? IEEE Software, 12(3):82–83, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1995:SLE

[Dak95d] Karl J. Dakin. Soft law: Establishing a fair price for software. IEEE Software, 12(6):105–106, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1995:WNT

[Dak95e] Karl J. Dakin. What’s new? tracing the origin of species. IEEE Software, 12(2):99– 100, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1996:SLK

[Dak96a] Karl Dakin. Soft law — keys to engineering a workable contract. IEEE Software, 13(1):99–100, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1996:SLD

[Dak96b] Karl Dakin. Soft law: Do we need an implied warranty of competence? IEEE Software, 13(6):97– 98, November 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1996:SLT

[Dak96c] Karl Dakin. Soft law: The transaction-based pricing trend. IEEE Software, 13 (3):96–97, May 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1996:WPI

[Dak96d] Karl Dakin. Warranties:

Promising the improbable. IEEE Software, 13(2):105– 106, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1996:SLL

[Dak96e] Karl J. Dakin. Soft law:

Legal implications of Bloun schedules. IEEE Software, 13(5):101–102, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1996:SDM

[Dak96f] Karl J. Dakin. Softlaw: Are developers morally challenged? IEEE Software, 13(4):20, 22, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1997:SSE

[Dak97a] Karl Dakin. Softlaw: Software “engineer”? time will tell. IEEE Software, 14(3): 105–106, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dakin:1997:SWI

[Dak97b] Karl Dakin. Softlaw: What if there were no software piracy? IEEE Software, 14(1):20–21, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Damian:2007:SGR

[Dam07] Daniela Damian. Stakeholders in global requirements engineering: Lessons learned from practice. IEEE Software, 24(2): 21–27, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Daniels:2002:MSP

[Dan02]  John Daniels.  Mod-

eling with a sense of purpose. IEEE Software, 19(1):8–10, January/ February 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1008abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1008.pdf.

Dance:2003:ER

[Dan03]  John Dance.  Extreme

reading. IEEE Software, 20(3):108–111, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3108.
pdf.

Daneva:2004:ERE

[Dan04]  Maya Daneva.  ERP

requirements engineering practice: Lessons learned. IEEE Software, 21(2):26– 33, March/April 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Daskalantonakis:1994:AHS

[Das94] Michael Daskalantonakis.

Achieving higher SEI levels. IEEE Software, 11(4): 17–24, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dybaa:2007:THB

[DAS^{+}07] Tore Dybå, Erik Arisholm, Dag I. K. Sjøberg, Jo E. Hannay, and Forrest Shull. Are two heads better than one? on the effectiveness of pair programming. IEEE Software, 24(6):12– 15, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1985:LPP

[Dav85] Ruth E. Davis. Logic programming and prolog: a tutorial. IEEE Software, 2 (5):53–62, September 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davidson:1986:PHL

[Dav86] Scott Davidson. Progress in high-level microprogramming. IEEE Software, 3(4):18–26, August 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1995:ERE

[Dav95a] Alan Davis. From the Editor: Report from the editorial trenches. IEEE Software, 12(4):4, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1995:ETS

[Dav95b] Alan Davis. From the Editor: Tracing: a simple necessity neglected. Winston Royce dies. correction. IEEE Software, 12(5):6, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1995:EFB

[Dav95c] Alan M. Davis. From the Editor: Filling big shoes: a proud legacy. IEEE Software, 12(1):4–13, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1995:ETB

[Dav95d] Alan M. Davis. From the Editor: Tomorrow’s blacksmiths? IEEE Software, 12 (3):4, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1995:ID

[Dav95e] Alan M. Davis. Information for decisionmakers. IEEE Software, 12(2):4, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1996:EST

[Dav96a] Al Davis. Eras of software technology transfer; McConnell named to editorial board; 1995 best article award. IEEE Software, 13 (2):4–7, March 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1996:EIF

[Dav96b] Al Davis. From the Editor:

It feels like Déja vu all over again. IEEE Software, 13 (4):4, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1996:EWP

[Dav96c] Al Davis. From the Editor: What’s practical.

IEEE Software, 13(5):4–5, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1996:AEO

[Dav96d] Alan M. Davis. Art or engineering, one more time. IEEE Software, 13(6):4–5, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1996:FAB

[Dav96e] Alan M. Davis. Farewell to Angela Burgess. IEEE Software, 13(3):5, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1996:PHT

[Dav96f] Alan M. Davis. Practitioner, heal thyself. IEEE Software, 13(3):4–5, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1997:EBS

[Dav97a] Al Davis. From the Editor: Between Scylla and Charybdis. IEEE Software, 14(5):4, 6, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5004.
pdf.

Davis:1997:EMM

[Dav97b] Al Davis. From the Editor: Making a mark on the world. IEEE Software, 14(6):4, 6, November/December 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6004.
pdf.

Davis:1997:EWW

[Dav97c] Al Davis. From the Editor:

Words of wisdom. openings, closings. errata. IEEE Software, 14(3):4, 6, 9, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1997:TOR

[Dav97d]  Al Davis.  Thanks to

our 1997 referees. IEEE Software, 14(6):125–126, November/December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s6125.pdf.

Davis:1997:ESF

[Dav97e] Alan M. Davis. From the Editor: Software’s future is in the cards. IEEE Software, 14(4):4–6, 8, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4004.
pdf.

Davis:1998:ERM

[Dav98a]  Al Davis.  From the

Editor: Recovering from method abuse. IEEE Software, 15(1):4, 6–7, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1004.
pdf.

Davis:1998:EMW

[Dav98b] Alan M. Davis. From the Editor: More words of wisdom. IEEE Software, 15(3): 6–8, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3006.
pdf.

Davis:1998:EPF

[Dav98c] Alan M. Davis. From the Editor: Predictions and farewells. IEEE Software, 15(4):6–9, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4006.
pdf.

Davis:1998:EHR

[Dav98d] Alan M. Davis. From the Editor; the harmony in rechoirments. IEEE Software, 15(2):6, 8, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2006.
pdf.

Davis:1998:IGT

[Dav98e] Alan M. Davis. Interview:

a golden thread in software’s tapestry. IEEE Software, 15(6):18–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6018.
pdf.

Davis:2000:SSC

[Dav00] Alan M. Davis. Soapbox: The software company machine. IEEE Software, 17(2):14–15, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2014.
pdf.

Davis:2003:SP

[Dav03] Alan M. Davis. System phenotypes. IEEE Software, 20(4):54–56, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s4054.
                                        pdf.

Dawkins:1995:FCC

[Daw95a] Spencer Dawkins. Focus on careful coding. IEEE Software, 12(2):119–121, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dawkins:1995:HSY

[Daw95b] Spencer Dawkins. How to supercharge your RISC engine. IEEE Software, 12 (2):114–115, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). Review of Kevin Dowd’s High Performance Computing.

Dribin:1996:LDD

[DB96] Lawrence S. Dribin and Maarten Boasson. Letters: Defining design terminology. IEEE Software, 13(3): 8, 10, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Deck:1997:LTR

[DB97] Michael Deck and Boris Beizer. Letters: a truce? reentering cleanroom. IEEE Software, 14(5):8, 10, 12, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5008.
pdf.

Dupuis:1999:GSE

[DB99] Robert Dupuis and Pierre Bourque. Guide to the software engineering body of knowledge project. IEEE Software, 16(4):61, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4061.
pdf.

Duggan:2004:TAO

[DBL04] Jim Duggan, Jason Byrne, and Gerard J. Lyons. A task allocation optimizer for software construction. IEEE Software, 21(3):76– 82, May/June 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dingsoyr:2009:VEW

[DBS09] Torgeir Dingsøyr, Finn Olav Bjørnson, and Forrest Shull. Voice of evidence: What do we know about knowledge management? practical implications for software engineering. IEEE Software, 26(3):100–103, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dorenbos:2004:ILL

[DC04] David Dorenbos and Annie Combelles. Introduction: Lessons learned around the world: Key success factors to enable process change. IEEE Software, 21(4):20– 21, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        04/s4020.pdf.

Durbin:2002:LAS

[DCF^{+}02] John P. Durbin, Reidar Conradi, Alfonso Fuggetta, Paul Freedman, Ferdinand Wagner, and Peter Wolstenholme. Letters: Not another SPI framework; pair programming à la Dijkstra; a radical paradigm shift? IEEE Software, 19(6):8–9, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s6008.
                                        pdf.

Denne:2004:IFM

[DCH04] Mark Denne and Jane Cleland-Huang. The incremental funding method:

Data-driven software development. IEEE Software, 21 (3):39–47, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dedene:1995:RSR

[DD95] Guido Dedene and Jean-Pierre De Vreese. Realities of off-shore reengineering. IEEE Software, 12(1):35– 45, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dybaa:2009:WDW

[DD09] Tore Dybå and Torgeir Dingsoyr. What do we know about agile software development? IEEE Software, 26(5):6–9, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Damm:1986:AAF

[DDMS86] Werner Damm, Gert Doehmen, Klaus Merkel, and Mathilde Sichelschmidt. AADL/S* approach to firmware design verification. IEEE Software, 3(4):27–37, August 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dorner:2009:EUB

[DDPW09] Christian Dorner, Sebastian Draxler, Volkmar Pipek, and Volker Wulf. End users at the bazaar: Designing next-generation enterprise resource planning systems. IEEE Software, 26(5):45–51, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DePasquale:1997:LWB

[DDS^{+}97] Mike DePasquale, Karl Dakin, Bryan Schmidt, Tom Adams, Al Davis, Ken Porter, and Denis C. Meredith. Letters: Who’s to blame for blown schedules? art of creating a discipline. analogy breakdown. why build software? A key question. putting code in its place. IEEE Software, 14 (3):11–14, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Deck:2001:MPD

[Dec01] Michael Deck. Managing process diversity while improving your practices. IEEE Software, 18(3):21– 27, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s3021abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s3021.pdf.

Dedeke:2009:LOL

[Ded09] Adenekan (Nick) Dedeke.

Loyal opposition: Is Linux better than Windows software? IEEE Software, 26(3):104, 103, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeMarco:1990:MDS

[DeM90] Tom DeMarco. Making a difference in the schools. IEEE Software, 7(6):78–82, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeMarco:1995:MWL

[DeM95a] Tom DeMarco. Manager:

What ‘lean and mean’ really means. IEEE Software, 12(6):101–102, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeMarco:1995:TTR

[DeM95b] Tom DeMarco. Thoughtful treatise on reusable objects: review of Design Patterns: Elements of Reusable Object-Oriented Software. IEEE Software, 12(3):101–102, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeMarco:1999:CIA

[DeM99] Tom DeMarco. Counterpoint: It ain’t broke, so don’t fix it. IEEE Software, 16(6):67, 69, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6066.
pdf.

DeMarco:2009:VSE

[DeM09] Tom DeMarco. Viewpoints: Software engineering: An idea whose time has come and gone? IEEE Software, 26(4):96, 95, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Denney:1991:TCG

[Den91] Richard Denney. Test-case generation from Prologbased specifications. IEEE Software, 8(2):49–57, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Desilets:2008:TMS

[Dés08] Alain Désilets. Tell me a story. IEEE Software, 25(2):14–15, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Damian:2000:UDC

[DESG00] Daniela E. Herlea Damian, Armin Eberlein, Mildred L. G. Shaw, and Brian R. Gaines. Using different communication media in requirements negotiation. IEEE Software, 17(3):28– 36, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3028abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s3028.pdf.

Deutsch:1988:FRT

[Deu88] Michael S. Deutsch. Focusing real-time systems analysis on user operations. IEEE Software, 5(5):39–50, September 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Duarte:2000:BJA

[DFL00] Carlos H. C. Duarte, Martin Fogarty, and Robert C. Larrabee. Bookshelf: Java application frameworks use case driven object: Modeling with UML: a practical approach: Chaos and complexity in software, challenging the industry and the new science. IEEE Software, 17 (5):120–123, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5120.
pdf.

Donnellan:2005:IOS

[DFLS05] Brian Donnellan, Brian Fitzgerald, Brian Lake, and John Sturdy. Implementing an Open Source knowledge base. IEEE Software, 22(6):92–95, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Debou:1993:SBT

[DFS93] Christophe Debou, Norbert Fuchs, and Heinz Saria. Selling believable technology. IEEE Software, 10 (6):22–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Delisle:1990:FSO

[DG90] Norman Delisle and David Garlan. A formal specification of an oscilloscope. IEEE Software, 7(5):29–36, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dohmke:2007:TDD

[DG07] Thomas Dohmke and Henrik Gollee. Test-driven development of a PID controller. IEEE Software, 24 (3):44–50, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeRoure:2009:SDE

[DG09a] David De Roure and Carole Goble. Software design for empowering scientists. IEEE Software, 26(1): 88–95, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Demirors:2009:CAF

[DG09b] Onur Demirors and Cigdem Gencel. Conceptual association of functional size measurement methods. IEEE Software, 26 (3):71–78, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Duenas:2007:AEC

[DGC^{+}07] Juan C. Dueñas, Hugo A. Parada G., Félix Cuadrado, Manuel Santillán, and José L. Ruiz. Apache and Eclipse: Comparing open source project incubators. IEEE Software, 24(6):90–98, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Diaz-Herrera:1993:ISS

[DH93] Jorge L. Diaz-Herrera. The importance of static structures in software construction. IEEE Software, 10(3): 75–87, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dugan:2001:RFT

[DHK^{+}01] Joanne Bechta Dugan, Les Hatton, Karama Kanoun, Jean-Claude Laprie, and Mladen A. Vouk. Roundtable: Fault tolerance. IEEE Software, 18(4):54–57, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4054.
pdf;   http://dlib.computer.
org/so/books/so2001/s4054abs.
htm.

Diaz-Herrera:2002:CPR

[DHMR02] Jorge Díaz-Herrera, Mike Murphy, and Dawn Ramsey. A collaborative program to retrain Lockheed Martin aero engineers. IEEE Software, 19(5):30–34, September/ October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5030.
pdf;   http://www.computer.
org/software/so2002/s5030abs.
htm.

DiGiacomo:2005:MLL

[Di 05] Mariella Di Giacomo. MySQL: Lessons learned on a digital library. IEEE Software, 22 (3):10–13, May/June 2005.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
03/s3010.pdf.

Dickson:1995:LW

[Dic95] Scott Dickson. Letters: Not to worry. IEEE Software, 12(5):14, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dick:2005:DT

[Dic05] Jeremy Dick. Design traceability. IEEE Software, 22(6):14–16, November/ December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Digre:1998:BOC

[Dig98] Tom Digre. Business object component architecture. IEEE Software, 15(5):60–68, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5060.
pdf;   http://www.computer.
org/computer/co1998/s5060abs.
htm.

Dion:1993:PIC

[Dio93] Raymond Dion. Process improvement and the corporate balance sheet. IEEE Software, 10(4):28–??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dozier:2005:MTE

[DJ05] Christopher Dozier and Peter Jackson. Mining text for expert witnesses. IEEE Software, 22(3):94– 100, May/June 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Deissenboeck:2008:TSC

[DJH^{+}08] Florian Deissenboeck, Elmar Juergens, Benjamin Hummel, Stefan Wagner, Benedikt Mas y Parareda, and Markus Pizka. Tool support for continuous quality control. IEEE Software, 25(5):60–67, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dieste:2004:HHE

[DJM04] Oscar Dieste, Natalia Juristo, and Ana M. Moreno. How higher-education systems influence software engineering degree programs. IEEE Software, 21(4):78– 85, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dieste:2008:UCW

[DJS08] Oscar Dieste, Natalia Juristo, and Forrest Shull. Understanding the customer: What do we know about requirements elicitation? IEEE Software, 25(2):11–13, March/ April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2008/extras/mso2008020011x1.
html.

Dart:1997:DAS

[DJSS97] Philip Dart, Lorraine Johnston, Cameron Schmidt, and Liz Sonenberg. Developing an accredited software engineering program. IEEE Software, 14(6):66–70, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6066.
pdf;   http://dlib.computer.
org/so/books/so1997/pdf/
s6066abs.htm.

Darcy:2005:OMP

[DK05] David P. Darcy and Chris F.

Kemerer. OO metrics in practice. IEEE Software, 22(6):17–19, November/ December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dybaa:2005:EBS

[DKJ05] Tore Dybå, Barbara A. Kitchenham, and Magne Jorgensen. Evidence-based software engineering for practitioners. IEEE Software, 22(1):58–65, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dongarra:1990:SRG

[DKKK90] Jack Dongarra, Alan H. Karp, Ken Kennedy, and David Kuck. Special report: 1989 Gordon Bell Prize. IEEE Software, 7(3):100–104, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dongarra:1991:SRG

[DKMS91] Jack Dongarra, Alan Karp, Ken Miura, and Horst Simon. Special report: 1990 Gordon Bell Prize. IEEE Software, 8(3):92–98, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeMarco:1998:PEH

[DL98] Tom DeMarco and Timothy Lister. Peopleware 2 excerpt: Human capital. IEEE Software, 15(6):103–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s6103abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s6103.pdf.

DeMarco:2003:RMD

[DL03] Tom DeMarco and Tim Lister. Risk management during requirements. IEEE Software, 20(5):99–101, September/October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/dl/mags/so/2003/05/
s5099.pdf.

deLemos:2006:FTA

[dLdCGR06] Rogério de Lemos, Paulo Asterio de Castro Guerra, and Cecília Mary Fischer Rubira. A faulttolerant architectural approach for dependable systems. IEEE Software, 23(2):80–87, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ducasse:2007:SFE

[DLR07] Stéphane Ducasse, Adrian Lienhard, and Lukas Renggli. Seaside: a flexible environment for building dynamic Web applications. IEEE Software, 24(5):56–63, September/ October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dangle:2005:SPI

[DLSZ05] Kathleen Coleman Dangle, Patricia Larsen, Michele Shaw, and Marvin V. Zelkowitz. Software process improvement in small organizations: a case study. IEEE Software, 22(6):68– 75, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dutta:1999:SEE

[DLV99] Soumitra Dutta, Michael Lee, and Luk Van Wassenhove. Software engineering in Europe: a study of best practices. IEEE Software, 16(3):82–81, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s3082abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s3082.pdf.

Diederich:1987:EPS

[DM87] Jim Diederich and Jack Milton. Experimental prototyping in Smalltalk. IEEE Software, 4(3):50– 64, May 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1993:OOD

[DM93] John Davis and Tom Morgan. Object-oriented development at Brooklyn Union Gas. IEEE Software, 10 (1):67–74, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeMarco:1996:MLS

[DM96] Tom DeMarco and Ann Miller. Managing large software projects. IEEE Software, 13(4):24–27, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Damian:2006:GEI

[DM06] Daniela Damian and Deependra Moitra. Guest Editors’ introduction: Global software development: How far have we come? IEEE Software, 23(5):17–19, September/October 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
05/s5017.pdf.

Davis:1998:LWR

[DMN^{+}98] Alan Davis, Steve Mc-Connell, Patrick Naughton, Dale Strok, and Michale G. Morris. Letters: Who is the real Patrick Naughton? IEEE Software, 15(2):11, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2010.
pdf.

Dawson:1997:ISE

[DN97] Ray Dawson and Ron Newsham. Introducing software engineers to the real world. IEEE Software, 14(6):37–43, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6037.
pdf;   http://dlib.computer.
org/so/books/so1997/pdf/
s6037abs.htm.

Drobka:2004:PXF

[DNR04] Jerry Drobka, David Noftz, and Rekha Raghu. Piloting XP on four mission-critical projects. IEEE Software, 21(6):70–75, November/ December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Deiss:2006:IDT

[DNS^{+}06] Thomas Deiß, Andreas J. Nyberg, Stephan Schulz, Risto Teittinen, and Colin Willcock. Industrial deployment of the TTCN-3 testing technology. IEEE Software, 23(4):48–54, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dominus:1998:PJW

[Dom98] Mark-Jason Dominus. Perl:

Not just for Web programming. IEEE Software, 15(1):69–74, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1069abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s1069.pdf.

Donahue:2001:UBL

[Don01] George M. Donahue. Usability and the bottom line. IEEE Software, 18(1): 31–37, January/February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s1031abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s1031.pdf.

Donzelli:2006:DSS

[Don06] Paolo Donzelli. A decision support system for software project management. IEEE Software, 23(4):67– 75, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dorchester:1999:SWL

[Dor99a] Dave Dorchester. Soapbox: Why license software engineers? IEEE Software, 16(2):101–103, March/April 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s2101.
                                        pdf.

Dorfman:1999:CWC

[Dor99b] Merlin Dorfman. Culture at work: Commercial vs. aerospace worlds: Comparing software engineering cultures. IEEE Software, 16(6):97–100, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s6097.
                                        pdf.

Douglass:1985:QAP

[Dou85] Robert J. Douglass. Qualitative assessment of parallelism in expert systems. IEEE Software, 2(3):70– 81, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dowson:1987:IPS

[Dow87] Mark Dowson. Integrated project support with IStar. IEEE Software, 4(6):6–15, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Davis:1994:GEI

[DP94a] Alan M. Davis and Hsia Pei. Guest Editors’ introduction: Giving voice to requirements engineering. IEEE Software, 11(2):12– 15, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See letter and response in March 1994 issue, pp. 6, 8–9.

Diaz:1994:EOU

[DP94b] Oscar Diaz and Norman W. Paton. Extending ODBMSs using metaclasses. IEEE Software, 11 (3):40–47, May 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dubois:2003:GEI

[DP03] Eric Dubois and Klaus Pohl. Guest Editors’ introduction: RE 02 — a major step toward a mature requirements engineering community. IEEE Software, 20(1):14–15, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1014.
pdf.

Dromey:1996:CC

[Dro96] R. Geoff Dromey. Cornering the Chimera. IEEE Software, 13(1):33–43, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dromey:2006:CNS

[Dro06] R. Geoff Dromey. Climbing over the “no silver bullet” brick wall. IEEE Software, 23(2):120, 118–119, March/April 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Decker:2007:WBS

[DRR^{+}07] Björn Decker, Eric Ras, Jörg Rech, Pascal Jaubert, and Marco Rieth. Wikibased stakeholder participation in requirements engineering. IEEE Software, 24(2):28–35, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dunsmore:2003:PCI

[DRW03] Alastair Dunsmore, Marc Roper, and Murray Wood. Practical code inspection techniques for objectoriented systems: An experimental comparison. IEEE Software, 20(4):21– 29, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4021.
pdf;   http://www.computer.
org/software/so2003/s4021abs.
htm.

Diaz:1997:HSP

[DS97] Michael Diaz and Joseph Sligo. How software process improvement helped Motorola. IEEE Software, 14(5):75–81, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5075.
pdf;   http://www.computer.
org/software/so1997/s5075abs.
htm.

Denger:2007:PAQ

[DS07] Christian Denger and Forrest Shull. A practical approach for quality-driven inspections. IEEE Software, 24(2):79–86, March/ April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

deLemos:1995:ASR

[dSA95] Rogério de Lemos, Amer Saeed, and Tom Anderson. Analyzing safety requirements for process-control systems. IEEE Software, 12 (3):42–53, May 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Durham:1999:BOO

[DSB99] George B. Durham, Diomidis Spinelis, and Mariá Bieliková. Bookshelf: Object-oriented software design and construction with C++: Decrypted secrets: Methods and maxims of cryptology: Software engineering: Theory and practice. IEEE Software, 16(4):114–117, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4114.
pdf.

Dance:1999:BRC

[DSBM99] John Dance, Rudy Schild, Mária Bieliková, and Deependra Moitra. Bookshelf: Ruminations on C++; a decade of programming insight and experience; migrating from Pascal to C++; designing the user interface; strategies for effective human-computer interaction; practical software reuse; strategies for introducing reuse concepts in your organization. IEEE Software, 16(2):91– 96, March/April 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2091.
pdf.

Dimitrov:2002:UBP

[DSD02] Evgeni Dimitrov, Andreas Schmietendorf, and Reiner Dumke. UML-based performance engineering possibilities and techniques. IEEE Software, 19(1):74–83, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1074abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1074.pdf.

Denno:2003:MDI

[DSLB03] Peter Denno, Michelle Potts Steves, Don Libes, and Edward J. Barkmeyer. Modeldriven integration using existing models. IEEE Software, 20(5):59–63, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5059.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5059abs.htm.

Dongarra:1985:ADD

[DSS85] J. J. Dongarra, B. T.

Smith, and D. Sorensen. Algorithm design for different computer architectures. IEEE Software, 2(4): 79–80, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

                                        tronic). URL http://www.
                                        netlib.org/utk/people/
                                        JackDongarra/PAPERS/Algorithm-
                                        Design-for-Different-
                                        Computer-Architectures.
                                        pdf.

deSouza:2009:GEI

[dSSS^{+}09] Cleidson R. B. de Souza, Helen Sharp, Janice Singer, Li-Te Cheng, and Gina Venolia. Guest Editors’ introduction: Cooperative and human aspects of software engineering. IEEE Software, 26(6):17– 19, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dori:1994:SMI

[DT94] Dov Dori and Erez Tatcher. Selective multiple inheritance. IEEE Software, 11 (3):77–85, May 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dogru:2003:PMC

[DT03] Ali H. Dogru and Murat M. Tanik. A process model for componentoriented software engineering. IEEE Software, 20(2):34–41, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s2034.

pdf;   http://www.computer.
org/software/so2003/s2034abs.
htm.

Duarte:2002:BCD

[Dua02] Carlos H. C. Duarte.

Brazil: Cooperative development of a software industry. IEEE Software, 19(3):84–87, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3084.
pdf;   http://www.computer.
org/software/so2002/s3084abs.
htm.

Duggan:1998:BFO

[Dug98]  Jim Duggan.  Book-

shelf: The future is objectoriented … for now. IEEE Software, 15(4):87– 88, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4087.
pdf.

Duke:1989:VFM

[Duk89] Eugene L. Duke. V&V of flight and mission-critical software. IEEE Software, 6(3):39–45, May 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dumais:1995:IFW

[Dum95]  Susan Dumais.  Inter-

face: Finding what you want: New tools and tricks. IEEE Software, 12(5):79– 81, September 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dunham:1989:VND

[Dun89] Janet R. Dunham. V&V in the next decade. IEEE Software, 6(3):47–53, May 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dutton:1993:CAP

[Dut93] James E. Dutton. Commonsense approach to process modeling. IEEE Software, 10(4):56–??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Diaz:1989:SDO

[DV89] Michel Diaz and Chris Vissers. SEDOS: Designing open distributed systems. IEEE Software, 6(6):24–33, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dittrich:2009:ECS

[DVG09] Yvonne Dittrich, Sebastien Vaucouleur, and Stephen Giff. ERP customization as software engineering:

Knowledge sharing and cooperation. IEEE Software, 26(6):41–47, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Du:1990:SBI

[DW90] Weichang Du and William W.

Wadge. A 3D spreadsheet based on intensional logic. IEEE Software, 7(3):78– 89, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dang:2000:BID

[DW00] Bruce Dang and Nathan Ward. Bookshelf: Internetworking: Devices powering the Internet. Interconnections, Second Edition: Bridges, Routers, Switches, and Internetworking Protocols; The Team Handbook, Second Edition. IEEE Software, 17(3):99–100, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3099.
pdf.

Dasgupta:1986:ASF

[DWH86] Subrata Dasgupta, Philip A.

Wilsey, and Juha Heinanen. Axiomatic specifications in firmware development systems. IEEE Software, 3(4):49–58, August 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dik:1993:FQP

[DY93] Lun Lee Dik and Yung Leung Yim. Fast query processing in deductive databases. IEEE Software, 10(6):66–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dyadkin:1995:MPN

[Dya95] Lev J. Dyadkin. Multibox parsers. no more handwritten lexical analyzers. IEEE Software, 12(5):61– 67, September 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dybaa:2000:ISS

[Dyb00] Tore Dybå. Improvisation in small software organizations. IEEE Software, 17(5):82–87, September/October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so/s5082abs.htm;   http:
                                        //dlib.computer.org/so/
                                        books/so2000/pdf/s5082.
                                        pdf.

Donzelli:2005:ECC

[DZB^{+}05] Paolo Donzelli, Marvin Zelkowitz, Victor Basili, Dan Allard, and Kenneth N. Meyer. Evaluating COTS component dependability in context. IEEE Software, 22(4):46– 53, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ehrlich:1994:GWT

[E^{+}94] Kate Ehrlich et al. Getting the whole team into usability testing; who’s on the team? IEEE Software, 11(1):89–91, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Eickelmann:2003:SPC

[EA03] Nancy Eickelmann and Animesh Anant. Statistical process control: What you don’t measure can hurt you! IEEE Software, 20(2):49–51, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2049.
pdf.

Ebert:2007:B

[EA07] Christof Ebert and Anthony Akins. Bookshelf. IEEE Software, 24(3):110– 112, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2007/
                                        03/s3110.pdf.

Everett:1994:SBG

[EAS94] William W. Everett, Mikio Aoyama, and Jawed Siddiqi. Software beyond 2001: a global vision. IEEE Software, 11(6):8–12, November 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2001:BHE

[Ebe01]  Christof Ebert.  Book-

shelf: Hands-on experience implementing a product line: Software Product-Line Engineering: A Family-Based Software Development Process. IEEE Software, 18(1):105–106, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s1105.
                                        pdf.

Ebert:2006:UPL

[Ebe06] Christof Ebert. Understanding the product life cycle: Four key requirements engineering techniques. IEEE Software, 23 (3):19–25, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2007:OSD

[Ebe07]  Christof Ebert.  Open

source drives innovation. IEEE Software, 24(3):105– 109, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2008:OSS

[Ebe08a]  Christof Ebert.  Open

source software in industry. IEEE Software, 25 (3):52–53, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2008/
03/mso2008030052.pdf.

Ebert:2008:STB

[Ebe08b] Christof Ebert. Software technology: a brief history of software technology. IEEE Software, 25(6):22– 25, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2009:OSS

[Ebe09]  Christof Ebert.  Open

source software: Guest Editor’s introduction: How open source tools can benefit industry. IEEE Software, 26(2):50–51, March/ April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Everett:2009:JDM

[ECW09] William Everett, James Cusick, and Laurie Williams. John D. Musa. IEEE Software, 26(5):102, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2001:SGS

[ED01] Christof Ebert and Philip De Neve. Surviving global software development. IEEE Software, 18(2):62–69, March/April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s2062abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2062.pdf.

Egyhazy:2001:DTP

[EEM01] Csaba Egyhazy, Scott Eyestone, and Janet Martino. Defining team processes using OO metaphors. IEEE Software, 18(3):74– 83, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s3074abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s3074.pdf.

Engelstad:1991:DCB

[EFM^{+}91] Steve Engelstad, Keith Falck, Warren Montgomery, Joe Neumann, Ralph Straubs, Jim Vandendorpe, and Mike Wilde. A dynamic Cbased object-oriented system for Unix. IEEE Software, 8(3):73–85, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2004:GEI

[EFS04] Hakan Erdogmus, John Favaro, and Wolfgang Strigel. Guest Editors’ introduction: Return on investment. IEEE Software, 21(3):18–22, May/ June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2004/03/s3018.
pdf.

ElEmam:1996:NST

[EG96] Khaled El Emam and Dennis R. Goldenson. In the news: Spice trials produce positive results. IEEE Software, 13(6), November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Egyed:2004:IRC

[EG04a] Alexander Egyed and Paul Grunbacher. Identifying requirements conflicts and cooperation: How quality attributes and automated traceability can help. IEEE Software, 21(6):50–58, November/ December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Eugster:2004:DPT

[EG04b] Patrick T. Eugster and Rachid Guerraoui. Distributed programming with typed events. IEEE Software, 21(2):56–64, March/ April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Emrath:1992:DNP

[EGP92] Perry A. Emrath, Sanjoy Ghosh, and David A. Padua. Detecting nondeterminacy in parallel programs. IEEE Software, 9(1):69–77, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Everett:1995:GEI

[EH95] William W. Everett and Shinichi Honiden. Guest Editors’ introduction: Reliability and safety of real time systems. IEEE Software, 12(3):13–15, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Eickelmann:2003:IVC

[Eic03] Nancy Eickelmann. An insider’s view of CMM level 5. IEEE Software, 20(4):79– 81, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4079.
pdf.

Embley:1995:OSA

[EJW95] David W. Embley, Robert B.

Jackson, and Scott N. Woodfield. OO systems analysis: is it or isn’t it? IEEE Software, 12(4):19– 33, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebner:2002:TAA

[EK02] Gerald Ebner and Hermann Kaindl. Tracing all around in reengineering. IEEE Software, 19(3):70– 77, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3070.
pdf;   http://www.computer.
org/software/so2002/s3070abs.
htm.

ElEmam:2008:RSI

[EK08] Khaled El Emam and A. Günes Koru. A replicated survey of IT software project failures. IEEE Software, 25(5):84–90, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Evans:2002:ISU

[EL02] David Evans and David Larochelle. Improving security using extensible lightweight static analysis. IEEE Software, 19(1): 42–51, January/February 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2002/s1042abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1042.pdf.

Eickelmann:2004:MMG

[EL04] Nancy Eickelmann and Motorola Labs. Measuring maturity goes beyond process. IEEE Software, 21(4): 12–13, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ellison:1999:SNS

[ELLM99] Robert J. Ellison, Richard C. Linger, Thomas Longstaff, and Nancy R. Mead. Survivable network system analysis: a case study. IEEE Software, 16(4):70– 77, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s4070abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s4070.pdf.

Ehrlich:1990:ARM

[ELM90] Willa K. Ehrlich, S. Keith Lee, and Rex H. Molisani. Applying reliability measurement: a case study. IEEE Software, 7(2):56– 64, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Etzioni:1995:SAI

[ELST95] Oren Etzioni, Henry M.

Levy, Richard B. Segal, and Chandramohan A. Thekkath. The Softbot approach to OS interfaces. IEEE Software, 12(4):42– 51, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Elshoff:1991:MEP

[EM91] James L. Elshoff and Michael Marcotty. Microtool: An environment for programming microprocessors. IEEE Software, 8 (1):42–48, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Eriksson:1993:MKA

[EM93] Henrik Eriksson and Mark Musen. Metatools for knowledge acquisition. IEEE Software, 10(3):23–29, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

ElEmam:1996:QTD

[EM96] Khaled El Emam and Nazim H. Madhavji. Quality time: Does organizational maturity improve quality? IEEE Software, 13(5):109–110, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2004:B

[EM04] Christof Ebert and Stephen Mellor. Bookshelf. IEEE Software, 21(4):92–94, July/ August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/so/2004/04/s4092.
                                        pdf.

Everett:2009:PC

[EM09] Gerald D. Everett and Bertrand Meyer. Point/ counterpoint. IEEE Software, 26(4):62–65, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Evans:1999:LQM

[EMB^{+}99] John Evans, David B. Moffett, Randall Brukardt, Richard Kuhn, Bill Dietrich, David Budgen, and Soheil Khajenoori. Letters: Quality musings; what about Ada?; the art of design; noting another body of knowledge. IEEE Software, 16(6):9–12, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6009.
pdf.

Emery:1995:LDR

[Eme95]  David Emery.  Letters:

“Defects” the result of a rigorous Ada culture. IEEE Software, 12(4):7–8, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Egyed:2005:GEI

[EMP05] Alexander Egyed, Hausi A.

Muller, and Dewayne E. Perry. Guest Editors’ introduction: Integrating COTS into the development process. IEEE Software, 22(4): 16–18, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
04/s4016.pdf.

Emrath:1985:XOS

[Emr85] Perry Emrath. Xylem: an operating system for the Cedar multiprocessor. IEEE Software, 2(4):30– 37, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:1997:RMN

[EMW^{+}97] Christof Ebert, Tomoo Matsubara, Todd Webb, Mauro Pezzé, and Olav W. Bertelsen. The road to maturity: Navigating between craft and science. IEEE Software, 14(6):77–88, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s6077.
                                        pdf;   http://dlib.computer.
                                        org/so/books/so1997/pdf/
                                        s6077abs.htm.

Endres:1993:LLI

[End93] Albert Endres. Lessons learned in an industrial software lab. IEEE Software, 10(5):58–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Engel:1999:PCS

[Eng99] Gerald L. Engel. Program criteria for software engineering accreditation programs. IEEE Software, 16(6):31–34, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6031abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6031.pdf.

Ehrlich:1993:DCS

[EPSJ93] Willa Ehrlich, Bala Prasanna, John Stampfel, and Wu Jar. Determining the cost of a stop-test decision. IEEE Software, 10(2):33–??, March 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2005:EIL

[Erd05] Hakan Erdogmus. The economic impact of learning and flexibility on process decisions. IEEE Software, 22(6):76–83, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2007:ACA

[Erd07a] Hakan Erdogmus. Agile’s coming of age … or not. IEEE Software, 24(6):2–5, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.

                                        org/comp/mags/so/2007/
                                        06/mso2007060002.pdf.

Erdogmus:2007:ISB

[Erd07b] Hakan Erdogmus. IEEE Software–2007 and beyond. IEEE Software, 24(1):5– 7, January/February 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2007/
                                        01/s1005.pdf.

Erdogmus:2007:NS

[Erd07c] Hakan Erdogmus. Novelty in sameness. IEEE Software, 24(3):5–7, May/ June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/so/2007/03/s3005.
                                        pdf.

Erdogmus:2007:DES

[Erd07d] Hakan Erdogmus. Ondemand enterprise services: Where’s the catch? IEEE Software, 24(4):5–7, July/August 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2007/
                                        04/s4005.pdf.

Erdogmus:2007:TSA

[Erd07e] Hakan Erdogmus. Tips for software authors. IEEE Software, 24(5):5–7, September/October 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
05/s5005.pdf.

Erdogmus:2007:WGS

[Erd07f] Hakan Erdogmus. What’s good software, anyway? IEEE Software, 24(2):5–7, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
02/s2005.pdf.

Erdogmus:2008:ESP

[Erd08a]  Hakan Erdogmus.  Es-

sentials of software process. IEEE Software, 25 (4):4–7, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2008/
04/mso2008040004.pdf.

Erdogmus:2008:EYS

[Erd08b] Hakan Erdogmus. From the Editor: 25 years of software. IEEE Software, 25(6):2–5, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2008:IRM

[Erd08c] Hakan Erdogmus. The infamous ratio measure. IEEE Software, 25(3):4– 7, May/June 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2008/
                                        03/mso2008030004.pdf.

Erdogmus:2008:MA

[Erd08d] Hakan Erdogmus. Measurement acquiescence. IEEE Software, 25(2):4–6, March/ April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/so/2008/02/mso2008020004.
                                        pdf.

Erdogmus:2008:MSR

[Erd08e] Hakan Erdogmus. Must software research stand divided? IEEE Software, 25(5):4–6, September/October 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2008/
                                        05/mso2008050004.pdf.

Erdogmus:2008:MLL

[Erd08f] Hakan Erdogmus. So many languages, so little time. IEEE Software, 25(1):4– 6, January/February 2008.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2008/
01/mso2008010004.pdf.

Erdogmus:2009:AMA

[Erd09a]  Hakan Erdogmus.  Ar-

chitecture meets agility. IEEE Software, 26(5):2–4, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2009:ETT

[Erd09b] Hakan Erdogmus. From the Editor: a tale of two conferences. IEEE Software, 26(1):4–7, January/ February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2009:ECC

[Erd09c] Hakan Erdogmus. From the Editor: Cloud computing: Does Nirvana hide behind the nebula? IEEE Software, 26(2):4–6, March/ April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2009:EDS

[Erd09d] Hakan Erdogmus. From the Editor: Diversity and software development. IEEE Software, 26(3):2–4, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2009:EST

[Erd09e] Hakan Erdogmus. From the Editor: The seven traits of superprofessionals. IEEE Software, 26(4):4–6, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2009:P

[Erd09f] Hakan Erdogmus. A process that is not. IEEE Software, 26(6):4–7, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2002:TT

[ERJI02] Christof Ebert, Howard Rubin, Margaret Johnson, and Susan Iventosch. Table typo. IEEE Software, 19(3):8–12, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s3008.

pdf. See [RJI02].

Erwig:2009:SES

[Erw09] Martin Erwig. Software engineering for spreadsheets. IEEE Software, 26(5):25–30, September/ October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ege:1992:DMR

[ES92] Raimund K. Ege and Christian Stary. Designing maintainable, reusable interfaces. IEEE Software, 9 (6):24–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2009:FES

[ES09] Christof Ebert and Jürgen Salecker. FOCUS 1: Embedded software: Guest Editors’ introduction: Embedded software. IEEE Software, 26(3):14–18, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Etlinger:1996:BAC

[Etl96a] Hank Etlinger. Bookshelf:

Applying communication principles to visual interface design. IEEE Software, 13(2):117–118, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Etlinger:1996:BNK

[Etl96b] Hank Etlinger. Bookshelf:

The new Karel is longer and clearer — but is it better? IEEE Software, 13(3):110– 112, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Everett:1995:ASS

[Eve95] William Everett. Article summaries: Safety and reliability. IEEE Software, 12 (3):14–15, May 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fach:2001:DRT

[Fac01] Peter W. Fach. Design reuse through frameworks and patterns. IEEE Software, 18(5):71–76, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s5071abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5071.pdf.

Fafchamps:1994:OFR

[Faf94] Danielle Fafchamps. Organizational factors and reuse. IEEE Software, 11 (5):31–41, September 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fairley:1994:RMS

[Fai94] Richard Fairley. Risk management for software projects. IEEE Software, 11(3): 57–67, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fairley:2002:MAE

[Fai02] Dick Fairley. Making accurate estimates. IEEE Software, 19(6):61–63, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6061.
pdf.

Fairley:2005:SRM

[Fai05] Richard E. Fairley. Software risk management. IEEE Software, 22(3):101, May/June 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
03/s3101.pdf.

Farrow:1984:GPC

[Far84] Rodney Farrow. Generating a production compiler from an attribute grammar. IEEE Software, 1(4):77–93, October 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Farmer:2006:GGH

[Far06] Eugene Farmer. The gatekeeper’s guide, or how to kill a tool. IEEE Software, 23(6):12–13, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Favaro:1996:QTW

[Fav96]  John Favaro.  Quality

time: When the pursuit of quality destroys value. IEEE Software, 13(3):93– 95, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Favaro:2002:MRB

[Fav02]  John Favaro.  Manag-

ing requirements for business value. IEEE Software, 19(2):15–17, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2015.
pdf;   http://www.computer.
org/software/so2002/s2015abs.
htm.

Franch:2003:UQM

[FC03] Xavier Franch and Juan Pablo Carvallo. Using quality models in software package selection. IEEE Software, 20(1):34–41, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1034.
pdf;   http://www.computer.
org/software/so2003/s1034abs.
htm.

Florac:2000:SPC

[FCB00] William A. Florac, Anita D.

Carleton, and Julie Barnard. Statistical process control: Analyzing a Space Shuttle onboard software process. IEEE Software, 17(4):97–106, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4097abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4097.pdf.

Feather:2008:BQM

[FCH^{+}08] Martin S. Feather, Steven L.

Cornford, Kenneth A. Hicks, James D. Kiper, and Tim Menzies. A broad, quantitative model for making early requirements decisions. IEEE Software, 25(2):49–56, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Feathers:2004:BC

[Fea04]  Michael Feathers.  Be-

fore clarity. IEEE Software, 21(6):86–88, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Feitelson:2007:ALR

[Fei07] Dror G. Feitelson. Asimov’s Laws of Robotics applied to software. IEEE Software, 24(4):112, 111, July/August 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://bell.computer.
                                        org/dlcomments/.

Fenick:1990:IMM

[Fen90] Stewart Fenick. Implementing management metrics: An Army program. IEEE Software, 7(2):65– 72, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fenton:1996:CDS

[Fen96] Norman Fenton. Counterpoint: Do standards improve product quality? IEEE Software, 13(1):23– 24, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Feord:1997:TTM

[Feo97] Michael Feord. Toolbox: Testing for millennium risk management. top drawer. IEEE Software, 14(3):126, 127, 130, 131, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ferne:1995:BSE

[Fer95] Peter Ferne. Bookshelf: Solid examination of distributed operating systems: review of Distributed Operating Systems. IEEE Software, 12(6):115, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ferdinandi:1998:CWF

[Fer98] Patricia L. Ferdinandi.

Culture at work: Facilitating communication. IEEE Software, 15(5):92– 96, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1998/pdf/
s5092.pdf.

Ferguson:2001:MCD

[Fer01a] Jack Ferguson. Manager:

Crouching dragon, hidden software: Software in DoD weapon systems. IEEE Software, 18(4):105–107, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4105.
pdf.

Ferrell:2001:BMG

[Fer01b] Jerry Ferrell. Bookshelf:

Making the grade: Distributed Applications with Microsoft Visual C++ 6.0. IEEE Software, 18(4):108– 109, July/August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2001/pdf/
s4108.pdf.

Floerkemeier:2008:RAI

[FF08] Christian Floerkemeier and Elgar Fleisch. RFID applications: Interfacing with readers. IEEE Software, 25 (3):67–70, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

France:2003:MAP

[FGSK03] Robert France, Sudipto Ghosh, Eunjee Song, and Dae-Kyoo Kim. A metamodeling approach to pattern-based model refactoring. IEEE Software, 20(5):52–58, September/ October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5052.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5052abs.htm.

Faegri:2007:CPC

[FH07] Tor Erlend Fægri and Geir Kjetil Hanssen. Collaboration, process control, and fragility in evolutionary product development. IEEE Software, 24(3):96– 104, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Frieder:1985:LSP

[FHMU85] Gideon Frieder, Gabor T.

Herman, Craig Meyer, and Jayaram Udupa. Large software problems for small computers: an example from medical imaging. IEEE Software, 2(5):37–47, September 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fayad:1993:USM

[FHRK93] Mohamed E. Fayad, Louis J.

Hawn, Mark A. Roberts, and Jerry R. Klatt. Using the Shlaer-Mellor objectoriented analysis method. IEEE Software, 10(2):43– ??, March 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Floch:2006:UAM

[FHS^{+}06] Jacqueline Floch, Svein Hallsteinsen, Erlend Stav, Frank Eliassen, Ketil Lund, and Eli Gjorven. Using architecture models for runtime adaptability. IEEE Software, 23(2):62– 70, March/April 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Frakes:1994:SFS

[FI94] William B. Frakes and Sadahiro Isoda. Success factors of systematic reuse.

IEEE Software, 11(5):15– 19, September 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fidge:1996:FDS

[Fid96] Colin Fidge. Fundamentals of distributed system observation. IEEE Software, 13 (6):77–83, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Finkelstein:1996:IPU

[Fin96] Anthony Finkelstein. Improving public understanding. IEEE Software, 13 (6):20–21, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fischer:1987:CVR

[Fis87] Gerhard Fischer. Cognitive view of reuse and redesign. IEEE Software, 4(4):60– 72, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fischer:1989:HCI

[Fis89] Gerhard Fischer. Humancomputer interaction software: Lessons learned, challenges ahead. IEEE Software, 6(1):44–52, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ferre:2001:UBS

[FJWC01] Xavier Ferré, Natalia Juristo, Helmut Windl, and Larry Constantine. Usability basics for software developers. IEEE Software, 18(1):22–29, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s1022abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s1022.pdf.

Fitzgerald:2004:DIS

[FK04] Brian Fitzgerald and Tony Kenny. Developing an information systems infrastructure with Open Source software. IEEE Software, 21(1):50–55, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Foley:1989:DIH

[FKKM89] James Foley, Won Chul Kim, Srdjan Kovacevic, and Kevin Murray. Defining interfaces at a high level of abstraction. IEEE Software, 6(1):25–32, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Flower:1991:FEA

[FKLW91] Jon Flower, Adam Kolawa, Timothy Liang, and Victor Weingarten. Finiteelement analysis on a PC. IEEE Software, 8(5):50–??, September 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fenton:2002:SMU

[FKN02] Norman Fenton, Paul Krause, and Martin Neil. Software measurement: Uncertainty and causal modeling. IEEE Software, 19(4):116–122, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4116.
pdf;   http://www.computer.
org/software/so2002/s4116abs.
htm.

Fidge:1997:FMB

[FKU97] C. Fidge, P. Kearney, and M. Utting. A formal method for building concurrent real-time software. IEEE Software, 14(2):99– 106, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Faustini:1986:TRT

[FL86] Antony A. Faustini and Edgar B. Lewis. Toward a real-time dataflow language. IEEE Software, 3(1):29–35, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fleming:1999:FPO

[Fle99] Ryan Fleming. A fresh perspective on old problems. IEEE Software, 16(1):106–??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1106.
pdf.

Funka-Lea:1991:IVM

[FLKMR91] Cynthia Funka-Lea, Tasos Kontogiorgos, Robert Morris, and Larry Rubin. Interactive visual modeling for performance analysis. IEEE Software, 8(5):58–??, September 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fogle:2001:BPO

[FLN01] Sam Fogle, Carol Loulis, and Bill Neuendorf. The benchmarking process: One team’s experience. IEEE Software, 18(5):40–47, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5040abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5040.pdf.

Fung:2004:EDE

[FLR04] Kam Hay Fung, Graham Low, and Pradeep Kumar Ray. Embracing dynamic evolution in distributed systems. IEEE Software, 21(2):49–55, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fraser:2008:NSB

[FM08] Steven Fraser and Dennis Mancl. No silver bullet: Software engineering reloaded. IEEE Software, 25(1):91–94, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fickas:1988:CSS

[FN88] Stephen Fickas and P. Nagarajan. Critiquing software specifications. IEEE Software, 5(6):37–47, November 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Feitelson:1999:STS

[FN99] Dror G. Feitelson and Michael Naaman. Selftuning systems. IEEE Software, 16(2):52–60, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s2052abs.htm;

http://dlib.computer.
org/so/books/so1999/pdf/
s2052.pdf.

Fernstrom:1992:SFP

[FNO92] Christer Fernström, Kjell-Håkan Närfelt, and Lennart Ohlsson. Software factory principles, architectures, and experiments. IEEE Software, 9(2):36– 44, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fischer:2009:MGS

[FNY09] Gerhard Fischer, Kumiyo Nakakoji, and Yunwen Ye. Metadesign: Guidelines for supporting domain experts in software development. IEEE Software, 26(5):37–44, September/ October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fridrich:1985:HAX

[FO85] Marek Fridrich and William Older. Helix: The architecture of the XMS distributed file system. IEEE Software, 2(3):21–29, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fitzgerald:1999:LSS

[FO99] Brian Fitzgerald and Tom O’Kane. A longitudinal study of software process improvement. IEEE Software, 16(3):37–45, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s3037abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s3037.pdf.

Fong:1996:ARI

[Fon96]  Joseph Fong.  Adding

a relational interface to a nonrelational database. IEEE Software, 13(5):89– 97, September 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ford:1988:CAR

[For88]  Ray Ford.  Concur-

rent algorithms for realtime memory management. IEEE Software, 5(5):10–23, September 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Forte:1992:TFL

[For92] Gene Forte. Tools fair: Out of the lab, onto the shelf. IEEE Software, 9(3): 70–??, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Forte:1997:MCR

[For97] G. Forte. Managing change for rapid development. IEEE Software, 14(2):120– 122, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fowler:2001:DAR

[Fow01a] Martin Fowler. Design:

Avoiding repetition. IEEE Software, 18(1):97–99, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1097.
pdf.

Fowler:2001:DRC

[Fow01b] Martin Fowler. Design:

Reducing coupling. IEEE Software, 18(4):102–104, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4102.
pdf.

Fowler:2001:DSU

[Fow01c] Martin Fowler. Design separating user interface code. IEEE Software, 18(2):96– 97, March/April 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so2001/pdf/s2096.
pdf.

Fowler:2001:E

[Fow01d] Martin Fowler. To be explicit. IEEE Software, 18(6):10–15, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6010abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6010.pdf.

Fowler:2002:PVP

[Fow02a] Martin Fowler. Public versus published interfaces. IEEE Software, 19(2):18– 19, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2018.
pdf;   http://www.computer.
org/software/so2002/s2018abs.
htm.

Fowler:2002:UM

[Fow02b]  Martin Fowler.  Using

metadata. IEEE Software, 19(6):13–17, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6013.
pdf.

Fowler:2002:YAO

[Fow02c] Martin Fowler. Yet another optimization article. IEEE Software, 19(3):20– 21, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3020.
pdf;   http://www.computer.
org/software/so2002/s3020abs.
htm.

Fowler:2003:DAR

[Fow03a] Martin Fowler. Data access routines. IEEE Software, 20(6):96–98, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6096.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6096abs.htm.

Fowler:2003:P

[Fow03b] Martin Fowler. Patterns.

IEEE Software, 20(2):56– 57, March/April 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2056.
pdf.

Fowler:2003:WMT

[Fow03c] Martin Fowler. When to make a type. IEEE Software, 20(1):12–13, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s1012.
                                        pdf.

Fowler:2003:WNA

[Fow03d] Martin Fowler. Who needs an architect? IEEE Software, 20(5):11–13, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/05/s5011.
                                        pdf.

Fowler:2004:MA

[Fow04] Martin Fowler. Module assembly. IEEE Software, 21(2):65–67, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fowler:2005:SD

[Fow05] Martin Fowler. The state of design. IEEE Software, 22 (6):12–13, 16, November/ December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fowler:2009:TTP

[Fow09] Martin Fowler. Tools of the trade: a pedagogical framework for domain-specific languages. IEEE Software, 26(4):13–14, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fenton:1994:SSC

[FPG94] Norman Fenton, Shari Lawrence Pfleeger, and Rober L. Glass. Science and substance: a challenge to software engineers. IEEE Software, 11(4):86–95, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Freeman:2007:PC

[FPK07] Steve Freeman, Nat Pryce, and Joshua Kerievsky. Point/counterpoint. IEEE Software, 24(3):80–83, May/ June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Favela:2001:ECS

[FPM01] Jesús Favela and Feniosky Peña-Mora. An experience in collaborative software engineering education. IEEE Software, 18(2):47–53, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2047abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2047.pdf.

Frailey:1995:PSE

[Fra95] Dennis J. Frailey. Proving software engineering’s legitimacy. IEEE Software, 12 (1):14, January 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Franz:1998:JVM

[Fra98] Michael Franz. The Java Virtual Machine: a passing fad? IEEE Software, 15(6):26–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1998/s6026abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s6026.pdf.

Frailey:1999:PSE

[Fra99] Dennis J. Frailey. Point: Software engineering grows up. IEEE Software, 16 (6):66, 68, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s6066.
                                        pdf.

Fritz:1997:LPL

[Fri97]  Robert Fritz.  Letters:

The programmer … fishing? IEEE Software, 14(5):14–15, September/ October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5008.
pdf.

Frost:2007:JEW

[Fro07] Randall Frost. Jazz and the Eclipse way of collaboration. IEEE Software, 24(6):114–117, November/ December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fruhauf:1999:BCO

[Frü99]  Karol Frühauf.  Book-

shelf: Customer oriented software quality assurance: Client/server programming with JAVA and CORBA. IEEE Software, 16(5):131–133, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s5131.
pdf.

Ferguson:1998:LYC

[FS98] JoAn Ferguson and Sarah Sheard. Leveraging your CMM efforts for IEEE/ EIA 12207. IEEE Software, 15(5):23–29, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s5023.
                                        pdf.

Frezza:2006:CAS

[FTB06] Stephen Frezza, Mei-Huei Tang, and Barry J. Brinkman. Creating an accreditable Software Engineering Bachelor’s program. IEEE Software, 23(6):27–35, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fayad:1994:AOO

[FTR^{+}94] Mohamed E. Fayad, Wei-Tek Tsai, Mark A. Roberts, Louis J. Hawn, and Jay W. Schooley. Adapting an object-oriented development method. IEEE Software, 11(3):68–76, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fujino:1999:STJ

[Fuj99] Terunobu Fujino. Soapbox: Traditional Japanese architecture blends beauty and rationale. IEEE Software, 16(6):101–103, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s6101.
                                        pdf.

Furey:1997:WWS

[Fur97] S. Furey. Why we should use function points [software metrics]. IEEE Software, 14(2):28–??, 30, March/April 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Frank:1997:LPM

[FV97] Karl Frank and Jeffrey Voas. Letters: Is pipe metaphor a wash out? IEEE Software, 14(5):15, 17, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s5008.pdf.

Ferguson:2007:PC

[FV07] Donald Ferguson and Richard Veryard. Point/counterpoint. IEEE Software, 24(6):78– 81, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fritzson:1995:HLM

[FVHF95] Peter Fritzson, Lars Viklund, Johan Herber, and Dag Fritzson. High-level mathematical modeling and programming. IEEE Software, 12(4):77–87, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fairley:2003:WVS

[FW03] Richard E. Fairley and Mary Jane Willshire. Why the Vasa sank: 10 problems and some antidotes for software projects. IEEE Software, 20(2):18–25, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2018.
pdf;   http://www.computer.
org/software/so2003/s2018abs.
htm.

Frank:1985:MCN

[FWB85] Ariel J. Frank, Larry D. Wittie, and Arthur J. Bernstein. Multicast communication on network computers. IEEE Software, 2(3): 49–61, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gacek:2004:MMO

[GA04] Cristina Gacek and Budi Arief. The many meanings of Open Source. IEEE Software, 21(1):34–40, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2005:EPD

[GA05] Greg Goth and Benjamin Alfonsi. Enter the protectionist dragon? China’s software and standards policies stir debate and reflection. IEEE Software, 22(2):83–87, March/ April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2005/02/s2083.
pdf.

Galin:2006:CPI

[GA06] Daniel Galin and Motti Avrahami. Are CMM program investments beneficial? analyzing past studies. IEEE Software, 23(6):81–87, November/ December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gait:1994:OUD

[Gai94] Jason Gait. Optimizing Unix database file operations. IEEE Software, 11 (3):48–56, May 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garlan:1995:AMW

[GAO95] David Garlan, Robert Allen, and John Ockerbloom. Architectural mismatch: why reuse is so hard. IEEE Software, 12(6):17– 26, November 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garlan:2009:ATP

[GAO09] David Garlan, Robert Allen, and John Ockerbloom. 25th-anniversary top picks: Architectural mismatch: Why reuse is still so hard. IEEE Software, 26(4):66– 69, July/August 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garr:2000:CWW

[Gar00] Dixie Garr. Culture at work: The work culture at Cisco Systems. IEEE Software, 17(6):16–18, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s6016.
                                        pdf.

Garcia:2005:HSE

[Gar05] Suzanne Garcia. How standards enable adoption of project management practice. IEEE Software, 22(5): 22–29, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2001:NOM

[GAS^{+}01] Greg Goth, Scott Lorenz Andresen, Gil Alexander Shif, Paula Powers, and Scott Lorenz Andresen. In the news: The open market woos open source; awards & recognition: Anup Ghosh named Young Engineer of the Year; next-generation entertainment consoles: Not all fun and games; Asia Pacific region poised to follow India in software development revolution. IEEE Software, 18(2):104–111, March/April 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2104.
pdf.

Gause:2005:WCM

[Gau05] Donald C. Gause. Why context matters—and what can we do about it? IEEE Software, 22(5):13– 15, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gabber:1993:PPP

[GAY93] Eran Gabber, Amir Averbuch, and Amiram Yehudai. Portable, parallelizing Pascal compiler. IEEE Software, 10(2):71– ??, March 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gordon:1995:RPL

[GB95] V. Scott Gordon and James M. Bieman. Rapid prototyping: Lessons learned. IEEE Software, 12(1):85– 95, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2000:NNG

[GB00] Greg Goth and Cheryl Baltes. In the news: The next gold rush: Application service providers stake their claims in a red-hot market: Are consumer rights at risk?: Standards-based training: Grand opening: SPC’s resource center. IEEE Software, 17(2):96–100, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2096.
pdf.

Gonzalez-Barahona:2000:LSC

[GBdlHQCG^{+}00] Jesús M. González-Barahona, Pedro de-las Heras-Quirós, José Centeno-González, Vicente Matellán-Olivera, and Francisco J. Ballesteros. Libre software for computer science classes. IEEE Software, 17(3):76– 79, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3076abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s3076.pdf.

Gammage:1985:XRB

[GC85] Neil Gammage and Liam Casey. XMS: a rendezvousbased distributed system software architecture. IEEE Software, 2(3):9– 20, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2005:Na

[GC05] Greg Goth and Terry Costlow. In the news. IEEE Software, 22(5):110–114, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
05/s5110.pdf.

Goth:2007:C

[GC07] Greg Goth and Terry Costlow. Currents. IEEE Software, 24(2):94–98, March/ April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gilb:2008:PC

[GC08] Tom Gilb and Alistair Cockburn. Point/counterpoint. IEEE Software, 25(2):64– 67, March/April 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2006:STI

[GCB^{+}06] Robert L. Glass, Ross Collard, Antonia Bertolino, James Bach, and Cem Kaner. Software testing and industry needs. IEEE Software, 23(4):55– 57, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gerhart:1994:CSM

[GCR94a] Susan Gerhart, Dan Craigen, and Ted Ralston. Case study: Multinet Gateway System. IEEE Software, 11(1):37–39, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gerhart:1994:CSP

[GCR94b] Susan Gerhart, Dan Craigen, and Ted Ralston. Case study: Paris Metro signaling system. IEEE Software, 11(1):32–35, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gerhart:1994:EFM

[GCR94c] Susan Gerhart, Dan Craigen, and Ted Ralston. Experience with formal methods in critical systems. IEEE Software, 11(1):21– 28, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2002:NNR

[GD02] Greg Goth and Angela Davids. In the news: NIST report takes a step toward better testing; national strategy to secure cyberspace proposed. IEEE Software, 19(6):97–100, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6097.
pdf.

Goth:2007:CTP

[GDD07] Greg Goth, Dave Dorenbos, and Virginia Dorenbos. Currents: Trends, people, projects. IEEE Software, 24(4):100–105, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Godefroid:2008:AST

[GdHN^{+}08] Patrice Godefroid, Peli de Halleux, Aditya V. Nori, Sriram K. Rajamani, Wolfram Schulte, Nikolai Tillmann, and Michael Y. Levin. Automating software testing using program analysis. IEEE Software, 25(5):30–37, September/ October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Guerrero:2004:ASC

[GE04] Felipe Guerrero and Yadran Eterovic. Adopting the SW-CMM in a small IT organization. IEEE Software, 21(4):29–35, July/ August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Geer:2006:SDP

[Gee06]  David Geer.  Software

developer profession expanding. IEEE Software, 23(2):112–115, March/ April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/so/2006/02/s2112.
                                        pdf.

Gemignani:1985:WOW

[Gem85] Michael C. Gemignani. Who owns what software products? IEEE Software, 2(5):48–52, September 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gerhart:1990:AFM

[Ger90] Susan L. Gerhart. Applications of formal methods: Developing virtuoso software. IEEE Software, 7(5):6–10, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gougher:1999:BJB

[GF99] Rib Gougher and Karol Frühauf. Bookshelf: Jim Blinn’s Corner: Dirty Pixels; Software Runaways– Lessons Learned from Massive Software Project Failures. IEEE Software, 16 (6):112–114, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6112.
pdf.

GormLarsen:1996:AFS

[GFB96] Peter Gorm Larsen, John Fitzgerald, and Tom Brookes. Applying formal specification in industry. IEEE Software, 13(3):48–56, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gall:2009:CAE

[GFP09] Harald C. Gall, Beat Fluri, and Martin Pinzger. Change analysis with Evolizer and ChangeDistiller. IEEE Software, 26(1):26–33, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goldman:1989:QPP

[GG89] Ron Goldman and Richard P.

Gabriel. Qlisp: Parallel processing in Lisp. IEEE Software, 6(4):51–59, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gamble:2008:MMI

[GG08] M. Todd Gamble and Rose Gamble. Monoliths to mashups: Increasing opportunistic assets. IEEE Software, 25(6):71– 79, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Guarna:1989:FIE

[GGJ^{+}89] Vincent A. Guarna, Jr., Dennis Gannon, David Jablonowski, Allen D. Malony, and Yogesh Gaur. Faust: An integrated environment for parallel programming. IEEE Software, 6(4):20–27, July 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gunter:2000:RMR

[GGJZ00] Carl A. Gunter, Elsa L. Gunter, Michael Jackson, and Pamela Zave. A reference model for requirements and specifications. IEEE Software, 17(3):37– 43, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3037abs.htm;
http://dlib.computer.

org/so/books/so2000/pdf/
s3037.pdf.

Gorschek:2006:MTT

[GGLW06] Tony Gorschek, Per Garre, Stig Larsson, and Claes Wohlin. A model for technology transfer in practice. IEEE Software, 23(6):88– 95, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Green:2000:SDI

[GH00] Gina C. Green and Alan R.

Hevner. The successful diffusion of innovations: Guidance for software development organizations. IEEE Software, 17(6):96–103, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6096abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6096.pdf.

Godfrey:2009:FMS

[GHH^{+}09] Michael W. Godfrey, Ahmed E.

Hassan, James Herbsleb, Gail C. Murphy, Martin Robillard, Prem Devanbu, Audris Mockus, Dewayne E. Perry, and David Notkin. Future of mining software archives: a roundtable. IEEE Software, 26(1):67–70, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Guttag:1985:LFS

[GHW85] John V. Guttag, James J. Horning, and Jeannette M. Wing. Larch family of specification languages. IEEE Software, 2(5):24–36, September 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ghosh:2002:BSS

[GHW02] Anup K. Ghosh, Chuck Howell, and James A. Whittaker. Building software securely from the ground up. IEEE Software, 19(1):14–16, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2002/s1014abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2002/pdf/
                                        s1014.pdf.

Gilb:1996:MLW

[Gil96a] Tom Gilb. Manager — level 6: Why we can’t get there from here. IEEE Software, 13(1):97–98, 103, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gillette:1996:PCM

[Gil96b]  Walt Gillette.  Point

counterpoint: Managing megaprojects: a focused approach. IEEE Software, 13(4):28, 30, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Grable:1999:MSP

[GJPD99] Ross Grable, Jacquelyn Jernigan, Casey Pogue, and Dale Divis. Metrics for small projects: Experiences at the SED. IEEE Software, 16(2):21–29, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s2021abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s2021.pdf.

Guntersdorfer:2002:HSP

[GK02] Michael S. Guntersdorfer and David G. Kay. How software patents can support COTS component business. IEEE Software, 19(3):78–83, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3078.
pdf;   http://www.computer.
org/software/so2002/s3078abs.
htm.

Gold:2004:USO

[GKMM04] Nicolas Gold, Claire Knight, Andrew Mohan, and Malcolm Munro. Understanding service-oriented software. IEEE Software, 21(2):71–77, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:1985:ST

[Gla85] Robert L. Glass. Software theft. IEEE Software, 2(4): 82–85, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:1994:MWD

[Gla94a] Robert L. Glass. Mailbox: Where does the rush hour metaphor lead us? IEEE Software, 11(1):6–7, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:1994:SRC

[Gla94b] Robert L. Glass. The software-research crisis. IEEE Software, 11(6):42– 47, November 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:1997:STG

[Gla97] Robert L. Glass. Soapbox: Telling good numbers from bad ones. IEEE Software, 14(4):15, 16, 19, July/August 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4015.
pdf.

Glass:1998:PP

[Gla98a] Robert L. Glass. In praise of practice. IEEE Software, 15(1):30–35, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1030abs.htm.

Glass:1998:LOD

[Gla98b] Robert L. Glass. Loyal opposition: Defining quality intuitively. IEEE Software, 15(3):103–104, 107, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3103.
pdf.

Glass:1998:LOT

[Gla98c] Robert L. Glass. Loyal opposition: Is there really a software crisis? IEEE Software, 15(1):104–105, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so1998/pdf/s1104.
                                        pdf.

Glass:1998:LOM

[Gla98d] Robert L. Glass. Loyal opposition: Maintenance: Less is not more. IEEE Software, 15(4):67–68, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s4067.
                                        pdf.

Glass:1998:LOR

[Gla98e] Robert L. Glass. Loyal opposition: Reuse: What’s wrong with this picture? IEEE Software, 15(2):57– 59, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s2057.
                                        pdf.

Glass:1999:CHP

[Gla99a] Robert L. Glass. Cobol: a historic past, a vital future? IEEE Software, 16(4):120, 118–119, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s4120.
                                        pdf.

Glass:1999:LOD

[Gla99b] Robert L. Glass. Loyal opposition: On design. IEEE Software, 16(2):104, March/April 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2104.
pdf.

Glass:1999:NBL

[Gla99c] Robert L. Glass. News briefs: Loyal opposition: Of open source, Linux . . . . . . and hype. IEEE Software, 16(1):128, 126–127, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1128.
pdf.

Glass:1999:SSD

[Gla99d] Robert L. Glass. A snapshot of systems development practice. IEEE Software, 16(3):112, 110–111, May/June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3112.
pdf.

Glass:2000:LON

[Gla00a] Robert L. Glass. Loyal opposition: New answer to “How Important Is Mathematics to the Software Practitioner?”. IEEE Software, 17(6): 136, 135, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6136.
pdf.

Glass:2000:LOP

[Gla00b] Robert L. Glass. Loyal opposition: Process diversity and a computing old wives’/husbands’ tale. IEEE Software, 17(4): 128–128, 127–127, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4128.
pdf.

Glass:2000:LOG

[Gla00c] Robert L. Glass. Loyal opposition: The generalization of an application domain. IEEE Software, 17(5):128–128, 127–127, September/October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2000/pdf/
s5128.pdf.

Glass:2000:LOS

[Gla00d] Robert L. Glass. Loyal opposition: The sociology of open source: Of cults and cultures. IEEE Software, 17(3):104–105, May/June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3104.
pdf.

Glass:2000:LOY

[Gla00e] Robert L. Glass. Loyal opposition: Y2K and other software noncrises. IEEE Software, 17(2): 104–105, 100, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2104.
pdf.

Glass:2001:EPG

[Gla01a] Robert L. Glass. Extreme Programming: The good, the bad, and the bottom line. IEEE Software, 18 (6):112, 111, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6112abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6112.pdf.

Glass:2001:LOF

[Gla01b] Robert L. Glass. Loyal opposition: Frequently forgotten fundamental facts about software engineering. IEEE Software, 18 (3):112, 110–111, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s3112.
                                        pdf.

Glass:2001:SAC

[Gla01c]  Robert L. Glass.  A

story about the creativity involved in software work. IEEE Software, 18(5):96, c3, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s5096abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5096.pdf.

Glass:2002:PFM

[Gla02a] Robert Glass. Predicting future maintenance cost, and how we’re doing it wrong. IEEE Software, 19 (6):112, 111, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/

books/so2002/pdf/s6112.
pdf.

Glass:2002:FLM

[Gla02b] Robert L. Glass. Failure is looking more like success these days. IEEE Software, 19(1):104, 103, January/ February 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1104abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1104.pdf.

Glass:2002:SMT

[Gla02c] Robert L. Glass. In search of meaning (a tale of two words). IEEE Software, 19(4):136, 134–135, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4136.
pdf.

Glass:2002:LOP

[Gla02d] Robert L. Glass. Loyal opposition: Project retrospectives, and why they never happen. IEEE Software, 19(5):112, 111, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so2002/pdf/s5112.
pdf.

Glass:2002:NOO

[Gla02e] Robert L. Glass. The naturalness of object orientation: Beating a dead horse? IEEE Software, 19(3):104, 103, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3104.
pdf.

Glass:2003:BPA

[Gla03a] Robert L. Glass. A big problem in academic software engineering and a potential outside-the-box solution. IEEE Software, 20(4):96, 94–95, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4096.
pdf.

Glass:2003:EFS

[Gla03b] Robert L. Glass. Errorfree software remains extremely elusive. IEEE Software, 20(1):104, 103, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1104.
pdf.

Glass:2003:GEI

[Gla03c] Robert L. Glass. Guest Editor’s introduction: The state of the practice of software engineering. IEEE Software, 20(6):20–21, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6020.
pdf.

Glass:2003:QSE

[Gla03d] Robert L. Glass. Questioning the software engineering unquestionables. IEEE Software, 20(3):120, 119, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3120.
pdf.

Glass:2004:AES

[Gla04a] Robert L. Glass. Anarchy and the effects of schedule pressure. IEEE Software, 21(5):112, 111, September/ October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2004:LDS

[Gla04b] Robert L. Glass. Learning to distinguish a solution from a problem. IEEE Software, 21(3):112, 111, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2004:MD

[Gla04c] Robert L. Glass. On modeling and discomfort. IEEE Software, 21(2):104, 102– 103, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2004:SHR

[Gla04d] Robert L. Glass. Some heresy regarding software engineering. IEEE Software, 21(4):104, 102–103, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2004:SSD

[Gla04e] Robert L. Glass. Sources for software development: a mugwumpish view. IEEE Software, 21(1):112, 111, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2005:FLS

[Gla05a] Robert L. Glass. A followthe-leader story with a strange ending. IEEE Software, 22(6):112, 111, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2005:NCT

[Gla05b] Robert L. Glass. Never the CS and IS twain shall meet? IEEE Software, 22(5):120, 119, September/ October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2005:SSS

[Gla05c] Robert L. Glass. A sad SAC story about the state of the practice. IEEE Software, 22(4):120, 119, July/ August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2005:VBG

[Gla05d] Robert L. Glass. Viruses are beginning to get to me! IEEE Software, 22 (1):104, 102–103, January/ February 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2006:GVR

[Gla06a] Robert L. Glass. Greece vs. Rome: Two very different software cultures. IEEE Software, 23(6):112, 111, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2006:HMS

[Gla06b] Robert L. Glass. How much of the software engineering old still remains new? IEEE Software, 23 (4):104, 103, July/August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2006:HTM

[Gla06c]  Robert L. Glass.  Of

health, trust, money … and integrity. IEEE Software, 23(1):120, 119, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Glass:2007:SEFa

[Gla07a] Robert L. Glass. Is software engineering fun? IEEE Software, 24(1):96, 95, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Glass:2007:SEFb

[Gla07b] Robert L. Glass. Is software engineering fun? part 2. IEEE Software, 24(2):104, 103, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Glass:2007:WBT

[Gla07c] Robert L. Glass. What’s with this blog thing? IEEE Software, 24(5):104, 103, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Glass:2008:IRD

[Gla08a]  Robert L. Glass.  In-

tuition’s role in decision making. IEEE Software, 25(1):96, 95, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Glass:2008:LOA

[Gla08b] Robert L. Glass. Loyal opposition: An ancient (but still valid?) look at the classification of testing. IEEE Software, 25(6):112, 111, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2008:NPW

[Gla08c] Robert L. Glass. Negative productivity and what to do about it. IEEE Software, 25(5):96, 95, September/October 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://bell.computer.
                                        org/dlcomments/;   http:
                                        //www.adobe.com/products/
                                        acrobat/readstep2.html.

Glass:2008:IEL

[Gla08d]  Robert L. Glass.  On

the impurity of the English language. IEEE Software, 25(2):96, 95, March/April 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://bell.computer.
                                        org/dlcomments/.

Glass:2008:SHZ

[Gla08e] Robert L. Glass. Software: Hero or zero? IEEE Software, 25(3):96, 95, May/June 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://bell.computer.
                                        org/dlcomments/.

Glass:2008:TME

[Gla08f] Robert L. Glass. Two mistakes and error-free software: a confession. IEEE Software, 25(4):96, 95, July/August 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://bell.computer.
                                        org/dlcomments/.

Glass:2009:DSS

[Gla09a] Robert L. Glass. Doubt and software standards. IEEE Software, 26(5):104, 103, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2009:G

[Gla09b] Robert L. Glass. Goodbye! IEEE Software, 26(6):95, 95, November/ December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2009:LOC

[Gla09c] Robert L. Glass. Loyal opposition: a classification system for testing, Part 2. IEEE Software, 26(1):104, 103, January/ February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2009:LOM

[Gla09d] Robert L. Glass. Loyal opposition: Making research more relevant while not diminishing its rigor. IEEE Software, 26(2):96, 95, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glinz:2008:RBV

[Gli08] Martin Glinz. A riskbased, value-oriented approach to quality requirements. IEEE Software, 25(2):34–41, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gray:1999:NCT

[GLS99] Jonathan Gray, Anna Liu, and Louise Scott. In the news: CASE tool conference roundup. IEEE Software, 16(6): 108, 111, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6104.
pdf.

Graaf:2003:ESE

[GLT03] Bas Graaf, Marco Lormans, and Hans Toetenel. Embedded software engineering: The state of the practice. IEEE Software, 20(6):61–69, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6061.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6061abs.htm.

Goth:2005:Nb

[GM05] Greg Goth and Bart Massey. In the news. IEEE Software, 22(6):98–102, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
06/s6098.pdf.

Gotel:2009:RMT

[GM09] Olly Gotel and Stephen Morris. Requirements: More than just “Lost in Translation”. IEEE Software, 26(2):7–9, March/ April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Grossman:1996:PTL

[GMS^{+}96] David Grossman, M. Catherine McCabe, Christopher Station, Bret Bailey, Ophir Frieder, and David C. Roberts. Performance testing a large finance application. IEEE Software, 13 (5):50–54, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Greenberg:1995:LIC

[GN95] Steven S. Greenberg and Jakob Nielsen. Letters: Interface colors are yours for the choosing. IEEE Software, 12(5):14, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gomaa:2006:RTS

[Gom06]  Hassan Gomaa.  Real-

time software engineering, part 2. IEEE Software, 23(5):102, September/October 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
05/s5102.pdf.

GonzalezRodriguez:2002:ADG

[Gon02] Martín González Rodríguez. Automatic data-gathering agents for remote navigability testing. IEEE Software, 19(6):78–85, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6078.
pdf;   http://www.computer.
org/software/so2002/s6078abs.
htm.

Gonzales:2005:DRD

[Gon05] Regina Gonzales. Developing the requirements discipline: Software vs. systems. IEEE Software, 22(2):59– 61, March/April 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gorton:1995:SRP

[Gor95] Ian Gorton. Soapbox: Real programmers do use Delphi. IEEE Software, 12(6): 8, 10, 12, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:1999:NSE

[Got99a]  Greg Goth.  In the

news: The shareware explosion: Certifying the certifiers. IEEE Software, 16(6):104–108, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6104.
pdf.

Gotterbarn:1999:HNS

[Got99b]  Don Gotterbarn.  How

the new software engineering code of ethics affects you. IEEE Software, 16(6):58–64, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6058abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6058.pdf.

Goth:2000:NNA

[Got00] Greg Goth. In the news:

New air traffic control software takes an incremental approach. IEEE Software, 17(4):108–111, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4108.
pdf.

Goth:2001:NDD

[Got01a] Greg Goth. In the news: Danger! danger! console game developers ready themselves for the next generation. IEEE Software, 18(4):96–100, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s4096.
                                        pdf.

Goth:2001:NMP

[Got01b] Greg Goth. In the news: Military PDAs: All handhelds on deck! IEEE Software, 18(5):81–85, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s5081abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5081.pdf.

Goth:2001:NNC

[Got01c]  Greg Goth.  In the

news: New center will help software development “grow up”. IEEE Software, 18(3):99–103, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s3099.
                                        pdf.

Goth:2001:NPS

[Got01d]  Greg Goth.  In the

news: Privacy, security concerns take on new meaning; DMCA “fixes” delayed. IEEE Software, 18(6):102–105, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6102abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6102.pdf.

Goth:2001:NLB

[Got01e]  Greg Goth.  NSDIR:

a legacy beyond failureearly benchmarking effort was ahead of its time. IEEE Software, 18(5):53–56, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5053abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5053.pdf.

Goth:2002:NFG

[Got02a] Greg Goth. In the news:

Federal Government calls for more secure software design; secure with Cyclone language; off-the-shelf software powers International Space Station troubleshooter. IEEE Software, 19(1):90–94, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1090abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1090.pdf;   http://www.
research.att.com/projects/
cyclone.

Goth:2002:NOO

[Got02b] Greg Goth. In the news: Has object-oriented programming delivered? IEEE Software, 19(5):104–107, September/October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2002/pdf/
s5104.pdf;   http://www.
computer.org/software/
so2002/s5104abs.htm.

Goth:2002:NWC

[Got02c] Greg Goth. In the news: Will the cyber-UL concept take hold? IEEE Software, 19(4):12–15, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4012.
pdf.

Goth:2002:LTD

[Got02d] Greg Goth. Long-term data preservation efforts ramping up. IEEE Software, 19(2):98–102, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2098.
pdf;   http://www.computer.
org/software/so2002/s2098abs.
htm.

Goth:2002:NER

[Got02e] Greg Goth. A new era for research and development? IEEE Software, 19(3):98– 100, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3098.
pdf.

Gottesdiener:2003:RCG

[Got03] Ellen Gottesdiener. Requirements by collaboration: Getting it right the first time. IEEE Software, 20(2):52–55, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2052.
pdf.

Gotel:2006:SSC

[Got06a] Olly Gotel. In search of the system concept. IEEE Software, 23(1):102–103, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2006:N

[Got06b] Greg Goth. In the news. IEEE Software, 23(4):93– 97, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        04/s4093.pdf.

Goth:2007:STO

[Got07]  Greg Goth.  Sprinting

toward open source development. IEEE Software, 24(1):88–91, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2008:GTP

[Got08a] Greg Goth. “Googling” test practices? Web giant’s culture encourages process improvement. IEEE Software, 25(2):92–94, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2008:USR

[Got08b] Greg Goth. Ultralarge systems: Redefining software engineering? IEEE Software, 25(3):91–94, May/ June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2009:CAT

[Got09a]  Greg Goth.  Currents:

Agile tool market growing with the philosophy. IEEE Software, 26(2):88– 91, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2009:TBI

[Got09b] Greg Goth. The task-based interface: Not your father’s desktop. IEEE Software, 26(6):88–91, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gougher:1997:BCC

[Gou97] Robert Gougher. Bookshelf: Civilizing cyberspace. IEEE Software, 14(5):117, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s5115.pdf.

Gougher:1998:BCN

[Gou98] Robert A. Gougher. Bookshelf: Crafting nonhuman champions. IEEE Software, 15(2):120–122, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s2120.
                                        pdf.

Gargaro:1987:RIA

[GP87] Anthony Gargaro and T. L. Pappas. Reusability issues and Ada. IEEE Software, 4 (4):43–51, July 1987. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garzas:2005:OMD

[GP05] Javier Garzas and Mario Piattini. An ontology for microarchitectural design knowledge. IEEE Software, 22(2):28–33, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garvey:1997:IAR

[GPW97] Paul R. Garvey, Douglas J. Phair, and John A. Wilson. An information architecture for risk assessment and management. IEEE Software, 14(3):25– 34, May/June 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Grady:1990:WPA

[Gra90] Robert B. Grady. Workproduct analysis: The philosopher’s stone of software? IEEE Software, 7(2):26–34, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gray:1997:PWY

[Gra97] Wayne D. Gray. Point:

Who ya gonna call? you’re on your own. IEEE Software, 14(4):26, 28, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4026.
pdf.

Graham:1999:SAQ

[Gra99a] Larry Graham. Softlaw:

Act quickly to avoid losing patents. IEEE Software, 16(2):33–35, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2033.
pdf.

Graham:1999:SLI

[Gra99b] Larry Graham. Softlaw:

Legal implications of operating systems. IEEE Software, 16(1):20–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so1999/pdf/s1020.
                                        pdf.

Graham:1999:STP

[Gra99c]  Larry Graham.  Soft-

law: Trademarks and potholes on the superhighway. IEEE Software, 16(5):19–21, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s5019.
                                        pdf.

Graham:2000:SDS

[Gra00a]  Larry Graham.  Soft-

Law: Debunking software patent myths. IEEE Software, 17(4):122–123, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s4122.
                                        pdf.

Graham:2000:SKY

[Gra00b] Larry Graham. SoftLaw: Keep your bots to yourself. IEEE Software, 17 (6):106–107, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s6106.
                                        pdf.

Graham:2000:SRP

[Gra00c]  Larry Graham.  Soft-

Law: Recent patent reform legislation. IEEE Software, 17(3):101–103, May/June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3101.
pdf.

Graham:2000:SLT

[Gra00d]  Larry Graham.  Soft-

law: The legal tortoise and the technology hare. IEEE Software, 17(5):18–19, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5018.
pdf.

Graham:2002:RRT

[Gra02]  Dorothy Graham.  Re-

quirements: Requirements and testing: Seven missinglink myths. IEEE Software, 19(5):15–17, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5015.
pdf;   http://www.computer.
org/software/so2002/s5015abs.
htm.

Graham:2003:CRA

[Gra03] Ian Graham. The compleat requirements analyste. IEEE Software, 20(6):99–101, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/06/s6099.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/so/2003/06/
                                        s6099abs.htm.

Gras:2004:EED

[Gra04] Jean-Jacques Gras. Endto-end defect modeling. IEEE Software, 21(5):98– 100, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Greenspan:1996:BJL

[Gre96] Sol J. Greenspan. Bookshelf: Jackson’s latest: More than a lexicon. IEEE Software, 13(2):116–117, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Grenning:2001:LEP

[Gre01] James Grenning. Launching Extreme Programming at a process-intensive company. IEEE Software, 18(6):27–33, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6027abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6027.pdf.

Griffin:1995:ILL

[Gri95] William G. Griffin. Insider:

Lessons learned in software reuse. IEEE Software, 12 (4):11, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2008:LSP

[GRM08] Robert L. Glass, Johann Rost, and Matthias S. Matook. Lying on software projects. IEEE Software, 25(6):90–95, November/ December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Groth:2004:SIP

[Gro04] Robert Groth. Is the software industry’s productivity declining? IEEE Software, 21(6):92–94, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
06/s6092.pdf.

Guilford:2002:PPP

[GRS02] Sheila Guilford, Gordon Rugg, and Niall Scott. Pleasure and pain: Perceptual bias and its implications for software engineering. IEEE Software, 19(3):63–69, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3063.
pdf;   http://www.computer.
org/software/so2002/s3063abs.
htm.

Garg:1990:HSM

[GS90a] Pankaj K. Garg and Walt Scacchi. A hypertext system to manage software life-cycle documents. IEEE Software, 7(3):90–98, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Geetha:1990:RNL

[GS90b] T. V. Geetha and R. K. Subramanian. Representing natural language with Prolog. IEEE Software, 7(2):85–92, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garousi:2007:B

[GS07] Vahid Garousi and Art Sedighi. Bookshelf. IEEE Software, 24(5):101–102, September/October 2007.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
05/s5101.pdf.

Goodbrand:2003:LAR

[GSB03] Alan Goodbrand, André Sintzoff, and Mark Bullock. Letters: Asking the right questions; distance learning not delivering — yet; not really pair programming. IEEE Software, 20(1):8–9, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1008.
pdf.

Griswold:2006:MSD

[GSS^{+}06] William G. Griswold, Kevin Sullivan, Yuanyuan Song, Macneil Shonle, Nishit Tewari, Yuanfang Cai, and Hridesh Rajan. Modular software design with crosscutting interfaces. IEEE Software, 23(1):51–60, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:2005:IFR

[GT05] Robert L. Glass and Computing Trends. IT failure rates—70% or 10–15%? IEEE Software, 22(3):112, 110–111, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gibbs:1987:MMF

[GTF^{+}87] Simon Gibbs, Dennis Tsichritzis, Akis Fitas, Dimitri Konstantas, and Yannis Yeorgaroudakis. Muse: a multimedia filing system. IEEE Software, 4(2): 4–15, March 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Guerrieri:1994:CSD

[Gue94] Ernesto Guerrieri. Case study: Digital’s application generator. IEEE Software, 11(5):95–96, September 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Guest:1995:LBI

[Gue95] Ronald A. Guest. Letters: Blame it on the market. IEEE Software, 12(6): 4, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gumm:2006:DDS

[Gum06] Dorina C. Gumm. Distribution dimensions in software development projects: a taxonomy. IEEE Software, 23(5):45–51, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gutknecht:1986:SCM

[Gut86] J. Gutknecht. Separate compilation in Modula-2: an approach to efficient symbol files. IEEE Software, 3(6):29–38, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Grady:1994:KLA

[GV94] Robert B. Grady and Tom Van Slack. Key lessons in achieving widespread inspection use. IEEE Software, 11(4):46–57, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glass:1995:CAD

[GV95] Robert L. Glass and Iris Vessey. Contemporary application-domain taxonomies. IEEE Software, 12(4):63–76, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Goth:2004:N

[GV04] Greg Goth and Danna Voth. In the news. IEEE Software, 21(4):96– 100, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Glinz:2007:GEI

[GW07] Martin Glinz and Roel J. Wieringa. Guest Editors’ introduction: Stakeholders in requirements engineering. IEEE Software, 24(2): 18–20, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gordijn:2006:SDU

[GYvdR06] Jaap Gordijn, Eric Yu, and Bas van der Raadt. e-service design using i* and e3value modeling. IEEE Software, 23(3):26– 33, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Horl:2000:VVC

[HA00] Johann Hörl and Bernhard K. Aichernig. Validating voice communication requirements using lightweight formal methods. IEEE Software, 17(3):21–27, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3021abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s3021.pdf.

Hafiz:2007:OSP

[HAJ07] Munawar Hafiz, Paul Adamczyk, and Ralph E. Johnson. Organizing security patterns. IEEE Software, 24(4):52–60, July/August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hudenpohl:1996:ESM

[HAK^{+}96] John P. Hudenpohl, Stephen J.

Aud, Tahoi M. Khoshooftaar, Edward B. Allen, and Jean Mayrand. Emerald: Software metrics and models on the desktop. IEEE Software, 13(5):56– 60, September 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hall:1990:SMF

[Hal90] J. Anthony Hall. Seven myths of formal methods. IEEE Software, 7(5):11–19, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hall:1995:NQE

[Hal95] Tracy Hall. No quality without equality. IEEE Software, 12(2):101–102, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Haley:1996:SPI

[Hal96a] Thomas J. Haley. Software process improvement at Raytheon. IEEE Software, 13(6):33–41, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hall:1996:UFM

[Hal96b] Anthony Hall. Using formal methods to develop an ATC information system. IEEE Software, 13(2):66– 76, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hall:2009:CDE

[Hal09] Duncan Hall. Career development: The ethical software engineer. IEEE Software, 26(4):9–10, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hamlet:1992:WTT

[Ham92]  Dick Hamlet.  Are we

testing for true reliability? IEEE Software, 9(4): 21–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hansen:2006:PVS

[Han06] Kai T. Hansen. Project visualization for software. IEEE Software, 23(4):84– 92, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harrison:1987:RFI

[Har87] William Harrison. RPDE^{3}:

a framework for integrating tool fragments. IEEE Software, 4(6):46–57, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harrington:1989:ACO

[Har89a] Paul Harrington. Applying customer-oriented quality metrics. IEEE Software, 6(6):71, 74–??, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hartson:1989:UIM

[Har89b]  Rex Hartson.  User-

interface management control and communication. IEEE Software, 6(1):62–70, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harrison:2003:EHI

[Har03a] Warren Harrison. From the Editor: How IEEE Software engineers its content. IEEE Software, 20(5):5–7, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5005.
pdf.

Harrison:2003:ESE

[Har03b] Warren Harrison. From the Editor: Is software engineering as we know it over the hill? IEEE Software, 20(3):5–7, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3005.
pdf.

Harrison:2003:EOA

[Har03c] Warren Harrison. From the Editor: Our 20th anniversary: Bruce Shriver, EIC 1983–1986; Ted Lewis, EIC 1987–1990; Carl Chang, EIC 1991–1994; Al Davis, EIC 1995–1998; Steve McConnell, EIC 1999– 2002; Warren Harrison, EIC 2003–. IEEE Software, 20(6):5–7, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6005.
pdf.

Harrison:2003:EMR

[Har03d] Warren Harrison. From the Editor: The marriage of research and practice. IEEE Software, 20(2):5–7, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

tronic). URL http://www.
computer.org/software/
so2003/s2005.pdf.

Harrison:2003:ESD

[Har03e] Warren Harrison. From the Editor: The software developer as movie icon. IEEE Software, 20(1):5–7, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
homepage/2003/s1eic.htm;
 http://dlib.computer.
org/so/books/so2003/pdf/
s1005.pdf.

Harrison:2003:EWI

[Har03f] Warren Harrison. From the Editor: Whose information is it anyway? IEEE Software, 20(4):5–7, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4005.
pdf.

Harrison:2004:EBP

[Har04a] Warren Harrison. From the Editor: Best practices— who says? IEEE Software, 21(1):8–11, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.

                                        org/comp/mags/so/2004/
                                        01/s1008.pdf.

Harrison:2004:ECO

[Har04b] Warren Harrison. From the Editor: Clueless—and oblivious. IEEE Software, 21(3):5–7, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        03/s3005.pdf.

Harrison:2004:EPS

[Har04c] Warren Harrison. From the Editor: Propaganda and software development. IEEE Software, 21(5):5–7, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        05/s5005.pdf.

Harrison:2004:EDE

[Har04d] Warren Harrison. From the Editor: The dangers of end-user programming. IEEE Software, 21(4):5–7, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        04/s4005.pdf.

Harrison:2004:LOS

[Har04e] Warren Harrison. Learning organizations and the software developer. IEEE Software, 21(2):5–7, March/ April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2004/02/s2005.
pdf.

Harrison:2005:CCJ

[Har05a] Warren Harrison. Constant connectivity: Just because you can doesn’t mean you should. IEEE Software, 22 (2):5–7, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
02/s2005.pdf.

Harrison:2005:DYL

[Har05b]  Warren Harrison.  Do

you learn just in time or just in case? IEEE Software, 22(1):5–7, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
01/s1005.pdf.

Harrison:2005:HTD

[Har05c] Warren Harrison. Hightech disasters. IEEE Software, 22(6):3–5, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
06/s6003.pdf.

Harrison:2005:SW

[Har05d]  Warren Harrison.  The

saboteur within. IEEE Software, 22(4):5–7, July/ August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2005/04/s4005.
pdf.

Harrison:2005:SWS

[Har05e] Warren Harrison. Skinner wasn’t a software engineer. IEEE Software, 22 (3):5–7, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
03/s3005.pdf.

Harrison:2005:WDS

[Har05f] Warren Harrison. What do software developers need to know about business? IEEE Software, 22(5):5–7, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.

org/comp/mags/so/2005/
05/s5005.pdf.

Harrison:2006:BRF

[Har06a] Warren Harrison. Building references for the future. IEEE Software, 23 (2):5–7, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
02/s2005.pdf.

Harrison:2006:CMS

[Har06b] Warren Harrison. Content mismanagement systems. IEEE Software, 23(1):5– 8, January/February 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
01/s1005.pdf.

Harrison:2006:EYO

[Har06c] Warren Harrison. Eating your own dog food. IEEE Software, 23(3):5– 7, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
03/s3005.pdf.

Harrison:2006:GB

[Har06d] Warren Harrison. Giving back. IEEE Software, 23(6):5–7, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
06/s6005.pdf.

Harrison:2006:PP

[Har06e] Warren Harrison. Passwords and passion. IEEE Software, 23(4):5–7, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2006/04/s4005.
pdf.

Harrison:2006:TTT

[Har06f] Warren Harrison. Technology transfer and the tech broker. IEEE Software, 23(5):5–7, September/October 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
05/s5005.pdf.

Hassler:2005:OSL

[Has05]  Vesna Hassler.  Open

Source libraries for information retrieval. IEEE Software, 22(5):78–82, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hatton:1995:QTS

[Hat95] Les Hatton. Quality time:

Static inspection: Tapping the wheels of software. IEEE Software, 12(3):85– 87, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See letter and response in July 1995 issue, pp. 7, 8, 10.

Hatton:1997:RFD

[Hat97a]  L. Hatton.  Reexamin-

ing the fault density component size connection. IEEE Software, 14(2):89– 97, March/April 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See full text of corrected article

at http://www.computer.
org/pubs/software/s2hat.
pdf.

Hatton:1997:VDV

[Hat97b]  Les Hatton.  N-version

design vs. one good version. IEEE Software, 14(6):71–76, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6071.
pdf;   http://dlib.computer.
org/so/books/so1997/pdf/
s6071abs.htm.

Hatton:1998:DOS

[Hat98] Les Hatton. Does OO sync with how we think? IEEE Software, 15(3):46– 54, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1998/s3046abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s3046.pdf.

Hatton:2001:ERD

[Hat01] Les Hatton. Exploring the role of diagnosis in software failure. IEEE Software, 18(4):34–39, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s4034abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s4034.pdf.

Hatton:2008:TVC

[Hat08] Les Hatton. Testing the value of checklists in code inspections. IEEE Software, 25(4):82–88, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hayes:2003:DYL

[Hay03] Jane Huffman Hayes. Do you like piña coladas? how improved communication can improve software quality. IEEE Software, 20(1):90–92, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1090.
pdf.

Hayes:2004:VK

[Hay04] Jane Huffman Hayes. On the virtues of not knowing. IEEE Software, 21 (3):74–75, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harrison:2007:UPC

[HAZ07] Neil B. Harrison, Paris Avgeriou, and Uwe Zdun. Using patterns to capture architectural decisions. IEEE Software, 24(4):38– 45, July/August 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Huang:2006:HMS

[HB06] LiGuo Huang and Barry Boehm. How much software quality investment is enough: a value-based approach. IEEE Software, 23(5):88–95, September/ October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hassenzahl:2001:EJ

[HBB01] Marc Hassenzahl, Andreas Beu, and Michael Burmester. Engineering joy. IEEE Software, 18(1):70–76, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s1070abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s1070.pdf.

Hevner:1992:ICC

[HBP92] Alan R. Hevner, Shirley A. Becker, and Lenard B. Pedowitz. Integrated CASE for cleanroom development. IEEE Software, 9(2):69– 76, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Humphrey:1991:CCL

[HC91] Watts S. Humphrey and Bill Curtis. Comment on “A Critical Look at Software Capability Evaluations”. IEEE Software, 8 (4):42–46, July 1991. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See [BM91].

Hall:2002:CCD

[HC02] Anthony Hall and Roderick Chapman. Correctness by construction: Developing a commercial secure system. IEEE Software, 19(1):18–25, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1018abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/

s1018.pdf. See correction [Ano02j].

Hardgrave:2000:COO

[HD00] Bill C. Hardgrave and E. Reed Doke. Cobol in an object-oriented world: a learning perspective. IEEE Software, 17(2):26– 29, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2026abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2026.pdf.

Higgins:2003:MRM

[HdLGG03] Stewart A. Higgins, Maurice de Laat, Paul M. C. Gieles, and Emilienne M. Geurts. Managing requirements for medical IT products. IEEE Software, 20(1):26–33, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so2003/pdf/s1026.
                                        pdf;   http://www.computer.
                                        org/software/so2003/s1026abs.
                                        htm.

Hailpern:1986:MRS

[HDR^{+}86] Brent Hailpern, Roger Dannenberg, Dean Rubine, Bjarne Stroustrup, Prabhaker Mateti, Frances Hunt, George W. Ernst, Gerald Radack, Yves Caseau, Mario Tokoro, Yutaka Ishikawa, Mark R. Laff, Yonathan Malachi, Zohar Manna, and Richard Waldinger. Multiparadigm research: a survey of nine projects. IEEE Software, 3(1):70–77, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hayes:2005:IAF

[HDS05] Jane Huffman Hayes, Alex Dekhtyar, and Senthil Karthekeyan Sundaram. Improving after-the-fact tracing and mapping: Supporting software quality predictions. IEEE Software, 22(6):30– 37, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Heath:1991:VPP

[HE91] Michael Heath and Jennifer Etheridge. Visualizing the performance of parallel programs. IEEE Software, 8 (5):29–39, September 1991.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hecker:1999:SSB

[Hec99] Frank Hecker. Setting up shop: The business of opensource software. IEEE Software, 16(1):45–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s1045abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s1045.pdf.

Hefley:1995:HUH

[Hef95]  William Hefley.  Help-

ing users help themselves. IEEE Software, 12(2):93– 95, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hayes:2006:SSQ

[HEH06] Jane Hayes, Nancy Eickelmann, and E. Ashlee Holbrook. Security and software quality: An interview with Frank Perry. IEEE Software, 23(3):12– 15, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hein:1998:CWB

[Hei98]  Fred Hein.  Culture

at work: The blame game. IEEE Software, 15(6):89–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6089.
pdf.

Heires:2001:WDL

[Hei01] James T. Heires. What I did last summer: a software development benchmarking case study. IEEE Software, 18(5):33–39, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5033abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5033.pdf.

Henninger:1994:UIR

[Hen94]  Scott Henninger.  Us-

ing iterative refinement to find reusable software. IEEE Software, 11(5):48– 59, September 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hendrickson:1999:CWS

[Hen99] Elisabeth Hendrickson. Culture at work: Signs of a healthy startup. IEEE Software, 16(3):13–14, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3013.
pdf.

Henry:1995:LSI

[HF95] Emmanuel Henry and Benoit Faller. Largescale industrial reuse to reduce cost and cycle time. IEEE Software, 12(5):47– 53, September 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hall:1997:IES

[HF97] T. Hall and N. Fenton. Implementing effective software metrics programs. IEEE Software, 14(2):55– 65, March/April 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hendrickson:2002:SEI

[HF02] Elizabeth Hendrickson and Martin Fowler. The software engineering of Internet software: Guest Editors’ introduction. IEEE Software, 19(2):23–24, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2023.
pdf;   http://www.computer.
org/software/so2002/s2023abs.
htm.

Herbsleb:1999:ACD

[HG99] James D. Herbsleb and Rebecca E. Grinter. Architectures, coordination, and distance: Conway’s law and beyond. IEEE Software, 16(5):63–70, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5063abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5063.pdf.

Hantos:2000:ISP

[HG00] Peter Hantos and Mario Gisbert. Identifying software productivity improvement approaches and risks: Construction industry case study. IEEE Software, 17(1):48–56, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s1048abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s1048.pdf.

Harel:2009:FET

[HGK09] David Harel and Michal Gordon-Kiwkowitz. FEA-TURE: Education: On teaching visual formalisms. IEEE Software, 26(3):87– 95, May/June 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hayhurst:2002:NAS

[HH02a] Kelly J. Hayhurst and C. Michale Holloway. In the news: Aviation software guidelines. IEEE Software, 19(5):107, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5104.
pdf;   http://www.computer.
org/software/so2002/s5104abs.
htm.

Hilburn:2002:ICS

[HH02b] Thomas B. Hilburn and Watts S. Humphrey. The impending changes in software education. IEEE Software, 19(5):22–24, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5022.
pdf;   http://www.computer.
org/software/so2002/s5022abs.
htm.

Hilburn:2002:TT

[HH02c] Thomas B. Hilburn and Watts S. Humphrey. Teaching teamwork. IEEE Software, 19(5):72–77, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s5072.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s5072abs.
                                        htm.

Hallen:1999:LW

[HHJC99] Jacob Hallén, Anders Hammarqvist, Fredrik Juhlin, and Anders Chrigström. Linux in the workplace. IEEE Software, 16(1): 52–??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://computer.org/software/
                                        so1999/s1052abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s1052.pdf.

Henry:1994:ISM

[HHM94] Joel Henry, Sallie Henry, and Lance Matheson. Improving software maintenance at Martin Marietta. IEEE Software, 11(4):67– 75, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hilburn:1997:SEM

[Hil97] Thomas B. Hilburn. SE education: a modest proposal. IEEE Software, 14(6):44–48, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6044.
pdf;   http://dlib.computer.
org/so/books/so1997/pdf/
s6044abs.htm.

Hirsh:2005:UPR

[Hir05] Barbara Hirsh. Using positive reinforcement as a quality tool. IEEE Software, 22(2):62–63, March/ April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hirschi:2007:TLL

[Hir07] Ashwin Hirschi. Traveling light, the Lua way. IEEE Software, 24(5):31– 38, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hix:1990:GUI

[Hix90] Deborah Hix. Generations of user-interface management systems. IEEE Software, 7(5):77–87, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hummel:2008:CCP

[HJA08] Oliver Hummel, Werner Janjic, and Colin Atkinson. Code conjurer: Pulling reusable software out of thin air. IEEE Software, 25(5):45–52, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Halang:1994:SAP

[HK94] Wolfgang A. Halang and Bernd J. Kramer. Safety assurance in process control. IEEE Software, 11(1):61– 67, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Honiden:1993:FSM

[HKK93] Shinichi Honiden, Nobuto Kotaka, and Yoshinori Kishimoto. Formalizing specification modeling in OOA. IEEE Software, 10(1):54–66, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Horowitz:1985:SAG

[HKN85] Ellis Horowitz, Alfons Kemper, and Balaji Narasimhan. Survey of application generators. IEEE Software, 2(1):40–54, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Heeks:2001:SSG

[HKNS01] Richard Heeks, S. Krishna, Brian Nicholson, and Sundeep Sahay. Synching or sinking: Global software outsourcing relationships. IEEE Software, 18(2):54– 60, March/April 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2054abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2054.pdf.

Horch:1996:BTL

[HKPS96] John W. Horch, Joseph M.

Kusmiss, Kasilingam Periyasamy, and Daniel J. Salomon. Bookshelf: Thorough look at software inspection; informal survey of software fault tolerance in distributed systems; highlevel text on higher order logic; achieving little goals in language design and implementation. IEEE Software, 13(1):116–119, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harrison:1990:TMC

[HKR^{+}90] Warren Harrison, Bernd Kramer, Walter Rudd, Sol Shatz, Carl Chang, Zary Segall, David Clemmer, Joel Williamson, Ben Peek, Bill Appelbe, Kevin Smith, and Adam Kolawa. Tools for multiple-CPU environments. IEEE Software, 7 (3):45–51, May 1990. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Helmbold:1985:DAT

[HL85] David Helmbold and David Luckham. Debugging Ada tasking programs. IEEE Software, 2(2):47– 57, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Holland:1999:CSF

[HL99] Christopher P. Holland and Ben Light. A critical success factors model for ERP implementation. IEEE Software, 16(3):30– 36, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s3030abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s3030.pdf.

Hofmann:2001:RES

[HL01] Hubert F. Hofmann and Franz Lehner. Requirements engineering as a success factor in software projects. IEEE Software, 18(4):58–66, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s4058abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s4058.pdf.

Hagge:2005:SRE

[HL05] Lars Hagge and Kathrin Lappe. Sharing requirements engineering experience using patterns. IEEE Software, 22(1):24–31, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harbert:1990:GSS

[HLS90] Andrew Harbert, William Lively, and Sallie Sheppard. A graphical specification system for userinterface design. IEEE Software, 7(4):12–20, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Horgan:1992:ATT

[HM92] Joseph R. Horgan and Aditya P. Mathur. Assessing testing tools in research and education. IEEE Software, 9(3):61–69, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Herbsleb:2001:GEI

[HM01a] James D. Herbsleb and Deependra Moitra. Guest Editors’ introduction: Global software development. IEEE Software, 18(2):16–20, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2016.
pdf;   http://dlib.computer.
org/so/books/so2001/s2016abs.
htm.

Holmberg:2001:SPF

[HM01b] Lena Holmberg and Lars Mathiassen. Survival patterns in fast-moving software organizations. IEEE Software, 18(6):51–55, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6051abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6051.pdf.

Hieatt:2002:GFT

[HM02a] Edward Hieatt and Robert Mee. Going faster: Testing the Web application. IEEE Software, 19(2):60– 65, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2060.
pdf;   http://www.computer.
org/software/so2002/s2060abs.
htm.

Hoglund:2002:PCS

[HM02b] Greg Hoglund and Gary McGraw. Point/counterpoint: Security band-aids: More cost-effective than “secure” coding; building secure software: Better than protecting bad software. IEEE Software, 19(6):56–59, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6056.
pdf.

Haase:1994:BFT

[HMK^{+}94] Volkmar Haase, Richard Messnarz, G. Koch, H. J. Kugler, and Paul Decrinis. Bootstrap: Finetuning process assessment. IEEE Software, 11(4):25– 35, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hess:2001:LTT

[HMP^{+}01] Narada Hess, Keith Miller, Bret Pettichord, Cem Kaner, Brian Marick, Keith Miller, Bob Ferguson, Bill Brykczynski, Watts S. Humphrey, Bill Krosky, Aditya Mathur, and Bob Glass. Letters: Testing troubles; art, craft, and cowboys; team training; math — a good but ineffective exercise. IEEE Software, 18(3):8– 13, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so2001/pdf/s3008.
                                        pdf.

Harandi:1990:KBP

[HN90] Mehdi T. Harandi and Jim Q. Ning. Knowledgebased program analysis. IEEE Software, 7(1):74–81, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Haibert:1987:UTI

[HO87] Daniel C. Haibert and Patrick D. O’Brien. Using types and inheritance in object-oriented programming. IEEE Software, 4 (5):71–79, September 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ho:1989:STT

[Ho89]  Irving T. Ho.  Soft-

ware technology in Taiwan. IEEE Software, 6(2):68– 73, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hoare:1984:PSS

[Hoa84] C. A. R. Hoare. Programming: Sorcery or science? IEEE Software, 1(2):5–12, 15–16, April 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hohmann:1999:MCR

[Hoh99]  Luke Hohmann.  Man-

ager: Coaching the rookie manager. IEEE Software, 16(1):16–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1016.
pdf.

Hohmann:2003:DBM

[Hoh03] Luke Hohmann. The difference between marketecture and tarchitecture. IEEE Software, 20(4):51– 53, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4051.
pdf.

Hohpe:2005:YCS

[Hoh05] Gregor Hohpe. Your coffee shop doesn’t use two-phase commit. IEEE Software, 22(2):64–66, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Holzmann:1992:PDR

[Hol92] Gerard J. Holzmann. Protocol design: Redefining the state of the art. IEEE Software, 9(1):17–22, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Holzinger:2004:RPV

[Hol04] Andreas Holzinger. Rapid prototyping for a virtual medical campus interface. IEEE Software, 21(1): 92–99, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Honekamp:2009:STA

[Hon09] Uwe Honekamp. Software technology: The Autosar XML schema and its relevance for Autosar tools. IEEE Software, 26(4):73– 76, July/August 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Horspool:1987:AGG

[Hor87] R. Nigel Horspool. Alternative to the Graham-Glanville code-generation method. IEEE Software, 4 (3):33–39, May 1987. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Horch:1995:TOH

[Hor95] John W. Horch. Two orientations on how to work with objects. IEEE Software, 12(2):117–118, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Horch:1996:BWB

[Hor96] John W. Horch. Bookshelf:

a whale of a book on analysis and how to make it work. IEEE Software, 13 (2):119, March 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Horowitz:1997:CAI

[Hor97]  Ellis Horowitz.  Call

for articles: IEEE Software special issue: Migrating software to the World Wide Web. IEEE Software, 14(4):50, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4100.
pdf.

Horch:1998:BSM

[Hor98a] John W. Horch. Bookshelf: a Sharpe modeling approach. IEEE Software, 15(4):88–89, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4087.
pdf.

Horowitz:1998:GEI

[Hor98b] Ellis Horowitz. Guest Editor’s introduction:: Migrating software to the World Wide Web. IEEE Software, 15(3):18–21, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3018abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3018.pdf.

Horstmann:2000:CCV

[Hor00] Cay S. Horstmann. Counterpoint: Cobol versus Java. IEEE Software, 17(2):23, 25, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2022.
pdf.

House:1995:LNQ

[Hou95] Ron House. Letters: No “quality” in twisted attack on men. IEEE Software, 12 (4):5–6, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Houlihan:2001:BDP

[Hou01] Peter Houlihan. Bookshelf: Design patterns for the enterprise: Enterprise Modeling with UML: Designing Successful Software Through Business Analysis. IEEE Software, 18(1):106–107, January/ February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1105.
pdf.

Howden:1985:TPF

[How85] William E. Howden. Theory and practice of functional testing. IEEE Software, 2(5):6–17, September 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hunter:2000:LEN

[HPG^{+}00] John A. Hunter, Bill Pyritz, Thomas Gilb, Walter Royce, and Ruth Buys. Letters: Exploring new avenues; inspection hype?; CMM in the small. IEEE Software, 17(6):6–8, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6006.
pdf.

Hausler:1990:UFA

[HPLH90] Philip A. Hausler, Mark G.

Pleszkoch, Richard C. Linger, and Alan R. Hevner. Using function abstraction to understand program behavior. IEEE Software, 7(1):55–63, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hale:2000:ECE

[HPSD00] Joanne Hale, Allen Parrish, Randy K. Smith, and Brandon Dixon. Enhancing the Cocomo estimation models. IEEE Software, 17(6):45–50, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s6045abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s6045.pdf.

Hartmann:1990:TSR

[HR90] Jean Hartmann and David J. Robson. Techniques for selective revalidation. IEEE Software, 7(1):31–36, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Huntbach:1995:PCL

[HR95] Matthew M. Huntbach and Graem A. Ringwood. Programming in concurrent logic languages. IEEE Software, 12(6):71– 82, November 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Helsen:2008:WMJ

[HRS08] Simon Helsen, Arthur Ryman, and Diomidis Spinellis. Where’s my jetpack? IEEE Software, 25(5):18–21, September/ October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Henry:1990:PSC

[HS90] Sallie Henry and Calvin Selig. Predicting sourcecode complexity at the design stage. IEEE Software, 7(2):36–44, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harrold:1991:SUD

[HS91] Mary Jean Harrold and Mary Lou Soffa. Selecting and using data for integration testing. IEEE Software, 8(2):58–65, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Henderson-Sellers:1997:MOP

[HS97] Brian Henderson-Sellers.

Manager: OO project management: The need for process. IEEE Software, 14(4):96–97, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so1997/pdf/s4096.
pdf.

Henderson-Sellers:2000:OFE

[HS00] Brian Henderson-Sellers. The OPEN framework for enhancing productivity. IEEE Software, 17(2):53– 58, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2053abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2053.pdf.

Hodges:1989:CSM

[HSA89] Matthew E. Hodges, Russell M. Sasnett, and Mark S. Ackerman. Construction set for multimedia applications. IEEE Software, 6(1):37–43, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hall:2008:WDW

[HSB^{+}08] Tracy Hall, Helen Sharp, Sarah Beecham, Nathan Baddoo, and Hugh Robinson. What do we know about developer motivation? IEEE Software, 25(4):92–94, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/

mags/so/2008/extras/mso2008040092x1.
pdf.

Hsia:1994:FAS

[HSG^{+}94] Pei Hsia, Jayarajan Samuel, Jerry Gao, David Kung, Yasufumi Toyoshima, and Cris Chen. Formal approach to scenario analysis. IEEE Software, 11(2):33– 41, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hsia:1993:LPL

[Hsi93] Pei Hsia. Learning to put lessons into practice. IEEE Software, 10(5):14– ??, September 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hsia:1996:MSD

[Hsi96] Pei Hsia. Making software development visible. IEEE Software, 13(3):23– 26, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Humphrey:1991:SPI

[HSW91] Watts S. Humphrey, Terry R.

Snyder, and Ronald R. Willis. Software process improvement at Hughes Aircraft. IEEE Software, 8(4): 11–23, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hunt:2002:SA

[HT02a] Andy Hunt and Dave Thomas. Software archaeology. IEEE Software, 19(2):20–22, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2020.
pdf;   http://www.computer.
org/software/so2002/s2020abs.
htm.

Hunt:2002:SCZ

[HT02b] Andy Hunt and Dave Thomas. Software construction: Zero-tolerance construction. IEEE Software, 19(5):100–102, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5100.
pdf.

Hunt:2002:UA

[HT02c] Andy Hunt and Dave Thomas. Ubiquitous automation. IEEE Software, 19(1):11–13, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1011abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1011.pdf.

Hunt:2003:PRM

[HT03a] Andy Hunt and Dave Thomas. Preparing the raw material. IEEE Software, 20(5):97–98, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5097.
pdf.

Hunt:2003:TPD

[HT03b] Andy Hunt and Dave Thomas. The trip-packing dilemma. IEEE Software, 20(3):106–107, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3106.
pdf.

Hunt:2003:AE

[HT03c] Andy Hunt and David Thomas. The art of enbugging. IEEE Software, 20(1):10–11, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1010.
pdf.

Hunt:2004:I

[HT04a] Andy Hunt and Dave Thomas. Imaginate.

IEEE Software, 21(5):96– 97, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hunt:2004:OOS

[HT04b] Andy Hunt and Dave Thomas. OO in one sentence: Keep it DRY, shy, and tell the other guy. IEEE Software, 21(3):101– 103, May/June 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hunt:2004:TLN

[HT04c] Andy Hunt and Dave Thomas. Three legs, no wobble. IEEE Software, 21(1):18–19, 22, January/ February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hen-Tov:2009:MWE

[HTLPS09] Atzmon Hen-Tov, David H. Lorenz, Assaf Pinhasi, and Lior Schachter. ModelTalk: When everything is a domain-specific language. IEEE Software, 26(4):39– 46, July/August 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hudak:1988:EPP

[Hud88] Paul Hudak. Exploring parafunctional programming: Separating the what from the how. IEEE Software, 5(1):54–61, January 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Humphrey:1988:CSP

[Hum88] Watts S. Humphrey. Characterizing the software process: a maturity framework. IEEE Software, 5(2):73–79, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Humphrey:1995:MMP

[Hum95] Watts S. Humphrey. Manager: Making process improvement personal. IEEE Software, 12(5):82–83, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Humphrey:1996:UDM

[Hum96] Watts S. Humphrey. Using a defined and measured personal software process. IEEE Software, 13(3):77– 88, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Humphrey:2000:PSP

[Hum00] Watts S. Humphrey. The personal software process: Status and trends: Guest Editor’s introduction. IEEE Software, 17(6):71–75, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6071.
pdf.

Humphrey:2001:EWT

[Hum01] Watts S. Humphrey. Engineers will tolerate a lot of abuse. IEEE Software, 18(5):13–15, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5013abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5013.pdf.

Harpold:2000:LCE

[HVF^{+}00] Jim Harpold, Daniel Vijay, Francesco G. Fantauzzi, Alan M. Davis, Thomas Allen, David B. Anderson, Robert L. Glass, David Jameson, and Greg A. Woods. Letters: Cobol exists!; where do we go from here?; leadership not management; opening the gates of discussion. IEEE Software, 17(4):8–11, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4008.
pdf.

Hatton:2009:PC

[HvG09] Les Hatton and Michiel van Genuchten. Point/ counterpoint. IEEE Software, 26(3):66–69, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Islam:1997:FSS

[IAJR97] Nayeem Islam, Rangachari Anand, Trent Jaeger, and Josyula R. Rao. A flexible security system for using Internet content. IEEE Software, 14(5):52–59, September/ October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5052.
pdf;   http://www.computer.
org/software/so1997/s5052abs.

htm. Discusses the FlexxGuard security system built using the Java Development Kit 1.1.

Ielse:1996:IIB

[Iel96] Jakob Ielse. Interface — the importance of being beautiful. IEEE Software, 13(1):92–94, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ichikawa:1990:IPW

[IH90] Tadao Ichikawa and Masahito Hirakawa. Iconic programming: Where to go? IEEE Software, 7(6):63–68, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ingalsbe:2008:TMD

[IKBM08] Jeffrey A. Ingalsbe, Louis Kunimatsu, Tim Baeten, and Nancy R. Mead. Threat modeling: Diving into the deep end. IEEE Software, 25(1):28–34, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ingebretsen:2008:CUC

[Ing08] Mark Ingebretsen. Currents: Unconferences catch on with developers. IEEE Software, 25(6):108–110, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Irvine:1997:NCC

[Irv97] Cynthia E. Irvine. In the news: Challenges in computer security education. IEEE Software, 14 (5):110–111, September/ October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5106.
pdf.

Isoda:1994:SEA

[IS94] Sadahiro Isoda and Motoshi Saeki. Software engineering in Asia. IEEE Software, 11(6):63–68, November 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Isoda:1987:VVD

[ISO87] Sadahiro Isoda, Takao Shimomura, and Yuji Ono. VIPS: a visual debugger. IEEE Software, 4(3):8– 19, May 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jaaksi:2002:DMB

[Jaa02] Ari Jaaksi. Developing mobile browsers in a product line. IEEE Software, 19(4):73–80, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4073.
pdf;   http://www.computer.
org/software/so2002/s4073abs.
htm.

Jacobson:1993:OTS

[Jac93] Ivar Jacobson. Is object technology software’s industrial platform? IEEE Software, 10(1):24–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jackson:1994:PMS

[Jac94] Michael Jackson. Problems, methods, and specialization. IEEE Software, 11 (6):57–62, November 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jackson:1995:QTC

[Jac95] Michael Jackson. Quality time: Critical reading for software developers (book excerpt). IEEE Software, 12(6):103–104, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jackson:1996:CRN

[Jac96] Daniel Jackson. Counterpoint: Requirements need form, maybe formality. IEEE Software, 13(2):21– 22, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jackman:1998:CWH

[Jac98a] Michele Jackman. Culture at work: Homeopathic remedies to team toxicity. IEEE Software, 15(4):43– 45, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/

books/so1998/pdf/s4043.
pdf.

Jackson:1998:LOD

[Jac98b] Michael Jackson. Loyal opposition: Defining a discipline of description. IEEE Software, 15(5):14– 16, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1998/pdf/
s5014.pdf.

Jackson:1998:WTE

[Jac98c]  Michael Jackson.  Will

there ever be software engineering? IEEE Software, 15(1):36–39, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1036abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s1036.pdf.

Jackson:1999:LOS

[Jac99] Michael Jackson. Loyal opposition: Specializing in software engineering. IEEE Software, 16(6): 120–121, 119, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so1999/pdf/s6120.
                                        pdf.

Jackson:2004:SMW

[Jac04]  Michael Jackson.  See-

ing more of the world. IEEE Software, 21(6):83– 85, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jakobsen:1998:BPI

[Jak98] Allan Baktoft Jakobsen. Bottom-up process improvement tricks. IEEE Software, 15(1):64–68, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1998/s1064abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s1064.pdf.

Jakobsen:2000:SPL

[Jak00] Allan Baktoft Jakobsen. Software processes: Live and let die. IEEE Software, 17(3):71–75, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s3071abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s3071.pdf.

Jalics:1989:RPP

[Jal89] Paul J. Jalics. Realizing the performance potential of Cobol. IEEE Software, 6 (5):70–79, September 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Janert:2004:ITD

[Jan04] Philipp K. Janert. Introducing test-driven software development. IEEE Software, 21(6):100–101, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
06/s6100.pdf.

Jarke:1992:SIC

[Jar92] Matthias Jarke. Strategies for integrating CASE environments. IEEE Software, 9(2):54–61, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Johnson:1996:RRS

[JB96] Donna L. Johnson and Judith G. Brodman. Realities and rewards of software process improvement. IEEE Software, 13(6):99– 101, November 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Johnson:2000:ACP

[JB00] Donna L. Johnson and Judith G. Brodman. Applying CMM project planning practices to diverse environments. IEEE Software, 17(4):40–47, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4040abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4040.pdf.

Jorgensen:2004:EUC

[JB04] Jens Baek Jorgensen and Claus Bossen. Executable use cases: Requirements for a pervasive health care system. IEEE Software, 21(2): 34–41, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jansen:2008:OSS

[JBHD08] Slinger Jansen, Sjaak Brinkkemper, Ivo Hunink, and Cetin Demir. Opportunistic software systems development: Pragmatic and opportunistic reuse in innovative startup companies. IEEE Software, 25(6):42–49, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jacobson:1999:EUS

[JBR99] Ivar Jacobson, Grady Booch, and James Rumbaugh. Excerpt from The Unified Software Development Process: The unified process. IEEE Software, 16(3):82–90, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s3082abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s3096.pdf.

Jorgensen:2009:VSD

[JBR09] Magne Jørgensen, Barry Boehm, and Stan Rifkin. Viewpoints: Software development effort estimation: Formal models or expert judgment? IEEE Software, 26(2):14–19, March/ April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jackson:2000:RAT

[JC00] Daniel Jackson and John Chapin. Redesigning air traffic control: An exercise in software design. IEEE Software, 17(3):63– 70, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3063abs.htm;

                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s3063.pdf.

Jones:1997:LDM

[JD97] Capers Jones and Al Davis. Letters: Did measurement issue miss the mark? IEEE Software, 14(5):8, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s5008.
                                        pdf.

Johnson:1998:QTP

[JD98] Philip M. Johnson and Anne M. Disney. Quality time: The personal software process: a cautionary case study. IEEE Software, 15(6):85–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s6085.
                                        pdf.

Jezequel:2001:ROL

[JDT01] Jean-Marc Jézéquel, Daniel Deveaux, and Yves Le Traon. Reliable objects: Lightweight testing for OO languages. IEEE Software, 18(4):76–83, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s4076abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s4076.pdf.

Jeffries:1997:IDI

[Jef97] Robin Jeffries. Interface:

Designing interfaces for programmers. IEEE Software, 14(6):89–91, November/December 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6089.
pdf.

Jorgensen:2008:AIM

[JG08] Magne Jørgensen and Stein Grimstad. Avoiding irrelevant and misleading information when estimating development effort. IEEE Software, 25(3):78– 83, May/June 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jenkins:1986:PSN

[JGM86] Michael A. Jenkins, Janice I. Glasgow, and Carl D. McCrosky. Programming styles in Nial. IEEE Software, 3(1):46–55, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Johnson:2007:IPT

[JHMW07] Michael J. Johnson, Chih-Wei Ho, E. Michael Maximilien, and Laurie Williams. Incorporating performance testing in test-driven development. IEEE Software, 24 (3):67–73, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jung:2004:MSP

[JKC04] Ho-Won Jung, Seung-Gweon Kim, and Chang-Shin Chung. Measuring software product quality: a survey of ISO/ IEC 9126. IEEE Software, 21(5):88–92, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jaspan:2009:SME

[JKM^{+}09] Ciera Jaspan, Michael Keeling, Larry Maccherone, Gabriel L. Zenarosa, and Mary Shaw. Software mythbusters explore formal methods. IEEE Software, 26(6):60–63, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Johnson:2005:ISD

[JKP^{+}05] Philip M. Johnson, Hongbing Kou, Michael Paulding, Qin Zhang, Aaron Kagawa, and Takuya Yamashita. Improving software development management through software project telemetry. IEEE Software, 22(4):76– 85, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Janert:2003:BAC

[JL03] Philipp K. Janert and Robert C. Larrabee. Bookshelf: Art, craft, science— how about profession?; history of Bell Labs. IEEE Software, 20(6):108–111, November/December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/dl/mags/so/2003/06/
s6108.pdf;   http://csdl.
computer.org/dl/mags/
so/2003/06/s6108abs.htm.

Review of Software Craftsmanship by Pete McBreen and Bell Labs: Life in the Crown Jewel by Narain Gehani.

Jirotka:2006:SRV

[JL06] Marina Jirotka and Paul Luff. Supporting requirements with video-based analysis. IEEE Software, 23 (3):42–44, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jackson:2003:LBT

[JLLP03] Adam Jackson, Michael Lutz, Phil Laplante, and Scott Preece. Letters: Benchmarks are a tricky thing; requirements aren’t the whole picture — or play. IEEE Software, 20(2):8–11, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s2008.
                                        pdf.

Juristo:2002:QTR

[JM02] Natalia Juristo and Ana M. Moreno. Quality time: Reliable knowledge for software development. IEEE Software, 19(5):98–99, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s5098.
                                        pdf.

Jeffries:2007:GEI

[JM07] Ron Jeffries and Grigori Melnik. Guest Editors’ introduction: TDD– the art of fearless programming. IEEE Software, 24 (3):24–30, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.

org/comp/mags/so/2007/
03/s3024.pdf.

Johnson:2000:EGS

[JMDB00] Philip M. Johnson, Carleton A. Moore, Joseph A. Dane, and Robert S. Brewer. Empirically guided software effort guesstimation. IEEE Software, 17(6):51–56, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6051abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6051.pdf.

Juristo:2000:HUL

[JML00] Natalia Juristo, Ana Maria Moreno, and Marta López. How to use linguistic instruments for objectoriented analysis. IEEE Software, 17(3):80–89, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3080abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s3080.pdf.

Janert:2004:B

[JML04] Philipp K. Janert, Shantha Mohan, and Robert C. Larrabee. Bookshelf. IEEE Software, 21(1):104–106, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Juristo:2002:EIM

[JMS02] Natalia Juristo, Ana Moreno, and Andrés Silva. Is the European industry moving toward solving requirements engineering problems? IEEE Software, 19(6):70–77, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s6070.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s6070abs.
                                        htm.

Juristo:2006:GEI

[JMS06] Natalia Juristo, Ana M. Moreno, and Wolfgang Strigel. Guest Editors’ introduction: Software testing practices in industry. IEEE Software, 23(4):19– 21, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        04/s4019.pdf.

Juristo:2006:SWW

[JMVS06] Natalia Juristo, Ana M. Moreno, Sira Vegas, and Martín Solari. In search of what we experimentally know about unit testing. IEEE Software, 23(6):72– 80, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Juristo:2009:VEL

[JMVS09] Natalia Juristo, Ana Moreno, Sira Vegas, and Forrest Shull. Voice of evidence: a look at 25 years of data. IEEE Software, 26(1): 15–17, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Johnson:1995:LML

[Joh95] Bruce Johnson. Letters:

More lessons from civil engineering. IEEE Software, 12(6):6, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Johnson:2000:LLL

[JOL^{+}00] Bob Johnson, Hideaki Okada, Marty Leisner, Bill Dietrich, Steve Mulaik, Merlin Dorfman, Warren Montgomery, Steve Mc-Connell, David Gelperin, Tom Adams, David Lorge Parnas, Pierre Bourque, Robert Dupuis, Alain Abran, and James W. Moore. Letters: To license or not to license; good design versus bad software; nuking Brooks’ Law; raising questions; design is not enough; straightening out the record. IEEE Software, 17(1):6–9, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s1006.
                                        pdf.

Jones:1994:GSS

[Jon94] Capers Jones. Globalization of software supply and demand. IEEE Software, 11 (6):17–24, November 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jones:1996:OWC

[Jon96] Capers Jones. Our worst current development practices. IEEE Software, 13 (2):102–104, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jones:1999:EYU

[Jon99] Capers Jones. The Euro, Y2K, and the US software labor shortage. IEEE Software, 16(3):55–61, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s3055abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s3055.pdf.

Jones:2003:VSD

[Jon03] Capers Jones. Variations in software development practices. IEEE Software, 20(6):22–27, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6022.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6022abs.htm.

Jones:2005:TAD

[Jon05] Steve Jones. Toward an acceptable definition of service. IEEE Software, 22 (3):87–93, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jones:2007:DPL

[Jon07] Peter Edward Jones. Do programming languages make software too soft? IEEE Software, 24(3):120, 118–119, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Joos:1994:SRM

[Joo94]  Rebecca Joos.  Soft-

ware reuse at Motorola. IEEE Software, 11(5):42– 47, September 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jorgensen:2005:PGE

[Jor05] Magne Jorgensen. Practical guidelines for expertjudgment-based software effort estimation. IEEE Software, 22(3):57–63, May/ June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jorgensen:2009:HAS

[Jør09] Magne Jørgensen. How to avoid selecting bids based on overoptimistic cost estimates. IEEE Software, 26 (3):79–84, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jacob:2003:SPC

[JP03] Alice Leslie Jacob and S. K. Pillai. Statistical process control to improve coding and code review. IEEE Software, 20(3):50– 55, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3050.
                                        pdf;   http://www.computer.
                                        org/software/so2003/s3050abs.
                                        htm.

Jimenez:2009:HDS

[JRS^{+}09] Manuel Jiménez, Francisca Rosique, Pedro Sánchez, Bárbara Álvarez, and Andrés Iborra. Habitation: a domain-specific language for home automation. IEEE Software, 26(4): 30–38, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jiang:1995:FPA

[JS95] T. Ming Jiang and Lakshmanan Sankaran. Fast, portable application mirroring. IEEE Software, 12(2):57–63, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Janzen:2008:DTD

[JS08] David Janzen and Hossein Saiedian. Does test-driven development really improve software design quality? IEEE Software, 25(2):77– 84, March/April 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ju:1998:ICB

[Ju98] Dehua Ju. Insider: China’s budding software industry. IEEE Software, 15(2):20– 21, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2020.
pdf.

Ju:2001:CRC

[Ju01] Dehua Ju. Country report: China’s budding software industry. IEEE Software, 18(3):92–95, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s3092.
                                        pdf.

Jullig:1993:AFS

[Jul93] Richard K. Jullig. Applying formal software synthesis. IEEE Software, 10(3): 17–22, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Jung:1998:OVC

[Jun98] Ho-Won Jung. Optimizing value and cost in requirements analysis. IEEE Software, 15(4):74–78, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        cg1998/s4074abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s4074.pdf.

Juristo:1997:BSN

[Jur97] Natalia Juristo. Bookshelf: Sculpting a new software design paradigm: Beyond Programming: To a New Era of Design, by Bruce Blum. IEEE Software, 14(4):116–117, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4114.
pdf.

Jurik:2009:STB

[JW09] Andrew D. Jurik and Alfred C. Weaver. Software technology: Body sensors: Wireless access to physiological data. IEEE Software, 26(1):71–73, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Juristo:2001:GEI

[JWC01] Natalia Juristo, Helmut Windl, and Larry Constantine. Guest Editors’ introduction: Introducing usability. IEEE Software, 18(1):20–21, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1020.
pdf.

Jepsen:1999:LUE

[JWKZ99] Thomas C. Jepsen, Stephen A.

Wright, Richard L. Klevans, and Ze Zhang. Linux update: An experimental ATM network. IEEE Software, 16(5):32–39, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s5032abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s5032.pdf.

Kazman:1996:SBA

[KABC96] Rick Kazman, Gregory Abowd, Len Bass, and Paul Clements. Scenario-based analysis of software architecture. IEEE Software, 13 (6):47–55, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ko:2009:GEI

[KABM09] Andrew J. Ko, Robin Abraham, Margaret M. Burnett, and Brad A. Myers. Guest Editors’ introduction: End-user software engineering. IEEE Software, 26(5):16–17, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kaindl:1999:DTO

[Kai99] Hermann Kaindl. Difficulties in the transition from OO analysis to design. IEEE Software, 16(5):94–102, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5094abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5094.pdf.

Khoshgoftaar:1996:EQP

[KAKG96] Taghi M. Khoshgoftaar, Edward B. Allen, Kalai S. Kalaichelvan, and Nishith Goel. Early quality prediction: a case study in telecommunications. IEEE Software, 13(1):65–71, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kaluzniacky:1994:BCR

[Kal94] Eugene Kaluzniacky. Bookshelf: Consciousness-Raising Perspective on Computers in Society (review of the social impact of computers). IEEE Software, 11 (1):117–118, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kamel:1987:EMS

[Kam87] Ragui F. Kamel. Effect of modularity on system evolution. IEEE Software, 4(1):48–54, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kant:1993:SMM

[Kan93] Elaine Kant. Synthesis of mathematical-modeling software. IEEE Software, 10(3):30–41, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kansala:1997:IRA

[Kän97] Kari Känsälä. Integrating risk assessment with cost estimation. IEEE Software, 14(3):61–67, May/ June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kanoun:2001:RWD

[Kan01] Karama Kanoun. Realworld design diversity: a case study on cost. IEEE Software, 18(4):29– 33, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s4029abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s4029.pdf.

Kandrup:2005:SC

[Kan05] Susanne Kandrup. On systems coaching. IEEE Software, 22(1):52–54, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kandt:2009:EIF

[Kan09] Ronald Kirk Kandt. Experiences in improving flight software development processes. IEEE Software, 26 (3):58–64, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kappelman:2000:SSY

[Kap00] Leon A. Kappelman. Some strategic Y2K blessings. IEEE Software, 17(2):42– 46, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2042abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2042.pdf.

Karam:1988:IBD

[Kar88] Gerald M. Karam. ICONbased design method for PROLOG. IEEE Software, 5(4):51–65, July 1988. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kajihara:1993:LB

[KAS93] Juichirou Kajihara, Goro Amamiya, and Tetsuo Saya. Learning from bugs. IEEE Software, 10(5):46– ??, September 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kautz:1999:MSM

[Kau99] Karlheinz Kautz. Making sense of measurement for small organizations. IEEE Software, 16(2):14– 20, March/April 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s2014abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s2014.pdf.

Kay:1998:IIF

[Kay98] Alan Kay. Interview: Inventing the future. IEEE Software, 15(2):22–24, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s2022.
                                        pdf.

Karp:1988:CPF

[KB88] Alan H. Karp and Robert G. Babb. A comparison of 12 parallel Fortran dialects. IEEE Software, 5(5):52–67, September 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kozaczynski:1998:GEI

[KB98] Wojtek Kozaczynski and Grady Booch. Guest Editors’ introduction. IEEE Software, 15(5):34–36, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5034.
pdf;   http://www.computer.
org/computer/co1998/s5034abs.
htm.

Kazman:2002:MAR

[KB02] Rick Kazman and Len Bass. Making architecture reviews work in the real world. IEEE Software, 19(1):67–73, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1067abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1067.pdf.

Krumme:1990:TPE

[KC90] David Krumme and Alva Couch. Triplex parallelexecution monitor. IEEE Software, 5(3):28, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). [Abstract].

Kuntzmann-Combelles:1996:ICW

[KC96] Annie Kuntzmann-Combelles.

Insider: Can we learn lessons better? IEEE Software, 13(3):21, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kerth:1997:UPI

[KC97] Norman L. Kerth and Ward Cunningham. Using patterns to improve our architectural vision. IEEE Software, 14(1):53–59, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kesavasamy:2001:LIY

[KCC^{+}01] K. Kesavasamy, Yuri Chernak, Harry Clarke, Keith Miller, Claude Y. Laporte, Michiel van Genuchten, Cor van Dijk, Henk Scholten, Doug Vogel, Henrik A. Schutz, and Padraic Moran. Letters: Improving your testing; ethics and product liability; experience or education?; erring about erring; Ireland’s indigenous industry. IEEE Software, 18(4):8–11, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s4008.
                                        pdf.

Chang:1995:VLS

[kCCP^{+}95] Shi kuo Chang, Gennaro Costagliola, Giuliano Pacini, Maurizio Tucci, Genoveffa Tortora, Bing Yu, and Jing-Sheng Yu. Visual-language system for user interfaces. IEEE Software, 12(2):33–44, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2009:DVR

[KCD09] Philippe Kruchten, Rafael Capilla, and Juan Carlos Dueñas. The decision view’s role in software architecture practice. IEEE Software, 26(2):36– 42, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kendall:2008:DWF

[KCF^{+}08] Richard Kendall, Jeffrey C.

Carver, David Fisher, Dale Henderson, Andrew Mark, Douglass Post, Clifford E. Rhoades, Jr., and Susan Squires. Development of a weather forecasting code: a case study. IEEE Software, 25(4):59–65, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kuntzmann-Combelles:1993:AEF

[KCL93] Annie Kuntzmann-Combelles and Luqi. Advancing Europe’s fortunes. IEEE Software, 10(6):14–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Knight:2002:OW

[KD02] Alan Knight and Naci Dai. Objects and the Web. IEEE Software, 19(2):51– 59, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2051.
pdf;   http://www.computer.
org/software/so2002/s2051abs.
htm.

Koschmann:1988:BGB

[KE88] Timothy Koschmann and Martha Walton Evens. Bridging the gap between object-oriented and logic programming. IEEE Software, 5(4 or 5?):36–42, July 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). A description is given of an interface that was developed between Loops and Xerox Quintus Prolog. Loops is an extension to the Xerox AI Environment to support object-oriented programming; Xerox Quintus Prolog is a version of Prolog that runs on Xerox Lisp machines. Such a bridge enables all the support tools of both environments to be accessed, and degradation of performance that occurs when one language is implemented on top of another is avoided. The interface has three layers. At the lowest level, a set of Prolog predicates gives the Prolog programmer access to Loops objects. This lowest level is the bridge from Prolog to Loops. At the next level, programming tools in the Loops environment let object methods be defined in Prolog. At the highest level, the Prolog programmer can treat Prolog clauses as Loops objects that can be manipulated outside the Prolog database. Each layer can be used independently. 9 Refs.

Keeni:2000:EQP

[Kee00] Gargi Keeni. The evolution of quality processes at Tata Consultancy Services. IEEE Software, 17(4):79– 88, July/August 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4079abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4079.pdf.

Kelly:2007:SCS

[Kel07] Diane F. Kelly. A software chasm: Software engineering and scientific computing. IEEE Software, 24(6): 120, 118–119, November/ December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kemmerer:1990:IFM

[Kem90] Richard A. Kemmerer. Integrating formal methods into the development process. IEEE Software, 7 (5):37–50, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kemerer:1992:HLC

[Kem92] Chris F. Kemerer. How the learning curve affects CASE tool adoption. IEEE Software, 9(3):23–28, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kernighan:1999:IWY

[Ker99] Brian Kernighan. Interview: What have you learned today? IEEE Software, 16(2):66–68, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s2066.
                                        pdf.

Kernighan:2008:TTS

[Ker08] Brian Kernighan. Tools of the trade: Sometimes the old ways are best. IEEE Software, 25(6):18– 19, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kaiser:1988:IAS

[KFP88] Gail E. Kaiser, Peter H.

Feiler, and Steven S. Popovich. Intelligent assistance for software development and maintenance. IEEE Software, 5(3):40– 49, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kaiser:1987:MSS

[KG87] Gail E. Kaiser and David Garlan. Melding software systems from reusable building blocks. IEEE Software, 4(4):17–24, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Karabuk:2007:CMP

[KG07] Suleyman Karabuk and F. Hank Grant. A common medium for programming operations-research models. IEEE Software, 24(5): 39–47, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Karp:2003:USC

[KGRB03] Alan H. Karp, Rajiv Gupta, Guillermo J. Rozas, and Arindam Banerji. Using split capabilities for access control. IEEE Software, 20(1):42–49, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1042.
pdf;   http://www.computer.
org/software/so2003/s1042abs.
htm.

Kosoresow:1997:IDS

[KH97] Andrew P. Kosoresow and Steven A. Hofmeyr. Intrusion detection via system call traces. IEEE Software, 14(5):24–42, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5035.
pdf;   http://www.computer.
org/software/so1997/s5035abs.
htm.

Kamatar:2000:ERP

[KH00] Jagadish Kamatar and Will Hayes. An experience report on the personal software process. IEEE Software, 17(6):85–89, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6085abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6085.pdf.

Khan:2002:CSA

[KH02] Khaled M. Khan and Jun Han. Composing securityaware software. IEEE Software, 19(1):34–41, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1034abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1034.pdf.

Koru:2003:LOS

[KHA^{+}03] A. Güneṡ Koru, Warren Harrison, Omar Alonso, Martin Fowler, and Holger Eichelberger. Letters: Open source research ethics; balancing both worlds; the agile sweet spot; more recent references. IEEE Software, 20(6):8–11, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6008.
pdf.

Kim:2008:VSB

[KIB^{+}08] Sangsoo Kim, Hoh Peter In, Jongmoon Baik, Rick Kazman, and Kwangsin Han. VIRE: Sailing a blue ocean with value-innovative requirements. IEEE Software, 25(1):80–87, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kiczales:1996:SBB

[Kic96] Gregor Kiczales. Soapbox: Beyond the black box: Open implementation. IEEE Software, 13(1):8–11, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kilpi:2001:ISM

[Kil01] Tapani Kilpi. Implementing a software metrics program at Nokia. IEEE Software, 18(6):72–77, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s6072abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s6072.pdf.

Kim:2004:PMS

[Kim04] Hee-Woong Kim. A process model for successful CRM system development. IEEE Software, 21(4):22– 28, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kindlund:1997:INA

[Kin97] Erika Kindlund. Interface: Navigating the appletbrowser divide. IEEE Software, 14(5):22–25, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5022.

pdf. Discusses design issues for Java applets.

Kishida:1996:CMM

[Kis96] Kouichi Kishida. Counterpoint: Managing megaprojects: a free-form approach. IEEE Software, 13(4):29– 30, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kitchenham:2007:MMU

[KJC07] Barbara Kitchenham, David Ross Jeffery, and Colin Connaughton. Misleading metrics and unsound analyses. IEEE Software, 24(2): 73–78, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kozaczynski:1993:WIT

[KKC93] Wojtek Kozaczynski and Annie Kuntzmann-Combelles. What it takes to make OO work. IEEE Software, 10 (1):20–23, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kishimoto:1995:AOC

[KKH95] Yoshinori Kishimoto, Nobuto Kotaka, and Shinichi Honiden. Adapting objectcommunication methods dynamically. IEEE Software, 12(3):65–74, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kenny:1991:MAR

[KKJ91] Kevin Kenny and Lin Kwei-Jay. Measuring and analyzing real-time performance. IEEE Software, 8(5):41–??, September 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kanoun:1997:QQR

[KKL97] K. Kanoun, M. Kaaniche, and J.-P. Laprie. Qualitative and quantitative reliability assessment. IEEE Software, 14(2):77–87, March/ April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kaiser:1987:MDL

[KKM87] Gail E. Kaiser, Simon E. Kaplan, and Josephine Micallef. Multiuser, distributed language-based environments. IEEE Software, 4(6):58–67, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kienle:2009:TTT

[KKM^{+}09] Holger M. Kienle, Adrian Kuhn, Kim Mens, Mark van den Brand, and Roel Wuyts. Tools of the trade: Tool building on the shoulders of others. IEEE Software, 26(1):22–23, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruatrachue:1988:GSD

[KL88] Boontee Kruatrachue and Ted Lewis. Grain size determination for parallel processing. IEEE Software, 5(1):23–32, January 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Knight:1994:CTW

[KL94] John C. Knight and Bev Littlewood. Critical task of writing dependable software. IEEE Software, 11 (1):16–20, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kitchenham:1996:SQE

[KL96] Barbara Kitchenham and Shari Lawrence Pfleeger. Software quality. the elusive target. IEEE Software, 13(1):12–21, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kaner:1997:SUC

[KL97a] Cem Kaner and Brian Lawrence. Softlaw: UCC changes pose problems for developers. IEEE Software, 14(2):139–143, March/April 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kitchenham:1997:EUR

[KL97b] Barbara Kitchenham and Stephen Linkman. Estimates, uncertainty, and risk. IEEE Software, 14 (3):69–74, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kitchenham:1998:QTV

[KL98] Barbara Kitchenham and Stephen Linkman. Quality time: Validation, verification, and testing: Diversity rules. IEEE Software, 15(4):46–49, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4046.
pdf.

Kambayashi:2004:SPP

[KL04] Yasushi Kambayashi and Henry F. Ledgard. The separation principle: a programming paradigm. IEEE Software, 21(2):78– 87, March/April 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Koru:2005:BED

[KL05] A. Günes Koru and Hongfang Liu. Building effective defect-prediction models in practice. IEEE Software, 22(6):23–29, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Klatt:1995:LEN

[Kla95]  Garth Klatt.  Letters:

An engineer by any name doesn’t do it as well. IEEE Software, 12(3):5, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Klatt:1997:LPM

[Kla97] Garth Klatt. Letters: Parable or mirror? in praise of Beethoven. IEEE Software, 14(5):12, 14, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5008.
pdf.

Kang:2002:FOP

[KLD02] Kyo C. Kang, Jaejoon Lee, and Patrick Donohoe. Feature-oriented product line engineering. IEEE Software, 19(4):58– 65, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s4058.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s4058abs.
                                        htm.

Kleist:1998:LUC

[Kle98] Dave Kleist. Letters: Ubiquitous computing will demand rigorous practices. IEEE Software, 15(3):14, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s3010.
                                        pdf.

Kasik:1989:RUU

[KLR89] David J. Kasik, Michelle A. Lund, and Henry W. Ramsey. Reflections on using a UIMS for complex applications. IEEE Software, 6(1):54–61, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kim:1986:CAI

[KM86] Jin Kim and John McDermott. Computer aids for IC design. IEEE Software, 3(2):38–47, March 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Keepence:1999:UPM

[KM99] Barry Keepence and Mike Mannion. Using patterns to model variability in product families. IEEE Software, 16(4):102–108, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s4102abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s4102.pdf.

Kalb:2002:CPI

[KM02] Goerge E. Kalb and Gerald M. Masson. Continuous process improvement and the risk to information assurance. IEEE Software, 19(1):84–86, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1084abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1084.pdf.

Knauber:2000:APL

[KMSW00] Peter Knauber, Dirk Muthig, Klaus Schmid, and Tanya Widen. Applying product line concepts in small and medium-sized companies. IEEE Software, 17(5):88–95, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/

                                        so/s5088abs.htm;   http:
                                        //dlib.computer.org/so/
                                        books/so2000/pdf/s5088.
                                        pdf.

Kelly:1991:IDD

[KMY91] John P. J. Kelly, Thomas I. McVittie, and Wayne I. Yamamoto. Implementing design diversity to achieve fault tolerance. IEEE Software, 8(4):61–71, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kavi:1996:STA

[KN96] Krishna Kavi and Ez Nahouraii. Software tools assessment. IEEE Software, 13(5):23–26, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kemp:2001:KMI

[KNR^{+}01] Linda Larson Kemp, Kenneth E. Nidiffer, Louis C. Rose, Robert Small, and Michael Stankosky. Knowledge management: Insights from the trenches. IEEE Software, 18(6):66– 68, November/December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://computer.org/software/
                                        so2001/s6066abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s6066.pdf.

Knudsen:1987:BEH

[Knu87] Jorgen Lindskov Knudsen.

Better exception-handling in block-structured systems. IEEE Software, 4(3): 40–49, May 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ko:2009:AWC

[Ko09] Andrew J. Ko. Automating the Web with Co-Scripter: An interview with Tessa Lau. IEEE Software, 26(5):52–53, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kochikar:1998:OPW

[Koc98] Vivekanand P. Kochikar.

The object-powered Web. IEEE Software, 15(3):57– 62, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3057abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3057.pdf.

Kocher:2000:MSP

[Koc00] Bryan Kocher. A model for software practices from the accounting profession. IEEE Software, 17(1): 57–60, January/February 2000. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://computer.org/software/
                                        so2000/s1057abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s1057.pdf.

Kodaganallur:2004:ILP

[Kod04] Viswanathan Kodaganallur. Incorporating language processing into Java applications: a JavaCC tutorial. IEEE Software, 21(4): 70–77, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2004/
                                        04/s4070s.pdf.

Kohl:2005:REC

[Koh05] Ronald J. Kohl. Requirements engineering changes for COTS-intensive systems. IEEE Software, 22(4):63–64, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Korth:1986:ESR

[Kor86] Henry F. Korth. Extending the scope of relational languages. IEEE Software, 3(1):19–28, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Korzeniowski:1996:NPO

[Kor96] Paul Korzeniowski. In the news: Plan 9: From outer space to kernel space. IEEE Software, 13(3):105, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2006:PPF

[KOS06] Philippe Kruchten, Henk Obbink, and Judith Stafford. The past, present, and future for software architecture. IEEE Software, 23(2): 22–30, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kotula:1998:UPC

[Kot98] Jeffrey Kotula. Using patterns to create component documentation. IEEE Software, 15(2):84–92, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s2084abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2084.pdf.

Kernighan:1999:EPP

[KP99] Brian Kernighan and Rob Pike. Excerpt from The Practice of Programming: Finding performance improvements. IEEE Software, 16(2):61–65, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s2061abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s2061.pdf.

Kelly:2009:WPD

[KP09] Steven Kelly and Risto Pohjonen. Worst practices for domain-specific modeling. IEEE Software, 26(4): 22–29, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kitchenham:1995:CSM

[KPP95] Barbara Kitchenham, Lesley Pickard, and Shari Lawrence Pfleeger. Case studies for method and tool evaluation. IEEE Software, 12(4): 52–62, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kelley:1988:MEH

[KR88] Keith L. Kelley and Marek Rusinkiewicz. Multikey, extensible hashing for relational databases. IEEE Software, 5(4):77–85, July 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Karlsson:1997:CVA

[KR97] Joachim Karlsson and Kevin Ryan. A cost-value approach for prioritizing requirements. IEEE Software, 14(5):67–74, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5067.
pdf;   http://www.computer.
org/software/so1997/s5067abs.
htm.

Kulkarni:2003:SCM

[KR03] Vinay Kulkarni and Sreedhar Reddy. Separation of concerns in model-driven development. IEEE Software, 20(5):64–69, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5064.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5064abs.htm.

Karlstrom:2005:CAM

[KR05] Daniel Karlstrom and Per Runeson. Combining agile methods with stagegate project management. IEEE Software, 22(3):43– 49, May/June 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kraterfield:1995:LWG

[Kra95] Susan E. Kraterfield. Letters: What’s the genderquality connection? IEEE Software, 12(4):5, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kobitzsch:2001:OI

[KRF01] Werner Kobitzsch, Dieter Rombach, and Raimund L. Feldmann. Outsourcing in India. IEEE Software, 18(2):78–86, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s2078abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s2078.pdf.

Kruchten:1995:VMA

[Kru95] Philippe B. Kruchten. The 4 + 1 view model of architecture. IEEE Software, 12 (6):42–50, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2005:CSD

[Kru05a] Philippe Kruchten. Casting software design in the function-behavior-structure framework. IEEE Software, 22(2):52–58, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2005:EIS

[Kru05b] Philippe Kruchten. Editor’s introduction: Software design in a postmodern era. IEEE Software, 22(2):16–18, March/ April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2005/02/s2016.
pdf.

Kruchten:2008:BHL

[Kru08a] Philippe Kruchten. The biological half-life of software engineering ideas. IEEE Software, 25(5):10– 11, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2008:CDL

[Kru08b] Philippe Kruchten. Career development: Licensing software engineers? IEEE Software, 25(6):35– 37, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2009:CDW

[Kru09a] Philippe Kruchten. Career development: When Robert rules. IEEE Software, 26(1):20–21, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2009:CDY

[Kru09b] Philippe Kruchten. Career development: You are what you read. IEEE Software, 26(2):10–11, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kramer:1989:DIE

[KS89] Bernd Kramer and Heinz-Wilhelm Schmidt. Developing integrated environments with ASDL. IEEE Software, 6(1):98– 107, January 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kirby:1996:LAD

[KS96] James Kirby, Jr. and Mary Shaw. Letters: Architecture? design? debating classifications. IEEE Software, 13(3):10, 12, 14, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kelly:2004:LKA

[KS04] Diane Kelly and Terry Shepard. A little knowledge about software. IEEE Software, 21(2):46–48, March/ April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kshetri:2004:ELA

[Ksh04] Nir Kshetri. Economics of Linux adoption in developing countries. IEEE Software, 21(1):74–81, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kshetri:2005:SSC

[Ksh05] Nir Kshetri. Structural shifts in the Chinese software industry. IEEE Software, 22(4):86–93, July/ August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Komi-Sirvio:2002:TPS

[KSMS02] Seija Komi-Sirviö, Annukka Mäntyniemi, and Veikko Seppänen. Toward a practical solution for capturing knowledge for software projects. IEEE Software, 19(3):60– 62, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3060.
pdf;   http://www.computer.
org/software/so2002/s3060abs.
htm.

Kruchten:1984:SPU

[KSS84] Philippe Kruchten, Edmond Schonberg, and Jacob Schwartz. Software prototyping using the Setl programming language. IEEE Software, 1(4):66–75, October 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Koskimies:1998:ASM

[KSTM98] Kai Koskimies, Tarja Systä, Jyrki Tuomi, and Tatu Männistö. Automated support for modeling OO software. IEEE Software, 15(1):87–94, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1998/s1087abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s1087.pdf.

Kallstrom:1988:PTP

[KT88] Marta Kallstrom and Shreekant S. Thakkar. Programming three parallel computers. IEEE Software, 5 (1):11–22, January 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Koru:2004:DHM

[KT04] A. Gunes Koru and Jeff Tian. Defect handling in medium and large Open Source projects. IEEE Software, 21(4):54–61, July/ August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Keil:2005:BCD

[KT05] Mark Keil and Amrit Tiwana. Beyond cost: The drivers of COTS application value. IEEE Software, 22(3):64–69, May/ June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kuhn:1998:SDS

[Kuh98] Sarah Kuhn. The software design studio: An exploration. IEEE Software, 15(2):65–71, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s2065abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2065.pdf.

Kusmiss:1995:REI

[Kus95] Joseph M. Kusmiss. Reassurance for the electronically insecure: review of Network and Internetwork Security: Principles and Practices. IEEE Software, 12(3):103, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kusmiss:1996:BCP

[Kus96] Joseph M. Kusmiss. Bookshelf: Computers and privacy: Crypto policy questions. IEEE Software, 13 (2):122, March 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kircher:2007:GEI

[KV07] Michael Kircher and Markus Völter. Guest Editors’ introduction: Software patterns. IEEE Software, 24(4):28–30, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2007/04/s4028.
pdf.

Karunanithi:1992:UNN

[KWM92] Nachimuthu Karunanithi, Darrell Whitley, and Yashwant K. Malaiya. Using neural networks in reliability prediction. IEEE Software, 9(4):53–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Krishna:2004:PDS

[KYM^{+}04] Arvind S. Krishna, Cemal Yilmaz, Atif M. Memon, Adam A. Porter, Douglas C. Schmidt, Aniruddha Gokhale, and Balachandran Natarajan. Preserving distributed systems’ critical properties: a model-driven approach. IEEE Software, 21(6):32– 40, November/December 2004. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Laddad:2003:QT

[LA03] Ramnivas Laddad and Roger Alexander. Quality time. IEEE Software, 20(6):90–93, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6090.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6090abs.htm.

Labudde:1994:WRE

[Lab94] Edward V. Labudde. Why is requirements engineering underused? IEEE Software, 11(2):6, 8, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See response in March 1994 issue, pp. 8–9.

Lai:1993:MMP

[Lai93] Robert Lai. The move to mature processes. IEEE Software, 10(4):14–??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lai:2008:JIA

[Lai08] Charlie Lai. Java insecurity: Accounting for subtleties that can compromise code. IEEE Software, 25(1):13–19, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lampson:1984:HCS

[Lam84] Butler W. Lampson. Hints for computer system design. IEEE Software, 1 (1):11–28, January 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lanubile:1996:QTW

[Lan96] Filippo Lanubile. Quality time: Why software reliability predictions fail. IEEE Software, 13(4):131–132, 137, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lanir:2006:N

[Lan06] Danna Voth Heather Lanir. In the news. IEEE Software, 23(5):96–98, September/October 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        05/s5096.pdf.

Lagos:2007:IBD

[LANC07] María Ester Lagos, Rosa Alarcón, Miguel Nussbaum, and Francisca Capponi. Interaction-based design for mobile collaborativelearning software. IEEE Software, 24(4):80–89, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Larman:2001:DPV

[Lar01a] Craig Larman. Design:

Protected variation: The importance of being closed. IEEE Software, 18(3):89– 91, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s3089.
pdf.

Larrabee:2001:BSP

[Lar01b] Robert C. Larrabee. Bookshelf: a (software) Pilgrim’s Progress: The Limits of Software: People, Projects, and Perspectives. IEEE Software, 18(2):102, March/April 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2102.
pdf.

Lauesen:1998:RLO

[Lau98] Søren Lauesen. Real-life object-oriented systems. IEEE Software, 15(2):76– 83, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/

so1998/s2076abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2076.pdf.

Lauesen:2003:TDF

[Lau03] Soren Lauesen. Task descriptions as functional requirements. IEEE Software, 20(2):58–65, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2058.
pdf;   http://www.computer.
org/software/so2003/s2058abs.
htm.

Lavazza:2000:PAS

[Lav00] Luigi Lavazza. Providing automated support for the GQM measurement process. IEEE Software, 17(3):56–62, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3056abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s3056.pdf.

Lawrence:1996:PDY

[Law96] Brian Lawrence. Point: Do you really need formal requirements? IEEE Software, 13(2):20, 22, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lawrence:1998:CDM

[Law98] Brian Lawrence. Counterpoint: Designers must do the modeling. IEEE Software, 15(2):31, 33, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2031.
pdf.

Luqi:1988:RPR

[LB88] Luqi and Valdis Berzins.

Rapidly prototyping realtime systems. IEEE Software, 5(5):25–36, September 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Laitinen:2000:PCS

[LB00] Mauri Laitinen and John Boddie. Point/counterpoint: Scaling down is hard to do vs. do we ever really scale down? IEEE Software, 17(5):78–81, September/October 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5078.
pdf.

Lutz:2006:GEI

[LB06] Michael J. Lutz and Donald Bagert. Guest Editors’ introduction: Software engineering curriculum development. IEEE Software, 23(6):16–18, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
06/s6016.pdf.

Larus:2004:RS

[LBD^{+}04] James R. Larus, Thomas Ball, Manuvir Das, Robert DeLine, Manuel Fahndrich, Jon Pincus, Sriram K. Rajamani, and Ramanathan Venkatapathy. Righting software. IEEE Software, 21(3):92–100, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

LeBlanc:1985:HLB

[LC85] Thomas J. LeBlanc and Robert P. Cook. Highlevel broadcast communication for local area networks. IEEE Software, 2(3):40– 48, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Leblang:1987:PSC

[LC87] David B. Leblang and Robert P. Chase, Jr. Parallel software configuration management in a network environment. IEEE Software, 4(6):28–35, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Li:2009:CBS

[LCB^{+}09] Jingyue Li, Reidar Conradi, Christian Bunse, Marco Torchiano, Odd Petter N. Slyngstad, and Maurizio Morisio. Componentbased software development: Development with off-the-shelf components: 10 facts. IEEE Software, 26(2):80–87, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lutz:2001:LMM

[LCE^{+}01] Mike Lutz, Stephen W.

Carson, Mike Ellims, Robert L. Baber, Robert Glass, Ambrose Li, and John Boddie. Letters: More on mathematics and software practitioners; scaling down. IEEE Software, 18(2):6–8, March/April 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2006.
pdf.

Lange:2006:PUS

[LCM06] Christian Lange, Michel Chaudron, and Johan Muskens. In practice: UML software architecture and design description. IEEE Software, 23(2):

40–46, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Leveson:1991:SVA

[LCS91] Nancy G. Leveson, Stephen S. Cha, and Timothy J. Shimeall. Safety verification of Ada programs using software fault trees. IEEE Software, 8(4):48–59, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lee:1997:DRV

[LCS97] D. L. Lee, Huei Chuang, and K. Seamons. Document ranking and the vector-space model. IEEE Software, 14(2):67–75, March/ April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Leonhard:1995:JSD

[LD95] Corey A. Leonhard and J. Steve Davis. Job-shop development model: a case study. IEEE Software, 12(2):86–92, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Leavens:1991:MSV

[Lea91] Gary T. Leavens. Modular specification and verification of object-oriented programs. IEEE Software, 8(4):72–80, July 1991. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lehman:1986:PDR

[Leh86] John A. Lehman. Program design and rhetoric. IEEE Software, 3(3):71– 73, May 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lehman:1998:SFM

[Leh98] M. M. Lehman. Software’s future: Managing evolution. IEEE Software, 15(1):40–44, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1040abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s1040.pdf.

Leibovitch:1999:BCL

[Lei99] Evan Leibovitch. The business case for Linux. IEEE Software, 16(1):40–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s1040abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s1040.pdf.

Lesiecki:2006:AAJ

[Les06] Nicholas Lesiecki. Applying AspectJ to J2EE application development. IEEE Software, 23(1):24–32, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Leveson:1990:CBP

[Lev90] Nancy G. Leveson. The challenge of building processcontrol software. IEEE Software, 7(6):55–62, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Levendel:1991:IQM

[Lev91] Ytzhak Levendel. Improving quality with a manufacturing process. IEEE Software, 8(2):13–25, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lewis:1984:WPM

[Lew84] Ted G. Lewis. Word processing for the masses: a review of Bank Street Writer. IEEE Software, 1(2):89– 92, April 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lewis:1998:BDN

[Lew98]  Ted Lewis.  Bookshelf:

The Deadline: A Novel About Project Management; ParseMe. 1st: SGML for Software Developers; Creating a Software Engineering Culture; The Year 2000 Software Crisis: Challenge of the Century. IEEE Software, 15(1):107–110, January/ February 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1107.
pdf.

Lang:2005:HSD

[LF05] Michael Lang and Brian Fitzgerald. Hypermedia systems development practices: a survey. IEEE Software, 22(2):68–75, March/ April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lewis:1990:CG

[LFL^{+}90] Ted Lewis, Glenn Funkhouser, Kit Hung Lou, Atsushi Ito, and Rolf Bahlke. Code generators. IEEE Software, 7 (3):67–70, May 1990. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Laitinen:1999:SES

[LFW99] Mauri Laitinen, Mohamed Fayad, and Robert Ward. Software engineering in the small. IEEE Software, 16(4):117, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s4117.
                                        pdf.

Laitinen:2000:SES

[LFW00] Mauri Laitinen, Mohamed Fayad, and Robert Ward. Software engineering in the small: Guest Editors’ introduction. IEEE Software, 17(5):75–77, September/October 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s5075.
                                        pdf.

Luqi:1997:FMP

[LG97] Luqi and Joseph A. Goguen. Formal methods: Promises and problems. IEEE Software, 14(1):73–85, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Liu:2003:ACM

[LG03] Anna Liu and Ian Gorton. Accelerating COTS middleware acquisition: The i-Mate process. IEEE Software, 20(2):72–79, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2072.
pdf;   http://www.computer.
org/software/so2003/s2072abs.
htm.

Lieberherr:1989:AGS

[LH89] Karl J. Lieberherr and Ian M. Holland. Assuring good style for object-oriented programs. IEEE Software, 6(5):38–48, September 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lauesen:2001:VWL

[LH01] Søren Lauesen and Morten Borup Harning. Virtual windows: Linking user tasks, data models, and interface design. IEEE Software, 18(4):67–75, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s4067abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s4067.pdf.

Lee:1990:UID

[LHB^{+}90] Ed Lee, Frank Hall, Andrea Bowers, Sherry Yang, Len Bass, Andreas Lemke, and Yen-Ping Shan. User-interface development tools. IEEE Software, 7(3): 31–36, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Liebowitz:2002:LNG

[Lie02] Jay Liebowitz. A look at NASA Goddard Space Flight Center’s knowledge management initiatives. IEEE Software, 19(3):40– 42, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s3040.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s3040abs.
                                        htm.

Lim:1994:ERQ

[Lim94] Wayne C. Lim. Effects of reuse on quality, productivity, and economics. IEEE Software, 11(5):23– 30, September 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lindquist:1985:AUH

[Lin85] Timothy E. Lindquist. Assessing the usability of human-computer interfaces. IEEE Software, 2 (1):74–82, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Linton:1987:DMS

[Lin87] Mark A. Linton. Distributed management of a software database. IEEE Software, 4(6):70–76, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lindstrom:1993:FWD

[Lin93] David R. Lindstrom. Five ways to destroy a software project. IEEE Software, 10 (5):55–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Linger:1994:CPM

[Lin94] Richard C. Linger. Cleanroom process model. IEEE Software, 11(2):50–58, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lister:1997:PRM

[Lis97] Tim Lister. Point: Risk management is project management for adults. IEEE Software, 14(3):20, 22, May/June 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lister:1998:MHF

[Lis98] Tim Lister. Manager: Hallucinations at 37,000 feet. IEEE Software, 15(3):105– 107, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so1998/pdf/s3105.
                                        pdf.

Little:2004:VCC

[Lit04] Todd Little. Value creation and capture: a model of the software development process. IEEE Software, 21 (3):48–53, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Little:2005:CAA

[Lit05]  Todd Little.  Context-

adaptive agility: Managing complexity and uncertainty. IEEE Software, 22 (3):28–35, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Little:2006:SEU

[Lit06] Todd Little. Schedule estimation and uncertainty surrounding the cone of uncertainty. IEEE Software, 23(3):48–54, May/ June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lindquist:1988:TCG

[LJ88] Timothy E. Lindquist and Joyce R. Jenkins. Testcase generation with IO-Gen. IEEE Software, 5 (1):72–79, January 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lawrence:1997:MPS

[LJ97] Brian Lawrence and Bob Johnson. Manager: The project scoping gamble. IEEE Software, 14(3):107– 109, May/June 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Luqi:1988:CAP

[LK88] Luqi and Mohammad A.

Ketabchi. A computeraided prototyping system. IEEE Software, 5(2):66– 72, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lawler:2003:MMT

[LK03] Jim Lawler and Barbara Kitchenham. Measurement modeling technology. IEEE Software, 20(3):68– 75, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3068.
pdf;   http://www.computer.
org/software/so2003/s3068abs.
htm.

Litwin:1987:NMF

[LL87] Witold Litwin and David B.

Lomet. A new method for fast data search with keys. IEEE Software, 4(2):16– 24, March 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). Also published in/as: Proc. IEEE CS Intl. Conf. No. 2 on Data Engineering, Los Angeles, Feb.1986.

Lin:1992:PVL

[LL92a] Fuchun Joseph Lin and Ming T. Liu. Protocol validation for large-scale applications. IEEE Software, 9(1):23–26, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lin:1992:RPE

[LL92b] Fuchun Joseph Lin and Ming T. Liu. The rise of protocol engineering. IEEE Software, 9(1):14–??, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lutz:2003:CNF

[LL03] Michael H. Lutz and Phillip A. Laplante. C# and the .NET framework: Ready for real time? IEEE Software, 20(1):74–80, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1074.
pdf;   http://www.computer.
org/software/so2003/s1074abs.
htm.

Liu:2003:IAM

[LLH^{+}03] Jun Liu, Khiang Wee Lim, Weng Khuen Ho, Kay Chen Tan, Rajagopalan Srinivasan, and Arthur Tay. The intelligent alarm management system. IEEE Software, 20(2):66–71, March/ April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2066.
pdf;   http://www.computer.
org/software/so2003/s2066abs.
htm.

Liu:2005:UOS

[LLH^{+}05] Jun Liu, Khiang Wee Lim, Weng Khuen Ho, Kay Chen Tan, Arthur Tay, and Rajagopalan Srinivasan. Using the OPC standard for real-time process monitoring and control. IEEE Software, 22(6):54–59, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lethbridge:2006:SRU

[LLS^{+}06] Timothy C. Lethbridge, Richard J. LeBlanc, Jr., Ann E. Kelley Sobel, Thomas B. Hilburn, and Jorge L. Díaz-Herrera. SE2004: Recommendations for undergraduate software engineering curricula. IEEE Software, 23(6):19– 25, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Larrabee:1984:GPA

[LM84] Tracy Larrabee and Chad Leland Mitchell. GAMB&IT: a prototyping approach to video game design. IEEE Software, 1(4):28–36, October 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Larrabee:2003:BPM

[LM03] Robert C. Larrabee and Shantha Mohan. Bookshelf: Project management: a working reference; the importance of software performance engineering. IEEE Software, 20(5):102–103, September/ October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/05/s5102.

pdf. Review of Quality Software Project Management by Robert Futrell, Donald Shafer, and Linda Shafer, and Performance Solutions: A Practical Guide to Creating Responsive, Scalable Software by Connie U. Smith and Lloyd G. Williams.

Lutz:2004:ORD

[LM04] Robyn R. Lutz and I. C. Mikulski. Ongoing requirements discovery in highintegrity systems. IEEE Software, 21(2):19–25, March/ April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lalanda:2007:DCD

[LM07] Philippe Lalanda and Cristina Marin. A domainconfigurable development environment for serviceoriented applications. IEEE Software, 24(6):31–38, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lyu:1992:ARM

[LN92] Michael R. Lyu and Allen Nikora. Applying reliability models more effectively. IEEE Software, 9(4): 43–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

LeMetayer:1998:ESD

[LNR98] Daniel Le Métayer, Valérie-Anne Nicolas, and Olivier Ridoux. Exploring the software development trilogy. IEEE Software, 15(6):75–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s6075abs.htm;

                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s6075.pdf.

Lewis:1990:CSD

[LO90] Ted G. Lewis and Paul Oman. The challenge of software development. IEEE Software, 7(6):9–14, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Long:2002:CPM

[Lon02]  Carol A. Long.  Can

project management be extreme too? IEEE Software, 19(3):90–92, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s3090.
                                        pdf.

Long:2003:LMG

[Lon03]  Carol A. Long.  Lost,

but making good use of time? IEEE Software, 20(1):96–99, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s1096.
                                        pdf.

Lott:1997:MBN

[Lot97] Christopher M. Lott. Manager: Breathing new life into the waterfall model. IEEE Software, 14(5):103–105, September/ October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5103.
pdf.

Lounama:1995:FBS

[Lou95] Pertti Lounama. The future belongs to the specialized tool. IEEE Software, 12(2):9, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2005:JUT

[Lou05] Panagiotis Louridas. JUnit: Unit testing and coding in tandem. IEEE Software, 22(4):12–15, July/ August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2006:SWS

[Lou06a] Panagiotis Louridas. SOAP and Web services. IEEE Software, 23(6):62–67, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2006:SCA

[Lou06b] Panagiotis Louridas. Static code analysis. IEEE Software, 23(4):58–61, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2006:UWS

[Lou06c] Panagiotis Louridas. Using Wikis in software development. IEEE Software, 23(2):88–91, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2006:VC

[Lou06d] Panagiotis Louridas. Version control. IEEE Software, 23(1):104–107, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2007:DGP

[Lou07] Panagiotis Louridas. Declarative GUI programming in Microsoft Windows. IEEE Software, 24(4):16– 19, July/August 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2008:OWS

[Lou08] Panagiotis Louridas. Orchestrating Web services with BPEL. IEEE Software, 25(2):85–87, March/ April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lovgren:1994:IHC

[Lov94]  John Lovgren.  Inter-

face: How to choose good metaphors. IEEE Software, 11(3):86–88, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lindvall:2000:GEI

[LR00] Mikael Lindvall and Ioana Rus. Guest Editors’ introduction: Process diversity in software development. IEEE Software, 17(4):14– 18, July/August 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4014.
pdf.

Long:2002:BSM

[LRO02] Carol A. Long, V. V. S.

Raveendra, and Tope Omitola. Bookshelf: Software manufacturing manager’s handbook; inside Java 2 platform security; so you want to build an operating system. IEEE Software, 19(4):131–133, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4131.

pdf. Review of Software Product Lines: Practices and Patterns by Paul Clements and Linda Northrop, Inside Java 2 Platform Security: Architecture, API Design and Implementation by Li Gong, and Applied Operating Systems Concepts by Avi Silberschatz, Peter B. Galvin, Abraham Silberschatz, and Greg Gagne.

Letovsky:1986:DPP

[LS86] Stanley Letovsky and Elliot Soloway. Delocalized plans and program comprehension. IEEE Software, 3(3): 41–49, May 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lumsden:2007:PC

[LS07] Larry Lumsden and Wolfgang Strigel. Point/ counterpoint. IEEE Software, 24(1):54–57, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Larrabee:2002:BSA

[LSB02] Robert C. Larrabee, Diomidis Spinellis, and Angelo Bellotti. Bookshelf: Software architecture: a maturing discipline; a gentle introduction to embedded systems; practical transactional systems in the Internet Age. IEEE Software, 19(1):100–102, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1100abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1100.pdf.

Lethbridge:2003:HSE

[LSF03] Timothy C. Lethbridge, Janice Singer, and Andrew Forward. How software engineers use documentation: The state of the practice. IEEE Software, 20(6):35–39, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6035.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6035abs.htm.

Lemahieu:2005:CCA

[LSG^{+}05] Wilfried Lemahieu, Monique Snoeck, Frank Goethals, Manu De Backer, Raf Haesen, Jacques Vandenbulcke, and Guido Dedene. Coordinating COTS applications via a business event layer. IEEE Software, 22(4):28– 35, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lewis:1984:IPD

[LSM84] Ted G. Lewis, Keith R.

Spitz, and Paul E. McKenney. Interleave principle for demonstrating concurrent programs. IEEE Software, 1(4):54–64, October 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lindland:1994:UQC

[LSS94] Odd Ivar Lindland, Guttorm Sindre, and Arne Sølvberg. Understanding quality in conceptual modeling. IEEE Software, 11(2):42–49, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Luckham:1991:TDP

[LST91] David Luckham, Sriram Sankar, and Shuzo Takahashi. Two-dimensional pinpointing: Debugging with formal specifications. IEEE Software, 8(1):74–84, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lenz:1987:SRT

[LSW87] Manfred Lenz, Hans Albrecht Schmid, and Peter F. Wolf. Software reuse through building blocks. IEEE Software, 4(4):34– 42, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Liggesmeyer:2009:TES

[LT09] Peter Liggesmeyer and Mario Trapp. Trends in embedded software engineering. IEEE Software, 26 (3):19–25, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lubars:1991:RDD

[Lub91] Mitchell D. Lubars. Representing design dependencies in an issue-based style. IEEE Software, 8(4):81– 89, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Luqi:1992:CAP

[Luq92a] Luqi. Computer-aided prototyping for a commandand-control system using CAPS. IEEE Software, 9(1):56–67, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Luqi:1992:SRC

[Luq92b]  Luqi.  Status report:

Computer-aided protoyping. IEEE Software, 9 (6):77–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lussier:2004:NTH

[Lus04] Stephane Lussier. New tricks: How Open Source changed the way my team works. IEEE Software, 21(1):68–72, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Luckham:1985:OAS

[Lv85] David C. Luckham and Friedrich W. von Henke. Overview of Anna, a specification language for Ada. IEEE Software, 2(2):9–22, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lammel:2001:CLP

[LV01] Ralf Lämmel and Chris Verhoef. Cracking the 500language problem. IEEE Software, 18(6):78–88, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6078abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6078.pdf.

Lopez:1992:EPM

[LVS92] Leonard A. Lopez, Karim A. Valimohamed, and Linda G. Schub. An environment for painless MIMD development. IEEE Software, 9 (6):67–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lawrence:2001:TRR

[LWE01] Brian Lawrence, Karl Wiegers, and Christof Ebert. The top risks of requirements engineering. IEEE Software, 18(6):62–63, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6062abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6062.pdf.

Lokan:2001:OBU

[LWHS01] Chris Lokan, Terry Wright, Peter R. Hill, and Michael Stringer. Organizational benchmarking using the ISBSG data repository. IEEE Software, 18(5):26– 32, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s5026abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5026.pdf.

Lutz:1990:TT

[LWN^{+}90] Mike Lutz, Douglas G.

Wildes, Edward J. Nieters, Rob Chang, K. C. Tai, Robert Poston, Rhonda Martin, Randall Neff, Mary Jean Harrold, and Mary Lou Soffa. Testing tools. IEEE Software, 7(3): 53–57, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lee:1999:AUR

[LX99] Jonathan Lee and Nien-Lin Xue. Analyzing user requirements by use cases: a goal-driven approach. IEEE Software, 16(4):92–101, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s4092abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s4092.pdf.

Lauesen:1998:SQV

[LY98] Søren Lauesen and Houman Younessi. Is software quality visible in the code? IEEE Software, 15(4):69– 73, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        cg1998/s4069abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s4069.pdf.

Musa:1989:QSV

[MA89] John D. Musa and A. Frank Ackerman. Quantifying software validation: when to stop testing? IEEE Software, 6(3):19–27, May 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Michener:2000:SDK

[MA00] John R. Michener and Tolga Acar. Security domains: Key management in large-scale systems. IEEE Software, 17(5):52–58, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so/s5052abs.htm;   http:
//dlib.computer.org/so/
books/so2000/pdf/s5052.
pdf.

McLaughlin:2004:Nb

[MA04] Laurianne McLaughlin and Benjamin Alfonsi. In the news. IEEE Software, 21(5):101–105, September/October 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
05/s5101.pdf.

McLaughlin:2005:N

[MA05] Laurianne McLaughlin and Benjamin Alfonsi. In the news. IEEE Software, 22(1):98–101, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2005/
                                        01/s1098.pdf.

Maiden:2007:HM

[MA07] Neil Maiden and Ian Alexander. From the horse’s mouth. IEEE Software, 24(6):21–23, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mackey:1996:WBT

[Mac96] Karen Mackey. Why bad things happen to good projects. IEEE Software, 13(3): 27–32, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mackey:1998:CWB

[Mac98] Karen Mackey. Culture at work: Beyond Dilbert: Creating cultures that work. IEEE Software, 15(1):48–49, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s1048.
                                        pdf.

Mackey:1999:CWC

[Mac99a] Karen Mackey. Culture at work: Conscious conflict. IEEE Software, 16 (5):112–113, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s5112.
pdf.

Mackey:1999:CWG

[Mac99b]  Karen Mackey.  Cul-

ture at work: Give me a break. IEEE Software, 16(1):26–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1026.
pdf.

Mackey:1999:STD

[Mac99c] Karen Mackey. Stages of team development. IEEE Software, 16(4):90–91, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4090.
pdf.

Mackey:2000:CWM

[Mac00] Karen Mackey. Culture at work: Mars versus Venus. IEEE Software, 17(3):14– 15, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s3014.
                                        pdf.

Madachy:1997:HRA

[Mad97] Raymond J. Madachy. Heuristic risk assessment using cost factors. IEEE Software, 14(3):51–59, May/ June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maginnis:2000:EDB

[Mag00] Terri Maginnis. Engineers don’t build. IEEE Software, 17(1):34–39, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s1034abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s1034.pdf.

Mair:2002:BSS

[Mai02] Steven Mair. A balanced scorecard for a small software group. IEEE Software, 19(6):21–27, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s6021.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s6021abs.
                                        htm.

Maiden:2005:WRR

[Mai05] Neil Maiden. What has requirements research ever done for us? IEEE Software, 22(4):104–105, July/ August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2006:IYR

[Mai06a]  Neil Maiden.  Improve

your requirements: Quantify them. IEEE Software, 23(6):68–69, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2006:SYR

[Mai06b] Neil Maiden. Servicing your requirements. IEEE Software, 23(5):14–16, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2007:MRW

[Mai07]  Neil Maiden.  My re-

quirements? well, that depends. IEEE Software, 24(1):86–87, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2008:RRY

[Mai08a]  Neil Maiden.  Require-

ments: Requirements 25 years on. IEEE Software, 25(6):26–28, November/ December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2008:URS

[Mai08b] Neil Maiden. User requirements and system requirements. IEEE Software, 25(2):90–91, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2009:OA

[Mai09a] Neil Maiden. Oi, analyst. IEEE Software, 26(6):13– 14, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2009:RCS

[Mai09b]  Neil Maiden.  Require-

ments: Card sorts to acquire requirements. IEEE Software, 26(3):85–86, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2009:WWH

[Mai09c]  Neil Maiden.  Where

are we? handling context. IEEE Software, 26(5): 75–76, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mili:1999:TED

[MAMY99] Ali Mili, Edward Addy, Hafedh Mili, and Sherif Yacoub. Toward an engineering discipline of software reuse. IEEE Software, 16(5):22–31, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5022abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5022.pdf.

Maples:1985:ASP

[Map85] Creve Maples. Analyzing software performance in a multiprocessor environment. IEEE Software, 2(4): 50–63, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martin:1985:NSR

[Mar85] Joanne L. Martin. National Supercomputer Research Centers. IEEE Software, 2 (6):55–67, November 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martin:1988:SGC

[Mar88] Charles F. Martin. Secondgeneration CASE tools: a challenge to vendors. IEEE Software, 5(2):46– 49, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maring:1996:OOD

[Mar96] Blayne Maring. Objectoriented development of large applications. IEEE Software, 13(3):33–40, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Marceau:1998:LWJ

[Mar98] Carla Marceau. Letters: Why is Java dangerous? IEEE Software, 15(3):10, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s3010.
                                        pdf.

Martin:2000:SEP

[Mar00] Robert C. Martin. Soapbox: eXtreme programming development through dialog. IEEE Software, 17(4):12–13, July/August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s4012.pdf.

Marchesin:2004:ULR

[Mar04] Armand Marchesin. Using Linux for real-time applications. IEEE Software, 21(5):18–20, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martin:2007:PTD

[Mar07] Robert C. Martin. Professionalism and test-driven development. IEEE Software, 24(3):32–36, May/ June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mastorakis:1998:BSG

[Mas98] Nikos E. Mastorakis. Bookshelf: Stylish guide to Modula-3. IEEE Software, 15(4):89–90, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4087.
pdf.

Matsubara:1995:S

[Mat95a] Tomoo Matsubara. Soapbox. IEEE Software, 12(1): 15, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Matsumoto:1995:IRS

[Mat95b] Masao J. Matsumoto. Insider: Research to shed light on software shadows. IEEE Software, 12(5):16, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mattsson:1996:IRI

[Mat96] Christer Mattsson. ICSE 18: Research and industry come together, more or less. IEEE Software, 13(5): 104–??, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Matsubara:2000:SPC

[Mat00a] Tomoo Matsubara. Soapbox: Process certification: Double-edged sword. IEEE Software, 17(6):104–105, November/December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2000/pdf/
s6104.pdf.

Mattison:2000:BNS

[Mat00b] Harry Mattison. Bookshelf: No secrets: Global keys to business (and software) success; job training: What you never learned: High Output Management; lasting insight: Built to Last: Successful Habits of Visionary Companies; a revolutionary idea: The Fifth Discipline: The Art & Practice of the Learning Organization. IEEE Software, 17(1):97–99, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s1097.
pdf.

Matsubara:2001:JHI

[Mat01] Tomoo Matsubara. Japan:

a huge IT consumption market. IEEE Software, 18(5):77–80, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5077abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5077.pdf.

Maurer:1990:GTD

[Mau90] Peter M. Maurer. Generating test data with enhanced context-free grammars. IEEE Software, 7(4): 50–55, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maxwell:2001:CDC

[Max01] Katrina D. Maxwell. Collecting data for comparability: Benchmarking software development productivity. IEEE Software, 18(5):22–25, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/

                                        so2001/s5022abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5022.pdf.

Mellor:2005:UDD

[MBE05] Stephen Mellor, Fernando Berzal, and Christof Ebert. UML distilled: From difficulties to assets. IEEE Software, 22(3):106–109, May/June 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2005/
                                        03/s3106.pdf.

Manolescu:2008:VDD

[MBL08] Dragos Manolescu, Brian Beckman, and Benjamin Livshits. Volta: Developing distributed applications by recompiling. IEEE Software, 25(5):53–59, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mamrak:1993:BUI

[MBO93] Sandra A. Mamrak, Julie Barnes, and Conleth S. O’Connell. The benefits of using the Integrated Chameleon Architecture for data translation. IEEE Software, 10(4):82– 88, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martin:1994:PCH

[MC94] Robert L. Martin and Carl K. Chang. Pointcounterpoint: How to solve the management crisis? IEEE Software, 11(6):14– 15, November 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Menzies:2000:WTL

[MC00] Tim Menzies and Bojan Cukic. When to test less. IEEE Software, 17 (5):107–112, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so/s5107abs.htm;   http:
//dlib.computer.org/so/
books/so2000/pdf/s5107.
pdf.

McLaughlin:2003:N

[MC03a] Laurianne McLaughlin and Terry Costlow. In the news. IEEE Software, 20 (5):104–108, September/ October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5104.
pdf.

McLaughlin:2003:NAE

[MC03b] Laurianne McLaughlin and Terry Costlow. In the news: After ElcomSoft: DMCA still worries developers, researchers; software language shoudl help protect networks from hackers. IEEE Software, 20(2):86– 91, March/April 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2086.
pdf.

McLaughlin:2004:Na

[MC04] Laurianne McLaughlin and Terry Costlow. In the news. IEEE Software, 21(1):100–103, January/ February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McHugh:2000:DYR

[MCA00] John McHugh, Alan Christie, and Julia Allen. Defending yourself: The role of intrusion detection systems: (available without subscription). IEEE Software, 17(5):42–51, September/October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so/s5042abs.htm;   http:
//dlib.computer.org/so/
books/so2000/pdf/s5042.
pdf.

McCabe:1996:TCC

[McC96a] Thomas McCabe. Toolbox: Cyclomatic complexity and the year 2000. IEEE Software, 13(3):115– 117, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1996:BPA

[McC96b] Steve McConnell. Best practices: Avoiding classic mistakes. IEEE Software, 13(5):112, 111, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1996:BPD

[McC96c] Steve McConnell. Best practices: Daily build and smoke test. IEEE Software, 13(4):144, 143, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1996:BPH

[McC96d] Steve McConnell. Best practices: How to defend an unpopular schedule. IEEE Software, 13(3): 120, 118–119, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1996:BPW

[McC96e] Steve McConnell. Best practices: Who cares about software construction? IEEE Software, 13 (1):128–129, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1996:MAI

[McC96f] Steve McConnell. Missing in action: Information hiding. IEEE Software, 13 (2):128, 127, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1997:BPAb

[McC97a] Steve McConnell. Best practices: Achieving leaner software. IEEE Software, 14(6):128, 127, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6128.
pdf.

McConnell:1997:BPAa

[McC97b] Steve McConnell. Best practices: Annualized software delivery. IEEE Software, 14(1):104–??, 103, January/February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1997:BPG

[McC97c] Steve McConnell. Best practices: Gauging software readiness with defect tracking. IEEE Software, 14(3):136, 135, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1997:BPP

[McC97d] Steve McConnell. Best practices: The programmer writing. IEEE Software, 14(4):120, 119, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4120.
pdf.

McConnell:1997:BPT

[McC97e] Steve McConnell. Best practices: Tool support for project tracking. IEEE Software, 14(5):120, 119, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s5120.pdf.

McCracken:1997:PCS

[McC97f] W. Michael McCracken.

Point/counterpoint: SE education: What academia can do. IEEE Software, 14(6):27, 29, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McConnell:1998:BPF

[McC98a] Steve McConnell. Best practices: Feasibility studies. IEEE Software, 15(3):120, 119, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s3120.
                                        pdf.

McConnell:1998:BPP

[McC98b] Steve McConnell. Best practices; problem programmers. IEEE Software, 15(2):128, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s2128.
                                        pdf.

McConnell:1998:BPA

[McC98c] Steve McConnell. Best practices: The art, science, and engineering of software development. IEEE Software, 15(1): 120, 118–119, January/ February 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s1120.
                                        pdf.

McConnell:1998:BPW

[McC98d] Steve McConnell. Best practices: Why you should use routines … routinely. IEEE Software, 15(4):96, 94–95, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4096.
pdf.

McConnell:1998:EBC

[McC98e] Steve McConnell. From the Editor: Building the community. IEEE Software, 15(5):128, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5128.
pdf.

McConnell:1998:EHR

[McC98f] Steve McConnell. From the Editor: How to read a technical article. IEEE Software, 15(6):128–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6128.
pdf.

McConnell:1999:EAG

[McC99a] Steve McConnell. From the editor: After the gold rush. IEEE Software, 16(1):6–??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1006.
pdf.

McConnell:1999:EBL

[McC99b] Steve McConnell. From the Editor: Brooks’ Law repealed. IEEE Software, 16(6):6–8, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6006.
pdf.

McConnell:1999:EOS

[McC99c] Steve McConnell. From the editor: Open-source methodology: Ready for prime time? IEEE Software, 16(4):6–8, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4006.
pdf.

McConnell:1999:ESE

[McC99d] Steve McConnell. From the Editor: Software engineering principles. IEEE Software, 16(2):6–8, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2006.
pdf.

McConnell:1999:EUP

[McC99e] Steve McConnell. From the editor: Update on professional development. IEEE Software, 16(5):6–8, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1999/pdf/
s5006.pdf.

McConnell:1999:EYC

[McC99f] Steve McConnell. From the Editor: Y2K countdown to certification. IEEE Software, 16(3):4, 6–7, May/June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3004.
pdf.

McConnell:2000:ECC

[McC00a] Steve McConnell. From the Editor: Cargo cult software engineering. IEEE Software, 17(2):11–13, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so2000/pdf/s2011.
                                        pdf.

McConnell:2000:EQS

[McC00b] Steve McConnell. From the editor: Quantifying soft factors. IEEE Software, 17(6):9–11, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s6009.
                                        pdf.

McConnell:2000:ESS

[McC00c] Steve McConnell. From the editor: Sitting on the suitcase. IEEE Software, 17(3): 5–7, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s3005.
                                        pdf.

McConnell:2000:EBI

[McC00d] Steve McConnell. From the Editor: The best influences on software engineering. IEEE Software, 17(1):10–17, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s1010.
                                        pdf.

McConnell:2000:ESM

[McC00e] Steve McConnell. From the editor: The software manager’s toolkit. IEEE Software, 17(4):5–7, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4005.
pdf.

McConnell:2000:EWN

[McC00f] Steve McConnell. From the editor: What’s in a name? IEEE Software, 17(5):7–9, September/October 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5007.
pdf.

McConnell:2000:TOR

[McC00g] Steve McConnell. Thanks to our 1999 referees. IEEE Software, 17(2):94– 95, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2094.
pdf.

McConnell:2001:EOP

[McC01a] Steve McConnell. From the editor: An ounce of prevention. IEEE Software, 18(3):

5–7, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s3005.
pdf.

McConnell:2001:EAS

[McC01b] Steve McConnell. From the editor: Art, science, and engineering. IEEE Software, 18(2):9–11, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2009.
pdf.

McConnell:2001:ECS

[McC01c] Steve McConnell. From the editor: Common sense. IEEE Software, 18(4):5–7, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4005.
pdf;   http://dlib.computer.
org/so/books/so2001/s4005abs.
htm.

McConnell:2001:ERW

[McC01d] Steve McConnell. From the Editor: Roundtable: Who needs software engineering? IEEE Software, 18(1):5–8, January/ February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1005.
pdf.

McConnell:2001:NDS

[McC01e]  Steve McConnell.  The

nine deadly sins of project planning. IEEE Software, 18(5):5–7, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5005abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5005.pdf.

McConnell:2001:RYS

[McC01f] Steve McConnell. Raising your software consciousness. IEEE Software, 18(6):7–9, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6007abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6007.pdf.

McConnell:2001:RT

[McC01g]  Steve McConnell.  Re-

viewer thanks. IEEE Software, 18(1):108–111, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s1108.
                                        pdf.

McConnell:2002:CG

[McC02a] Steve McConnell. Closing the gap. IEEE Software, 19(1):3–5, January/ February 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2002/s1003abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2002/pdf/
                                        s1003.pdf.

McConnell:2002:ECG

[McC02b] Steve McConnell. From the Editor: Changing of the guard. IEEE Software, 19(6):5–7, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s6005.
                                        pdf.

McConnell:2002:EHW

[McC02c] Steve McConnell. From the Editor: How to write a good technical article. IEEE Software, 19(5):5–7, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/

books/so2002/pdf/s5005.
pdf.

McConnell:2002:EKW

[McC02d] Steve McConnell. From the Editor: I know what I know. IEEE Software, 19(3):5–7, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3005.
pdf;   http://www.computer.
org/software/so2002/s3005abs.
htm.

McConnell:2002:EBS

[McC02e] Steve McConnell. From the Editor: The business of software improvement. IEEE Software, 19(4):5–7, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4005.
pdf.

McConnell:2002:RQR

[McC02f] Steve McConnell. Real quality for real engineers. IEEE Software, 19(2):5–7, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2005.
pdf;   http://www.computer.
org/software/so2002/s2005abs.
htm.

McConnell:2002:RTS

[McC02g]  Steve McConnell.  Re-

viewer thanks: Saluting our 2001 referees. IEEE Software, 19(1):98–99, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2002/s1098abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2002/pdf/
                                        s1098.pdf.

McConnell:2006:SCP

[McC06] Steve McConnell. Software construction, part 1. IEEE Software, 23(1):99, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        01/s1099.pdf.

Mellor:2003:GEI

[MCF03] Stephen J. Mellor, Anthony N. Clark, and Takao Futagami. Guest Editors’ introduction: Model-driven development. IEEE Software, 20(5):14–18, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/05/s5014.
                                        pdf.

McKinney:2002:STB

[McK02] Dorothy McKinney. Six translations between softwarespeak and managementspeak. IEEE Software, 19(6):50–52, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6050.
pdf;   http://www.computer.
org/software/so2002/s6050abs.
htm.

Mead:1997:GEI

[MCL97] Nancy R. Mead, David Carter, and Michael Lutz. Guest Editors’ introduction: The state of SE education and training. IEEE Software, 14(6):22– 25, November/December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s6022.pdf.

Mclendon:1998:CWH

[Mcl98] Jean Mclendon. Culture at work: Healing the ailing workplace. IEEE Software, 15(3):97–99, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3097.
pdf.

Mylopoulos:2001:EAD

[MCL^{+}01] John Mylopoulos, Lawrence Chung, Stephen Liao, Huaiqing Wang, and Eric Yu. Exploring alternatives during requirements analysis. IEEE Software, 18(1):92–96, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s1092abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s1092.pdf.

McLaughlin:2005:ISP

[McL05] Laurianne McLaughlin. Inside the software patents debate: Some good news for Open Source developers. IEEE Software, 22(3): 102–105, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2005/
                                        03/s3102.pdf.

McLaughlin:2006:APN

[McL06]  Laurianne  McLaughlin.

Automated programming: The next wave of developer power tools. IEEE Software, 23(3):91–93, May/ June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        csdl.computer.org/comp/

mags/so/2006/03/s3091.
pdf.

Mcsharry:1996:SPA

[Mcs96] Maureen Mcsharry. SEW panelists ask, is process enough? IEEE Software, 13(2):115, 121, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mueller:1986:FMM

[MD86] Robert A. Mueller and Michael R. Duda. Formal methods of microcode verification and synthesis. IEEE Software, 3(4):38–48, August 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mills:1990:UMQ

[MD90] Harlan D. Mills and Peter B. Dyson. Using metrics to quantify development. IEEE Software, 7(2):14– 16, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Morris:1997:HUP

[MD97] Michael G. Morris and Andrew Dillon. How user perceptions influence software use. IEEE Software, 14(4):58–65, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4058.
pdf;   http://www.computer.
org/software/so1997/s4058abs.
htm.

McGarry:2002:ALC

[MD02] Frank McGarry and Bill Decker. Attaining level 5 in CMM process maturity. IEEE Software, 19(6):87–96, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6087.
pdf;   http://www.computer.
org/software/so2002/s6087abs.
htm.

Madanmohan:2004:OSR

[MD04] T. R. Madanmohan and Rahul De’. Open Source reuse in commercial firms. IEEE Software, 21(6):62– 69, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Moe:2009:OBS

[MDD09] Nils Brede Moe, Torgeir Dingsoyr, and Tore Dybå. Overcoming barriers to selfmanagement in software teams. IEEE Software, 26(6):20–26, November/ December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meier:1994:HRD

[MDM^{+}94] Andreas Meier, Rolf Dippold, Jacky Mercerat, Alex Muriset, Jean-Claude Untersinger, Robert Eckerlin, and Flavio Ferrara. Hierarchical to relational database migration. IEEE Software, 11(3):21–27, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Murru:2003:AXE

[MDM03] Orlando Murru, Roberto Deias, and Giampiero Mugheddu. Assessing XP at a European Internet company. IEEE Software, 20(3):37–43, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3037.
pdf;   http://www.computer.
org/software/so2003/s3037abs.
htm.

Mashayekhi:1993:DCS

[MDWTR93] Vahid Mashayekhi, Janet M.

Drake, Tsai Wei-Tek, and John Riedl. Distributed, collaborative software inspection. IEEE Software, 10(5):66–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Musa:1990:SRE

[ME90] John D. Musa and William W.

Everett. Software-reliability engineering: Technology for the 1990s. IEEE Software, 7(6):36–43, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Matsubara:2000:BAC

[ME00] Tomoo Matsubara and Christof Ebert. Benefits and applications of crosspollination: Guest Editors’ introduction. IEEE Software, 17(1):24–26, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s1024.
pdf.

Miller:2002:GEI

[ME02] Ann Miller and Christof Ebert. Guest Editors’ introduction: Software engineering as a business. IEEE Software, 19(6):18–20, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6018.
pdf.

Medvidovic:2005:SAE

[Med05] Nenad Medvidovic. Software architectures and embedded systems: a match made in heaven? IEEE Software, 22(5):83– 86, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Melford:1995:SIW

[Mel95] Robert Melford. Soapbox:

If we don’t license ourselves, who will? IEEE Software, 12(3):11, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mellor:2005:EIA

[Mel05] Stephen J. Mellor. Editor’s introduction: Adapting agile approaches to your project needs. IEEE Software, 22(3):17–20, May/ June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2005/03/s3017.
pdf.

Meyer:1985:FS

[Mey85] Bertrand Meyer. On formalism in specifications. IEEE Software, 2(1):6–26, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meyer:1987:RCO

[Mey87] Bertrand Meyer. Reusability: the case for objectoriented design. IEEE Software, 4(2):50–64, March 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meyers:1991:DIM

[Mey91] Scott Meyers. Difficulties in integrating multiview development systems. IEEE Software, 8(1):49–57, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meyers:1996:DMK

[Mey96a] Ware Meyers. Decision making: The key to successful project managment: Interview: Ron LaFleur. IEEE Software, 13(4):139– 140, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meyers:1996:NIC

[Mey96b] Ware Meyers. In the news: Interview with Christopher Meyer. IEEE Software, 13(3):100–101, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meyers:1996:INK

[Mey96c] Ware Meyers. Interview with Norm Kerth: Leading from a powerless position. IEEE Software, 13(5): 104–??, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meyers:2004:MID

[Mey04] Scott Meyers. The most important design guideline? IEEE Software, 21(4):14– 16, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maxwell:2000:BSD

[MF00] Katrina D. Maxwell and Pekka Forselius. Benchmarking software-development productivity: Applied research results. IEEE Software, 17(1):80–88, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s1080abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s1080.pdf.

Moitra:2004:B

[MF04] Deependra Moitra and Martin Fogarty. Bookshelf. IEEE Software, 21 (2):100–101, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2001:WDR

[MG01] Neil Maiden and Alexis Gizikis. Where do requirements come from?

IEEE Software, 18(5):10– 12, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2001/s5010abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5010.pdf.

Massey:2006:Nb

[MG06] Bart Massey and Greg Goth. In the news. IEEE Software, 23(6):99–103, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
06/s6099.pdf.

Merlo:1995:RUI

[MGG^{+}95] Ettore Merlo, Pierre-Yves Gagne, Jean-Francois Girard, Kostas Kontogiannis, Laurie Hendren, Prakash Panangaden, and Renato De Mori. Reengineering user interfaces. IEEE Software, 12(1):64–73, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2004:PCI

[MGR04] Neil Maiden, Alexis Gizikis, and Suzanne Robertson. Provoking creativity: Imagine what your requirements could be like. IEEE Software, 21(5):68–75, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Menzies:2006:EBC

[MH06] Tim Menzies and Jairus Hihn. Evidence-based cost estimation for betterquality software. IEEE Software, 23(4):64–66, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martin:2007:OSA

[MH07] Ken Martin and Bill Hoffman. An open source approach to developing software in a small organization. IEEE Software, 24(1): 46–53, January/February 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Murphy-Hill:2008:RTF

[MHB08] Emerson Murphy-Hill and Andrew P. Black. Refactoring tools: Fitness for purpose. IEEE Software, 25(5):38–44, September/ October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McLaughlin:2003:NEI

[MHC03]  Laurianne  McLaughlin,

Joan Hong, and Terry Costlow. In the news: An eye on India: Outsourcing debate continues; ISO approves Cosmic-FFP method; Web developers continue to focus on interoperability. IEEE Software, 20(3):114– 118, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3114.
pdf.

Malony:1991:TTV

[MHJ91] Allen D. Malony, David H. Hammerslag, and David J. Jablonowski. Traceview: a trace visualization tool. IEEE Software, 8(5):19–28, September 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Michener:1999:SII

[Mic99] John Michener. System insecurity in the Internet age. IEEE Software, 16(4):62– 69, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s4062abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s4062.pdf.

Mickel:2000:CWT

[Mic00] Sue Mickel. Culture at work: There’s a merger in your future. IEEE Software, 17(1):89–91, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s1089.
pdf.

Miles:1984:CCS

[Mil84] Dana E. Miles. Copyrighting computer software after Apple v. Franklin. IEEE Software, 1(2):84– 87, April 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mills:1986:SPR

[Mil86] Harlan D. Mills. Structured programming: Retrospect and prospect. IEEE Software, 3(6):58–66, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Miller:2001:GEI

[Mil01a] Ann Miller. Guest Editor’s introduction: Organizational change. IEEE Software, 18(3):18–20, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s3018.
pdf.

Miller:2001:ITI

[Mil01b] Ann Miller. Interview: Two interviews, two perspectives (with Eric Schmidt and Mary E. Boone). IEEE Software, 18(3):56– 59, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s3056.
pdf.

Miller:2001:QTM

[Mil01c] Keith W. Miller. Quality time: a modest proposal for software testing. IEEE Software, 18(2):98–100, March/April 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2098.
pdf.

Milicev:2002:DMU

[Mil02] Dragan Milicev. Domain mapping using Extended UML object diagrams. IEEE Software, 19(2):90– 97, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2090.
pdf;   http://www.computer.
org/software/so2002/s2090abs.
htm.

Miranda:2001:ISE

[Mir01] Eduardo Miranda. Improving subjective estimates using paired comparisons. IEEE Software, 18(1):87–91, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s1087abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s1087.pdf.

McConnell:1998:RWS

[MIZ^{+}98] Steve McConnell, Jim Isaak, Lin Zucconi, Norman F. Schneidewind, Val Varaart, and Maarten Boasson. Roundtable: Weighing in on standards. IEEE Software, 15(6):92–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6092.
pdf.

Misra:1988:TGV

[MJ88] Santosh K. Misra and Paul J. Jalics. Thirdgeneration versus fourthgeneration software development. IEEE Software, 5 (4):8–14, July 1988. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mellor:1997:WEO

[MJ97] Stephen J. Mellor and Ralph Johnson. Why explore object methods, patterns, and architectures? IEEE Software, 14(1): 27–??, January/February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Moore:2000:LKW

[MJP^{+}00] John E. Moore, Michael Jackson, David Lorge Parnas, Matt Edgar, Robert J. Schaaf, and Steve Mc-Connell. Letters: Knowing when to quit; the physics of education; the cult of software engineering. IEEE Software, 17(3): 8–10, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s3008.
                                        pdf.

Munson:1992:MDP

[MK92] John C. Munson and Taghi M. Khoshgoftaar. Measuring dynamic program complexity. IEEE Software, 9(6):48–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Murugappan:2003:BCS

[MK03] Mala Murugappan and Gargi Keeni. Blending CMM and six sigma to meet business goals. IEEE Software, 20(2):42– 48, March/April 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2042.
pdf;   http://www.computer.
org/software/so2003/s2042abs.
htm.

MacCormack:2003:TOB

[MKCC03] Alan MacCormack, Chris F.

Kemerer, Michael Cusumano, and Bill Crandall. Tradeoffs between productivity and quality in selecting software development practices. IEEE Software, 20(5):78–85, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5078.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5078abs.htm.

Murphy:2006:HJS

[MKF06] Gail C. Murphy, Mik Kersten, and Leah Findlater. How are Java software developers using the Eclipse IDE? IEEE Software, 23(4):76–83, July/August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mannion:1998:UVD

[MKH98] Mike Mannion, Barry Keepence, and David Harper. Using viewpoints to define domain requirements. IEEE Software, 15(1):95–102, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1998/s1095abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s1095.pdf.

Malloy:2006:IPA

[MKHV06] Brian A. Malloy, Nicholas A. Kraft, Jason O. Hallstrom, and Jeffrey M. Voas. Improving the predictable assembly of service-oriented architectures. IEEE Software, 23(2):12–15, March/ April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Monroe:1997:ASD

[MKMG97] Robert T. Monroe, Andrew Kompanek, Ralph Melton, and David Garlan. Architectural styles, design patterns, and objects. IEEE Software, 14(1):43–52, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Manolescu:2007:GDP

[MKMH07] Dragos Manolescu, Wojtek Kozaczynski, Ade Miller, and Jason Hogg. The growing divide in the patterns world. IEEE Software, 24(4):61–67, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Middleton:2004:WCS

[MLI04] Peter Middleton, Ho Woo Lee, and Shahrukh A. Irani. Why culling software colleagues is popular. IEEE Software, 21(5): 28–32, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mohan:2003:BCC

[MLL03] Shantha Mohan, Robert C.

Larrabee, and Dharmendra Lingaiah. Bookshelf: Commercial components: Panacea or Pandora’s Box?; business metrics for software; software radio: RF engineering’s new era. IEEE Software, 20(4):84–86, 95, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4084.

pdf. Review of Building Systems fro Commercial Components by Kurt C.

Wallnau, Scott A. Hissam, and Robert C. Seacord, Practical Software Measurement: Objective Information for Decision Makers Software Radio: A Modern Approach to Radio Engineering by Jeffrey H. Reed.

McGraw:2000:AMC

[MM00] Gary McGraw and Greg Morrisett. Attacking malicious code: a report to the Infosec Research Council. IEEE Software, 17(5):33–41, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so/s5033abs.htm;   http:
                                        //dlib.computer.org/so/
                                        books/so2000/pdf/s5033.
                                        pdf.

Martin:2005:TBI

[MM05] Robert C. Martin and Object Mentor. The test bus imperative: Architectures that support automated acceptance testing. IEEE Software, 22(4):65– 67, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Massey:2006:Na

[MM06] Bart Massey and Laurianne McLaughlin. In the news. IEEE Software, 23(1):108–113, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
01/s1108.pdf.

Mahaux:2008:TIK

[MM08a] Martin Mahaux and Neil Maiden. Theater improvisers know the requirements game. IEEE Software, 25(5):68–69, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martin:2008:TRR

[MM08b] Robert C. Martin and Grigori Melnik. Tests and requirements, requirements and tests: a Möbius strip. IEEE Software, 25(1): 54–59, January/February 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McCusker:2001:LMD

[MMM^{+}01] Patrick McCusker, William E.

McMillan, Alexey Martyushenko, Robert Glass, Francisco Torres, and David F. Rico. Letters: Math debates; vague references; PSP, again. IEEE Software, 18(1):13–15, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s1013.
                                        pdf.

Mathe:2009:MIG

[MMM^{+}09] Janos L. Mathe, Jason B. Martin, Peter Miller, Ákos Lédeczi, Liza M. Weavind, Andras Nadas, Anne Miller, David J. Maron, and Janos Sztipanovits. A modelintegrated, guideline-driven, clinical decision-support system. IEEE Software, 26(4):54–61, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mannucci:1989:GSD

[MMN^{+}89] Stefano Mannucci, Bruno Mojana, Maria Cristina Navazio, Valerio Romano, Maria Carla Terzi, and Piero Torrigiani. Graspin: a structured development environment for analysis and design. IEEE Software, 6(6):35–43, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Miller:1988:ADA

[MMS88] Keith W. Miller, Larry J. Morell, and Fred Stevens. Adding data abstraction to FORTRAN software. IEEE Software, 5(6):50–58, November 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martinez:2009:UBM

[MMSM09] Jesús Martínez, Pedro Merino, Alberto Salmerón, and Francisco Malpartida. UML-based model-driven development for HSDPA design. IEEE Software, 26 (3):26–33, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Malachi:1986:TFR

[MMW86] Yonathan Malachi, Zohar Manna, and Richard Waldinger. Tablog: Functional and relational programming in one framework. IEEE Software, 3 (1):75–76, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:1998:ACS

[MN98] Neil A. Maiden and Cornelius Ncube. Acquiring COTS software selection requirements. IEEE Software, 15(2):46–56, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s2046abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2046.pdf.

Mathiassen:2005:MCS

[MNA05] Lars Mathiassen, Ojelanki K. Ngwenyama, and Ivan Aaen. Managing change in software process improvement. IEEE Software, 22(6):84–91, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McGregor:2002:GEI

[MNJP02] John D. McGregor, Linda M. Northrop, Salah Jarrad, and Klaus Pohl. Guest Editors’ introduction: Initiating software product lines. IEEE Software, 19(4):24– 27, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s4024.
                                        pdf.

Moen:1990:DDT

[Moe90]  Sven Moen.  Drawing

dynamic trees. IEEE Software, 7(4):21–28, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mohan:2001:NIS

[Moh01]  Kalpana Mohan.  In

the news: The international software connection. IEEE Software, 18(1):100–103, January/ February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1100.
pdf.

Moitra:2001:CRI

[Moi01] Deependra Moitra. Country report: India’s software industry. IEEE Software, 18(1):77–80, January/February 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1077.
pdf.

Moore:1999:ICS

[Moo99] James W. Moore. An integrated collection of software engineering standards. IEEE Software, 16(6):51–57, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6051abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6051.pdf.

Moore:2001:YTCb

[Moo01a] Melody M. Moore. From your Technical Council: Focus on software reliability engineering. IEEE Software, 18(3):98, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s3098.
                                        pdf.

Moore:2001:YTCa

[Moo01b] Melody M. Moore. From your Technical Council: TCSE: Incubator of software engineering programs. IEEE Software, 18(1):104, January/February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s1104.pdf.

Moore:2001:SRS

[Moo01c] Melody M. Moore. Software reuse: Silver bullet? IEEE Software, 18(5):86, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s5086abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5086.pdf.

Moore:2002:RER

[Moo02a] Melody M. Moore. Reverse engineering and reengineering: After Y2K. IEEE Software, 19(2):109, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2109.
pdf;   http://www.computer.
org/software/so2002/s2109abs.
htm.

Moore:2002:TSE

[Moo02b] Melody M. Moore. TCSE:

Software engineering education. IEEE Software, 19(5):103, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5103.
pdf.

Moore:2003:LPS

[Moo03] Melody M. Moore. A license to practice software engineering. IEEE Software, 20(3):112–113, May/June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3112.
pdf.

Morisio:2000:API

[Mor00]  Maurizio Morisio.  Ap-

plying the PSP in industry. IEEE Software, 17(6):90–95, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s6090abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s6090.pdf.

Mosley:1992:HAT

[Mos92] Vicky Mosley. How to assess tools efficiently and quantitatively. IEEE Software, 9(3):29–32, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Motoyama:2006:ISD

[Mot06] Tetsuro Motoyama. Improving software development through three stages. IEEE Software, 23(5):81– 87, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Moustaki:2001:SHC

[Mou01]  Nikki Moustaki.  Self-

healing computers more than a band-aid. IEEE Software, 18(4):99, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2001/pdf/s4102.
                                        pdf.

Moynihan:1997:HEP

[Moy97] Tony Moynihan. How experienced project managers assess risk. IEEE Software, 14(3):35–41, May/ June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

www.computer.org/pubs/
software/1997/s3035abs.
htm.

Meyers:1988:MPW

[MP88] Richard J. Meyers and Jeff W. Parrish. Macintosh programmer’s workshop. IEEE Software, 5(3): 59–66, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mariani:2007:DDC

[MP07] Leonardo Mariani and Mauro Pezzè. Dynamic detection of COTS component incompatibility. IEEE Software, 24(5):76– 85, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Madhavji:1986:MML

[MPT86] Nazim H. Madhavji, Luc Pinsonneault, and Kamel Toubache. Modula-2/MUPE-2: Language and environment interactions. IEEE Software, 3(6):7–17, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Malony:1990:PAV

[MR90] Allen Malony and Daniel Reed. Performance analysis, visualization with hyperview. IEEE Software, 5 (3):26, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). [Abstract].

Madhyastha:1995:DSD

[MR95] Tara M. Madhyastha and Daniel A. Reed. Data sonification: Do you see what I hear? IEEE Software, 12(2):45–56, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McDaid:2009:TDD

[MR09] Kevin McDaid and Alan Rust. Test-driven development for spreadsheet risk management. IEEE Software, 26(5):31–36, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Messer:1998:BJU

[MRA98] Robert Messer, Steven R. Rakitin, and Charles Ashbacher. Bookshelf: Java, the universe, and everything: Java How to Program with an Introduction to J++; Software Quality Assurance and Measurement: A Worldwide Perspective; Object Solutions: Managing the Object-Oriented Process. IEEE Software, 15(3):115– 118, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3115.
pdf.

Maiden:2005:GEI

[MRE05] Neil Maiden, Suzanne Robertson, and Christof Ebert. Guest Editors’ introduction: Shake, rattle, and requirements. IEEE Software, 22(1):13–15, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
01/s1013.pdf.

McLellan:1998:BMU

[MRTS98] Samuel G. McLellan, Alvin W.

Roesler, Joseph T. Tempest, and Clay I. Spinuzzi. Building more usable APIs. IEEE Software, 15(3):78– 86, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3078abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3078.pdf.

Maranzano:2005:ARP

[MRZ^{+}05] Joseph F. Maranzano, Sandra A. Rozsypal, Gus H. Zimmerman, Guy W. Warnken, Patricia E. Wirth, and David M. Weiss. Architecture reviews: Practice and experience. IEEE Software, 22(2):34–43, March/ April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mi:1992:PIC

[MS92] Peiwei Mi and Walt Scacchi. Process integration in CASE environments. IEEE Software, 9(2):45– 53, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Millington:1995:SRD

[MS95] Don Millington and Jennifer Stapleton. Special report: Developing a RAD standard. IEEE Software, 12(5):54–55, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Muller:1999:HHA

[MS99] Gilles Muller and Ulrik Pagh Schultz. Harissa: a hybrid approach to Java execution. IEEE Software, 16(2):44–51, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s2044abs.htm;
                                        http://dlib.computer.

org/so/books/so1999/pdf/
s2044.pdf.

Messerschmitt:2004:MIS

[MS04] David G. Messerschmitt and Clemens Szyperski. Marketplace issues in software planning and design. IEEE Software, 21(3):62– 70, May/June 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Murphy:2006:GEI

[MS06] Gail Murphy and Christa Schwanninger. Guest Editors’ introduction: Aspectoriented programming. IEEE Software, 23(1):20–23, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
01/s1020.pdf.

Macala:1996:MDS

[MSG96] Randall R. Macala, Lynn D.

Stuckey, Jr., and David C. Gross. Managing domainspecific, product-line development. IEEE Software, 13 (3):57–67, May 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2007:DSN

[MSG^{+}07] Neil Maiden, Norbert Seyff, Paul Grünbacher, Omo O. Otojare, and Karl Mitteregger. Determining stakeholder needs in the workplace: How mobile technologies can help. IEEE Software, 24(2):46– 52, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Macaulay:2009:UUC

[MSJ^{+}09] Catriona Macaulay, David Sloan, Xinyi Jiang, Paula Forbes, Scott Loynton, Jason R. Swedlow, and Peter Gregor. Usability and user-centered design in scientific software development. IEEE Software, 26(1):96–102, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Moitra:2000:BAS

[MSK00] Deependra Moitra, Diomidis Spinellis, and Robert Bruce Kelsey. Bookshelf: Adaptive Software Development: a Collaborative Approach to Managing Complex Systems; eXtreme Programming Explained: Embrace Change; Software Process Quality: Management and Control. IEEE Software, 17(4):112–115, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so2000/pdf/s4112.
pdf.

McConnell:1999:PSE

[MT99] Steve McConnell and Leonard Tripp. Professional software engineering: Fact or fiction? IEEE Software, 16(6):13–18, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6013.
pdf.

Morisio:2000:DRP

[MTE00] Maurizio Morisio, Colin Tully, and Michel Ezran. Diversity in reuse processes. IEEE Software, 17(4):56–63, July/August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2000/s4056abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4056.pdf.

Mojdehbakhsh:1994:RSS

[MTKE94] Ramin Mojdehbakhsh, Wei-Tek Tsai, Shekhar Kirani, and Lynn Elliott. Retrofitting software safety in an implantable medical device. IEEE Software, 11 (1):41–50, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mubarak:2008:DFS

[Mub08] Hisham Mubarak. Developing flexible software using agent-oriented software engineering. IEEE Software, 25(5):12–15, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mugridge:2008:MAP

[Mug08] Rick Mugridge. Managing Agile project requirements with Storytest-driven development. IEEE Software, 25(1):68–75, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Muller:1986:MPS

[Mul86] Carlo Muller. Modula-Prolog: a software development tool. IEEE Software, 3(6):39–45, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Munson:1998:PSL

[Mun98] John C. Munson. Point: Software lives too long. IEEE Software, 15(4):18, 20, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s4018.
                                        pdf.

Murphy:1995:EIE

[Mur95] Melissa J. Murphy. Enterprise integration extends to people. IEEE Software, 12 (1):16, January 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Musa:1985:SEF

[Mus85] John D. Musa. Software engineering: the future of a profession. IEEE Software, 2(1):55–62, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Musa:1993:OPS

[Mus93] John D. Musa. Operational profiles in softwarereliability engineering. IEEE Software, 10(2):14–??, March 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McCracken:1999:LSR

[MVS^{+}99] Daniel D. McCracken, Robert Vanderwall, Wolfgang Strigel, Ambrose Li, Roger Sessions, Martin Larsson, Eric Raymond, and Terry Bollinger. Letters: Setting the record straight; to train or not to train; quality vs. hype; getting quantitative results. IEEE Software, 16(4):12– 14, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4012.
pdf.

Moher:1991:OHL

[MW91] Thomas G. Moher and Paul R. Wilson. Offsetting human limitations with debugging technology. IEEE Software, 8(3):11– 13, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McLendon:1996:BBC

[MW96] Jean McLendon and Gerald M. Weinberg. Beyond blaming: congruence in large systems development projects. IEEE Software, 13(4):33–42, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mockus:2001:GCQ

[MW01] Audris Mockus and David M. Weiss. Globalization by chunking: a quantitative approach. IEEE Software, 18(2):30–37, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2030abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2030.pdf.

Mignolet:2009:EMS

[MW09] Jean-Yves Mignolet and Roel Wuyts. Embedded multiprocessor systems-on-chip programming. IEEE Software, 26(3):34–41, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1989:UIT

[Mye89a] Brad A. Myers. Userinterface tools: Introduction and survey. IEEE Software, 6(1):15–23, January 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1989:APT

[Mye89b] Ware Myers. Allow plenty of time for large-scale software. IEEE Software, 6(4): 92–99, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1994:PWE

[Mye94]  Ware Myers.  Proceed-

ings: Workshop explores large-grained reuse. IEEE Software, 11(1):108–109, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1995:IKH

[Mye95a]  Ware Myers.  ICSE

keynotes hit on manmonth, Microsoft. IEEE Software, 12(4):105–108, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1995:IYP

[Mye95b]  Ware Myers.  Improve

your personal software process: review of A Discipline for Software Engineering. IEEE Software, 12(3):104– 105, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1995:NDM

[Mye95c] Ware Myers. In the news: Does Microsoft portend the future? interview with Michael Cusumano. IEEE Software, 12(6):110–111, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1995:PPP

[Mye95d] Ware Myers. Pushing process with a personal approach: Interview with Watts Humphrey, Software Engineering Institute. IEEE Software, 12(4):102– 103, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Myers:1998:LSP

[Mye98]  Ware Myers.  Letters:

Scenarios prove useful in many practical contexts. IEEE Software, 15(3):14, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3010.
pdf.

Myers:2001:IDN

[Mye01]  Ware Myers.  Inter-

view: Developers need some slack. IEEE Software, 18(4):90–92, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4090.
pdf.

Myers:2002:IJS

[Mye02] Ware Myers. Ivar Jacobson: Shaping software development. IEEE Software, 19(3):93–95, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3093.
pdf.

Nakakoji:1996:BLT

[Nak96] Kumiyo Nakakoji. Beyond language translation: Crossing the cultural divide. IEEE Software, 13 (6):42–46, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Narain:1985:MIE

[Nar85] Sanjai Narain. Mycin: Implementing the expert system in Loglisp. IEEE Software, 2(3):83–88, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Naughton:1998:BJA

[Nau98] Patrick Naughton. Basic to Java: Assembling a career. IEEE Software, 15(1):14–17, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s1014.
                                        pdf.

Naveda:1995:BST

[Nav95] J. Fernando Naveda. Bookshelf: Solid text on software engineering: review of Software Engineering with Abstractions. IEEE Software, 12(4):113, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Norman:1992:WTI

[NC92] Ronald J. Norman and Minder Chen. Working together to integrate CASE. IEEE Software, 9(2):12– ??, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nunes:2000:WSE

[NC00] Nuno J. Nunes and João F.

Cunha. Wisdom: a software engineering method for small software development companies. IEEE Software, 17(5):113–119, September/October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so/s5113abs.htm;   http:
//dlib.computer.org/so/
books/so2000/pdf/s5113.
pdf.

Nyman:2002:LHH

[NCC^{+}02] Jeff Nyman, William Carroll, Bob Corrick, Martin Fowler, Brian Henderson-Sellers, Jh Boddie, and T. Scott Ankrum. Letters: Handling the hurdles; to be, or not to be, explicit; not RUP, not XP; what is architecture?; bad news. IEEE Software, 19(2):8–10, March/April 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2008.
pdf;   http://www.computer.
org/software/so2002/s2008abs.
htm.

Narayana:1990:FSL

[ND90] K. T. Narayana and S. Dharap. Formal specification of a look manager. IEEE Transactions on Software Engineering, 16(9):1089–1103, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Nidiffer:2005:EDP

[ND05] Kenneth E. Nidiffer and Dana Dolan. Evolving distributed project management. IEEE Software, 22(5):63–72, September/ October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nichols:1990:PT

[NDJ^{+}90] Kathleen M. Nichols, Donald Dubois, Carol Janczewski, Jon Flower, David Flanagan, Jerry Yan, Allen Malony, Daniel Reed, Nakul Saraiya, Larry Snyder, David Krumme, Alva Couch, and Tokuda Hideyuki. Performance tools. IEEE Software, 7(3):21–30, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Niu:2007:YTY

[NE07] Nan Niu and Steve Easterbrook. So, you think you know others’ goals? A repertory grid study. IEEE Software, 24(2):53– 61, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nerurkar:2001:WUI

[Ner01] Uttara Nerurkar. Web user interface design: Forgotten lessons. IEEE Software, 18(6):69–71, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6069abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6069.pdf.

Nesi:1998:MOP

[Nes98] Paolo Nesi. Managing OO projects better. IEEE Software, 15(4):50–60, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
cg1998/s4050abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s4050.pdf.

Newcomb:1995:TWB

[New95] Philip Newcomb. Toolbox: Web-based businessprocess reengineering. IEEE Software, 12(6):116–117, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

NattochDag:2005:LEA

[NGBR05] Johan Natt och Dag, Vincenzo Gervasi, Sjaak Brinkkemper, and Bjorn Regnell. A linguisticengineering approach to large-scale requirements management. IEEE Software, 22(1):32–39, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nicinski:1992:ETA

[Nic92] Tom Nicinski. Extending Teamwork for architecture diagrams. IEEE Software, 9(3):54–60, May 1992. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1995:IHP

[Nie95a] Jakob Nielsen. Interface: a home-page overhaul using other Web sites. IEEE Software, 12(3):75–78, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1995:IUP

[Nie95b]  Jakob Nielsen.  Inter-

face: Using paper prototypes in home-page design. IEEE Software, 12(4):88– 89, 97, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). See letter and response in September 1995 issue, p. 14.

Nielsen:1996:IDF

[Nie96a] Jacob Nielsen. The impending demise of the file system. IEEE Software, 13(2):100–101, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1996:IPU

[Nie96b] Jacob Nielsen. Interface:

Putting the user in userinterface testing. IEEE Software, 13(3):89–90, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1996:IDS

[Nie96c] Jakob Nielsen. Interface:

Designing to seduce the user. IEEE Software, 13 (5):18–20, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1996:IGG

[Nie96d] Jakob Nielsen. Interface:

Going global with user testing. IEEE Software, 13(4): 129–130, July 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1996:UMT

[Nie96e] Jakon Nielsen. Usability metrics: Tracking interface improvements. IEEE Software, 13(6):12– 13, November 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1997:IUM

[Nie97a] Jacob Nielsen. Interface: The use and misuse of focus groups. IEEE Software, 14(1):94–95, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1997:CSB

[Nie97b] Jakob Nielsen. Counterpoint: Something is better than nothing. IEEE Software, 14(4):27–28, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4026.
pdf.

Nielsen:1997:ILR

[Nie97c]  Jakob Nielsen.  Inter-

face: Learning from the real world. IEEE Software, 14(4):98–99, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4098.
pdf.

Nielsen:1997:ILA

[Nie97d] Jakob Nielsen. Interface: Let’s ask the users. IEEE Software, 14(3):110–111, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nisse:2000:CWL

[Nis00]  David Nisse.  Culture

at work: Leadership, Army style. IEEE Software, 17(2):92–94, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2092.
pdf.

Nissen:1996:MMR

[NJJ^{+}96] Hans W. Nissen, Manfred A. Jeusfeld, Matthias Jarke, Georg V. Zemanek, and Harald Huber. Managing multiple requirements perspectives with metamodels. IEEE Software, 13(2):37–48, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Norris:2004:MCD

[NK04] Jeffrey S. Norris and Poul-Henning Kamp. Missioncritical development with Open Source software: Lessons learned. IEEE Software, 21(1):42–49, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Neill:2003:RES

[NL03] Colin J. Neill and Phillip A. Laplante. Requirements engineering: The state of the practice. IEEE Software, 20(6):40–45, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/06/s6040.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/so/2003/06/
                                        s6040abs.htm.

Nielsen:1996:SSS

[NM96] Jim Nielsen and Ann Miller. Selecting software subcontractors. IEEE Software, 13(4):104–109, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nichols:1991:NCA

[NO91] Kathleen Nichols and Paul Oman. Navigating complexity to achieve high performance. IEEE Software, 8(5):12–??, September 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ncube:2008:GEI

[NOK08] Cornelius Ncube, Patricia Oberndorf, and Anatol W. Kark. Guest Editors’ introduction: Opportunistic software systems development: Making systems from what’s available. IEEE Software, 25(6):38– 41, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nolan:1999:LS

[Nol99] Andrew J. Nolan. Learning from success. IEEE Software, 16(1):97–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1097.
pdf.

Northrop:2002:SSP

[Nor02] Linda M. Northrop. SEI’s software product line tenets. IEEE Software, 19(4):32– 40, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4032.
pdf;   http://www.computer.
org/software/so2002/s4032abs.
htm.

Nordberg:2003:MCO

[Nor03] Martin E. Nordberg III.

Managing code ownership. IEEE Software, 20(2):26– 33, March/April 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2026.

                                        pdf;   http://www.computer.
                                        org/software/so2003/s2026abs.
                                        htm.

Norden:2007:SRE

[Nor07] Barbara Norden. Screenwriting for requirements engineers. IEEE Software, 24(4):26–27, July/August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nielsen:1989:ELU

[NR89] Jakob Nielsen and John T. Richards. Experience of learning and using Smalltalk. IEEE Software, 6(3):73–77, May 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Naik:1992:TCP

[NS92] Kshirasagar Naik and Behcet Sarikaya. Testing communication protocols. IEEE Software, 9(1):27–37, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Noack:1999:IOD

[NS99] Jörg Noack and Bruno Schienmann. Introducing OO development in a large banking organization. IEEE Software, 16(3):71– 81, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s3071abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s3071.pdf.

Naveda:2005:PCS

[NS05] J. Fernando Naveda and Stephen B. Seidman. Professional certification of software engineers: The CSDP program. IEEE Software, 22(5):73–77, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Neto:2008:IEA

[NSV^{+}08] Arilo Dias Neto, Rajesh Subramanyan, Marlon Vieira, Guilherme Horta Travassos, and Forrest Shull. Improving evidence about software technologies: a look at model-based testing. IEEE Software, 25 (3):10–13, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2008/
extras/mso2008030010x1.
pdf.

Nelson:1992:UBS

[NT92] Peter C. Nelson and Anestis A. Toptsis. Unidirectional and bidirectional search algorithms. IEEE Software, 9(2):77– 83, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nejmeh:2002:BDP

[NT02] Brian Nejmeh and Ian Thomas. Business-driven product planning using feature vectors and increments. IEEE Software, 19(6):34–42, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s6034.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s6034abs.
                                        htm.

Nord:2006:SAC

[NT06] Robert L. Nord and James E. Tomayko. Software architecturecentric methods and agile development. IEEE Software, 23(2):47–53, March/ April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nusenoff:1996:BSS

[Nus96]  Ron Nusenoff.  Book-

shelf: Software success: Microsoft’s method. IEEE Software, 13(3):107–109, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nuseibeh:1997:SAW

[Nus97] Bashar Nuseibeh. Soapbox: Ariane 5: Who dunnit? IEEE Software, 14 (3):15–16, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Newkirk:2002:DHN

[NV02] James Newkirk and Alexei A.

Vorontsov. Design: How .NET’s custom attributes affect design. IEEE Software, 19(5):18–20, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5018.
pdf.

Nano:2007:GEI

[NZ07] Olivier Nano and Andrea Zisman. Guest Editors’ introduction: Realizing service-centric software systems. IEEE Software, 24(6):28–30, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
06/mso2007060028.pdf.

Nagappan:2009:MSA

[NZZ09]  Nachiappan  Nagappan,

Andreas Zeller, and Thomas Zimmermann. Mining software archives: Guest Editors’ introduction: Mining software archives. IEEE Software, 26(1):24–25, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oates:1998:TED

[Oat98]  Joe Oates.  Toolbox:

Evaluating data warehouse toolkits. IEEE Software, 15(1):52–54, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1052.
pdf.

Oman:1990:CAD

[OBM^{+}90] Paul W. Oman, Adrian J. Bowles, Ron Mount, Gerald Karam, David Kalinsky, Mikko Tervonen, Victoria Bundonis, Herm Fischer, Michael Fish, Debbie Longshore, and Nobby Akiha. CASE: Analysis and design tools. IEEE Software, 7(3):37–43, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ofuonye:2008:HDW

[OBR^{+}08] Ejike Ofuonye, Patricia Beatty, Ian Reay, Scott Dick, and James Miller. How do we build trust into E-commerce Web sites?

IEEE Software, 25(5):7–9, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ozkaya:2008:MPU

[OBSN08] Ipek Ozkaya, Len Bass, Raghvinder S. Sangwan, and Robert L. Nord. Making practical use of quality attribute information. IEEE Software, 25(2):25– 33, March/April 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oldehoeft:1988:APS

[OC88] Rodney R. Oldehoeft and David C. Cann. Applicative parallelism on a shared-memory multiprocessor. IEEE Software, 5 (1):62–70, January 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oman:1990:BPI

[OC90a] Paul W. Oman and Curtis R. Cook. The book paradigm for improved maintenance. IEEE Software, 7(1):39–45, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Osborne:1990:FPM

[OC90b] Wilma M. Osborne and Elliot J. Chikofsky. Fitting pieces to the maintenance puzzle. IEEE Software, 7(1):11–??, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Olsson:1991:SDH

[OCHW91] Ronald A. Olsson, Richard H. Crawford, W. Wilson Ho, and Christopher E. Wee. Sequential debugging at a high level of abstraction. IEEE Software, 8(3):27– 36, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Organick:1984:TAP

[OCM^{+}84] E. I. Organick, T. M. Carter, M. P. Maloney, A. Davis, A. B. Hayes, D. Klass, G. Lindstrom, B. E. Nelson, and K. F. Smith. Transforming an Ada program unit to silicon and verifying its behavior in an Ada environment: a first experiment. IEEE Software, 1(1):31–38, 40, 42–49, January 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

ODay:1998:OH

[O’D98] Dan O’Day. This old house. IEEE Software, 15(2):72– 75, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/

so1998/s2072abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2072.pdf.

Offutt:2002:QAW

[Off02] Jeff Offutt. Quality attributes of Web software applications. IEEE Software, 19(2):25–32, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2025.
pdf;   http://www.computer.
org/software/so2002/s2025abs.
htm.

Obrenovic:2007:OSS

[OG07] Željko Obrenovic and Dragan Gaševic. Open source software: All you do is put it together. IEEE Software, 24(5):86–95, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Obrenovic:2008:SCT

[OGE08] Zeljko Obrenovic, Dragan Gasevic, and Anton Eliëns. Stimulating creativity through opportunistic software development. IEEE Software, 25(6):64– 70, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oren:2007:FIF

[OHM^{+}07] Eyal Oren, Armin Haller, Cédric Mesnage, Manfred Hauswirth, Benjamin Heitmann, and Stefan Decker. A flexible integration framework for Semantic Web 2.0 applications. IEEE Software, 24(5):64–71, September/ October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Offen:1997:ESM

[OJ97] R. J. Offen and R. Jeffery. Establishing software measurement programs. IEEE Software, 14(2):45–53, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

OLeary:2009:MCI

[O’L09] Daniel O’Leary. The most cited IEEE Software articles. IEEE Software, 26(1): 12–14, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ortin:2004:SAP

[OLPS04] Francisco Ortin, Benjamin Lopez, and J. Baltasar Garcia Perez-Schofield. Separating adaptable persistence attributes through computational reflection. IEEE Software, 21(6):41– 49, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Olson:1985:PPM

[Ols85] Robert Olson. Parallel processing in a message-based operating system. IEEE Software, 2(4):39–49, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Olsen:1993:SRH

[Ols93] Neil C. Olsen. The software rush hour. IEEE Software, 10(5):29–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Olsen:1995:SFI

[Ols95]  Neil C. Olsen.  Sur-

vival of the fastest: improving service velocity. IEEE Software, 12(5):28– 38, September 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oman:1989:OLC

[Oma89] Paul Oman. Objective look at C++ environments. IEEE Software, 6(2):100– 103, 105, March 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

OConnor:1994:RCC

[OMTHC94] James O’Connor, Catharine Mansour, Jerri Turner-Harris, and Grady H. Campbell, Jr. Reuse in command-and-control systems. IEEE Software, 11 (5):70–79, September 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oberg:1992:ERG

[ON92] Bruce Oberg and David Notkin. Error reporting with graduated color. IEEE Software, 9(6):33–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

ONeill:1997:SIS

[O’N97] Don O’Neill. Soapbox: Issues in software inspection. IEEE Software, 14(1): 18–19, January/February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oman:1990:MT

[ONR^{+}90] Paul Oman, Andrew Novobilski, Vaclav Rajlich, Joel Harband, Thomas Mc-Cabe, Jr., James Cross, Leonard Vanek, Linda Davis, Keith Gallagher, and Norman Wilde. Maintenance tools. IEEE Software, 7(3):59–65, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Olson:1992:IDM

[OO92] Gary M. Olson and Judith Olson, S. Interface — defining a metaphor for group work. IEEE Software, 9(3): 93–95, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Orr:2004:ARO

[Orr04] Ken Orr. Agile requirements: Opportunity or oxymoron? IEEE Software, 21(3):71–73, May/ June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oshana:1998:TCI

[Osh98] Robert S. Oshana. Tailoring cleanroom for industrial use. IEEE Software, 15(6):46–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s6046abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s6046.pdf.

Onoma:1995:PST

[OY95] Akira K. Onoma and Tsuneo Yamaura. Practical steps toward quality development. IEEE Software, 12 (5):68–77, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Orme:2006:CMO

[OYE06] Anthony M. Orme, Haining Yao, and Letha H. Etzkorn. Coupling metrics for ontology-based systems. IEEE Software, 23 (2):102–108, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pressman:1998:RCI

[PAD^{+}98] Roger S. Pressman, Ben Adida, Tom DeMarco, Thomas Gilb, Brent Gorda, Watts Humphrey, Ray Johnson, Ted Lewis, and Ellen Ullman. Roundtable: Can Internet-based applications be engineered? IEEE Software, 15(5):104– 109, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s5104.pdf.

Paech:2008:WRE

[Pae08] Barbara Paech. What is a requirements engineer? IEEE Software, 25(4):16– 17, July/August 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Paige:1986:PI

[Pai86] Robert Paige. Programming with invariants. IEEE Software, 3(1):56–69, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Palmer:1996:SBL

[Pal96] John Palmer. Soapbox: Beware the lure of software slogans. IEEE Software, 13 (3):18–20, May 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Palma:2001:BMG

[Pal01a] Pedro Sánchez Palma.

Bookshelf: Making the grade: Modular Development of Object-Oriented Programs. IEEE Software, 18(4):109, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4108.
pdf.

Palshikar:2001:AFS

[Pal01b] Girish Keshav Palshikar.

Applying formal specifications to real-world software development. IEEE Software, 18(6):89–97, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6089abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6089.pdf.

Parke:1996:BOR

[Par96a] Tom Parke. Bookshelf: Objectifying real-time systems: a down-to-earth guide. IEEE Software, 13 (2):123, March 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Parris:1996:IA

[Par96b] Kathy V. C. Parris. Implementing accountability. IEEE Software, 13(4):83– 93, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Parnas:1999:SEP

[Par99] David Lorge Parnas. Software engineering programs are not computer science programs. IEEE Software, 16(6):19–30, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6019abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6019.pdf.

Parsons:2003:CWC

[Par03] Rebecca Parsons. Components and the world of chaos. IEEE Software, 20(3):83–85, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3083.
pdf.

Parsons:2005:EAJ

[Par05] Rebecca J. Parsons. Enterprise architects join the team. IEEE Software, 22(5):16–17, September/ October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Patton:2007:UUC

[Pat07] Jeff Patton. Understanding user centricity. IEEE Software, 24(6):9–11, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Patton:2008:ABV

[Pat08a] Jeff Patton. Ambiguous business value harms software products. IEEE Software, 25(1):50–51, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Patton:2008:CMS

[Pat08b] Jeff Patton. Consider multiple solutions. IEEE Software, 25(5):72–73, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Patton:2008:GSR

[Pat08c] Jeff Patton. Getting software RITE. IEEE Software, 25(3):20–21, May/ June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Patton:2008:UCC

[Pat08d] Jeff Patton. User centric: a conversation with Alan Cooper: The origin of interaction design. IEEE Software, 25(6):15–17, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Patton:2009:UCL

[Pat09] Jeff Patton. User centric: Leah Buley: Toward collaborative, pragmatic userexperience work. IEEE Software, 26(4):93–94, July/ August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Paulk:1985:ANC

[Pau85] Mark C. Paulk. ARC network: a case study.

IEEE Software, 2(3):62– 69, May 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Paulk:1995:HIC

[Pau95] Mark C. Paulk. How ISO 9001 compares with the CMM. IEEE Software, 12(1):74–83, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Paulk:2001:EPC

[Pau01] Mark C. Paulk. Extreme Programming from a CMM perspective. IEEE Software, 18(6):19–26, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6019abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6019.pdf.

Pawson:2002:NO

[Paw02] Richard Pawson. Naked objects. IEEE Software, 19(4):81–83, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4081.
pdf.

Page:1998:LRW

[PB98] John Page and Robert Binder. Letters: Reuse: What’s wrong with this word? IEEE Software, 15(3):13–14, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s3010.
                                        pdf.

Pfleeger:1997:GEI

[PC97] Charles P. Pfleeger and Deborah M. Cooper. Guest Editor’s introduction: Security and privacy: Promising advances. IEEE Software, 14(5):27–32, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Parnas:2009:PC

[PC09] David Lorge Parnas and Bill Curtis. Point/ counterpoint. IEEE Software, 26(6):56–59, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Paulk:1993:CMM

[PCCW93] Mark C. Paulk, Bill Curtis, Mary Beth Chrissis, and Charles V. Weber. Capability maturity model, version 1.1. IEEE Software, 10(4):

18–??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Puketza:1997:SPT

[PCOM97] Nicholas Puketza, Mandy Chung, Ronald A. Olsson, and Biswanath Mukherjee. A software platform for testing intrusion detection systems. IEEE Software, 14(5):43–51, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5043.
pdf;   http://www.computer.
org/software/so1997/s5043abs.
htm.

Plaisant:1995:IBT

[PCS95] Catherine Plaisant, David Carr, and Ben Shneiderman. Image-browser taxonomy and guidelines for developers. IEEE Software, 12(2):21–32, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Prieto-Diaz:1993:SRS

[PD93] Ruben Prieto-Diaz. Status report: Software reusability. IEEE Software, 10(3): 61–66, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Prieto-Diaz:1987:CSR

[PDF87] Ruben Prieto-Diaz and Peter Freeman. Classifying software for reusability. IEEE Software, 4(1):6– 16, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Powell:1997:ASC

[PDHT97] Gerald M. Powell, Jorge L. Diaz-Herrera, and Dennis J. Turner. Achieving synergy in collaborative education. IEEE Software, 14(6):58–65, November/December 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s6058.
                                        pdf;   http://dlib.computer.
                                        org/so/books/so1997/pdf/
                                        s6058abs.htm.

Pei:1993:SRR

[PDK93] Hsia Pei, Alan Davis, and David Kung. Status report: Requirements engineering. IEEE Software, 10 (6):75–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Paech:2005:IRE

[PDK05] Barbara Paech, Jorg Dorr, and Mathias Koehler. Improving requirements engineering communication in multiproject environments.

IEEE Software, 22(1): 40–47, January/February 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pemberton:1987:ASL

[Pem87] Steven Pemberton. Alternative simple language and environment for PCs. IEEE Software, 4(1):56–64, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Perlman:1990:TUI

[Per90] Gary Perlman. Teaching user-interface development. IEEE Software, 7(6):85–86, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Periyasamy:1995:AOB

[Per95] Kasilingam Periyasamy.

Apples and oranges are both fruit. IEEE Software, 12(2):118–119, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Perkins:1999:CCR

[Per99]  Greg Perkins.  Culture

clash and the road to world domination. IEEE Software, 16(1):80–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s1080abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s1080.pdf.

Pescio:1997:SWP

[Pes97]  Carlo Pescio.  Soap-

box: When past solutions cause future problems. IEEE Software, 14(5):19–21, September/ October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5019.
pdf.

Pescio:2006:LYT

[Pes06] Carlo Pescio. Listen to your tools and materials. IEEE Software, 23(5): 74–80, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Petschenik:1985:PPS

[Pet85] Nathan H. Petschenik. Practical priorities in system testing. IEEE Software, 2(5):18–23, September 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Petersen:1999:MMD

[Pet99]  Sue Petersen.  Man-

ager: The manager’s dance.

IEEE Software, 16(3):91– 93, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3091.
pdf.

Palmer:1992:IER

[PF92] James D. Palmer and N. Ann Fields. An integrated environment for requirements engineering. IEEE Software, 9(3):80– 85, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:1993:LLB

[Pfl93] Shari Lawrence Pfleeger.

Lessons learned in building a corporate metrics program. IEEE Software, 10 (3):67–74, May 1993. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:1995:ULC

[Pfl95a]  Charles Pfleeger.  Un-

cryptic look at cryptography. IEEE Software, 12 (1):121–123, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). Review of Bruce Schneier’s Applied Cryptography.

Pfleeger:1995:MRW

[Pfl95b] Shari Lawrence Pfleeger. Making reuse work at Hewlett–Packard. IEEE Software, 12(1):105–107, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:1996:MRC

[Pfl96] Shari Lawrence Pfleeger. Measuring reuse: a cautionary tale. IEEE Software, 13(4):118–127, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:1997:QTF

[Pfl97a] Charles P. Pfleeger. Quality time: The fundamentals of information security. IEEE Software, 14 (1):15–16, 60, January/ February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:1997:BEM

[Pfl97b] Shari Lawrence Pfleeger. Bookshelf: Enduring monument to software architecture: Software Architecture: Perspectives on an Emerging Discipline, by Mary Shaw and David Garlan. IEEE Software, 14(4):114–118, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4114.
pdf.

Pfleeger:1997:QTG

[Pfl97c] Shari Lawrence Pfleeger.

Quality time: Guidelines for applying research results. IEEE Software, 14(3):102–104, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:1998:QTN

[Pfl98] Shari Lawrence Pfleeger.

Quality time: The nature of system change. IEEE Software, 15(3):87– 90, May/June 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3087.
pdf.

Pfleeger:2002:WSE

[Pfl02] Shari Lawrence Pfleeger.

What software engineering can learn from soccer. IEEE Software, 19(6):64–65, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6064.
pdf.

Pfleeger:2005:SAR

[Pfl05] Shari Lawrence Pfleeger. Soup or art? the role of evidential force in empirical software engineering. IEEE Software, 22(1):66–73, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:2008:VES

[Pfl08] Shari Lawrence Pfleeger. Voice of evidence: Software metrics: Progress after 25 years? IEEE Software, 25(6):32–34, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pinheiro:1996:OOT

[PG96] Francisco A. C. Pinheiro and Joseph A. Goguen. An object oriented tool for tracing requirements. IEEE Software, 13(2):52– 64, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pulli:1993:CER

[PH93] Petri J. Pulli and Marko P. Heikkinen. Concurrent engineering for real-time systems. IEEE Software, 10 (6):39–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Poole:2001:UEP

[PH01] Charles Poole and Jan Willem Huisman. Using Extreme Programming in a maintenance environment. IEEE Software, 18(6):42–50, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6042abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6042.pdf.

Phillips:1996:MPM

[Phi96] Dwayne Phillips. Manager:

Project management: Filling in the gaps. IEEE Software, 13(4):17–18, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ponamgi:1991:DMP

[PHK91] M. Krish Ponamgi, Wenwey Hseush, and Gail E. Kaiser. Debugging multithreaded programs with MPD. IEEE Software, 8(3): 37–43, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ploski:2007:IVC

[PHRS07] Jan Ploski, Wilhelm Hasselbring, Jochen Rehwinkel, and Stefan Schwierz. Introducing version control to database-centric applications in a small enterprise. IEEE Software, 24(1): 38–44, January/February 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pittman:1993:LLM

[Pit93] Matthew Pittman. Lessons learned in managing object-oriented development. IEEE Software, 10(1):43–??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pitterman:2000:TTJ

[Pit00] Bill Pitterman. Telcordia technologies: The journey to high maturity. IEEE Software, 17(4):89– 96, July/August 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4089abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4089.pdf.

Pfleeger:1997:SRS

[PJCK97] S. L. Pfleeger, R. Jeffery, B. Curtis, and B. Kitchenham. Status report on software measurement. IEEE Software, 14(2):33– 43, March/April 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pankratius:2009:PBC

[PJT09] Victor Pankratius, Ali Jannesari, and Walter F. Tichy. Parallelizing Bzip2: a case study in multicore software engineering. IEEE Software, 26(6):70–77, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Parnas:2003:GEI

[PL03] David L. Parnas and Mark Lawford. Guest Editors’ introduction: Inspection’s role in software quality assurance. IEEE Software, 20(4):16–20, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4016.
pdf.

Putnam:1997:MHS

[PM97] Lawrence H. Putnam and Ware Myers. Manager: How solved is the cost estimation problem? IEEE Software, 14(6):105–107, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6105.
pdf.

Pietrucha:1998:NBI

[PM98] Bill Pietrucha and Craig Menefee. In the news: Bumping Internet packets to first class: NSF grant funds National Network Engineering Center: New spec has multimedia developers smiling: Surveys seek input on software standards. IEEE Software, 15(3):108–110, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3108.
pdf.

Perrochon:1999:ID

[PM99a] Louis Perrochon and Walter Mann. Inferred designs. IEEE Software, 16(5):46–51, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5046abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5046.pdf.

Putnam:1999:YWC

[PM99b] Larry Putnam and Ware Myers. Y2K work comes down to the wire. IEEE Software, 16(1):90–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1090.
pdf.

Pfleeger:2000:MTS

[PM00] Shari Lawrence Pfleeger and Winifred Menezes. Marketing technology to software practitioners. IEEE Software, 17(1):27–33, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s1027abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s1027.pdf.

Pepa:2007:BS

[PMD07] Caroline Pepa, Naseem Mariam, and Kevin C. Desouza. Book shelf. IEEE Software, 24(2):91– 93, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
02/s2091.pdf.

Poore:1993:PCS

[PMM93] J. H. Poore, Harlan D.

Mills, and David Mutchler. Planning and certifying software system reliability. IEEE Software, 10(1):88– ??, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pont:1998:SWJ

[Pon98] Michael J. Pont. Soapbox: Why Java is dangerous. IEEE Software, 15(1):20–22, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s1020.
                                        pdf.

Poston:1985:PSR

[Pos85] Robert M. Poston. Preventing software requirements specification errors with IEEE 830. IEEE Software, 2(1):83–86, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Poston:1995:TTC

[Pos95] Robert M. Poston. Testing tools combine best of new and old. IEEE Software, 12(2):122–126, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Potts:1993:SER

[Pot93]  Colin Potts.  Software-

engineering research revisited. IEEE Software, 10 (5):19–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Poulin:1995:BST

[Pou95] Jeffrey S. Poulin. Broad but shallow treatment of OO databases: review of Object-oriented Databases: Technology, Applications, and Products. IEEE Software, 12(4):112–113, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Poulin:1999:BHP

[Pou99] Jeffrey S. Poulin. Bookshelf: High Performance Oracle8 Object-Oriented Design; Object-Oriented Client/Server Internet Environments. IEEE Software, 16(3):105–106, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3105.
pdf.

PinheirodaSilva:2003:UIM

[PP03] Paulo Pinheiro da Silva and Norman W. Paton. User interface modeling in UMLi. IEEE Software, 20(4):62– 69, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4062.
pdf;   http://www.computer.
org/software/so2003/s4062abs.
htm.

Pfleeger:1994:MBP

[PR94] Shari Lawrence Pfleeger and Hans Dieter Rombach. Measurement-based process improvement. IEEE Software, 11(4):8–11, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pfleeger:2008:CEI

[PR08] Shari Lawrence Pfleeger and Rachel Rue. Cybersecurity economic issues: Clearing the path to good practice. IEEE Software, 25(1):35–42, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pratt:1985:PEP

[Pra85] Terrence W. Pratt. Pisces: an environment for parallel scientific computation. IEEE Software, 2(4): 7–20, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pressman:1995:SAN

[Pre95] Roger Pressman. Software according to Niccolo Machiavelli. IEEE Software, 12(1):101–102, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pressman:1996:MSPa

[Pre96a] Roger Pressman. Manager:

Software process impediment. IEEE Software, 13 (5):16–17, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pressman:1996:MSPb

[Pre96b] Roger S. Pressman. Manager: Software process perceptions. IEEE Software, 13(6):16, 18, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pressman:1998:MFT

[Pre98] Roger Pressman. Manager: Fear of trying: The plight of rookie project managers. IEEE Software, 15(1):50–51, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1050.
pdf.

Pressman:2000:MWT

[Pre00] Roger S. Pressman. Manager: What a tangled Web we weave. IEEE Software, 17(1):18–21, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so2000/pdf/s1018.
pdf.

Prechelt:2001:ALE

[Pre01] Lutz Prechelt. Accelerating learning from experience: Avoiding defects faster. IEEE Software, 18(6):56–61, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6056abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6056.pdf.

Plaisant:1997:LEH

[PRSV97] Catherine Plaisant, Anne Rose, Ben Shneiderman, and Ajit J. Vanniamparampil. Low-effort, highpayoff user interface reengineering. IEEE Software, 14(4):66–72, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4066.
pdf;   http://www.computer.
org/software/so1997/s4066abs.
htm.

Prell:1984:BQP

[PS84] Edward M. Prell and Alan P. Sheng. Building quality and productivity into a large software system. IEEE Software, 1(3):

47–54, July 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Porter:1990:EGS

[PS90] Adam A. Porter and Richard W. Selby. Empirically guided software development using metricbased classification trees. IEEE Software, 7(2):46– 54, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Poston:1992:EST

[PS92] Robert M. Poston and Michael P. Sexton. Evaluating and selecting testing tools. IEEE Software, 9(3): 33–42, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pooley:2002:CAW

[PSC02] Rob Pooley, Dave Senior, and Duncan Christie. Collecting and analyzing Web-based project metrics. IEEE Software, 19(1):52–58, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1052abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1052.pdf.

Parrish:2004:FSD

[PSHH04] Allen Parrish, Randy Smith, David Hale, and Joanne Hale. A field study of developer pairs: Productivity impacts and implications. IEEE Software, 21(5):76–79, September/ October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Perry:1994:POP

[PSV94] Dewayne E. Perry, Nancy A. Staudenmeyer, and Lawrence G. Votta. People, organizations, and process improvement: Two experiments to discover how developers spend their time. IEEE Software, 11(4):36–45, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pyster:2005:GEI

[PT05] Arthur B. Pyster and Richard H. Thayer. Guest Editors’ introduction: Software engineering project management 20 years later. IEEE Software, 22(5):18– 21, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2005/
                                        05/s5018.pdf.

Pothier:2009:BFO

[PT09] Guillaume Pothier and Eric Tanter. Back to the future: Omniscient debugging. IEEE Software, 26(6):78–85, November/ December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Potts:1994:IBR

[PTA94] Colin Potts, Kenji Takahashi, and Annie I. Anton. Inquiry — based requirements analysis. IEEE Software, 11(2):21–32, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pyster:2009:MDS

[PTH^{+}09] Arthur Pyster, Richard Turner, Devanandham Henry, Kahina Lasfer, and Larry Bernstein. Master’s Degrees in Software Engineering: An analysis of 28 university programs. IEEE Software, 26(5):94– 101, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Puncello:1988:AKB

[PTP^{+}88] P. Paolo Puncello, Piero Torrigiani, Francesco Pietri, Riccardo Burlon, Bruno Cardile, and Mirella Conti. ASPIS: a knowledge-based CASE environment. IEEE Software, 5(2):58–65, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Puder:2004:MOS

[Pud04] Arno Puder. MICO: An Open Source CORBA implementation. IEEE Software, 21(4):17–19, July/ August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Puerta:1997:MBI

[Pue97]  Angel R. Puerta.  A

model-based interface development environment. IEEE Software, 14(4):40– 47, July/August 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s4040.
                                        pdf;   http://www.computer.
                                        org/software/so1997/s4040abs.
                                        htm.

Putnam:1991:QTT

[Put91] Lawrence H. Putnam. Quality time — trends in measurement, estimation, and control. IEEE Software, 8(2):105–107, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Porter:1997:QTW

[PV97] Adam Porter and Lawrence Votta. Quality time:

What makes inspections work? IEEE Software, 14(6):99–102, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6099.
pdf.

Pfleeger:1995:BSP

[PWS^{+}95] Shari Lawrence Pfleeger, Walter Wilf, Martin Sheppard, Jacques L. Schlesinger, Spencer Dawkins, Henry A. Etlinger, and John Baumert. Bookshelf: Software process design: Out of the tarpit debugging the development process. advances in human-computer interaction. user interface software. developing user interfaces. integrated software reuse: Management and techniques. zen of code optimization. design paradigms: Case histories of error and judgement in engineering. user interface design. ISO 9000 for software developers. IEEE Software, 12(5):105–111, 114, September 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Putrycz:2005:PTC

[PWW05] Erik Putrycz, Murray Woodside, and Xiuping Wu. Performance techniques for COTS systems. IEEE Software, 22(4):36– 44, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Quinn:2006:LSE

[QBN^{+}06] Brendan Quinn, Leonor Barroca, Bashar Nuseibeh, Juan Fernández-Ramil, Lucia Rapanotti, Pete Thomas, and Michel Wermelinger. Learning software engineering at a distance. IEEE Software, 23(6):36–43, November/ December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Quinn:1990:DPP

[QH90] Michael J. Quinn and Philip J. Hatcher. Dataparallel programming on multicomputers. IEEE Software, 7(5):69–76, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2007:AE

[Rai07a] J. B. Rainsberger. Ask for examples. IEEE Software, 24(4):22–23, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2007:AD

[Rai07b] J. B. Rainsberger. Avoiding defects. IEEE Software, 24(2):14–15, March/April 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2007:JTI

[Rai07c] J. B. Rainsberger. Just try it. IEEE Software, 24(6):16–17, November/ December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2007:PP

[Rai07d] J. B. Rainsberger. Personal planning. IEEE Software, 24(1):16–17, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2007:RC

[Rai07e] J. B. Rainsberger. Rescuing code. IEEE Software, 24(5):26–27, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2007:SE

[Rai07f] J. B. Rainsberger. Ship effortlessly. IEEE Software, 24(3):16–17, May/ June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2008:IO

[Rai08]  J. B. Rainsberger.  In

or out? IEEE Software, 25(5):16–17, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rajlich:2000:IRU

[Raj00] Václav Rajlich. Incremental redocumentation using the Web. IEEE Software, 17(5):102–106, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so/s5102abs.htm;   http:
//dlib.computer.org/so/
books/so2000/pdf/s5102.
pdf.

Ramsey:1994:LPS

[Ram94] Norman Ramsey. Literate programming simplified. IEEE Software, 11(5): 97–105, September 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ramaswamy:2000:HSB

[Ram00] Ramkumar Ramaswamy. How to staff businesscritical maintenance projects. IEEE Software, 17 (3):90–94, May/June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so2000/s3090abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s3090.pdf.

Ramaswamy:2001:MOO

[Ram01] Ramkumar Ramaswamy.

Mentoring object-oriented projects. IEEE Software, 18(3):36–40, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s3036abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s3036.pdf.

Ramesh:2002:PKM

[Ram02] Balasubramaniam Ramesh.

Process knowledge management with traceability. IEEE Software, 19(3):50– 52, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3050.
pdf;   http://www.computer.
org/software/so2002/s3050abs.
htm.

Rasmusson:2003:IXG

[Ras03] Jonathan Rasmusson. Introducing XP into Greenfield Projects: Lessons learned. IEEE Software, 20(3):21–28, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3021.
                                        pdf;   http://www.computer.
                                        org/software/so2003/s3021abs.
                                        htm.

Runeson:2006:WDW

[RAT^{+}06] Per Runeson, Carina Andersson, Thomas Thelin, Anneliese Andrews, and Tomas Berling. What do we know about defect detection methods? IEEE Software, 23(3):82– 90, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rooijmans:1996:SQC

[RAv96] Jan Rooijmans, Hans Aerts, and Michiel van Genuchten. Software quality in consumer electronics products. IEEE Software, 13(1):55–64, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Raymond:1999:ILO

[Ray99]  Eric Raymond.  Inter-

view: Linux and opensource success. IEEE Software, 16(1):85–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/

books/so1999/pdf/s1085.
pdf.

Raybould:2006:EOD

[Ray06] Andrew Raybould. Explanationoriented design. IEEE Software, 23(5):104, 103, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Reifer:2003:ELL

[RBBC03] Donald J. Reifer, Victor R. Basili, Barry W. Boehm, and Betsy Clark. Eight lessons learned during COTS-based systems maintenance. IEEE Software, 20(5):94–96, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5094.
pdf.

RadydeAlmeida:2003:BPC

[RBBP03] Jorge Rady de Almeida, Jr., João Batista Camargo, Jr., Bruno Abrantes Basseto, and Sérgio Miranda Paz. Best practices in code inspection for safety-critical software. IEEE Software, 20(3):56– 63, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3056.
                                        pdf;   http://www.computer.
                                        org/software/so2003/s3056abs.
                                        htm.

Rech:2007:UWT

[RBH07] Jörg Rech, Christian Bogner, and Volker Haas. Using Wikis to tackle reuse in software projects. IEEE Software, 24(6):99–104, November/December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Roberts:1997:DSU

[RBIM97] Dave Roberts, Dick Berry, Scott Isensee, and John Mullaly. Developing software using OVID. IEEE Software, 14(4):51–57, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s4051.
                                        pdf;   http://www.computer.
                                        org/software/so1997/s4051abs.
                                        htm.

Raghavan:1989:DSE

[RC89] Sridhar A. Raghavan and Donald R. Chand. Diffusing software-engineering methods. IEEE Software, 6 (4):81–90, July 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reilly:1995:PCD

[RC95] John P. Reilly and Erran Carmel. Point counterpoint: Does RAD live up to the hype? IEEE Software, 12(5):25–26, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ruddock:1996:MPG

[RD96] David E. Ruddock and Balakrishnan Dasarathy. Multithreading programs: Guidelines for DCE applications. IEEE Software, 13(1):80–90, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rodriguez-Dapena:1999:SSC

[RD99] Patricia Rodríguez-Dapena.

Software safety certification: a multidomain problem. IEEE Software, 16(4):31–38, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s4031abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s4031.pdf.

Rothermel:2003:PYB

[RE03] Gregg Rothermel and Sebastian Elbaum. Putting your best tests forward. IEEE Software, 20(5):74– 77, September/October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/dl/mags/so/2003/05/

s5074.pdf. See correction [Ano03l].

Ruffin:2004:UOS

[RE04] Michel Ruffin and Christof Ebert. Using Open Source software in product development: a primer. IEEE Software, 21(1):82–86, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reeves:1991:PPC

[Ree91] Anthony Reeves. Parallel programming for computer vision. IEEE Software, 8 (6):51–??, November 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reel:1999:CSF

[Ree99] John S. Reel. Critical success factors in software projects. IEEE Software, 16(3):18–23, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s3018abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s3018.pdf.

Reed:2000:YTC

[Ree00] Karl Reed. From your technical council: Software engineering — a new millennium? IEEE Software, 17(4):107, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4107.
pdf.

Reed:2003:GEG

[Ree03] Karl Reed. Good enough is not good enough. IEEE Software, 20(5):109, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5109.
pdf.

Reed:2005:SEP

[Ree05] Karl Reed. Software engineering: The past, the future, and your TCSE. IEEE Software, 22(4):106– 107, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reed:2006:EPR

[Ree06a] Karl Reed. EQUITY and the problem of return on IT investment. IEEE Software, 23(1):114–115, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2006/
                                        01/s1114.pdf.

Reed:2006:ESE

[Ree06b] Karl Reed. Experimental software engineering: a new conference. IEEE Software, 23(4):98–99, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reiss:1987:WGE

[Rei87] Steven P. Reiss. Working in the garden environment for conceptual programming. IEEE Software, 4(6):16–27, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reiss:1990:CTU

[Rei90] Steven P. Reiss. Connecting tools using message passing in the Field environment. IEEE Software, 7 (4):57–66, July 1990. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1994:EUI

[Rei94a] Sorel Reisman. End user: Intermedia pauses for reality check. IEEE Software, 11(3):128, 127, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1994:EUM

[Rei94b] Sorel Reisman. End user:

Multimedia stars at a slightly calmer Comdex. IEEE Software, 11(1):128– 127, January 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1995:COC

[Rei95a] Sorel Reisman. Come one, come all to Comdex swap meet. IEEE Software, 12 (1):127–128, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1995:EUF

[Rei95b] Sorel Reisman. End user:

Fuzzyspeak threatens computing’s future. IEEE Software, 12(4):120, 119, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1995:EUS

[Rei95c] Sorel Reisman. End user:

I have seen the future: It is flooded with email. IEEE Software, 12(3):112, 111, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1995:EUWa

[Rei95d] Sorel Reisman. End user: We may be smart, but are we savvy? IEEE Software, 12(5):120, 119, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1995:EUWb

[Rei95e] Sorel Reisman. End user: Why CD-ROM is a deadend technology. IEEE Software, 12(6):119–120, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1995:PCF

[Rei95f]  Sorel Reisman.  Partly

cloudy forecasts for the year in computing. IEEE Software, 12(2):128, 127, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1996:IWM

[Rei96a] Sorel Reisman. Industry watch: MacWorld: Lots of light, sound, and audio — little strategy. IEEE Software, 13(2):113–114, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1996:IWO

[Rei96b]  Sorel Reisman.  Indus-

try watch: Old-timers and upstarts chime in with new telephony products. IEEE Software, 13(3):103– 104, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1996:TST

[Rei96c] Sorel Reisman. Too small, too hot, and still too cool. IEEE Software, 13(4):141– 142, July 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reifer:2000:MRM

[Rei00a] Donald J. Reifer. Manager: Requirements management: The search for nirvana. IEEE Software, 17(3):45–47, May/ June 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s3045.
pdf.

Reifer:2000:MSM

[Rei00b] Donald J. Reifer. Manager:

Software management: The good, the bad, and the ugly. IEEE Software, 17(2):73– 75, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2073.
pdf.

Reifer:2000:WDE

[Rei00c] Donald J. Reifer. Web development: Estimating quick-to-market software. IEEE Software, 17(6):57–64, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6057abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6057.pdf.

Reifer:2001:MSM

[Rei01] Donald J. Reifer. Manager: Software management’s seven deadly sins. IEEE Software, 18(2):12– 15, March/April 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s2012.
pdf;   http://dlib.computer.
org/so/books/so2001/s2012abs.
htm.

Reifer:2002:TDR

[Rei02a]  Donald Reifer.  Ten

deadly risks in Internet and intranet software development. IEEE Software, 19(2):12–14, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s2012.

pdf;   http://www.computer.
org/software/so2002/s2012abs.
htm.

Reifer:2002:HGA

[Rei02b] Donald J. Reifer. How good are agile methods? IEEE Software, 19(4):16– 18, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4016.
pdf.

Reifer:2002:LBK

[Rei02c] Donald J. Reifer. A little bit of knowledge is a dangerous thing. IEEE Software, 19(3):14–15, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3014.
pdf;   http://www.computer.
org/software/so2002/s3014abs.
htm.

Reifer:2003:SES

[Rei03a]  Donald J. Reifer.  Is

the software engineering state of the practice getting closer to the state of the art? IEEE Software, 20(6):78–83, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6078.

                                        pdf;   http://csdl.computer.
                                        org/dl/mags/so/2003/06/
                                        s6078abs.htm.

Reifer:2003:XC

[Rei03b] Donald J. Reifer. XP and the CMM. IEEE Software, 20(3):14–15, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3014.
                                        pdf.

Reifer:2004:CBR

[Rei04a] Donald J. Reifer. Catching the brass ring. IEEE Software, 21(3):12–14, May/ June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reifer:2004:SHO

[Rei04b] Donald J. Reifer. Seven hot outsourcing practices. IEEE Software, 21(1): 14–16, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Remillard:2005:SCR

[Rem05] Jason Remillard. Source code review systems. IEEE Software, 22(1):74–77, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Riehle:2009:OCW

[REM^{+}09] Dirk Riehle, John Ellenberger, Tamir Menahem, Boris Mikhailovski, Yuri Natchetoi, Barak Naveh, and Thomas Odenwald. Open collaboration within corporations using software forges. IEEE Software, 26(2):52–58, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Riedemann:2009:STM

[RF09] Catharina Riedemann and Regine Freitag. Software technology: Modeling usage: Techniques and tools. IEEE Software, 26(2):20– 24, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reddy:1986:KBS

[RFHM86] Y. V. Ramana Reddy, Mark S. Fox, Nizwer Husain, and Malcolm McRoberts. Knowledgebased simulation system. IEEE Software, 3(2):26– 37, March 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Roe:2002:SSD

[RG02] Colleen Roe and Sergio Gonik. Server-side design principles for scalable Internet systems. IEEE Software, 19(2):34–41, March/ April 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s2034.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s2034abs.
                                        htm.

Rajlich:2004:ICO

[RG04] Vaclav Rajlich and Prashant Gosavi. Incremental change in object-oriented programming. IEEE Software, 21(4):62–69, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rubin:1985:TPG

[RGR85] Robert V. Rubin, Eric J. Golin, and Steven P. Reiss. Think Pad: a graphical system for programming by demonstration. IEEE Software, 2(2):73–79, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ramanathan:1989:GIT

[RH89] Jay Ramanathan and Ronald L. Hartung. A generic iconic tool for viewing databases. IEEE Software, 6(5):50–57, September 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Robertson:2005:PC

[RH05] James Robertson and Connie Heitmeyer. Point/ counterpoint. IEEE Software, 22(1):48–51, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ramger:2003:LIW

[RHK^{+}03] David Ramger, Paul W.

Horstmann, Dan Kalcevic, Ben Linders, Yusuf Altunel, Ali H. Dogru, and Murat M. Tanik. Letters: It was Winston, not Walker, Royce; pick the approach that fits; what you don’t measure; a component-oriented process model. IEEE Software, 20(4):8–11, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4008.
pdf.

Rifkin:2001:LOW

[Rif01a] Stan Rifkin. Loyal opposition: Why software process innovations are not adopted. IEEE Software, 18(4):112, 110–111, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4112.
pdf.

Rifkin:2001:WMM

[Rif01b]  Stan Rifkin.  What

makes measuring software so hard? IEEE Software, 18(3):41–45, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s3041abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s3041.pdf.

Rifkin:2009:FSE

[Rif09]  Stan Rifkin.  FOCUS

2: Software estimation: Guest Editor’s introduction: Software measurement. IEEE Software, 26(3):70, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rinn:1997:LAL

[Rin97]  Steve Rinn.  Letters:

Another look at software “essentials”. IEEE Software, 14(4):14, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s4010.
                                        pdf.

Repenning:2001:UCR

[RIP^{+}01] Alexander Repenning, Andri Ioannidou, Michele Payton, Wenming Ye, and Jeremy Roschelle. Using components for rapid distributed software development. IEEE Software, 18(2):38–45, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2038abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2038.pdf.

Rising:2007:UPA

[Ris07] Linda Rising. Understanding the power of abstraction in patterns. IEEE Software, 24(4):46–51, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rising:2000:SSD

[RJ00] Linda Rising and Norman S. Janoff. The Scrum software development process for small teams. IEEE Software, 17(4):26– 32, July/August 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4026abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4026.pdf.

Ramasubramanian:2002:KMI

[RJ02] Shivram Ramasubramanian and Gokulakrishnan Jagadeesan. Knowledge management at Infosys. IEEE Software, 19(3):53– 55, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3053.
pdf;   http://www.computer.
org/software/so2002/s3053abs.
htm.

Rubin:2002:USI

[RJI02] Howard Rubin, Margaret Johnson, and Susan Iventosch. The US software industry. IEEE Software, 19(1):95–97, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1095abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/

s1095.pdf. See errata [ERJI02].

Rossak:1997:GMS

[RKJ^{+}97] Wilhelm Rossak, Vassilka Kirova, Leon Jololian, Harold Lawson, and Tamar Zemel. A generic model for software architectures. IEEE Software, 14(4):84– 92, July/August 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so1997/pdf/s4084.
pdf.

Ramasubbu:2005:LGR

[RKK05] Narayan Ramasubbu, M. S.

Krishnan, and Prasad Kompalli. Leveraging global resources: a process maturity framework for managing distributed development. IEEE Software, 22(3):80–86, May/ June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rao:2008:SOM

[RKP08] Uma Sudhakar Rao, Srikanth Kestur, and Chinmay Pradhan. Stochastic optimization modeling and quantitative project management. IEEE Software, 25(3):29–36, May/ June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rus:2002:GEI

[RL02] Ioana Rus and Mikael Lindvall. Guest Editors’ introduction: Knowledge management in software engineering. IEEE Software, 19(3):26–38, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3026.

pdf;   http://www.computer.
org/software/so2002/s3026abs.
htm.

Russ:2000:SDP

[RM00] Melissa L. Russ and John D. McGregor. A software development process for small projects: (available without subscription). IEEE Software, 17(5):96–101, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so/s5096abs.htm;   http:
//dlib.computer.org/so/
books/so2000/pdf/s5096.
pdf.

Raymond:2004:PC

[RM04] Eric S. Raymond and David G. Messerschmitt. Point/counterpoint. IEEE Software, 21(1):88–91, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reifer:2003:SAM

[RME03] Donald J. Reifer, Frank Maurer, and Hakan Erdogmus. Scaling agile methods. IEEE Software, 20(4):12– 14, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4012.
pdf.

Robertson:1995:QTV

[Rob95] Suzanne Robertson. Quality time: Visibility: The key to quality improvement. IEEE Software, 12 (4):95–97, July 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Roberts:1999:MWC

[Rob99] Tom Roberts. Manager:

Why can’t we implement this SDM? IEEE Software, 16(6):70–71, 75, November/December 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6070.
pdf.

Robertson:2001:RWA

[Rob01a] Suzanne Robertson. Requirements: Are we afraid of the dark? IEEE Software, 18(4):12–15, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4012.
pdf.

Robinson:2001:LOR

[Rob01b] Hugh Robinson. Loyal opposition: Reflecting on research and practice. IEEE Software, 18 (1):112, 110–111, January/ February 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s1112.
pdf.

Robertson:2002:EWA

[Rob02] James Robertson. Eureka! why analysts should invent requirements. IEEE Software, 19(4):20–22, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4020.
pdf.

Rob:2003:PFS

[Rob03] Mohammad A. Rob. Project failures in small companies. IEEE Software, 20(6):94–95, November/ December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6094.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6094abs.htm.

Robertson:2004:RBC

[Rob04] Suzanne Robertson. Requirements and the business case. IEEE Software, 21(5):93–95, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Robertson:2005:LOD

[Rob05a] Suzanne Robertson. Learning from other disciplines. IEEE Software, 22(3):54– 56, May/June 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Robillard:2005:OPS

[Rob05b] Pierre N. Robillard. Opportunistic problem solving in software engineering. IEEE Software, 22(6):60– 67, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Robillard:2009:WMA

[Rob09a] Martin P. Robillard. What makes APIs hard to learn? answers from developers. IEEE Software, 26(6):27– 34, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Robinson:2009:SQT

[Rob09b] William N. Robinson. Seeking quality through usergoal monitoring. IEEE Software, 26(5):58–65, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rogers:2000:BBP

[Rog00] Wm. Paul Rogers. Bookshelf: a best practices approach to software requirements: Software Requirements; software for use: Your guide to user interface design: Software for Use: A Practical Guide to the Models and Methods of Usage-Centered Design: a requirements guide worthy of your collection: Mastering the Requirements Process. IEEE Software, 17(2):101–103, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2101.
pdf.

Rokosz:2003:LTT

[Rok03] Vaughn T. Rokosz. Longterm testing in a shortterm world. IEEE Software, 20(3):64–67, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3064.
pdf;   http://www.computer.
org/software/so2003/s3064abs.
htm.

Rugaber:1990:RDD

[ROL90] Spencer Rugaber, Stephen B. Ornburn, and Richard J. LeBlanc, Jr. Recognizing design decisions in programs. IEEE Software, 7 (1):46–54, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rombach:1990:DMS

[Rom90] H. Dieter Rombach. Design measurement: Some lessons learned. IEEE Software, 7(2):17–25, March 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rosenblum:1985:MDA

[Ros85] David S. Rosenblum. Methodology for the design of Ada transformation tools in a Diana environment. IEEE Software, 2(2):24– 33, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ross:1990:UMQ

[Ros90] Niall Ross. Using metrics in quality management. IEEE Software, 7(4):80–81, 85, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rosenblum:1991:SCS

[Ros91] David S. Rosenblum. Specifying concurrent systems with TSL. IEEE Software, 8(3):52–61, May 1991. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rosenberg:1998:QTF

[Ros98a] Jarrett Rosenberg. Quality time: Five easy steps to systematic data handling. IEEE Software, 15(1):75–77, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1075.
pdf.

Rossmeyer:1998:BIL

[Ros98b] Brian Rossmeyer. Bookshelf: Intermediate level Linux guide. IEEE Software, 15(4):90, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4087.
pdf.

Ross:2000:MDF

[Ros00] Nancy Ross. Manager: The decline and fall of high-tech corporate culture. IEEE Software, 17(6):12–15, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6012.
pdf.

Rost:2004:PRF

[Ros04] Johann Rost. Political reasons for failed software projects. IEEE Software, 21(6):104, 102–103, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rost:2005:SET

[Ros05] Johann Rost. Software engineering theory in practice. IEEE Software, 22 (2):96, 94–95, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rost:2006:BPR

[Ros06] Johann Rost. Are “best practices” requirements documents a myth? IEEE Software, 23(3):104, 103, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://bell.computer.
org/dlcomments/.

Rothman:1999:CWR

[Rot99] Johanna Rothman. Culture at work: Retrain your code czar. IEEE Software, 16(2):86–88, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2086.
pdf.

Rothman:2000:CWM

[Rot00] Johanna Rothman. Culture at work: Making telecommuting work. IEEE Software, 17(5):124–126, September/October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s5124.pdf.

Rovner:1986:EMB

[Rov86]  Paul Rovner.  Extend-

ing Modula-2 to build large, integrated systems. IEEE Software, 3(6):46–57, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rowe:1996:NPA

[Row96]  Alison Rowe.  In the

news: Professional awareness in software engineering. IEEE Software, 13(3): 102, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Royce:2000:MSM

[Roy00] Walker Royce. Manager: Software management renaissance. IEEE Software, 17(4):116–118, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s4116.
pdf.

Royce:2005:SSM

[Roy05] Walker Royce. Successful software management style: Steering and balance. IEEE Software, 22(5):40–47, September/ October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rafii:1995:ISC

[RP95] Farshad Rafii and Sam Perkins. Internationalizing software with concurrent engineering. IEEE Software, 12(5):39–46, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rapin:1998:LTY

[RP98] Charles Rapin and Michael Pont. Letters: Teach your students well. IEEE Software, 15(3):10–12, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3010.
pdf.

Radjenovic:2006:ADL

[RP06] Alek Radjenovic and Richard Paige. Architecture description languages for high-integrity real-time systems. IEEE Software, 23(2):71–79, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ruiz:2007:TDG

[RP07] Alex Ruiz and Yvonne Wang Price. Test-driven GUI development with TestNG and Abbot. IEEE Software, 24(3):51–57, May/ June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//csdl2.computer.org/
comp/mags/so/2007/03/
s3051_Web_Extra_2007.
pdf.

Robinson:2009:RSS

[RP09] William N. Robinson and Sandeep Purao. Requirements specification: Specifying and monitoring interactions and commitments in open business processes. IEEE Software, 26(2):72– 79, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rainsberger:2008:YDY

[RR08] J. B. Rainsberger and Johanna Rothman. Are you done yet? IEEE Software, 25(1):52–53, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rech:2007:IAG

[RRD07] Jörg Rech, Eric Ras, and Björn Decker. Intelligent assistance in German software development: a survey. IEEE Software, 24(4): 72–79, July/August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ramamritham:1984:DTS

[RS84] Krithivasan Ramamritham and John A. Stankovic. Dynamic task scheduling in hard real-time distributed systems. IEEE Software, 1 (3):65–75, July 1984. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Raynal:2001:MAP

[RS01] Michel Raynal and Mukesh Singhal. Mastering agreement problems in distributed systems. IEEE Software, 18(4):40–47, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s4040abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s4040.pdf.

Rugaber:2004:MDR

[RS04] Spencer Rugaber and Kurt Stirewalt. Model-driven reverse engineering. IEEE Software, 21(4):45–53, July/ August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ruhe:2005:ASS

[RS05] Günther Ruhe and Moshood Omolade Saliu. The art and science of software release planning. IEEE Software, 22(6):47– 53, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ratanotayanon:2008:ITU

[RS08] Sukanya Ratanotayanon and Susan Elliott Sim. Inventive tool use to comprehend big code. IEEE Software, 25(5):91–92, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reisman:1990:MIT

[RSE^{+}90] Sorel Reisman, Ian Sommerville, Susan Englert, Jennifer Arme, Guy Latge, Peter Voletsky, Edward Miller, Thomas Radi, and Ellis Horowitz. Management and integrated tools. IEEE Software, 7(3):71– 77, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Regnell:2008:SRQ

[RSO08] Björn Regnell, Richard Berntsson Svensson, and Thomas Olsson. Supporting roadmapping of quality requirements. IEEE Software, 25(2):42–47, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rugg:1993:UCE

[Rug93] David Rugg. Using a capability evaluation to select a contractor. IEEE Software, 10(4):36–??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Runeson:2006:SUT

[Run06] Per Runeson. A survey of unit testing practices. IEEE Software, 23(4):22– 29, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rupp:2002:RP

[Rup02] Chris Rupp. Requirements and psychology. IEEE Software, 19(3):16–18, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3016.

pdf;   http://www.computer.
org/software/so2002/s3016abs.
htm.

Russell:1991:EIU

[Rus91] Glen W. Russell. Experience with inspection in ultralarge-scale developments. IEEE Software, 8(1):25–31, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Richardson:2007:GEI

[RvW07] Ita Richardson and Christiane Gresse von Wangenheim. Guest Editors’ introduction: Why are small software organizations different? IEEE Software, 24(1):18–22, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
01/s1018.pdf.

Rich:1990:RPD

[RW90] Charles Rich and Linda M. Wills. Recognizing a program’s design: a graphparsing approach. IEEE Software, 7(1):82–89, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rudolf:1992:CJI

[RW92] Jim Rudolf and Cathy Waite. Completing the job of interface design. IEEE Software, 9(6):11–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rugaber:1998:RLL

[RW98] Spencer Rugaber and Jim White. Restoring a legacy: Lessons learned. IEEE Software, 15(4):28– 33, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
cg1998/s4028abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s4028.pdf.

Robinson:2004:FRU

[RW04] William N. Robinson and Han G. Woo. Finding reusable UML sequence diagrams automatically. IEEE Software, 21(5):60–67, September/ October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ranka:1988:PHM

[RWS88] Sanjay Ranka, Youngju Won, and Sartaj Sahni. Programming a hypercube multicomputer. IEEE Software, 5(5):69–77, September 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ryskin:1995:BFB

[Rys95] Alexander Ryskin. Bookshelf: Fortran becomes a language for the nineties: review of Migrating to Fortran 90. IEEE Software, 12(4):110–111, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sharon:1997:CSE

[SA97] D. Sharon and T. Anderson. A complete software engineering environment. IEEE Software, 14 (2):123–125, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schatz:2005:PGA

[SA05] Bob Schatz and Ibrahim Abdelshafi. Primavera gets agile: a successful transition to agile development. IEEE Software, 22(3):36– 42, May/June 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Serrano:2007:AFI

[SA07] Nicolás Serrano and Juan Pablo Aroztegi. Ajax frameworks in interactive Web apps. IEEE Software, 24(5):12–14, September/ October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stytz:1997:RPD

[SAG^{+}97] Martin R. Stytz, Terry Adams, Brian Garcia, Steven M. Sheasby, and Brian Zurita. Rapid prototyping for distributed virtual environments. IEEE Software, 14(5):83–92, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5083.
pdf;   http://www.computer.
org/software/so1997/s5083abs.
htm.

Salomon:1997:BOO

[Sal97a] Daniel J. Salomon. Bookshelf: Object-oriented compiler construction. IEEE Software, 14(5):117–118, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s5115.pdf.

Salomon:1997:BPO

[Sal97b] Daniel J. Salomon. Bookshelf: Pitfalls of objectoriented development. IEEE Software, 14(5):115–116, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1997/pdf/
                                        s5115.pdf.

Salit:2003:RCA

[Sal03] Richard Salit. Requirements are corporate assets. IEEE Software, 20(3):86– 88, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3086.
                                        pdf.

Samuelson:1988:CLS

[Sam88] Pamela Samuelson. Is copyright law steering the right course? IEEE Software, 5 (5):78–86, September 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Samuelson:1990:RES

[Sam90] Pamela Samuelson. Reverseengineering someone else’s software: Is it legal? IEEE Software, 7(1):90–96, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sammet:2000:RCC

[Sam00] Jean E. Sammet. The real creators of Cobol. IEEE Software, 17(2):30– 32, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2030abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2030.pdf.

Sanders:1995:NMR

[San95a]  James L. Sanders.  In

the news: Microsoft releases a work in progress. US Congress slashes software R and D funds. interview: James Sims, Cambridge Technology Partners. IEEE Software, 12 (5):90–91, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Santalesa:1995:NTR

[San95b] Rich Santalesa. In the news: Taligent readies a new development paradigm. IEEE Software, 12(2):103–109, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sanden:1997:MCS

[San97a] Bo I. Sanden. Modeling concurrent software. IEEE Software, 14(5):93–100, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1997/pdf/
s5093.pdf;   http://www.

                                        computer.org/software/
                                        so1997/s5093abs.htm.

Sandoe:1997:SSE

[San97b] Kent Sandoe. Soapbox: Split ends: Labor shortage and the CS-IS divide. IEEE Software, 14(6):19–21, November/ December 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s6019.
                                        pdf.

Sanders:1998:ELO

[San98a]  James Sanders.  Edge-

watch: Linux, open source, and software’s future. IEEE Software, 15(5):88– 91, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s5088.pdf.

Sanders:1998:EOO

[San98b]  James Sanders.  Edge-

watch: Opportunities online: Fact or fantasy? IEEE Software, 15(2):62– 64, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s2062.
                                        pdf.

Sanders:1998:EYD

[San98c]  James Sanders.  Edge-

watch: Y2K: Don’t play it again, Sam. IEEE Software, 15(3):100–102, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3100.
pdf.

Sanders:1998:PP

[San98d] James Sanders. Products for practitioners. IEEE Software, 15(3):111–112, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3111.
pdf.

Sanders:1998:PPC

[San98e] James Sanders. Products for practitioners: Caché; SQml Web 1.0; Temporal Rover; Solutions: Schedule; Installshield Java Edition. IEEE Software, 15(1):103, 102, January/ February 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1103.
pdf.

Sanders:1998:PPW

[San98f] James Sanders. Products for practitioners: Windows DLLs; Windows memory manager; middleware support; remote server; Java decompiler; Java connectivity; Java quick reference; terminal emulator. IEEE Software, 15(4):83– 86, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4083.
pdf.

Sanders:1998:SS

[San98g] James Sanders. Software synergies. IEEE Software, 15(1):13, January/ February 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s1013.
pdf.

Sanden:2003:ELM

[San03] Bo I. Sandén. Entitylife modeling: Modeling a thread architecture on the problem environment. IEEE Software, 20(4):70– 78, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4070.
pdf;   http://www.computer.

org/software/so2003/s4070abs.
htm.

Santos:2008:VUP

[San08] Carlos Santos, Jr. Viewpoint: Understanding partnerships between corporations and the open source community: a research gap. IEEE Software, 25(6):96– 97, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Savor:2008:TFR

[Sav08]  Tony Savor.  Testing

feature-rich reactive systems. IEEE Software, 25(4):74–81, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Srivas:1990:FVP

[SB90] Mandayam Srivas and Mark Bickford. Formal verification of a pipelined microprocessor. IEEE Software, 7(5):52–64, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sharon:1995:TBC

[SB95a] David Sharon and Rodney Bell. Tools that bind: Creating integrated environments. IEEE Software, 12(2):76–85, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Strigini:1995:LFS

[SB95b] Lorenzo Strigini and Antonia Bertolino. Letters: The flip side of designing for testability. IEEE Software, 12(5):12–13, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stevens:1996:LLD

[SB96] Robert J. Stevens and Stephen Barlas. Letters: Loral and DSR: Congress was (and is) watching. IEEE Software, 13(3):20, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schaffert:2008:SW

[SBBK08] Sebastian Schaffert, François Bry, Joachim Baumeister, and Malte Kiesel. Semantic Wikis. IEEE Software, 25(4):8–11, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stefik:1986:IAO

[SBK86] Mark J. Stefik, Daniel G. Bobrow, and Kenneth M. Kahn. Integrating accessoriented programming into a multiparadigm environment. IEEE Software, 3 (1):10–18, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Saiedian:2002:SEP

[SBM02] Hossein Saiedian, Donald J. Bagert, and Nancy R. Mead. Software engineering programs: Dispelling the myths and misconceptions. IEEE Software, 19(5):35–41, September/October 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5035.
pdf;   http://www.computer.
org/software/so2002/s5035abs.
htm.

Shetterly:1996:CPO

[SC96] Rick Shetterly and Alan Chmura. Creating the paperless office. IEEE Software, 13(2):124–126, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2000:BPP

[SC00] Diomidis Spinellis and Gerry Coleman. Bookshelf: The Pragmatic Programmer: From Journeyman to Master; Introduction to the Team Software Process. IEEE Software, 17(6):108–110, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s6108.
pdf.

Stark:2003:TSP

[SC03] John A. Stark and Ron Crocker. Trends in software process: The PSP and agile methods. IEEE Software, 20(3):89–91, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3089.
                                        pdf.

Serrano:2004:AAP

[SC04] Nicolas Serrano and Ismael Ciordia. Ant: Automating the process of building applications. IEEE Software, 21(6):89–91, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Serrano:2005:BIO

[SC05] Nicolas Serrano and Ismael Ciordia. Bugzilla, ITracker, and other bug trackers. IEEE Software, 22(2):11–13, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shaw:2006:GAS

[SC06] Mary Shaw and Paul Clements. The golden age of software architecture. IEEE Software, 23(2): 31–39, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Scanlan:1989:SFO

[Sca89] David A. Scanlan. Structured flowcharts outperform pseudocode: An experimental comparison. IEEE Software, 6(5):28–36, September 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Scacchi:2004:FOS

[Sca04] Walt Scacchi. Free and Open Source development practices in the game community. IEEE Software, 21(1):59–66, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Serrano:2004:AMM

[SCCA04] Nicolas Serrano, Ismael Ciordia, Sonia Calzada, and Fernando Alonso. Automated management of multicustomer code bases. IEEE Software, 21(6):26– 31, November/December 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schneeman:1992:PMA

[Sch92] Richard D. Schneeman.

Porting multimedia applications to the open system environment. IEEE Software, 9(6):39–??, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schamp:1995:TCT

[Sch95a] Alan Schamp. Toolbox: CM-tool evaluation and selection. IEEE Software, 12(4):114–118, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schlesinger:1995:OLS

[Sch95b] Jacques Schlesinger. Offbeat look at software engineering: review of Software Engineering Principles and Practice. IEEE Software, 12(3):102, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schmid:1996:CAC

[Sch96a] Hans Albrecht Schmid. Creating applications from components: a manufacturing framework design. IEEE Software, 13(6):67– 75, November 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schneidewind:1996:PDS

[Sch96b] Norman Schneidewind. Point: Do standards improve product quality? IEEE Software, 13(1):22, 24, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schneidewind:1998:HEL

[Sch98a] Norman F. Schneidewind.

How to evaluate legacy system maintenance. IEEE Software, 15(4):34–42, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
cg1998/s4034abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s4034.pdf.

Schultz:1998:MYP

[Sch98b] James E. Schultz. Managing a Y2K projectstarting now. IEEE Software, 15(3):63–71, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3063abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3063.pdf.

Schreiner:1999:NMS

[Sch99] Keri Schreiner. In the news: Malaysia’s Silicon Valley: European software markets: Software engineering licensing: Mozilla.org. IEEE Software, 16(5):126–130, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s5126.pdf.

Schricker:2000:CNM

[Sch00]  Don Schricker.  Cobol

for the next millennium. IEEE Software, 17(2):48– 52, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s2048abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s2048.pdf.

Schuh:2001:RRE

[Sch01] Peter Schuh. Recovery, redemption, and Extreme Programming. IEEE Software, 18(6):34–41, November/December 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s6034abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s6034.pdf.

Schrage:2004:NGC

[Sch04] Michael Schrage. Never go to a client meeting without a prototype. IEEE Software, 21(2):42–45, March/ April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schneidewind:2007:QAS

[Sch07] Norman Schneidewind. A quantitative approach to software development using IEEE 982.1. IEEE Software, 24(1):65–72, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Serrano:2004:POS

[SCSC04] Nicolas Serrano, Sonia Calzada, Jose Mari Sarriegui, and Ismael Ciordia. From proprietary to Open Source tools in information systems development. IEEE Software, 21(1):56–58, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sherman:1988:TSE

[SD88] Mark Sherman and Robert L.

Drysdale. Teaching software engineering in a workstation environment. IEEE Software, 5(3):68–76, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:B

[SDE^{+}05] Diomidis Spinellis, John R.

Dance, David Arthur Eatough, Kevin C. Desouza, and Yukika Awazu. Bookshelf.

IEEE Software, 22(1): 94–97, January/February 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
01/s1094.pdf.

Segura-Devillechaise:2006:DAS

[SDML^{+}06] Marc Ségura-Devillechaise, Jean-Marc Menaud, Nicolas Loriant, Rémi Douence, Mario Südholt, Thomas Fritz, and Egon Wuchner. Dynamic adaptation of the Squid Web cache with Arachne. IEEE Software, 23(1):34–41, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shrivastava:1991:OAD

[SDP91] Santosh K. Shrivastava, Graeme N. Dixon, and Graham D. Parrington. An overview of the Arjuna distributed programming system. IEEE Software, 8 (1):66–73, January 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schneidewind:1998:GEI

[SE98] Norman F. Schneidewind and Christof Ebert. Guest Editors’ introduction: Preserve or redesign legacy systems? IEEE Software, 15(4):14–17, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
cg1998/s4014abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s4014.pdf.

Seffah:1999:TDC

[Sef99] Ahmed Seffah. Training developers in critical skills. IEEE Software, 16(3):66– 70, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s3066abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s3066.pdf.

Seidewitz:2003:WMM

[Sei03] Ed Seidewitz. What models mean. IEEE Software, 20(5):26–32, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5026.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5026abs.htm.

Selby:1990:EIF

[Sel90] Richard W. Selby. Extensible integration frameworks for measurement. IEEE Software, 7(6):83–84, 112, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Selic:2003:PMD

[Sel03] Bran Selic. The pragmatics of model-driven development. IEEE Software, 20(5):19–25, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/05/s5019.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/so/2003/05/
                                        s5019abs.htm.

Selby:2009:ADD

[Sel09a] Richard W. Selby. Analyticsdriven dashboards enable leading indicators for requirements and designs of large-scale systems. IEEE Software, 26(1):41–49, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Selic:2009:ADA

[Sel09b] Bran Selic. Agile documentation, anyone? IEEE Software, 26(6):11–12, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Seviora:1987:KBP

[Sev87] Rudolph E. Seviora. Knowledgebased program debugging systems. IEEE Software, 4 (3):20–32, May 1987. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Segal:1993:FPM

[SF93] Mark E. Segal and Ophir Frieder. On-the-fly program modification: Systems for dynamic updating. IEEE Software, 10(2):53– 65, March 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sorel:1994:DDA

[SFG94] Pierre Sorel, Mariano G.

Fernandez, and Sumit Ghosh. A dynamic debugger for asynchronous distributed algorithms. IEEE Software, 11(1):69–76, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sierra:2008:DAU

[SFVFM08] José Luis Sierra, Alfredo Fernández-Valmayor, and Baltasar Fernández-Manjón. From documents to applications using markup languages. IEEE Software, 25(2):68– 76, March/April 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Seyff:2006:TYM

[SGM06] Norbert Seyff, Paul Grünbacher, and Neil Maiden. Take your mobile device out from behind the requirements desk. IEEE Software, 23(4):16– 18, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Slimick:2002:LEP

[SGW02] John Slimick, Patrick Georges, and Laurie Williams. Letters: Extremely pleased; one more deadly sin; try it, you’ll like it. IEEE Software, 19(1):6–7, January/ February 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2002/s1006abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2002/pdf/
                                        s1006.pdf.

Stoyenko:1993:EPI

[SH93] Alexander D. Stoyenko and Wolfgang A. Halang. Extending Pearl for industrial real-time applications. IEEE Software, 10(4):65– ??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shapiro:2002:EPD

[SH02] Jonathan S. Shapiro and Norm Hardy. EROS: a principle-driven operating system from the ground up. IEEE Software, 19(1):26–33, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1026abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1026.pdf.

Shaw:1984:ATM

[Sha84] Mary Shaw. Abstraction techniques in modern programming languages. IEEE Software, 1(4):10–24, 26, October 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shaw:1990:PED

[Sha90] Mary Shaw. Prospects for an engineering discipline of software. IEEE Software, 7 (6):15–24, November 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sharon:1995:TB

[Sha95a] David Sharon. Tool box.

IEEE Software, 12(1):124– 126, January 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shaw:1995:CAD

[Sha95b] Mary Shaw. Comparing architectural design styles. IEEE Software, 12(6):27– 41, November 1995. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sharon:1996:TMC

[Sha96]  David Sharon.  Tool-

box: Meeting the challenge of software maintenance. IEEE Software, 13 (1):122–126, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sharon:1997:TTD

[Sha97a] David Sharon. Toolbox: Top drawer. IEEE Software, 14(4):111, 113, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4107.
pdf.

Sharon:1997:TYT

[Sha97b] David Sharon. Toolbox: Year 2000 tool classification scheme. IEEE Software, 14(4):107–111, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4107.
pdf.

Sha:2001:USC

[Sha01] Lui Sha. Using simplicity to control complexity. IEEE Software, 18(4):20– 28, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s4020abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s4020.pdf.

Shannon:2007:C

[Sha07] Laurianne McLaughlin Meg Shannon. Currents. IEEE Software, 24(3):113–117, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
03/s3113.pdf.

Shaw:2009:CPE

[Sha09] Mary Shaw. Continuing prospects for an engineering discipline of software. IEEE Software, 26(6):64– 67, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sherer:1991:CEA

[She91] Susan A. Sherer. A costeffective approach to testing. IEEE Software, 8(2): 34–40, March 1991. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Surendran:2002:SSE

[SHM02] Ken Surendran, Helen Hays, and Andrew Macfarlane. Simulating a software engineering apprenticeship. IEEE Software, 19(5):49–56, September/ October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s5049.
                                        pdf;   http://www.computer.
                                        org/software/so2002/s5049abs.
                                        htm.

Shneiderman:1991:ITS

[Shn91] Ben Shneiderman. Interface — touch screens now offer compelling uses. IEEE Software, 8(2):93–94, 107, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shneiderman:1994:DQV

[Shn94] Ben Shneiderman. Dynamic queries for visual information seeking. IEEE Software, 11(6):70– 77, November 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shneiderman:1997:FSI

[Shn97] B. Shneiderman. A framework for search interfaces.

IEEE Software, 14(2):18– 20, March/April 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shore:2004:CD

[Sho04a] Jim Shore. Continuous design. IEEE Software, 21(1): 20–22, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shore:2004:FF

[Sho04b]  Jim Shore.  Fail fast.

IEEE Software, 21(5):21– 25, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shopyrin:2006:MCU

[Sho06] Danil Shopyrin. Multimethods in C++ using recursive deferred dispatching. IEEE Software, 23 (3):62–73, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2007:WNE

[Shu07a]  Forrest Shull.  Who

needs evidence, anyway? IEEE Software, 24(5):10– 11, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shum:2007:TNL

[Shu07b] Simon Buckingham Shum. There’s nothing like a good argument …. IEEE Software, 24(5):21–23, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schwartz:1998:LRH

[SHZ^{+}98] Barry Schwartz, Les Hatton, Nicholas Zvegintzov, Lee Aydelotte, and Danijel Rebolj. Letters: Rethinking how we construct language; shedding light on error correction; let’s review reuse; IT without borders? IEEE Software, 15(4):10– 13, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4010.
pdf.

Shimomura:1991:LLV

[SI91] Takao Shimomura and Sadahiro Isoda. Linkedlist visualization for debugging. IEEE Software, 8(3): 44–51, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Siddiqi:1994:CUT

[Sid94] Jawed Siddiqi. Challenging universal truths of requirements engineering.

IEEE Software, 11(2):18– 19, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Siegel:1992:WWN

[Sie92] Stan Siegel. Why we need checks and balances to assure quality. IEEE Software, 9(1):102–103, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sims:1995:NCA

[Sim95a] David Sims. In the news:

China agrees to fight copyright violations. IEEE Software, 12(3):88–89, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sims:1995:NIP

[Sim95b]  David Sims.  In the

news: Internet providers face life without NSFNet. IEEE Software, 12(4):100– 101, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sims:1995:PES

[Sim95c] David Sims. Proceedings:

Exploring solutions to the people problem. IEEE Software, 12(3):94–95, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Simmons:1996:QOD

[Sim96] Pamela Simmons. Quality outcomes: determining business value. IEEE Software, 13(1):25–32, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Simmons:2004:RTW

[Sim04] Erik Simmons. Requirements triage: What can we learn from a “medical” approach? IEEE Software, 21(4):86–88, July/ August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Simmons:2006:UMD

[Sim06] Erik Simmons. The usage model: Describing product usage during design and development. IEEE Software, 23(3):34–41, May/ June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schwan:1986:SRA

[SJ86] Karsten Schwan and Anita K. Jones. Specifying resource allocation for the Cm* multiprocessor. IEEE Software, 3(3):60–70, May 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stachura:2002:BUI

[SJM02] Victor Stachura, Philipp K.

Janert, and Shantha Mohan. Bookshelf: Useful ideas to improve requirements practices; software acquisition: a high-level view; software management for executives. IEEE Software, 19(5):108–110, September/October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2002/pdf/

s5108.pdf. Review of Effective Requirements Practices by Ralph R. Young, Managing Software Acquisition: Open Systems and COTS Products by B. Craig Meyers and Patricia Oberndorf, and Winning with Software: An Executive Strategy by Watts S. Humphrey.

Schneidewind:1992:ARM

[SK92] Norman F. Schneidewind and Ted W. Keller. Applying reliability models to the space shuttle. IEEE Software, 9(4):28–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sotirovski:1995:IDI

[SK95] Drasko M. Sotirovski and Philippe B. Kruchten. Implementing dialogue independence. IEEE Software, 12(6):61–70, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sendall:2003:MTH

[SK03] Shane Sendall and Wojtek Kozaczynski. Model transformation: The heart and soul of model-driven software development. IEEE Software, 20(5):42–45, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /csdl.computer.org/dl/
                                        mags/so/2003/05/s5042.
                                        pdf;   http://csdl.computer.
                                        org/dl/mags/so/2003/05/
                                        s5042abs.htm.

Sanders:2008:DRS

[SK08] Rebecca Sanders and Diane Kelly. Dealing with risk in scientific software development. IEEE Software, 25(4):21–28, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Setliff:1993:PSS

[SKC93] Dorothy Setliff, Elaine Kant, and Tom Cain. Practical software synthesis. IEEE Software, 10(3):1– 18, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sheldon:1992:RMT

[SKT^{+}92] Frederick T. Sheldon, Krishna M. Kavi, Robert C. Tausworthe, James T. Yu, Ralph Brettschneider, and William W. Everett. Reliability measurement: From theory to practice. IEEE Software, 9(4):13–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Skulikaris:1999:SES

[Sku99] Yannis Skulikaris. Softlaw: European software patents. IEEE Software, 16(6):109–111, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s6109.
pdf.

Sears:1997:CEU

[SL97] Andrew Sears and Arnold M.

Lund. Creating effective user interfaces. IEEE Software, 14(4):21–24, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4021.
pdf;   http://www.computer.
org/software/so1997/s4021abs.
htm.

Shere:2000:LPQ

[SLG^{+}00] David D. Shere, Ambrose Li, Robert Glass, Francisco Torres, Dick Botting, Natalia Juristo, Dick B. Simmons, and Jim Highsmith. Letters: The pursuit of quality; scratching an itch; maintaining morale; natural language dispute; Brooks’ Law review; a good blend. IEEE Software, 17(5):10–13, September/October 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5010.
pdf.

Schoenrank:2001:LVW

[SLH01] Rainder Schoenrank, Søren Lauesen, and Morten Borup Harning. Letters: Virtual windows debate. IEEE Software, 18(5):8–9, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5008abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5008.pdf.

Sockut:1988:FSF

[SM88] Gary H. Sockut and Ashok Malhotra. A fullscreen facility for defining relational and entityrelationship schemas. IEEE Software, 5(6):68–78, November 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Srimani:1992:SPR

[SM92] Pradip K. Srimani and Yashwant K. Malaiya. Steps to practical reliability measurement. IEEE Software, 9(4):10–??, July 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shlaer:1995:LWI

[SM95] Sally Shlaer and Stephen J.

Mellor. Letters: Was it or wasn’t it a valid comparison? IEEE Software, 12 (5):8, 10, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shlaer:1997:RDA

[SM97] Sally Shlaer and Stephen J.

Mellor. Recursive design of an application-independent architecture. IEEE Software, 14(1):61–72, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shimeall:1999:SSI

[SM99] Timothy J. Shimeall and John J. McDermott. Software security in an Internet world: An executive summary. IEEE Software, 16(4):58–61, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s4058abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s4058.pdf.

Symons:2001:PCS

[SM01] Charles Symons and John McGarry. Point/counterpoint: Software benchmarking: Serious management tool or a joke? vs. when it comes to measuring software, every project is unique. IEEE Software, 18(5):18–21, September/ October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s5018abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s5018.pdf.

Seetharaman:2006:TOU

[SM06a] Swaminathan Seetharaman and Krishna Murthy. Test optimization using software virtualization. IEEE Software, 23(5):66–69, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stribrny:2006:WPO

[SM06b] Scott Stribrny and Fran Boehme Mackin. When politics overshadow software quality. IEEE Software, 23(5): 72–73, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sandberg:2008:PMM

[SM08a] Anna Børjesson Sandberg and Lars Mathiassen. Project management: Managing slowdown in improvement projects. IEEE Software, 25(6):84– 89, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Segal:2008:DSS

[SM08b] Judith Segal and Chris Morris. Developing scientific software. IEEE Software, 25(4):18–20, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2008/04/mso2008040018.
pdf.

Sawyer:2009:RHU

[SM09a] Peter Sawyer and Neil Maiden. Requirements: How to use Web services in your requirements process. IEEE Software, 26(1): 76–78, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Segal:2009:DSS

[SM09b] Judith Segal and Chris Morris. Developing scientific software: Guest Editors’ introduction: Developing scientific software, Part 2. IEEE Software, 26 (1):79, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Smedinghoff:1992:SCE

[Sme92] Thomas J. Smedinghoff. Is source-code escrow a viable compromise? IEEE Software, 9(1):83–85, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Smith:2009:MAA

[SMHT09] Michael Smith, James Miller, Lily Huang, and Albert Tran. A more agile approach to embedded system development. IEEE Software, 26(3):50– 57, May/June 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Smith:1991:MIS

[Smi91]  Connie Smith.  Man-

ager — improving service while controlling costs. IEEE Software, 8(2):95– 96, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Suitor:2000:LTR

[SMK^{+}00] Terry Suitor, Steve Mc-Connell, Ulf Kurella, Steve McConnell, Steve Mattingley, David Lorge Parnas, Dick Botting, and Hermann Kaindl. Letters: Time to refocus? McConnell’s law; but MIS is not engineering; designing the real world. IEEE Software, 17(2):6–10, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2006.
pdf.

Spinellis:2007:B

[SMP07] Diomidis Spinellis, Naseem Mariam, and Caroline Pepa. Bookshelf. IEEE Software, 24(4):107–109, July/August 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2007/
04/s4107.pdf.

Shadbolt:1993:CKB

[SMR93] Nigel Shadbolt, Enrico Motta, and Alain Rouge. Constructing knowledgebased systems. IEEE Software, 10(6):34–??, November 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sprinkle:2009:FDS

[SMTS09] Jonathan Sprinkle, Marjan Mernik, Juha-Pekka Tolvanen, and Diomidis Spinellis. FOCUS: Domain-specific modeling: Guest Editors’ introduction: What kinds of nails need a domain-specific hammer? IEEE Software, 26(4):15– 18, July/August 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Squillante:1989:IHL

[SN89] Mark S. Squillante and David Notkin. Integrating heterogeneous local mail systems. IEEE Software, 6 (6):59–67, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sommerlad:2007:PC

[SN07] Peter Sommerlad and James Noble. Point/ counterpoint. IEEE Software, 24(4):68–71, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sneed:1984:SRC

[Sne84] Harry M. Sneed. Software renewal: a case study.

IEEE Software, 1(3):56– 63, July 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sneed:1995:PRL

[Sne95] Harry M. Sneed. Planning the reengineering of legacy systems. IEEE Software, 12(1):24–34, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Snyder:1993:EOC

[Sny93] Alan Snyder. The essence of objects: Concepts and terms. IEEE Software, 10 (1):31–42, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Smith:1990:STC

[SO90] Dennis B. Smith and Paul W. Oman. Software tools in context. IEEE Software, 7(3):14–17, 19, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Song:1992:TOS

[SO92] Xiping Song and Leon J.

Osterweil. Toward objective, systematic designmethod comparisons. IEEE Software, 9(3):43–53, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schamp:1997:TSI

[SO97] Alan Schamp and Heidi Owens. Toolbox: Successfully implementing configuration management. IEEE Software, 14(1):98–101, January/February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sommerville:2005:IRE

[Som05] Ian Sommerville. Integrated requirements engineering: a tutorial. IEEE Software, 22(1):16–23, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Song:1997:SID

[Son97] Xiping Song. Systematic integration of design methods. IEEE Software, 14 (2):107–117, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sotirovski:2001:HIS

[Sot01] Drasko Sotirovski. Heuristics for iterative software development. IEEE Software, 18(3):66–73, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s3066abs.htm;
http://dlib.computer.

org/so/books/so2001/pdf/
s3066.pdf.

Speed:1999:WDY

[Spe99] John R. Speed. What do you mean I can’t call myself a software engineer? IEEE Software, 16(6):45–50, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s6045abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s6045.pdf.

Spivey:1990:SRT

[Spi90] J. Michael Spivey. Specifying a real-time kernel. IEEE Software, 7(5):21–28, September 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:1998:LOC

[Spi98] Diomidis Spinellis. Loyal opposition: The computer’s new clothes. IEEE Software, 15(6):14–??, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s6014.
pdf.

Spinellis:2003:DSM

[Spi03] Diomidis Spinellis. On the declarative specification of models. IEEE Software, 20(2):96, 94–95, March/April 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s2096.
                                        pdf.

Spinellis:2005:DE

[Spi05a] Diomidis Spinellis. Dear editor. IEEE Software, 22(2): 14–15, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:JMS

[Spi05b] Diomidis Spinellis. Java makes scripting languages irrelevant? IEEE Software, 22(3):70–71, May/ June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:TWF

[Spi05c] Diomidis Spinellis. Tool writing: a forgotten art? IEEE Software, 22(4):9–11, July/August 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:TH

[Spi05d] Diomidis Spinellis. The tools at hand. IEEE Software, 22(1):10–12, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:VCP

[Spi05e] Diomidis Spinellis. Version control, Part I. IEEE Software, 22(5):107, September/October 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
05/s5107.pdf.

Spinellis:2005:VCS

[Spi05f] Diomidis Spinellis. Version control systems. IEEE Software, 22(5):108–109, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:WUT

[Spi05g] Diomidis Spinellis. Working with Unix tools. IEEE Software, 22(6):9–11, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:BB

[Spi06a] Diomidis Spinellis. Bug busters. IEEE Software, 23(2):92–93, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:CPL

[Spi06b] Diomidis Spinellis. Choosing a programming language. IEEE Software, 23(4):62–63, July/August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:DLF

[Spi06c] Diomidis Spinellis. Debuggers and logging frameworks. IEEE Software, 23 (3):98–99, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:OSP

[Spi06d] Diomidis Spinellis. Open source and professional advancement. IEEE Software, 23(5):70–71, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:PAP

[Spi06e] Diomidis Spinellis. Project asset portability. IEEE Software, 23(1):100–101, January/February 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:AV

[Spi07a] Diomidis Spinellis. Abstraction and variation. IEEE Software, 24(5):24– 25, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:CSR

[Spi07b] Diomidis Spinellis. Cracking software reuse. IEEE Software, 24(1):12–13, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:S

[Spi07c] Diomidis Spinellis. I spy.

IEEE Software, 24(2):16– 17, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:P

[Spi07d] Diomidis Spinellis. On paper. IEEE Software, 24(6):24–25, November/ December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:SBO

[Spi07e] Diomidis Spinellis. Silver bullets and other mysteries. IEEE Software, 24 (3):22–23, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:TWU

[Spi07f] Diomidis Spinellis. The tools we use. IEEE Software, 24(4):20–21, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2008:RM

[Spi08a] Diomidis Spinellis. Rational metaprogramming. IEEE Software, 25(1): 78–79, January/February 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2008:SB

[Spi08b] Diomidis Spinellis. Software builders. IEEE Software, 25(3):22–23, May/ June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2008:UAX

[Spi08c] Diomidis Spinellis. Using and abusing XML. IEEE Software, 25(2):88– 89, March/April 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2008:WWP

[Spi08d] Diomidis Spinellis. The way we program. IEEE Software, 25(4):89–91, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:BET

[Spi09a] Diomidis Spinellis. Basic etiquette of technical communication. IEEE Software, 26(6):86–87, November/December 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:JS

[Spi09b] Diomidis Spinellis. Job security. IEEE Software, 26(5):14–15, September/ October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:TTD

[Spi09c] Diomidis Spinellis. Tools of the trade: Drawing tools. IEEE Software, 26 (3):12–13, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:TTS

[Spi09d] Diomidis Spinellis. Tools of the trade: Start with the most difficult part. IEEE Software, 26(2):70– 71, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sprague:1995:SLW

[Spr95] Robert D. Sprague. Soft law: At what point does liability hit home? IEEE Software, 12(4):93–94, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sustar:2008:REO

[SPZ08] Helena Sustar, Ulrike Pfeil, and Panayiotis Zaphiris. Requirements elicitation with and for older adults. IEEE Software, 25(3):16– 17, May/June 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Segall:1985:PPI

[SR85] Zary Segall and Larry Rudolph. PIE: a programming and instrumentation environment for parallel processing. IEEE Software, 2(6):22–37, November 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stankovic:1991:SKN

[SR91a] John A. Stankovic and Krithi Ramamritham. The Spring kernel: a new paradigm for real-time systems. IEEE Software, 8(3): 62–72, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Su:1991:ETM

[SR91b] Jason Su and Paul R. Ritter. Experience in testing the Motif interface.

IEEE Software, 8(2):26– 33, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sarkar:2009:MLS

[SRK^{+}09] Santonu Sarkar, Shubha Ramachandran, G. Sathish Kumar, Madhu K. Iyengar, K. Rangarajan, and Saravanan Sivagnanam. Modularization of a large-scale business application: a case study. IEEE Software, 26(2):28–35, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sriram:1986:KBE

[SRS^{+}86] Duvvuru Sriram, Michael D.

Rychener, Stuart C. Shapiro, Sargur N. Srihari, James Geller, Ming-Ruey Taie, Michael J. Pazzani, Anne F. Brindle, Zhigang Xiang, J. L. Rogers, J. F. M. Barthelemy, Paul A. Fishwick, Bob E. Parry, Mona A. Yousry, and Mary Lou Maher. Knowledge — based engineering systems — research in progress. IEEE Software, 3(2):48– 60, March 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sharp:2000:SEC

[SRW00] Helen Sharp, Hugh Robinson, and Mark Woodman. Software engineering: Community and culture. IEEE Software, 17(1):40–47, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s1040abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s1040.pdf.

Slater:1995:LES

[SS95] Paul Slater and Christine Steven. Letters: Evolving a software curriculum. IEEE Software, 12 (3):8, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Siddiqi:1996:REE

[SS96] Jawed Siddiqi and M. Chandra Shekaran. Requirements engineering: The emerging wisdom. IEEE Software, 13(2):15–19, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2004:GEI

[SS04] Diomidis Spinellis and Clemens Szyperski. Guest Editors’ introduction: How is Open Source affecting software development? IEEE Software, 21(1): 28–33, January/February 2004. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
01/s1028.pdf.

Serrano:2006:OSS

[SS06a] Nicolás Serrano and José María Sarriegi. Open source software ERPs: a new alternative for an old need. IEEE Software, 23(3):94– 97, May/June 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sharma:2006:PC

[SS06b] Sharad Sharma and Girish Seshagiri. Point/counterpoint. IEEE Software, 23(5):62– 65, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2008:IHI

[SS08] Forrest Shull and Carolyn Seaman. Inspecting the history of inspections: An example of evidence-based technology diffusion. IEEE Software, 25(1):88–90, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sinha:2006:ECD

[SSC06] Vibha Sinha, Bikram Sengupta, and Satish Chandra. Enabling collaboration in distributed requirements management. IEEE Software, 23(5):52–61, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shapiro:1986:FDS

[SSGT86] S. C. Shapiro, S. N. Srihari, J. Geller, and M.- R. Taie. A fault diagnosis system based on an integrated knowledge base. IEEE Software, 3(2):48– 49, March 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shneiderman:1986:DSP

[SSSW86] Ben Shneiderman, Philip Shafer, Roland Simon, and Linda Weldon. Display strategies for program browsing: Concepts and experiment. IEEE Software, 3(3):7–15, May 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sawyer:1999:CBR

[SSV99] Pete Sawyer, Ian Sommerville, and Stephen Viller. Capturing the benefits of requirements engineering. IEEE Software, 16(2):78–85, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/

so1999/s2078abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s2078.pdf.

Shull:2006:VRB

[SSZ06] Forrest Shull, Carolyn Seaman, and Marvin Zelkowitz. Victor R. Basili’s contributions to software quality. IEEE Software, 23(1): 16–18, January/February 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Staringer:1994:CAR

[Sta94] Werner Staringer. Constructing applications from reusable components. IEEE Software, 11(5):61–68, September 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stallings:1997:LRR

[Sta97] William Stallings. Letters:

Review revisited. IEEE Software, 14(4):14, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4010.
pdf.

Statz:1999:MLY

[Sta99]  Joyce Statz.  Manager:

Leverage your lessons. IEEE Software, 16(2):30– 32, March/April 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2030.
pdf.

Stal:2006:UAP

[Sta06a] Michael Stal. Using architectural patterns and blueprints for serviceoriented architecture. IEEE Software, 23(2):54–61, March/ April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

State:2006:B

[Sta06b] Anthony Akins Radu State. Bookshelf. IEEE Software, 23(6):104–105, November/ December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2006/
06/s6104.pdf.

Sedighi:2005:B

[STBC05] Art Sedighi, Martyn Thomas, Paula M. Bach, and Mark A. Cohen. Bookshelf. IEEE Software, 22(5):115–118, September/October 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
05/s5115.pdf.

Stein:1995:TTA

[Ste95a]  Larry Stein.  Toolbox:

Trends in automating document generation. IEEE Software, 12(5):116–118, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stern:1995:PCS

[Ste95b] Max Stern. Power, comfort, and speed. IEEE Software, 12(4):6, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Steven:2002:PST

[Ste02]  John Steven.  Putting

software terminology to the test. IEEE Software, 19(3):88–89, May/ June 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3088.
pdf;   http://www.computer.
org/software/so2002/s3088abs.
htm.

Sorenson:1988:MSM

[STM88] Paul G. Sorenson, Jean-Paul Tremblay, and Andrew J. McAllister. The Metaview system for many specification environments. IEEE Software, 5(2):30– 38, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stolper:1999:SDA

[Sto99] Steven A. Stolper. Streamlined design approach lands Mars Pathfinder. IEEE Software, 16(5):52– 62, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://computer.org/software/
so1999/s5052abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5052.pdf.

Stone:2003:NVS

[Sto03] Adam Stone. In the news: E-voting: Should we pull the lever? IEEE Software, 20(6):12–14, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6012.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6012abs.htm.

Stroustrup:1988:WOO

[Str88] Bjarne Stroustrup. What is object-oriented programming? IEEE Software, 5 (3):10–20, May 1988. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL https:

//ieeexplore.ieee.org/
abstract/document/2020.

Strigini:1996:QTL

[Str96] Lorenzo Strigini. Quality time — limiting the dangers of intuitive decision making. IEEE Software, 13 (1):101–103, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Strigel:1999:GEI

[Str99] Wolfgang Strigel. Guest Editor’s introduction: What’s the problem: Labor shortage or industry practices? IEEE Software, 16(3):52– 54, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3052.
pdf.

Strigel:2001:GEI

[Str01] Wolfgang Strigel. Guest Editor’s introduction: Reports from the field — using Extreme Programming and other experiences. IEEE Software, 18(6):17–18, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6017abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6017.pdf.

Strigel:2002:CSI

[Str02]  Wolfgang Strigel.  The

Canadian software industry. IEEE Software, 19(4):86–89, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s4086.
                                        pdf.

Strok:2009:RTO

[Str09]  Dale Strok.  Reviewer

thanks: Our 2008 reviewers. IEEE Software, 26 (2):??, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shiratori:1992:UAI

[STSK92] Norio Shiratori, Kaoru Takahashi, Kenji Sugawara, and Tetsuo Kinoshita. Using artificial intelligence in communication system design. IEEE Software, 9(1):38–46, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sutton:2000:RPS

[Sut00] Stanley M. Sutton, Jr. The role of process in a software start-up. IEEE Software, 17(4):33–39, July/ August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4033abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4033.pdf.

Schmid:2002:EIP

[SV02] Klaus Schmid and Martin Verlage. The economic impact of product line adoption and evolution. IEEE Software, 19(4):50– 57, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4050.
pdf;   http://www.computer.
org/software/so2002/s4050abs.
htm.

Stone:2003:NSF

[SV03] Adam Stone and Danna Voth. In the news: Software flaws: To tell or not to tell?; open source in the US government. IEEE Software, 20(1):70–73, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1070.
pdf.

Schneider:2002:EIL

[SvHB02] Kurt Schneider, Jan-Peter von Hunnius, and Victor Basili. Experience in implementing a learning software organization. IEEE Software, 19(3):46– 49, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3046.
pdf;   http://www.computer.
org/software/so2002/s3046abs.
htm.

Singer:2009:PC

[SVSC09] Janice Singer, Mark Vigder, Judith Segal, and Steven Clarke. Point/counterpoint. IEEE Software, 26(5):54– 57, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sangwan:2008:SEC

[SVSL08] Raghvinder S. Sangwan, Pamela Vercellone-Smith, and Phillip A. Laplante. Structural epochs in the complexity of software over time. IEEE Software, 25(4):66–73, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sessions:1999:PCP

[SY99] Roger Sessions and Robert Young. Point counterpoint: Point: a lesson from Palm Pilot. IEEE Software, 16(1):36–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1036.
pdf;   http://dlib.computer.
org/so/books/so1999/pdf/
s1037.pdf.

Symonds:1988:CSE

[Sym88] Andrew J. Symonds. Creating a software-engineering knowledge base. IEEE Software, 5(2):50–56, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tyree:2005:ADD

[TA05] Jeff Tyree and Art Akerman. Architecture decisions: Demystifying architecture. IEEE Software, 22(2):19–27, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tarvo:2009:MSH

[Tar09] Alexander Tarvo. Mining software history to improve software maintenance quality: a case study. IEEE Software, 26(1):34–40, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tausworthe:1995:EMR

[Tau95] Robert C. Tausworthe. Experience more reliable than theory at 5th ISSRE. IEEE Software, 12(2):110–113, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Trammell:1997:LCC

[TB97] Carmen Trammell and Boris Beizer. Letters: Clearing the cleanroom air. IEEE Software, 14(4):10, 12, July/August 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4010.
pdf.

Tucker:2002:PCB

[TB02] Allen Tucker and Barry Boehm. Point/counterpoint: On the balance between theory and practice; software engineering is a value-based contact sport. IEEE Software, 19(5):94– 97, September/October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2002/pdf/
s5094.pdf;   http://www.
computer.org/software/
so2002/s5094abs.htm.

Tsai:1996:DTC

[TBY96] Jeffrey J. P. Tsai, Yao-Dong Bi, and Steve Jennhwa Yang. Debugging for timing constraint violations. IEEE Software, 13(2):89– 99, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thompson:1986:QMS

[TC86] Timothy F. Thompson and William J. Clancey. Qualitative modeling shell for process diagnosis. IEEE Software, 3(2):6–15, March 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tepfenhart:1997:UOT

[TC97] William M. Tepfenhart and James J. Cusick. A unified object topology. IEEE Software, 14(1):31–35, January/February 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Teitelman:1984:TTC

[Tei84] Warren Teitelman. Tour through Cedar. IEEE Software, 1(2):44–52, 54–62, 64–73, April 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tervonen:1996:SQB

[Ter96] Ilkka Tervonen. Support for quality based design and inspection. IEEE Software, 13(1):44–54, January 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Terekhov:2001:RSI

[Ter01] Andrey A. Terekhov. The Russian software industry. IEEE Software, 18(6):98–101, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6098abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6098.pdf.

Thiel:2002:MUP

[TH02a] Steffen Thiel and Andreas Hein. Modeling and using product line variability in automotive systems. IEEE Software, 19(4):66– 72, July/August 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4066.
pdf;   http://www.computer.
org/software/so2002/s4066abs.
htm.

Thomas:2002:SM

[TH02b] Dave Thomas and Andy Hunt. State machines. IEEE Software, 19(6):10– 12, November/December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2002/pdf/
s6010.pdf.

Thomas:2002:MO

[TH02c] Dave Thomas and Andy Hunte. Mock objects. IEEE Software, 19(3):22– 24, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3022.
pdf;   http://www.computer.
org/software/so2002/s3022abs.
htm.

Tveito:2002:RMD

[TH02d] Aase Tveito and Per Hasvold. Requirements in the medical domain: Experiences and prescriptions. IEEE Software, 19(6):66–69, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6066.
pdf.

Thomas:2003:P

[TH03a] Dave Thomas and Andy Hunt. Pipelines. IEEE Software, 20(6):17–19, November/December 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/06/s6017.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/06/
s6017abs.htm.

Thomas:2003:VN

[TH03b] Dave Thomas and Andy Hunt. Verbing the noun. IEEE Software, 20(4):82– 83, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s4082.
                                        pdf.

Thomas:2004:NR

[TH04a] Dave Thomas and Andy Hunt. Nurturing requirements. IEEE Software, 21(2):13–15, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thomas:2004:OSE

[TH04b] Dave Thomas and Andy Hunt. Open Source ecosystems. IEEE Software, 21(4):89–91, July/August 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thomas:2004:P

[TH04c] Dave Thomas and Andy Hunt. Practice. IEEE Software, 21(6):12–13, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thayer:2003:SEGa

[Tha03a] Richard H. Thayer. Software engineering glossary. IEEE Software, 20(3):c3, 85, 88, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s30c3.
pdf.

Thayer:2003:SEGb

[Tha03b] Richard H. Thayer. Software engineering glossary. IEEE Software, 20(4): c3, 53, 56, 93, July/ August 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s40c3.
pdf.

Talby:2006:AST

[THDK06] David Talby, Orit Hazzan, Yael Dubinsky, and Arie Keren. Agile software testing in a largescale project. IEEE Software, 23(4):30–37, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thimbleby:1988:DC

[Thi88] Harold Thimbleby. Delaying commitment. IEEE Software, 5(3):78–86, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thomas:1989:PIS

[Tho89] Ian Thomas. PCTE interfaces: Supporting tools in software-engineering environments. IEEE Software, 6(6):15–23, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thompson:1994:WYN

[Tho94] Jess Thompson. What you need to manage requirements. IEEE Software, 11 (2):115–118, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thompson:1997:TAM

[Tho97]  Jess Thompson.  Tool-

box: Avoiding a middleware muddle. IEEE Software, 14(6):92–95, November/December 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1997/pdf/s6092.
                                        pdf.

Thomsett:1998:MYB

[Tho98] Rob Thomsett. Manager: The Year 2000 bug: a forgotten lesson. IEEE Software, 15(4):91–93, 95, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4091.
pdf.

Thomas:2004:MRM

[Tho04] Dave Thomas. MDA: Revenge of the modelers or UML utopia? IEEE Software, 21(3):15–17, May/ June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thomas:2005:APD

[Tho05]  Dave Thomas.  Ag-

ile programming: Design to accommodate change. IEEE Software, 22(3):14– 16, May/June 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tang:2009:SAD

[THV09] Antony Tang, Jun Han, and Rajesh Vasa. Software architecture design reasoning: a case for improved methodology support. IEEE Software, 26(2): 43–49, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tian:2004:QEM

[Tia04]  Jeff Tian.  Quality-

evaluation models and measurements. IEEE Software, 21(3):84–91, May/ June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tick:1989:CTP

[Tic89] Evan Tick. Comparing two parallel logic-programming architectures. IEEE Software, 6(4):71–80, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tiwana:2004:BBB

[Tiw04] Amrit Tiwana. Beyond the black box: Knowledge overlaps in software outsourcing. IEEE Software, 21(5):51–58, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tondel:2008:SRR

[TJM08] Inger Anne Tøndel, Martin Gilje Jaatun, and Per Håkon Meland. Security requirements for the rest of us: a survey. IEEE Software, 25(1):20–27, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thomas:1989:MEG

[TK89] John C. Thomas and Wendy A. Kellogg. Minimizing ecological gaps in interface design. IEEE Software, 6(1):78–86, January 1989. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Tausworthe:1996:GTS

[TL96] Robert C. Tausworthe and Michael R. Lyu. A generalized technique for simulating software reliability. IEEE Software, 13(2):77– 88, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tan:1998:CPS

[TL98] Hee Beng Kuan Tan and Tok Wang Ling. Correct program slicing of database operations. IEEE Software, 15(2):105–112, March/April 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s2105abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2105.pdf.

Thomas:1994:TVB

[TM94] Martyn Thomas and Frank McGarry. Top-down vs. bottom-up process improvement. IEEE Software, 11(4):12–13, July 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tartalja:1997:CSB

[TM97] Igor Tartalja and Veljko Milutinović. Classifying software-based cache coherence solutions. IEEE Software, 14(3):90–101, May/ June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Torchiano:2004:OAC

[TM04] Marco Torchiano and Maurizio Morisio. Overlooked aspects of COTS-based development. IEEE Software, 21(2):88–93, March/April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thomas:1992:DTI

[TN92] Ian Thomas and Brian A. Nejmeh. Definitions of tool integration for environments. IEEE Software, 9(2):29–35, March 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tockey:1997:MLS

[Toc97] Steve Tockey. A missing link in software engineering. IEEE Software, 14(6):31–36, November/December 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6031.
pdf;   http://dlib.computer.
org/so/books/so1997/pdf/
s6031abs.htm.

Tomijima:1987:HJR

[Tom87] Althea Uhlman Tomijima. How Japan’s recently amended copyright law affects software. IEEE Software, 4(1):17–21, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tomayko:2003:SMM

[Tom03]  Jim Tomayko.  Scien-

tific management meets the personal software process. IEEE Software, 20(2):12– 14, March/April 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s2012.
pdf.

Toth:2006:EOS

[Tot06] Kal Toth. Experiences with open source software engineering tools. IEEE Software, 23(6):44–52, November/December 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tu:1986:FFA

[TP86] Hai-Chen Tu and Alan J.

Perlis. FAC: a functional APL language. IEEE Software, 3(1):36–45, January 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tian:1998:AIR

[TP98] Jeff Tian and Joe Palma. Analyzing and improving reliability: a tree-based approach. IEEE Software, 15(2):97–104, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s2097abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2097.pdf.

Tsai:1988:ISS

[TR88] Jeffrey J.-P. Tsai and Joel C. Ridge. Intelligent support for specifications transformation. IEEE Software, 5(6):28–35, November 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thompson:2005:USE

[TR05] J. Barrie Thompson and Karl Reed. Undergraduate software engineering education: The mark of a discipline. IEEE Software, 22(6):96–97, November/December 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tribble:2002:SS

[Tri02] Alan C. Tribble. Software safety. IEEE Software, 19(4):84–85, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4084.
pdf.

Trubow:1991:HLY

[Tru91] George B. Trubow. How liable are you for your software? IEEE Software, 8 (4):94–95, 101, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Trubow:1994:LRC

[Tru94] George Trubow. Law review: Confronting sleaze on the superhighway; how the law views obscenity. IEEE Software, 11(1):96– 97, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thelin:2003:PUC

[TRW03] Thomas Thelin, Per Runeson, and Claes Wohlin. Prioritized use cases as a vehicle for software inspections. IEEE Software, 20(4):30– 33, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4030.
pdf;   http://www.computer.
org/software/so2003/s4030abs.
htm.

Thomas:2001:USB

[TS01] Gareth C. Thomas and Howard R. Smith. Using structured benchmarking to fast-track CMM process improvement. IEEE Software, 18(5):48–52, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s5048abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5048.pdf.

Tillmann:2006:UTR

[TS06] Nikolai Tillmann and Wolfram Schulte. Unit tests reloaded: Parameterized unit testing with symbolic execution. IEEE Software, 23(4):38–47, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thew:2009:RES

[TSP^{+}09] Sarah Thew, Alistair Sutcliffe, Rob Procter, Oscar de Bruijn, John McNaught, Colin C. Venters, and Iain Buchan. Requirements engineering for E-science: Experiences in epidemiology. IEEE Software, 26(1): 80–87, January/February 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thomas:2002:LFS

[TTT^{+}02] Martyn Thomas, Todd Thiessen, Péter Törok, J. Edralin, Louis Bastarache, and Gary Chatters. Letters: Failure or success?; crossing barriers. IEEE Software, 19(3):8– 12, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3008.
pdf.

Turner:2003:SPA

[Tur03] Richard Turner. Seven pitfalls to avoid in the hunt for best practices. IEEE Software, 20(1):67–69, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1067.
pdf.

Tackett:1999:PCE

[TV99] Buford D. Tackett and Buddy Van Doren. Process control for error-free software: a software success story. IEEE Software, 16(3):24–29, May/ June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s3024abs.htm;

                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s3024.pdf.

Terekhov:2000:RLC

[TV00] Andrey A. Terekhov and Chris Verhoef. The realities of language conversions. IEEE Software, 17(6):111–124, November/ December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2000/s6111abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2000/pdf/
                                        s6111.pdf.

Tratt:2007:GEI

[TW07] Laurence Tratt and Roel Wuyts. Guest Editors’ introduction: Dynamically typed languages. IEEE Software, 24(5):28– 30, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://csdl.computer.
                                        org/comp/mags/so/2007/
                                        05/s5028.pdf.

Tsai:2005:RES

[TYZP05] Wei-Tek Tsai, Lian Yu, Feng Zhu, and Ray Paul. Rapid embedded system testing using verification patterns. IEEE Software, 22(4):68–75, July/August 2005. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Uhl:2003:PCM

[UA03] Axel Uhl and Scott W. Ambler. Point/counterpoint: Model driver architecture is ready for prime time; agile model driven development is good enough. IEEE Software, 20(5):70–73, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5070.
pdf.

Ungar:2007:PC

[UE07] David Ungar and Erik Ernst. Point/counterpoint. IEEE Software, 24(5):72– 75, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Umphress:2002:SPC

[UHC02] David Umphress, T. Dean Hendrix, and James H. Cross. Software process in the classroom: The Capstone Project experience. IEEE Software, 19(5):78–85, September/ October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s5078.
pdf;   http://www.computer.

                                        org/software/so2002/s5078abs.
                                        htm.

Uhl:2008:MDD

[Uhl08] Axel Uhl. Model-driven development in the enterprise. IEEE Software, 25(1): 46–49, January/February 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Upchurch:2001:UCS

[URK01] Linda Upchurch, Gordon Rugg, and Barbara Kitchenham. Using card sorts to elicit Web page quality attributes. IEEE Software, 18(4):84– 89, July/August 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s4084abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s4084.pdf.

Ulkuniemi:2004:CCA

[US04] Pauliina Ulkuniemi and Veikko Seppanen. COTS component acquisition in an emerging market. IEEE Software, 21(6):76–82, November/December 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Umarji:2009:MDA

[US09] Medha Umarji and Forrest Shull. Measuring developers: Aligning perspectives and other best practices. IEEE Software, 26(6):92– 94, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vardi:1988:URD

[Var88] Moshe Y. Vardi. Universalrelation data model for logical independence. IEEE Software, 5(2):80–85, March 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Varadan:1995:ITR

[Var95] Gulyam Varadan. Insider:

Trends in reliability and test strategies. IEEE Software, 12(3):10, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Valaer:1997:CUI

[VB97] Laura A. Valaer and Robert G. Babb, II. Choosing a user interface development tool. IEEE Software, 14(4):29–39, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4029.
pdf;   http://www.computer.
org/software/so1997/s4029abs.
htm.

Varma:2002:BSB

[VBM^{+}02] Tathagat Varma, Fernando Berzal, Nicolás Marín, Tope Omitola, and Matt Edgar. Bookshelf: Strategies for building better people; no bulletproof vest for your system; learning from the past; a treasure for software designers. IEEE Software, 19 (6):101–104, November/ December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6101.

pdf. Review of The Peopleware Papers: Notes on the Human Side of Software by Larry L. Constantine, Fundamentals of Network Security by John E. Canavan, Classic Operating Systems by Per Brinch Hansen, and Software Fundamentals: Collected Papers by David L. Parnas by David M. Weiss and Daniel M. Hoffman, eds.

vanSolingen:1998:IJM

[vBv98] Rini van Solingen, Egon Berghout, and Frank van Latum. Interrupts: Just a minute never is. IEEE Software, 15(5):97–103, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.

org/so/books/so1998/pdf/
s5097.pdf.

Voas:1997:PHB

[VCM^{+}97] Jeffrey Voas, Frank Charron, Gary McGraw, Keith Miller, and Michael Friedman. Predicting how badly “good” software can behave. IEEE Software, 14(4):73–83, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4073.
pdf.

Valet:1996:LBB

[VD96] John Valet and Alan Davis.

Letters: “back to basics” or back to the Dark Ages? IEEE Software, 13(3):14, 16, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanderLinden:2002:SPF

[vdL02] Frank van der Linden. Software product families in Europe: The Esaps & Café Projects. IEEE Software, 19(4):41–49, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s4041.
pdf;   http://www.computer.

                                        org/software/so2002/s4041abs.
                                        htm.

vanderLinden:2009:OSD

[vdLLM09] Frank van der Linden, Björn Lundell, and Pentti Marttiin. Open source development: Commodification of industrial software: a case for open source. IEEE Software, 26(4):77– 83, July/August 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Verner:2005:HSD

[VE05] June M. Verner and William M. Evanco. In-house software development: What project management practices lead to success? IEEE Software, 22(1):86–93, January/February 2005. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vegdahl:1986:MOE

[Veg86] Steven R. Vegdahl. Microcode optimization: Examples and approaches. IEEE Software, 3(4):59– 68, August 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Venugopal:2005:SGS

[Ven05]  C. Venugopal.  Single

goal set: a new paradigm for IT megaproject success. IEEE Software, 22(5):

48–53, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanGenuchten:2008:PC

[vGB08] Michiel van Genuchten and Robert Baillargeon. Point/ counterpoint. IEEE Software, 25(6):80–83, November/December 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanGenuchten:2001:UGS

[vGvDSV01] Michiel van Genuchten, Cor van Dijk, Henk Scholten, and Doug Vogel. Using group support systems for software inspections. IEEE Software, 18(3):60– 65, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s3060abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s3060.pdf.

Vickers:2007:SBP

[Vic07]  Andy Vickers.  Satis-

fying business problems. IEEE Software, 24(3):18– 20, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Visaggio:1994:PIT

[Vis94] Giuseppe Visaggio. Process improvement through data reuse. IEEE Software, 11 (4):76–85, July 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Viswanathan:2008:TRW

[Vis08] Viswa Viswanathan. Tutorial: Rapid Web application development: a Ruby on Rails tutorial. IEEE Software, 25(6):98– 106, November/December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vitter:1984:UNF

[Vit84] Jeffrey Scott Vitter. US&R: a new framework for redoing. IEEE Software, 1 (4):39–52, October 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vodde:2007:LTD

[VK07] Bas Vodde and Lasse Koskela. Learning testdriven development by counting lines. IEEE Software, 24(3):74–79, May/ June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanderLinden:1995:CAB

[vM95a] Frank J. van der Linden and Jurgen K. Muller. Creating architectures with building blocks. IEEE Software, 12(6):51–60, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Voas:1995:STN

[VM95b] Jeffrey M. Voas and Keith W.

Miller. Software testability: The new verification. IEEE Software, 12(3):17– 28, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Voas:2000:GEI

[VM00] Jeffrey Voas and Nancy Mead. Guest Editors’ introduction: Malicious IT. IEEE Software, 17(5):23–25, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s5023.
pdf.

vanZijl:1992:TGN

[vMC92] Lynette van Zijl, Deon Mitton, and Simon Crosby. A tool for graphical network modeling and analysis. IEEE Software, 9 (1):47–54, January 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Voas:1991:PWF

[VMM91] Jeffrey Voas, Larry Morell, and Keith Miller. Predicting where faults can hide from testing. IEEE Software, 8(2):41–48, March 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Viega:2000:SSJ

[VMMF00] John Viega, Tom Mutdosch, Gary McGraw, and Edward W. Felten. Statically scanning Java code: Finding security vulnerabilities. IEEE Software, 17(5):68–74, September/ October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so/s5068abs.htm;   http:
                                        //dlib.computer.org/so/
                                        books/so2000/pdf/s5068.
                                        pdf.

Vestli:1994:MCR

[VNS94] Mette Vestli, Inge Nordbo, and Arne Solvberg. Modeling control in rule-based systems. IEEE Software, 11(2):77–81, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Voth:2004:N

[VNS04] Danna Voth, David Nicholls, and Dale Strok. In the news. IEEE Software, 21(3):107–110, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vo:1990:ITB

[Vo90] Kiem-Phong Vo. Ifs: a tool to build application systems. IEEE Software, 7(4): 29–36, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Voas:1997:HAC

[Voa97a] J. Voas. How assertions can increase test effectiveness. IEEE Software, 14(2): 118–119, 122, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Voas:1997:QTC

[Voa97b] Jeffrey Voas. Quality time:

Can clean pipes produce dirty water? IEEE Software, 14(4):93–95, July/ August 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s4093.
pdf.

Voas:1998:MCB

[Voa98a] Jeffrey Voas. Maintaining component-based systems. IEEE Software, 15(4):22– 27, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
cg1998/s4022abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s4022.pdf.

Voas:1998:MCS

[Voa98b] Jeffrey Voas. Manager: COTS software: The economical choice? IEEE Software, 15(2):16–19, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2016.
pdf.

Voas:1999:CSH

[Voa99a] Jeffrey Voas. Certifying software for high-assurance environments. IEEE Software, 16(4):48–54, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s4048abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s4048.pdf.

Voas:1999:GEI

[Voa99b] Jeffrey Voas. Guest Editor’s introduction: Certification: Reducing the hidden costs of poor quality. IEEE Software, 16(4):22– 25, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s4022.
pdf.

Voas:1999:QTU

[Voa99c] Jeffrey Voas. Quality time:

a US government-regulated software industry. IEEE Software, 16(3):94–102, May/June 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s3094.
pdf.

Voas:1999:QTD

[Voa99d] Jeffrey Voas. Quality time:

Does untested software threaten infrastructures? IEEE Software, 16(2):89– 90, March/April 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s2089.
pdf.

Voas:1999:QTP

[Voa99e] Jeffrey Voas. Quality time:

Protecting against what? the Achilles heel of information assurance. IEEE Software, 16(1):28–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

                                        books/so1999/pdf/s1028.
                                        pdf.

Voas:1999:QTS

[Voa99f]  Jeffrey Voas.  Quality

time: Software quality’s eight greatest myths. IEEE Software, 16(5):118–120, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s5118.pdf.

Voas:2000:QTN

[Voa00a]  Jeffrey Voas.  Quality

time: a new generation of software quality conferences. IEEE Software, 17(1):22–23, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s1022.
                                        pdf.

Voas:2000:QTC

[Voa00b] Jeffrey Voas. Quality time: Can chaotic methods improve software quality predictions? IEEE Software, 17(5):20–22, September/October 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s5020.
                                        pdf.

Voas:2000:QTW

[Voa00c] Jeffrey Voas. Quality time:

Will the real operational profile please stand up? IEEE Software, 17(2):87– 89, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s2087.
pdf.

Voas:2001:QTC

[Voa01a] Jeffrey Voas. Quality time:

Composing software component “ilities”. IEEE Software, 18(4):16–17, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4016.
pdf.

Voas:2001:QTF

[Voa01b] Jeffrey Voas. Quality time:

Faster, better, and cheaper. IEEE Software, 18(3):96– 97, May/June 2001. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s3096.
pdf.

Voas:2001:SFT

[Voa01c] Jeffrey Voas. Software fault tolerance: Making software behave: Guest Editor’s introduction. IEEE Software, 18(4):18–19, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2001/pdf/s4018.
pdf.

Voas:2003:GEI

[Voa03] Jeffrey Voas. Guest Editor’s introduction: Assuring software quality assurance. IEEE Software, 20(3):48–49, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3048.
pdf.

Voas:2004:SER

[Voa04a] Jeffrey Voas. Software engineering’s role in business. IEEE Software, 21(5):26–27, September/ October 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2004/
05/s5026.pdf.

Voas:2004:SSS

[Voa04b] Jeffrey Voas. Software’s secret sauce: The “-ilities”. IEEE Software, 21(6):14– 15, November/December 2004. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

vanSolingen:2004:MRS

[vS04] Rini van Solingen. Measuring the ROI of software process improvement. IEEE Software, 21(3):32– 38, May/June 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanSolingen:2009:FRS

[vS09]  Rini van Solingen.  A

follow-up reflection on software process improvement ROI. IEEE Software, 26(5): 77–79, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Verner:1988:ESE

[VT88] June Verner and Graham Tate. Estimating size and effort in fourth-generation development. IEEE Software, 5(4):15–22, July 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Viega:2000:QTC

[VV00] John Viega and Jeffrey Voas. Quality time: Can aspect-oriented programming lead to more reliable software? IEEE Software, 17(6):19–21, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s6019.
                                        pdf.

vanVliet:2006:RSE

[vV06] Hans van Vliet. Reflections on software engineering education. IEEE Software, 23 (3):55–61, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Verheecke:2006:UCC

[VVJ06] Bart Verheecke, Wim Vanderperren, and Viviane Jonckers. Unraveling crosscutting concerns in Web services middleware. IEEE Software, 23(1):42–50, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ven:2008:SYA

[VVM08] Kris Ven, Jan Verelst, and Herwig Mannaert. Should you adopt open source software? IEEE Software, 25 (3):54–59, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vigder:2008:SSE

[VVS^{+}08] Mark Vigder, Norman G. Vinson, Janice Singer, Darlene Stewart, and Keith Mews. Supporting scientists’ everyday work: Automating scientific workflows. IEEE Software, 25(4):52–58, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanLatum:1998:AGB

[vvSO^{+}98] Frank van Latum, Rini van Solingen, Markku Oivo, Barbara Hoisl, Dieter Rombach, and Günther Ruhe. Adopting GQM-based measurement in an industrial environment. IEEE Software, 15(1):78–86, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1078abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s1078.pdf.

vonWangenheim:2006:HSC

[vWAS06] Christiane Gresse von Wangenheim, Alessandra Anacleto, and Clenio F. Salviano. Helping small companies assess software processes. IEEE Software, 23(1):91–98, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vonWangenheim:2009:EOS

[vWHvW09] Christiane Gresse von Wangenheim, Jean Carlo Rossa Hauck, and Aldo von Wangenheim. Enhancing open source software in alignment with CMMI-DEV. IEEE Software, 26(2):59– 67, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vonWangenheim:2009:VEG

[vWS09] Christiane Gresse von Wangenheim and Forrest Shull. Voice of evidence: To game or not to game? IEEE Software, 26(2):92– 94, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vitvar:2007:SET

[VZM^{+}07] Tomas Vitvar, Michal Zaremba, Matthew Moran, Maciej Zaremba, and Dieter Fensel. SESA: Emerging technology for servicecentric environments. IEEE Software, 24(6):56–67, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Whittaker:2002:SEE

[WA02] James A. Whittaker and Steven Atkin. Software engineering is not enough. IEEE Software, 19(4):108–115, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/

books/so2002/pdf/s4108.
pdf;   http://www.computer.
org/software/so2002/s4108abs.
htm.

Williams:1995:L

[WAGT95] Douglas E. Williams, Tom Adams, Robert Glass, and Jacques Truong. Letters. IEEE Software, 12(2):5– 6, March 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wales:1996:NEA

[Wal96] Elspeth Wales. In the news: ESI aims for competitive edge. IEEE Software, 13(5):104–??, September 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wang:1995:LDR

[Wan95] Rong-Shuang Wang. Letters: Design/requirement dilemma revisited. IEEE Software, 12(6):6, November 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ward:1989:HIO

[War89] Paul T. Ward. How to integrate object orientation with structured analysis and design. IEEE Software, 6(2):74–82, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wasserman:1996:TDS

[Was96] Anthony I. Wasserman. Toward a discipline of software engineering. IEEE Software, 13(6):23–31, November 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Watanabe:1997:IPT

[Wat97] Sadakazu Watanabe. Insider: Professionalism through OO and reuse. IEEE Software, 14(1): 26–??, January/February 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2006:DR

[WB06a] Rebecca Wirfs-Brock. Designing for recovery. IEEE Software, 23(4):11–13, July/ August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2006:CC

[WB06b] Rebecca J. Wirfs-Brock. Characterizing classes. IEEE Software, 23(2):9–11, March/ April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2006:EYD

[WB06c] Rebecca J. Wirfs-Brock. Explaining your design.

IEEE Software, 23(6):96– 98, November/December 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2006:LPA

[WB06d] Rebecca J. Wirfs-Brock.

Looking for powerful abstractions. IEEE Software, 23(1):13–15, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2006:RP

[WB06e] Rebecca J. Wirfs-Brock.

Refreshing patterns. IEEE Software, 23(3):45–47, May/ June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2006:TEH

[WB06f] Rebecca J. Wirfs-Brock.

Toward exception-handling best practices and patterns. IEEE Software, 23(5):11–13, September/ October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2007:DEC

[WB07a] Rebecca J. Wirfs-Brock.

Designing extensible classes. IEEE Software, 24(5):15– 17, September/October 2007. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2007:DBC

[WB07b] Rebecca J. Wirfs-Brock. Does beautiful code imply beautiful design? IEEE Software, 24(6):18–20, November/December 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2007:DLD

[WB07c] Rebecca J. Wirfs-Brock. Driven to … discovering your design values. IEEE Software, 24(1):9–11, January/February 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2007:GDA

[WB07d] Rebecca J. Wirfs-Brock. Giving design advice. IEEE Software, 24(4):13–15, July/ August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2007:HDC

[WB07e] Rebecca J. Wirfs-Brock. Handling design criticism. IEEE Software, 24(3):12– 14, May/June 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2007:TDS

[WB07f] Rebecca J. Wirfs-Brock.

Toward design simplicity. IEEE Software, 24(2):9–11, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2008:CDC

[WB08a] Rebecca J. Wirfs-Brock.

Connecting design with code. IEEE Software, 25(2):20–21, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2008:DDT

[WB08b] Rebecca J. Wirfs-Brock.

Design: Designing then and now. IEEE Software, 25(6):29–31, November/ December 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2008:DS

[WB08c] Rebecca J. Wirfs-Brock.

Design strategy. IEEE Software, 25(3):14–15, May/ June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2008:EC

[WB08d] Rebecca J. Wirfs-Brock.

Enabling change. IEEE Software, 25(5):70–71, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2008:FD

[WB08e] Rebecca J. Wirfs-Brock. Up-front design. IEEE Software, 25(4):12–13, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2008:VDR

[WB08f] Rebecca J. Wirfs-Brock. Valuing design repair. IEEE Software, 25(1): 76–77, January/February 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2009:DCS

[WB09a] Rebecca J. Wirfs-Brock. Design: Creating sustainable designs. IEEE Software, 26(3):5–7, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2009:DDF

[WB09b] Rebecca J. Wirfs-Brock. Design: Designing in the future. IEEE Software, 26(1):18–19, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2009:DDA

[WB09c] Rebecca J. Wirfs-Brock.

Design: Designing with an agile attitude. IEEE Software, 26(2):68–69, March/ April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2009:DT

[WB09d] Rebecca J. Wirfs-Brock.

Design for test. IEEE Software, 26(5):92–93, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2009:DPP

[WB09e] Rebecca J. Wirfs-Brock.

Design: Principles in practice. IEEE Software, 26(4): 11–12, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wirfs-Brock:2009:RD

[WB09f] Rebecca J. Wirfs-Brock.

The responsible designer. IEEE Software, 26(6):9–10, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Winner:1986:AVM

[WC86] Robert I. Winner and Edward M. Carter. Automated vertical migration to dynamic microcode: an overview and example. IEEE Software, 3(4):6–16, August 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wu:1989:STS

[WC89] Chisu Wu and Youseek Chun. Software technology in South Korea: an emerging industry. IEEE Software, 6(2):56–60, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Winn:2002:P

[WC02] Tiffany Winn and Paul Calder. Is this a pattern? IEEE Software, 19(1):59–66, January/February 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2002/s1059abs.htm;
http://dlib.computer.
org/so/books/so2002/pdf/
s1059.pdf.

Wang:2006:PCC

[WC06] Zhaoqing Wang and Harry H. Cheng. Portable C/C++ code for portable XML data. IEEE Software, 23 (1):76–81, January/February 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wu:2009:CCE

[WC09] Harris Wu and Lan Cao.

Community collaboration for ERP implementation. IEEE Software, 26(6):48– 55, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weller:2008:PC

[WCCR08] Ed Weller, David Card, Bill Curtis, and Bob Raczynski. Point/counterpoint. IEEE Software, 25(3):48– 51, May/June 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wolf:1985:ABS

[WCW85] Alexander L. Wolf, Lori A.

Clarke, and Jack C. Wileden. Ada-Based support for programming-in-the-large. IEEE Software, 2(2):58– 71, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wieringa:2004:GEI

[WE04] Roel Wieringa and Christof Ebert. Guest Editors’ introduction: RE’03— Practical Requirements Engineering Solutions. IEEE Software, 21(2):16–18, March/ April 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2004/02/s2016.
pdf.

Weide:2001:LOM

[WEF01] Bruce W. Weide, Matt Edgar, and Martin Fowler. Letters: Our mark on the world; informal technical writing. IEEE Software, 18(6):6, 16, November/ December 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s6006abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s6006.pdf.

Wegner:1984:CIS

[Weg84] Peter Wegner. Capitalintensive software technology. IEEE Software, 1(3): 7–45, July 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weinberg:1998:LTB

[Wei98a] Gerald Weinberg. Letters: Thumbs up to bottomup process improvement. IEEE Software, 15(3):12, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s3010.
pdf.

Weisz:1998:MDU

[Wei98b] Mark Weisz. Manager: Dilbert University. IEEE Software, 15(5):18–21, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5018.
pdf.

Weinberg:1999:EPC

[Wei99] Gerald M. Weinberg. Excerpt from the Psychology of Computer Programming and Silver Anniversary Edition, Egoless Programming. IEEE Software, 16(1):118–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s1118abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s1118.pdf.

Weller:1993:LTY

[Wel93] Edward F. Weller. Lessons from three years of inspection data. IEEE Software, 10(5):38–??, September 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weller:2000:PAS

[Wel00] Edward F. Weller. Practical applications of statistical process control.

IEEE Software, 17(3):48– 55, May/June 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s3048abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s3048.pdf.

Wallin:2002:IBS

[WEL02] Christina Wallin, Fredrik Ekdahl, and Stig Larsson. Integrating business and software development models. IEEE Software, 19(6):28–33, November/December 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s6028.
pdf;   http://www.computer.
org/software/so2002/s6028abs.
htm.

Woodfield:1987:CPR

[WES87] Scott N. Woodfield, David W. Embley, and Del T. Scott. Can programmers reuse software? IEEE Software, 4(4):52–59, July 1987. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wesselius:2008:BIC

[Wes08]  Jacco Wesselius.  The

bazaar inside the cathedral: Business models for internal markets. IEEE Software, 25(3):60–66, May/ June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weyuker:1998:TCB

[Wey98] Elaine J. Weyuker. Testing component-based software: a cautionary tale. IEEE Software, 15(5):54– 59, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so1998/pdf/
s5054.pdf;   http://www.
computer.org/computer/
co1998/s5054abs.htm.

Weyuker:2003:MSY

[Wey03] Elaine Weyuker. I have my state, you have yours. IEEE Software, 20(6):112, November/December 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/dl/mags/so/2003/06/
s6112.pdf;   http://csdl.
computer.org/dl/mags/
so/2003/06/s6112abs.htm.

Wallace:1989:SVV

[WF89] Dolores R. Wallace and Roger U. Fujii. Software verification and validation: an overview. IEEE Software, 6(3):10–17, May 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wright:1986:ESR

[WGFC86] M. Lattimer Wright, Milton W. Green, Gudrun Fiegl, and Perry F. Cross. Expert system for real-time control. IEEE Software, 3(2):16–24, March 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wall:2002:LSP

[WGK^{+}02] David Wall, Michael Guntersdorfer, David G. Kay, Brett Pettichord, John Steven, Mike Llewellyn, Robert Glass, and Roger Jack. Letters: Software patents pro and con. IEEE Software, 19(4):8–11, July/ August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2002/pdf/s4008.
                                        pdf.

Weiser:1985:CST

[WGM85] M. D. Weiser, J. D. Gannon, and P. R. Mc-Mullin. Comparison of structural test coverage metrics. IEEE Software, 2(2):80–85, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wei:2002:DEK

[WHC02] Chih-Ping Wei, Paul Jen-Hwa Hu, and Hung-Huang Chen. Design and evaluation of a knowledge management system. IEEE Software, 19(3):56– 59, May/June 2002. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2002/pdf/s3056.
pdf;   http://www.computer.
org/software/so2002/s3056abs.
htm.

White:2009:IDS

[WHG^{+}09] Jules White, James H. Hill, Jeff Gray, Sumant Tambe, Aniruddha S. Gokhale, and Douglas C. Schmidt. Improving domain-specific language reuse with software product line techniques. IEEE Software, 26(4):47–53, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Whitehouse:1996:NEN

[Whi96] Karen Whitehouse. Naming each needle in the haystack. IEEE Software, 13(4):138, July 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Whittaker:2000:WST

[Whi00] James A. Whittaker. What is software testing? and why is it so hard?: Practice tutorial. IEEE Software, 17(1):70–79, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2000/pdf/s1070.
                                        pdf.

Whittaker:2001:SIU

[Whi01] James A. Whittaker. Software’s invisible users. IEEE Software, 18(3):84–88, May/ June 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s3084abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s3084.pdf.

Wiederhold:1984:KDM

[Wie84] Gio Wiederhold. Knowledge and database management. IEEE Software, 1(1):63–73, January 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wiegers:1998:SRM

[Wie98a] Karl E. Wiegers. Soapbox: Read my lips: No new models! IEEE Software, 15(5):10–13, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1998/pdf/s5010.
                                        pdf.

Wiener:1998:WYL

[Wie98b] Richard Wiener. Watch your language! IEEE Software, 15(3):55–56, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s3055abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s3055.pdf.

Wiegers:1999:SPI

[Wie99] Karl Wiegers. Software process improvement in Web time. IEEE Software, 16(4):78–86, July/ August 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s4078abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s4078.pdf.

Williams:1994:ASC

[Wil94] Lloyd G. Williams. Assessment of safety-critical specifications. IEEE Software, 11(1):51–60, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wilson:1996:BOH

[Wil96] Gregory V. Wilson. Bookshelf: An overview of highperformance computing for the uninitiated. IEEE Software, 13(2):118, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wilson:1997:BPT

[Wil97a]  Greg Wilson.  Book-

shelf: Programming with threads. IEEE Software, 14(5):116, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5115.
pdf.

Wilson:1997:BTP

[Wil97b]  Greg Wilson.  Book-

shelf: Threads primer: a guide to multithreaded programming. IEEE Software, 14(5):116, September/October 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s5115.
pdf.

Wilson:1997:BFY

[Wil97c] Gregory V. Wilson. Bookshelf: First-year C++ text outdated by Java: Programming with Class: A C++ Introduction. the craft of software testing. operating systems. executive guide to preventing IT disasters. managing a programming projectprocesses and people. IEEE Software, 14(3):122–124, May/June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Willerton:1999:ACW

[Wil99a]  Don Willerton.  After

the curtain was pulled away. IEEE Software, 16 (5):114–117, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5114abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5114.pdf.

Wilson:1999:SOS

[Wil99b] Greg Wilson. Soapbox:

Is the open-source community setting a bad example? IEEE Software, 16(1):23–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1999/pdf/s1023.
pdf.

Williams:2003:GEI

[Wil03] Laurie Williams. Guest Editor’s introduction: The XP programmer—the fewminutes programmer. IEEE Software, 20(3):16–20, May/ June 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so2003/pdf/s3016.
                                        pdf.

Wilson:2009:CDS

[Wil09] Greg Wilson. Career development: Not on the shelves. IEEE Software, 26(3):8–9, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wing:1988:SSL

[Win88] Jeannette M. Wing. Study of 12 specifications of the library problem. IEEE Software, 5(4):66–76, July 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Winder:1990:MDT

[Win90a] Ann Winder. Managing data through naming standards. IEEE Software, 7(4): 84–85, July 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wing:1990:SIF

[Win90b] J. M. Wing. A specifier’s introduction to formal methods. Computer, 23(9):8–24, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Winters:2005:UCT

[Win05] Geri Winters. Use case terminology. IEEE Software, 22(2):67, March/ April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

csdl.computer.org/comp/
mags/so/2005/02/s2067.
pdf.

Withrow:1990:EDS

[Wit90] Carol Withrow. Error density and size in Ada software. IEEE Software, 7 (1):26–30, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wells:1997:BWB

[WJS97] Timothy D. Wells, Natalia Juriso, and Nan C. Schaller. Bookshelf: Wonderful book, but a sad commentary; why the dwarf king died; class act in parallel programming. IEEE Software, 14 (6):103–104, 107, November/December 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1997/pdf/s6103.
pdf.

West:2005:B

[WJSP05] Joel West, Angela Jury, Shubhashis Sengupta, and Stratton Penberthy. Bookshelf. IEEE Software, 22 (4):114–118, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/comp/mags/so/2005/
04/s4114.pdf.

Wood:2003:EXS

[WK03] William A. Wood and William L. Kleb. Exploring XP for scientific research. IEEE Software, 20(3):30– 36, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3030.
pdf;   http://www.computer.
org/software/so2003/s3030abs.
htm.

WayneSherer:1996:EUC

[WKA96] S. Wayne Sherer, Ara Kouchakdjian, and Paul G. Arnold. Experience using cleanroom software engineering. IEEE Software, 13 (3):69–76, May 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Williams:2000:SCP

[WKCJ00] Laurie Williams, Robert R. Kessler, Ward Cunningham, and Ron Jeffries. Strengthening the case for pair programming. IEEE Software, 17(4):19– 25, July/August 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s4019abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s4019.pdf.

Wyatt:1992:POO

[WKH92] Barbara B. Wyatt, Krishna Kavi, and Steve Hufnagel. Parallelism in object-oriented languages: a survey. IEEE Software, 9 (6):56–66, November 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wakahara:1989:EES

[WKUI89] Yasushi Wakahara, Yoshiaki Kakuda, Eiji Utsunomiya, and Atsushi Ito. Escort: an environment for specifying communication requirements. IEEE Software, 6(2):38–43, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wakid:1999:TCI

[WKW99] Shukri A. Wakid, D. Richard Kuhn, and Dolores R. Wallace. Toward credible IT testing and certification. IEEE Software, 16(4):39– 47, July/August 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so1999/s4039abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s4039.pdf.

Witten:2001:DOS

[WLC01] Brian Witten, Carl Landwehr, and Michael Caloyannides. Does open source improve system security? IEEE Software, 18(5):57– 61, September/October 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://computer.org/software/
                                        so2001/s5057abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s5057.pdf.

Woollard:2008:SSW

[WMGM08] David Woollard, Nenad Medvidovic, Yolanda Gil, and Chris A. Mattmann. Scientific software as workflows: From discovery to distribution. IEEE Software, 25(4):37–43, July/ August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wilde:1993:MOO

[WMH93] Norman Wilde, Paul Matthews, and Ross Huitt. Maintaining object-oriented software. IEEE Software, 10 (1):75–80, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Woolridge:2007:SRA

[WMH07] Richard W. Woolridge, Denise J. McManus, and Joanne E. Hale. Stakeholder risk assessment: An outcome-based approach. IEEE Software, 24(2):36– 45, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Walker:1987:SGP

[WMWM87] Janet H. Walker, David A.

Moon, Daniel L. Weinreb, and Mike McMahon. Symbolics Genera programming environment. IEEE Software, 4(6):36–45, November 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weyuker:2000:CCP

[WOBP00] Elaine J. Weyuker, Thomas J.

Ostrand, JoAnne Brophy, and Rathna Prasad. Clearing a career path for software testers. IEEE Software, 17(2):76–81, March/ April 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s2076abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s2076.pdf.

Wolfe:1988:MSC

[Wol88] Michael Wolfe. Multiprocessor synchronization for concurrent loops. IEEE Software, 5(1):34–42, January 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wolf:1989:PCT

[Wol89] Wayne Wolf. A practical comparison of two object-oriented languages. IEEE Software, 6(5):61–68, September 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weide:1994:RAE

[WOS94] Bruce W. Weide, William F. Ogden, and Murali Sitaraman. Recasting algorithms to encourage reuse. IEEE Software, 11(5):80– 88, September 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wartik:1986:FRT

[WP86] Steven P. Wartik and Maria H. Penedo. Fillin: a reusable tool for formoriented software. IEEE Software, 3(2):61–69, March 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weidenhaupt:1998:SSD

[WPJH98] Klaus Weidenhaupt, Klaus Pohl, Matthias Jarke, and Peter Haumer. Scenarios in system development: Current practice. IEEE Software, 15(2):34–45, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s2034abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s2034.pdf.

Wiegers:2000:MSP

[WS00] Karl E. Wiegers and Doris C. Sturzenberger. A modular software process mini-assessment method: Tool report. IEEE Software, 17(1):62–69, January/February 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s1062abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s1062.pdf.

Wolf:2009:MTB

[WSD^{+}09] Timo Wolf, Adrian Schröter, Daniela Damian, Lucas D. Panjer, and Thanh H. D. Nguyen. Mining taskbased social networks to explore collaboration in software teams. IEEE Software, 26(1):58–66, January/February 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Warboys:2005:AAA

[WSG^{+}05] Brian Warboys, Bob Snowdon, R. Mark Greenwood, Wykeen Seet, Ian Robertson, Ron Morrison, Dharini Balasubramaniam, Graham Kirby, and Kath Mickan. An activearchitecture approach to COTS integration. IEEE Software, 22(4):20–27, July/ August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wall:1999:LSL

[WSL^{+}99] David A. E. Wall, Roger Sessions, Marty Leisner, Michael Stiefel, Robin F. Goldsmith, and Suresh Peram. Letters: Sounding off on Linux; (re)building the community; real success stories; improving performance by improving design. IEEE Software, 16(3): 8–12, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s3008.
                                        pdf.

Wong:1995:SRC

[WTMS95] Kenny Wong, Scott R. Tilley, Hausi A. Muller, and Margaret-Anne D. Storey. Structural redocumentation: a case study. IEEE Software, 12(1):46– 54, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weis:2003:MM

[WUG03] Torben Weis, Andreas Ulbrich, and Kurt Geihs. Model metamorphosis. IEEE Software, 20(5):46–51, September/October 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/csdl.computer.org/dl/
mags/so/2003/05/s5046.
pdf;   http://csdl.computer.
org/dl/mags/so/2003/05/
s5046abs.htm.

Wang:2001:OSS

[WW01] Huaiqing Wang and Chen Wang. Open source software adoption: a status report. IEEE Software, 18(2):90–95, March/ April 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2001/s2090abs.htm;
http://dlib.computer.
org/so/books/so2001/pdf/
s2090.pdf.

Weber:2003:REA

[WW03] Matthias Weber and Joachim Weisbrod. Requirements engineering in automotive development: Experiences and challenges. IEEE Software, 20(1):16–24, January/February 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s1016.
pdf;   http://www.computer.
org/software/so2003/s1016abs.
htm.

Williams:1997:PRM

[WWD97] Ray C. Williams, Julie A. Walker, and Audrey J. Dorofee. Putting risk management into practice. IEEE Software, 14(3):75– 82, May/June 1997. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Xu:2003:MST

[Xu03]  Jia Xu.  Making soft-

ware timing properties easier to inspect and verify. IEEE Software, 20(4):34– 41, July/August 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s4034.
pdf;   http://www.computer.
org/software/so2003/s4034abs.
htm.

Yacoub:2003:AQD

[Yac03]  Sherif Yacoub.  Auto-

mated QA for document understanding systems.

IEEE Software, 20(3):76– 82, May/June 2003. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2003/pdf/s3076.
pdf;   http://www.computer.
org/software/so2003/s3076abs.
htm.

Yamaura:1998:HDP

[Yam98a] Tsuneo Yamaura. How to design practical test cases. IEEE Software, 15(6):30–??, November/ December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s6030abs.htm;
http://dlib.computer.
org/so/books/so1998/pdf/
s6030.pdf.

Yamaura:1998:SWJ

[Yam98b] Tsuneo Yamaura. Soapbox: Why Johnny can’t test. IEEE Software, 15(2):113–115, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2113.
pdf.

Yamamura:1999:SPI

[Yam99] George Yamamura. Soapbox: Process improvement satisfies employees. IEEE Software, 16(5):83–85, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

                                        /dlib.computer.org/so/
                                        books/so1999/pdf/s5083.
                                        pdf.

Yang:2005:VBP

[YBBP05] Ye Yang, Jesal Bhuta, Barry Boehm, and Daniel N. Port. Value-based processes for COTS-based applications. IEEE Software, 22(4):54–62, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yurcik:2001:AFT

[YD01] William Yurcik and David Doss. Achieving faulttolerant software with rejuvenation and reconfiguration. IEEE Software, 18(4):48–52, July/ August 2001. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        computer.org/software/
                                        so2001/s4048abs.htm;
                                        http://dlib.computer.
                                        org/so/books/so2001/pdf/
                                        s4048.pdf.

Yin:1994:TTI

[Yin94] Sandra Giger Yin. Toolbox: Tool integration through repositories. IEEE Software, 11(1):123–126, January 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yang:2007:UET

[YJ07] Zhihui Yang and Michael Jiang. Using eclipse as a tool-integration platform for software development. IEEE Software, 24(2):87– 89, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yoo:2009:FMV

[YJC09] Junbeom Yoo, Eunkyoung Jee, and Sungdeok (Steve) Cha. Formal modeling and verification of safetycritical software. IEEE Software, 26(3):42–49, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yamazaki:1993:OOD

[YKIY93] Seiichi Yamazaki, Kiyohiko Kajihara, Mitsutaka Ito, and Ryuichi Yasuhara. Object-oriented design of telecommunication software. IEEE Software, 10(1):81–87, January 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yu:1994:CDC

[YM94a] Clement Yu and Weiyi Meng. Confronting database complexities. IEEE Software, 11(3):7–10, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yu:1994:PDS

[YM94b] Clement Yu and Weiyi Meng. Progress in database search strategies. IEEE Software, 11(3):11–19, May 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yeh:1991:CMM

[YNM^{+}91] Raymond Yeh, David Naumann, Roland Mittermeir, Reinhard Schlemmer, William Gilmore, George Sumrall, and John LeBaron. A commonsense management model. IEEE Software, 8(6):23–??, November 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yourdon:1995:MWG

[You95] Edward Yourdon. Manager: When good enough software is best. IEEE Software, 12(3):79–81, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Yourdon:1998:TTF

[You98] Ed Yourdon. A tale of two futures. IEEE Software, 15(1):23–29, January/February 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1998/s1023abs.htm.

Yourdon:2000:SID

[You00] Edward Yourdon. Soapbox: IT departments: Battling the Y2K backlash. IEEE Software, 17(1):100–102, January/ February 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so2000/pdf/s1100.
pdf.

Yourdon:2007:CPA

[You07] Ed Yourdon. Celebrating Peopleware’s 20th anniversary. IEEE Software, 24(5):96–100, September/ October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zadeh:1994:SCF

[Zad94] Lotfi A. Zadeh. Soft computing and fuzzy logic. IEEE Software, 11(6):48– 56, November 1994. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zanger:1992:LRS

[Zan92] Larry M. Zanger. Law review — still no agreement on repossession issues. IEEE Software, 9(3): 97–98, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zanger:1994:LRS

[Zan94] Larry M. Zanger. Law review: Speculating on this year’s issues. IEEE Software, 11(2):88, 91, March 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zave:1989:CAM

[Zav89] Pamela Zave. Compositional approach to multiparadigm programming. IEEE Software, 6(5):15–25, September 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zagal:2002:MOD

[ZAV02] José Pablo Zagal, Raúl Santelices Ahués, and Miguel Nussbaum Voehl. Maintenanceoriented design and development: a case study. IEEE Software, 19(4):100– 106, July/August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.
org/so/books/so2002/pdf/
s4100.pdf;   http://www.
computer.org/software/
so2002/s4100abs.htm.

Zdun:2009:CDK

[Zdu09] Uwe Zdun. Capturing design knowledge: Guest Editor’s introduction: Capturing design knowledge. IEEE Software, 26(2):25– 27, March/April 2009. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zhao:1999:MRC

[ZF99] Liping Zhao and Ted Foster. Modeling roles with Cascade. IEEE Software, 16(5):86–93, September/ October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so1999/s5086abs.htm;
http://dlib.computer.
org/so/books/so1999/pdf/
s5086.pdf.

Zhang:2004:TDM

[Zha04] Yuefeng Zhang. Test-driven modeling for model-driven development. IEEE Software, 21(5):80–86, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zorn:1989:MES

[ZHL^{+}89] Benjamin Zorn, Kinson Ho, James Larus, Luigi Semenzato, and Paul Hilfinger. Multiprocessing extensions in Spur Lisp.

IEEE Software, 6(4):41– 49, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zhou:1989:SRD

[Zho89] Bosheng Zhou. Software research and development in China. IEEE Software, 6(2):53–55, March 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zave:1998:CBA

[ZJ98] Pamela Zave and Michael Jackson. A componentbased approach to telecommunication software. IEEE Software, 15(5):70–78, September/October 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s5070.
pdf;   http://www.computer.
org/computer/co1998/s5070abs.
htm.

Zeiger:2000:LOE

[ZJ00] Paul Zeiger and Joe Jeffrey. Loyal opposition: Ending the Holy War between academia and business. IEEE Software, 17 (1):104, 102–103, January/ February 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/

books/so2000/pdf/s1104.
pdf.

Zhu:1995:SHR

[ZLJW95] Jiang Zhu, Ted G. Lewis, Weldon Jackson, and Russel L. Wilson. Scheduling in hard real-time applications. IEEE Software, 12(3):54– 63, May 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zhong:2000:CFA

[ZME00] Xiamong Zhong, Nazim H.

Madhavji, and Khaled El Emam. Critical factors affecting personal software processes. IEEE Software, 17(6):76–84, November/December 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

computer.org/software/
so2000/s6076abs.htm;
http://dlib.computer.
org/so/books/so2000/pdf/
s6076.pdf.

Zachos:2005:RMS

[ZMT05] Konstantinos Zachos, Neil Maiden, and Amit Tosar. Rich-media scenarios for discovering requirements. IEEE Software, 22(5):89– 97, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zhang:2006:MSR

[ZS06] Yuefeng Zhang and Dhaval Sheth. Mining software repositories for modeldriven development. IEEE Software, 23(1):82–90, January/February 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zavidovique:1991:MCC

[ZSF91] Bertrand Zavidovique, Veronique Serfaty, and Christian Fortunel. Mechanism to capture and communicate image-processing expertise. IEEE Software, 8 (6):37–??, November 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zernik:1992:UVT

[ZSM92] Dror Zernik, Marc Snir, and Dalia Malki. Using visualization tools to understand concurrency. IEEE Software, 9(3):87–92, May 1992. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zvegintsov:1998:CSS

[Zve98a] Nicholas Zvegintsov. Counterpoint: Software should live longer. IEEE Software, 15(4):19, 21, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s4019.
pdf.

Zvegintzov:1998:QTF

[Zve98b] Nicholas Zvegintzov. Quality time: Frequently begged questions and how to answer them. IEEE Software, 15(2):93–96, March/ April 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/
books/so1998/pdf/s2093.
pdf.

Zvegintzov:2003:DWK

[Zve03] Nicholas Zvegintzov. Do we know enough to teach software engineering? IEEE Software, 20(5):112, 110– 111, September/October 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://csdl.computer.
org/dl/mags/so/2003/05/
s5112.pdf.

Zage:1993:EDM

[ZZ93] Wayne M. Zage and Dolores M. Zage. Evaluating design metrics on large-scale software. IEEE Software, 10(4):75–??, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).